### v1

# Containerization in Multi-Cloud Environment: Roles, Strategies, Challenges, and Solutions for Effective Implementation

Abstract.
Containerization in a multi-cloud environment facilitates workload portability and optimized resource utilization. Containerization in multi-cloud environments has received significant attention in recent years both from academic research and industrial development perspectives. However, there exists no effort to systematically investigate the state of research on this topic. The aim of this research is to systematically identify and categorize the multiple aspects of containerization in multi-cloud environment. We conduct the Systematic Mapping Study (SMS) on the literature published between January 2013 and March 2023. Eighty-six studies were finally selected and the key results are: (1) Four leading themes on cloud computing and network systems research were identified: ‘Scalability and High Availability’, ‘Performance and Optimization’, ‘Security and Privacy’, and ‘Multi-Cloud Container Monitoring and Adaptation’. (2) Seventy-four patterns and strategies for containerization in multi-cloud environment were classified across 10 subcategories and 4 categories. (3) Ten quality attributes considered were identified with 47 associated tactics. (4) Four frameworks were introduced from the analysis of identified challenges and solutions: security, automation, deployment, and monitoring challenge-solution frameworks. The results of this SMS will assist researchers and practitioners in pursuing further studies on containerization in multi-cloud environment and developing specialized solutions.
Containerization, Multi-Cloud Environment, Challenges, Solutions, Frameworks, Systematic Mapping Study
journal: TOSEM
1. Introduction
The use of containers in multi-cloud environment has been widespread in the industry for many years (Naik, 2016b). Containers are standalone and executable packages of software that include everything needed to run an application, such as code, system tools, libraries, and settings (Merkel et al., 2014). Containers allow the packaging of an application along with its required dependencies, which makes it easy to move the application across different environments with little or no modification (Docker, 2021). On the other hand, multi-cloud environment is the distribution of cloud assets (Petcu, 2013). Using containers in multi-cloud environment allows organizations to achieve flexibility, agility, and cost efficiency. Developers can build applications in a consistent environment and easily move them between multiple cloud platforms, thereby leveraging the unique strengths (e.g., extensive infrastructure, seamless integration, advanced data analytic) of each platform.
Containerization in multi-cloud is illustrated in Figure 1. This figure provides an overview of applications deployed within various cloud configurations including public, private, and hybrid models. Each cloud hosts multiple applications encapsulated in containers, and completes different types of tasks with their required binaries and libraries, emphasizing the portability and isolation that containerization offers. A container platform layer, which could be exemplified by systems like Kubernetes or Docker, manages and orchestrates these containers across the different cloud environments. This architecture enables a flexible and scalable approach, allowing efficient application deployment and operations in a multi-cloud setting, while maintaining consistency and resilience across platforms.
Figure 1. Context: Containerization in multi-cloud environment
However, utilizing containerization in multi-cloud environment is not without challenges that organizations may face (Saif et al., 2022). These challenges may arise during different phases and activities of container-based application development including architectural design phase, system implementation, or when establishing an automated development infrastructure. Furthermore, the challenges can also present themselves during system testing, the coding process, and the deployment phase. For instance, during the architectural design phase, challenges may involve ensuring seamless integration of containerized components with existing systems and selecting appropriate container orchestration patterns, strategies, and tools that can effectively manage containers across multiple clouds (Raj et al., 2018) (Baby and Vysala, 2015), and while establishing an automated development infrastructure, organizations may encounter difficulties in creating efficient Continuous Integration/Continuous Deployment (CI/CD) pipelines that handle multi-cloud deployment scenarios while maintaining consistent performance and security standards (Debroy et al., 2018). Similarly, In the system implementation phase, challenges may include dealing with potential compatibility issues between containers and various cloud platforms, as well as optimizing resource utilization to control costs effectively (Helali and Omri, 2021).
Recent research (e.g., see Table 8) has highlighted the growing significance of container utilization in multi-cloud environment. These studies have explored various aspects, including container roles and strategies, architectural patterns, Quality Attributes (QAs), and the tools and frameworks employed. Additionally, these studies have explored the challenges associated with automation, deployment, monitoring, and security of containerization applications in multi-cloud environment. Despite the breadth of knowledge available, such valuable information is dispersed across different publications, spanning scientific research papers and the grey literature. This fragmentation of knowledge about containerization applications in multi-cloud environment presents a significant navigational challenge for practitioners tasked with real-world implementations. They need to thoroughly scrutinize numerous aspects, such as patterns and strategies for containerization applications in multi-cloud environment, to locate specific information relevant to their use cases, whether it be a pattern, a challenge, or a solution to an existing problem. Consequently, this spreading of knowledge restricts the development of a holistic understanding of the containerization applications in multi-cloud environment, leaving practitioners underprepared to devise comprehensive solutions.
Furthermore, this state of affairs of containerization in multi-cloud also poses challenges for the academic community. Researchers bear the obligation of navigating through this vast array of information to uncover the precise aspects they seek - be it a unique pattern, an unexplored challenge, or an innovative solution. Moreover, the distributed nature of this knowledge delays the formation of a cohesive and unified understanding of the subject matter. Recent studies (e.g., (Bordeleau et al., 2020) (Mihai et al., 2022) (Sarker, 2022)) have notably highlighted that challenges in design, development, monitoring, and testing of containerization applications in multi-cloud environment are deeply interconnected with the software development life cycle.
To address this knowledge gap, we established the objective of this research is to systematically identify and categorize various facets of container utilization in multi-cloud environment. These facets encompass the roles of containers, their implementation strategies, architectural patterns, quality attributes, as well as the motivations of the customers. Additionally, we aspire to explore the challenges and solutions related to automation, deployment, monitoring, and security of containerization applications in multi-cloud environment. To achieve this objective, we employed a systematic approach to conduct a mapping study. We finally selected 86 studies that focus on containerized applications in multi-cloud environment for further analysis.
The empirical findings and frameworks of this Systematic Mapping Study (SMS) offer significant potential benefits for developers of container-based applications. By exploring a diverse range of implementation strategies, architectural patterns, and quality attributes, practitioners can gain valuable insights to enhance their practices. Moreover, a detailed examination of challenges and solutions related to automation, deployment, monitoring, and security further provides practitioners with the knowledge to address problems effectively and improve dedicated quality attributes. Additionally, these findings and frameworks offer valuable support to practitioners in comprehending various patterns, strategies, and tactics specific to containerized applications in multi-cloud environment. The proposed frameworks can act as a facilitator of knowledge sharing, training, and ensure a shared understanding among practitioners regarding the complexities and solutions within containerized applications in multi-cloud environment. Ultimately, this SMS serves to empower software developers and practitioners, enabling them to optimize their approaches and elevate the quality and efficiency of containerized applications in multi-cloud environment.
Our SMS makes the following key contributions:
  • We developed four frameworks to address challenges and solutions in security, automation, deployment, and monitoring for containerization in multi-cloud environment.
  • We systematically identified and classified an extensive list of patterns and strategies for container-based applications in multi-cloud environment.
  • We systematically identified and classified quality attributes and tactics for containerized applications in multi-cloud environment.
  • We identified and classified a wide range of tools and frameworks for containerized applications in multi-cloud environment.
  • We have publicly released a dataset, which is available online (Waseem et al., 2024), to enable researchers and practitioners to access all collected data, replicate, and validate our SMS. This dataset includes detailed hierarchies of the developed frameworks, identified and classified container implementation strategies, and the roles of containers in multi-cloud environment. It also encompasses patterns, strategies, quality attributes, tactics, tools, and frameworks for containerized-based applications in multi-cloud environment.
The paper is structured as follows: Section 2 describes the research methodology employed. Section 3 presents the results and findings of this SMS. Section 4 discusses the implications of the SMS results. Section 5 clarifies the potential threats to the validity of this SMS. Section 6 reviews relevant prior work and highlights the differences between our work and previous research. Section 7 draws conclusions and outlines future research directions.
2. Methodology
We conducted a Systematic Mapping Study (SMS) by following the guidelines in (Petersen et al., 2008a) and augmenting them with SLR strategies (Keele et al., 2007). Our SMS consists of three phases: specifying research questions and search string, conducting the literature search, and performing data analysis and documentation. Figure 2 illustrates the SMS execution process.
Figure 2. Schematic representation of the research methodology implemented in this study
2.1. Research Questions
To conduct this SMS, we formulated six Research Questions (RQs), as outlined in Table 1, aligning with the objective of our SMS. The table provides a structured view of the RQs, their corresponding rationale, and categorization.
Table 1. Research questions and their rationale
# Research Questions Rationale
Container Roles and Implementation
RQ1
What roles do containers play, and what strategies can be employed for implementing container-based applications in multi-cloud environment?
To investigate the role of containers and explore implementation strategies in multi-cloud environment, with the objective of enhancing understanding and facilitating efficient deployment and management of applications across multiple cloud environments.
RQ2
What architectural patterns and strategies are utilized in the implementation of container-based applications in multi-cloud environment?
To explore and document the various architectural patterns and strategies that are utilized in implementing container-based applications in multi-cloud environment, thus providing a roadmap for efficient and scalable application development and deployment.
RQ3
What are the quality attributes and the tactics associated with their implementation in container-based applications in multi-cloud environment?
To examine and analyze the quality attributes required for container-based applications in multi-cloud environment and the corresponding tactics employed in their implementation, enabling the identification of best practices and guidelines for developing robust and reliable applications in such environments.
Challenges and Solutions
RQ4
What are the challenges related to security, automation, deployment, and monitoring in container-based applications in multi-cloud environment?
To comprehensively evaluate and gain insights into the challenges surrounding security, automation, deployment, and monitoring concerning container-based applications in multi-cloud environment, this question aims to provide valuable insights into the obstacles and intricacies that organizations encounter while implementing and managing these applications in such a dynamic setting.
RQ5
How to address security, automation, deployment, and monitoring in container-based applications within multi-cloud environment?
This question seeks to investigate solutions aimed at tackling the challenges of security, automation, deployment, and monitoring within container-based applications operating in multi-cloud environment. By providing actionable guidance and recommendations, the goal is to elevate the overall management and performance of these applications in the context of multi-cloud environment.
Tools and Frameworks
RQ6
What tools and frameworks are used to implement container-based multi-cloud applications?
To investigate and identify the specific tools and frameworks utilized in the implementation of container-based multi-cloud applications, providing an overview of the technological landscape and aiding developers and organizations in making informed decisions when selecting appropriate tools for their application deployment and management.
2.2. Study Search and Selection Process
The study search and selection process for this study is divided into two phases. In the first phase, we conducted a primary search using a specific search string on various databases (see Table 2 and the Summary sheet in (Waseem et al., 2024)). In the second phase, we applied the snowballing technique on primary studies that were selected from the first phase.
Table 2. Search string and targeted search area for databases
Search String
((multicloud OR multi-cloud OR multi cloud) AND
(container*) AND (automation) And (security) AND (monitoring) AND (deploy*))
Databases
Database Targeted search area
ACM Digital Library Paper title, abstract
IEEE Xplore Paper title, keywords, abstract
Scopus Paper title, keywords, abstract
SpringerLink Paper title, abstract
ScienceDirect Paper title, keywords, abstract
Wiley Online Library Paper title, abstract
Engineering Village Paper title, abstract
Web of Science Paper title, keywords, abstract
2.2.1. Primary Search
The primary search involved querying digital databases (see Table 2) using customized search string. The search strings were executed concurrently on eight databases, as illustrated in Figure 2, spanning between January 2013 and March 2023 when we started this SMS. We chose 2013 as the starting point because it is the year when container technologies like Docker emerged (Docker, 2023), marking a key phase in containerization and cloud computing. This period also saw a rise in relevant research to containers in multi-cloud environment, making it ideal to understand the development of containerized applications in multi-cloud environments over the past decade. We followed the steps mentioned below to select the relevant studies.
Table 3. Inclusion and exclusion criteria for selecting primary studies in this SMS
Selection Criteria Inclusion Criteria Exclusion Criteria
Language
English
Non-English
Study Type
Primary studies including peer-reviewed journal articles, book chapters, conference papers, workshops, and symposium papers
Secondary studies or non-peer-reviewed content (e.g., blogs, webpages, videos, white papers, technical reports, grey literature)
Study Focus
Studies addressing container roles and strategies, architectural patterns, quality attributes, adoption drivers, automation, deployment, monitoring, security challenges, solutions, and tools/frameworks for multi-cloud environment.
Studies that focus solely on traditional cloud computing, single-cloud environments, or unrelated topics
Study Duration
Studies published between January 2013 and March 2023
Studies published before January 2013 or after March 2023
  • Step 1: Extraction of Studies: We ran custom search strings in the selected databases (see Table 2) to retrieve study titles, author names, publication years, venues, publication types, and abstracts. This initial search yielded 5,676 studies from eight databases.Following initial retrieval, we did not apply Cohen’s Kappa as this step is automated and does not involve subjective decision-making.
  • Step 2: Title and Keyword Screening: This step involves reviewing the titles and keywords of the collected studies to assess their relevance to the research topic. Initially, we removed any duplicated studies obtained from different databases (such as IEEE Xplore and Scopus) by sorting them in ascending order. This resulted in the elimination of several hundred duplicate studies. We divided the remaining studies between two authors (R1, R2), who independently applied the inclusion and exclusion criteria to identify studies relevant to our research objectives, as detailed in Table 3. To assess the consistency of study selection and mitigate bias, a Cohen’s Kappa analysis (Cohen, 1968) was performed on a random sample of 40 studies, as illustrated in Formula 1:
    (1) $\kappa=\frac{P_{o}-P_{e}}{1-P_{e}}$
    where $P_{o}$ represents the relative observed agreement among authors, and $P_{e}$ the hypothetical probability of chance agreement. The results, recorded in Table 4.a, indicated a Cohen’s Kappa of 0.531. This value reflects a moderate level of agreement between R1 and R2. Any disagreements were systematically resolved through discussions among all authors, leading to a consensus. Following this process, the number of studies was reduced to 809 for further analysis.
  • Step 3: Abstract-based Screening: During this step, we conducted a thorough review of the abstracts of the collected studies to determine their relevance to our research topic. Two researchers (R1 and R2) independently examined each abstract, assigning a status of ”relevant,” ”irrelevant,” or ”doubtful” with respect to our research goals. To reduce subjective bias and to quantify the consensus between the researchers, we conducted a Cohen’s Kappa analysis on a subset of 30 randomly selected studies. The outcomes of this assessment are depicted in Table 4.b, which show substantial agreement between R1 and R2, with 25 studies being unanimously considered relevant and only 4 studies disagree (3 with one researcher favoring relevance and the other not, and 1 where both disagreed) The Kappa value, calculated based on the observed agreements and disagreements was approximately K= 0.268 which suggests a fair agreement between the two researchers. Consequently, upon the compilation of step, we got 303 studies for more detailed scrutiny.
  • Step 4: Full-Text Screening: In this step, we reviewed the full texts of 303 studies that had been shortlisted through abstract-based screening. To ensure consistency in our review process, we randomly chose 20 studies (see Table 4.c) for a Cohen’s Kappa analysis to measure the agreement between two researchers. The resulting Kappa value was K=0.318, which represents a fair level of agreement. This measure confirms that our selection criteria were applied consistently during the full-text screening studies.
Table 4. Kappa Agreement across different screening steps.
(a) Title and Keyword Screening on 40 Randomly Selected Studies
R1 Yes R1 No
R2 Yes 35 2
R2 No 1 2
Total 36 4
(b) Abstract-based Screening on 30 Randomly Selected Studies
R1 Yes R1 No
R2 Yes 25 3
R2 No 1 1
Total 26 4
(c) Full-Text Screening on 20 Randomly Selected Studies
R1 Yes R1 No
R2 Yes 16 2
R2 No 1 1
Total 17 3
2.2.2. Snowballing
In Phase 2, we utilized the snowballing technique, as described in (Wohlin, 2014), to examine references within primary studies for identifying additional relevant studies. This strategy was augmented by forward snowballing, where we gathered studies that cited the selected studies, and backward snowballing, which involved using references within the selected studies. Notably, we encountered several dozen studies during the forward and backward snowballing that had been excluded in primary search. This phase resulted in the addition of 8 more studies, bringing the final count to 86.
2.3. Data Extraction and Analysis
Data Extraction: The data extraction form was designed based on predefined data items (see Table 5) that were formulated to address the RQs specified in Table 1. To ensure the reliability of the extracted data items, the pilot data extraction was conducted on ten studies by the first author, and all the other authors assessed the extracted data. Subsequently, the first author employed a revised set of data items (e.g., D11, D12, D16), determined after evaluating the extracted data items, for the formal data extraction from the selected studies. To mitigate the personal bias and ambiguity, all authors engaged in discussions regarding the extracted data. The data items labeled as D1 to D11 present a summary of the demographics of the primary studies selected, while D12 to D26 are specifically employed to address RQ1 to RQ6. A concise description of each data item is also presented in Table 5. Finally, Google Sheets were used to record and further analysis of the extracted data.
Data Analysis: We employed a descriptive statistics to analyze the quantitative data from data items D1, D5, D7, D8, and D9. For the remainder of the data, which primarily consist of qualitative free-text descriptions (e.g., study aim, roles of containers, challenges, and solutions), we conducted a thematic analysis in accordance with the guidelines outlined in (Braun and Clarke, 2006). Our thematic analysis process consists of the following steps:
  1. (1)
    Data Familiarization: We conducted a thorough review of the selected studies by repeatedly reading through them and meticulously noting key points related to study aims (D7), contributions (D8), roles of containers (D9), implementation strategies (D10), architectural patterns (D11), quality attributes (D12), tactics (D13), motivations (D14), and challenges and solutions (D15-D21) in automation, deployment, monitoring, and security challenges and solutions, as well as the tools, languages, and frameworks (D22) employed.
  2. (2)
    Generation of Initial Codes: After developing a thorough understanding of the data, we created an initial set of codes derived from the information extracted concerning the data items identified in the previous step.
  3. (3)
    Identification of Types and Emerging Themes: In this step, we conducted a two-tiered analysis. Initially, we examined the codes to ascertain their types. Subsequently, we developed subcategories based on these types, and then formulated overarching categories that encompass the related subcategories.
  4. (4)
    Critical Evaluation of Types, Subcategories, and Categories: All authors actively participated in rigorously reviewing and refining the coded data, including types, subcategories, and categories. During this collaborative process, we redefined, merged, or dropped certain themes based on collective input and discussion.
  5. (5)
    Defining and Naming Categories: At this point, we provided explicit definitions for each of the identified themes and further refined them, ensuring that the terminology used for the categories was precise and unambiguous.
Table 5. Data items to be extracted in this SMS
Code Data Item Description RQ
D1 Index The ID of the study Demographics
D2 Publication Year Publication year of the study
D3 Publisher The publisher of the study
D4 Venue The name of the publishing venue
D5 Publication Type Journal, conference, workshop, and book chapter
D6 Authors’ Affiliation The affiliation of the authors
D7 Study Aim The aim of the paper
D8 Study Contribution The contributions made by the paper
D9 Role of Containers Role of containers in the study RQ1
D10 Container Implementation strategy Strategy used for container implementation
D11 Architectural Pattern and Strategies Architectural patterns and strategies reported in the study RQ2
D12 Quality Attributes Quality attributes discussed in the study RQ3
D13 Tactics Tactics used in the study
D14 Automation Challenges Challenges in automation discussed in the study RQ4
D15 Automation Solutions Solutions for automation challenges discussed in the study RQ5
D16 Deployment Challenges Challenges in deployment discussed in the study RQ4
D17 Deployment Solutions Solutions for deployment challenges discussed in the study RQ5
D18 Monitoring Challenges Challenges in monitoring discussed in the study RQ4
D19 Monitoring Solutions Solutions for monitoring challenges discussed in the study RQ6
D20 Security Challenges Challenges in security discussed in the study RQ4
D21 Security Solutions Solutions for security challenges discussed in the study RQ5
D22 Tools, Languages, and Frameworks Tools, programming languages, and frameworks RQ6
The process of obtaining research themes executed in this SMS is shown in Figure 2. Two researchers participated in the process in order to reduce personal bias. The most important activity of this process is brainstorming sessions that were mainly conducted while reviewing, defining, and naming the research themes. In these sessions, both researchers discussed and validated the research themes found.
Concerning data item D9 (testing approaches) and D10 (testing challenges), we used the open coding and constant comparison techniques from Grounded Theory (Glaser and Strauss, 1967) to analyze the qualitative data extracted from the selected studies.
Finally, we provided a replication package (Waseem et al., 2024) containing the results of each phase of the study selection process (e.g., Phase 1, Phase 2) and detailed results (e.g., Contributions, Patterns, QAs, Challenges, Solutions) for verification and validation purposes of this SMS.
3. Results
This section presents the results of the SMS,derived from analyzing data from chosen studies. Section 3.1 provide a summary of the demographic data, classifications, and the mapping of research themes identified in the studies. Section 3.2 presents the results concerning container roles and implementation strategies. Section 3.3 reports on the patterns and strategies employed in containerized-based applications in multi-cloud environment. Section 3.4 presents the QAs and related tactics for containerized applications in multi-cloud environment. Section 3.5 presents the security challenges and solution framework, while Section 3.6 presents the automation challenges and solution framework. Section 3.7 presents the deployment challenges and solution framework, and Section 3.8 presents the monitoring challenges and solution framework. Lastly, the tools and frameworks are reported in Section 3.9.
3.1. Demographics, Research Aims, and Contributions
Studies Yearly Distribution: Figure 3-a illustrates the annual distribution of studies found in eight databases from January 2013 to March 2023. After applying our selection criteria, the final dataset includes studies published from 2014 to 2022. We found no primary studies on containerization in multi-cloud environments meeting our criteria in the first three months of 2023. The bars displays yearly study counts, with a trend showing the overall direction. A peak in 2017, followed by a decrease and occasional fluctuations, suggests an early spike in interest then more targeted research.
Figure 3. Demographics of the Selected Studies, Including: a) Frequency and Categories of Publications; b) Publisher of the Research; c) Affiliations of the Authors
Publishers Distribution: The pie chart presented in Figure 3-b illustrates the distribution of identified studies among different publishers. The largest portion of the studies, accounting for 72.09%, is published by IEEE, indicating its significant contribution to the field. Scopus follows with 11.62%, while Engineering Village, ACM, and Springer Link represent smaller portions at 6.97%, 5.81%, and 3.48%, respectively. The chart emphasizes the dominance of IEEE in the identified studies, suggesting its importance as a primary source of relevant literature.
Study Types: Figure 3-a provides a breakdown of the types of the selected studies. It is apparent that conferences dominate as the most common venue, comprising a substantial 72.09% of the total publications. Journals, which are often considered as venues for more mature and rigorously peer-reviewed research, make up 19.76%. This indicates a strong preference for disseminating information through conferences, which may be due to the faster publication process and opportunities for networking and collaboration. Symposiums and workshops account for 6.97% and 1.16% respectively, suggesting that these are less common avenues, possibly due to their more specialized or focused nature.
Authors’ Affiliations: Figure 3-c illustrates the distribution of authors’ affiliations in the collected publications. A staggering 93.02% of the affiliations are associated with academia. This overwhelmingly high percentage emphasizes the central role that academic institutions play in the production and dissemination of research in the field. Industry affiliations represent a mere 4.65%, which could imply that industry professionals may be more focused on practical applications and might not publish frequently in the venues considered. Mixed affiliations, at 2.32%, indicate a small proportion of collaborations between academia and industry. This low rate of cross-affiliation collaboration may suggest that there is potential for increased synergy between academic and industrial entities in future research endeavors.
Mapping of Research Themes: The thematic analysis of selected studies aims on containerization in multi-cloud environment, as illustrated in Figure 4, indicating various significant insights.The prominent research themes identified include “Scalability and High Availability” with 20 studies, and three other themes: ’Security and Privacy”m “Multi-Cloud Container Monitoring and Adaptation”, and “Performance and Optimization”, each attracting attention in 15 studies. This highlights that the core areas of interest in multi-cloud environment are proficient container management, scalability, and performance optimization. “Security and Privacy’ also stands as a considerable theme, as demonstrated by the 15 studies focusing on areas such as secure container handling, data security, and blockchain applications. This emphasizes the necessity of protecting data and applications in the context of containers spreading across multi-cloud environment.
Moreover, it is noted that several studies encapsulate more than one theme. For instance, study S19 is catalogued under both ‘Multi-cloud Orchestration and Management’ and ‘Application Deployment in Multi-cloud Container’, signifying that it encompasses aspects of both Kubernetes and microservices architecture.
In essence, the exploration into containerization in multi-cloud environment is substantially inclined towards, scalability, performance optimization security, and orchestration, with several studies showcasing inter-thematic relevance. This is suggestive of the diverse challenges and opportunities present in deploying containers within multi-cloud contexts.
Major Contributions of the Selected Studies: Table 6 provides an overview of the contributions made by the selected studies. The contributions are primarily categorized into five main categories and 22 subcategories (see the Contributions sheet in (Waseem et al., 2024)), through thematic analysis. The selected studies primarily focused on proposing and validating reference architectures, security approaches, cloud management and interoperability frameworks, orchestration and deployment approaches, as well as approaches and tools for optimization and deployment of containerized applications in multi-cloud environment.
The studies related to the reference architecture category mainly contributed to the understanding and design of architectures specifically tailored for multi-cloud containerization. They aimed to address the complexities associated with containerization in multi-cloud environment. On the other hand, the studies in the security approaches category discuss concerns related to cloud analysis, security and performance, cybersecurity, and access control within multi-cloud contexts.
Figure 4. Selected studies research aims in containerization for multi-cloud environment.
The studies on cloud management and interoperability frameworks primarily aim to improve resource and service management, monitoring, and deployment in multi-cloud environment. Their goal is to enhance the performance and efficiency of managing resources and services across multiple clouds. Meanwhile, the research in the area of orchestration and deployment approaches introduces methods for bench-marking, optimizing, and simulating container orchestration in multi-cloud settings. The objective of these studies is to streamline the process of deploying and orchestrating containers within multi-cloud environment.
Lastly, the research on optimization and deployment approaches and tools focuses on improving container deployment and creating efficient management solutions for multi-cloud environment. These studies underscore the importance of optimizing resource allocation and developing practical management tools for use in multi-cloud settings. Overall, these studies contribute to proposing reference architectures, security approaches, cloud management, and interoperability frameworks. These studies also offer strategies and tools for orchestration, deployment operations, and optimization in multi-cloud containerization.
Table 6. Selected Studies’ Contributions on Containerization in Multi-Cloud Environment
Category Subcategory Study #
Reference Architectures and Models for Cloud General Cloud Architectures S26, S33, S35, S42, S76
Fault Tolerance & Self-Healing S12, S55, S67, S74
Migration & Testing Architecture S17, S79
Domain-Specific Language S80, S82
Performance Modeling S28, S62
Optimization & Reliability Model S50, S66
IoT & Service Mesh Architecture S05, S59
Security Approaches Cloud Analysis Approach S18, S24, S37, S65, S69, S86
Security and Performance Solution S10, S23, S38, S75, S77
Cybersecurity & Access Control S11, S30, S34, S71, S73
Security Approach S06, S25, S40, S83
Cloud Management and Interoperability Framework Resource & Service Management Framework S48, S49, S47, S15, S58
Monitoring & Deployment Framework S19, S45, S46, S08
Interoperability Framework S03, S81
Evaluation & Modeling Framework S20, S85
Orchestration and Deployment Approach Benchmarking & Optimization Approach S01, S02, S41, S51, S68
Service Management Approach S07, S14, S29
Simulation & Integration Approach S27, S31, S56
Storage & Services Solution S64, S75
Optimization and Deployment Approaches and Tool Optimization & Scheduling Approach S22, S32, S36, S39
Multi-Cloud Platform S09, S16, S43, S60, S72
Deployment Approach and Tool S21, S52, S57, S61
[…]
3.2. Container Roles and Implementation Strategies (RQ1)
The next two subsections describe the roles and implementation strategiesS related to containers in multi-cloud environment.
3.2.1. Container Roles
Table 7 presents a systematic breakdown of the roles of containers in various aspects of multi-cloud environment. It is organized by the “Category Name” and the “Role of Containers” which detail the specific context or usage of containers in that category.
1. Container-based Application Deployment: This category discusses the use of containers for the deployment of various types of applications. This includes web apps, Platform as a Service (PaaS) solutions, data-intensive applications, and Internet of Things (IoT) apps, among others. Containers provide a means of packaging and distributing applications, including their dependencies, across various environments.
2. Container-based Service Hosting: This category refers to the use of containers as a hosting environment for different services. Container-based service hosting allow for scaling and isolation of these services, creating an efficient multi-cloud environment.
3. Container-Based Resource Management: This category discusses the use of containers to manage computational resources. Containers allow for better multi-tenancy solutions, resource sharing, and reduction in complexity, among other benefits.
4. Container Technology Features: This category discusses the specific features and advantages of container technology, such as lightweight virtualization, use in edge computing, and compatibility with orchestration tools like Kubernetes.
5. Containerization in App Development and IoT: The focus of this category is on the use of containers for app development, particularly in the context of IoT. Containers provide a means of encapsulating software units, increasing portability, improving efficiency, and integration with DevOps practices, among other benefits.
6. Performance and Efficiency through Containers: This category discusses how containers can enhance performance and efficiency in a cloud environment. Containers can improve scalability, isolation, and cost-effectiveness, and also aid in platform optimization.
7. Workload Management via Containers: This category is about the use of containers to manage computational workloads across multiple cloud environments. Containers facilitate the migration and execution of these workloads across different platforms.
8. Security and Compliance via Containers: In this category, containers are discussed as a tool for enhancing security and compliance in multi-cloud environment.
9. Other Container Uses: This category includes various other uses of containers, such as data storage, support in architecture, blockchain reliability, scheduling algorithm evaluation, and virtual network composition.
Each of these categories demonstrates a dedicated aspect of how containers can contribute to efficient and effective multi-cloud management. The studies listed in each row (e.g., S01, S02) provide a more in-depth discussion and evidence for the roles and applications of containers.
Table 7. Roles of containers in multi-cloud environment
Category Role of Containers ID
Container-based Application Deployment Containerized Web-App Deployment S01
Container-Driven App Deployment S05, S08
Containerized PaaS Prototyping S09
Data-Intensive App Deployment via Containers S28
Containerized App Management S29
Container-Based Aneka Service Deployment S51
Container-Driven Microservices Deployment & Scaling S62
Container-Based Multi-Cloud App Deployment & Scaling S74
Containerized IoT App Deployment S85
Container-based Service Hosting Container-Based App Component Hosting S02
Containerized Service Hosting S15
Container-Driven MicroCloud Scaling & Isolation S42
Container-Based Resource Management Container-Based Multi-Tenancy Solution S04
Container-Driven Resource Sharing & Multi-Tenancy S06
Multi-Tenant Service Management via Containers S17
Container-Enabled OS-Level Virtualization S40
Container-Based Complexity Reduction S41
Container Technology Features Container Usage in Edge Computing S22
Container-Enabled Lightweight Virtualization S24
Key Features of Container Technology S25
Container-Based Edge Computing Virtualization S35
Container Leverage in Kubernetes S70
Containerization in App Development & IoT App Development & IoT Integration via Containers S16
Container-Based OpenVAS & Integration of Chef S30
Containers as Standardized Software Units S32
Container Placement & Orchestration in Fog Computing S33
Container Deployment Speed & DevOps Integration S36
Containerizing Apps for Privacy & Compliance S37
Container-Driven Portability & Efficiency Improvement S39
Preventing Vendor Lock-In via Containers S63
Container-Based Extensible App Development S76
Performance & Efficiency via Containers Scalable App Building via Containers S18
Container-Driven SaaS Performance Isolation S23
Cloud Performance Enhancement via Containers S48
Evaluating Performance Impact of Containers S59
Cost Efficiency of Microservices via Containers S68
Container-Based Cloud Platform Optimization S69
Workload Management via Containers Container-Based Virtual Cluster Implementation S21
Multi-Cloud Workload Migration via Containers S60, S65
Seamless Workload Migration via Containers S65
Container-Based Multi-Cloud Platform Execution S79
Other Container Uses Container-Based Azure Data Storage S54
Container & VM Support in Architecture S55
Blockchain-Based Multi-Cloud Reliability via Containers S66
Evaluating Scheduling Algorithm via Containers S67
Container-Based Virtual Network Composition S72
[…]
3.2.2. Implementation Strategies
Table 8 provides an overview of various implementation strategies for container technology in multi-cloud environment. The strategies are grouped into several broad categories, each with a unique focus and varying subcategories. Within these subcategories, each implementation strategy is listed with a concise and revised name for reference.
1. Container Deployment and Management: This category focuses on how containers are deployed and managed in multi-cloud environment. An example from the “Deployment Strategies” subcategory is “Automated Orchestration for Web Applications”, where container deployment is automated for web applications. From the “Container Management” subcategory, the strategy “Container Image Optimization” highlights how the size and efficiency of a container can be optimized.
2. Container Orchestration: In this category, the focus is on coordinating and managing the life cycle of containers in a multi-cloud setup. An instance from the “Cloud Orchestration” subcategory is “SLA-Based Container Strategies”, illustrating how Service Level Agreements (SLAs) can guide the orchestration of containers. In the “Microservices Orchestration” subcategory, “Microservices Architecture in IoT Cloud Apps” represents a way to orchestrate microservices in IoT applications using containers.
3. Container Security: This category focus on the security of containers. An example from the “Container Placement” subcategory is “Secure Placement with Deep Learning”, which employs deep learning techniques to secure the placement of containers. From the “Security Policies” subcategory, the strategy “Access Control and Defense for Container Security” suggests ways to control access and defend containers from security threats.
4. Performance and Scaling: This category addresses the optimization of container performance and the scaling of container-based applications. For example, from the “Performance Optimization” subcategory, “LXC Virtual Cluster for Performance” shows how Linux Containers (LXC) can be used to enhance the performance of virtual clusters. An example from the “Scaling Strategies” subcategory is “Edge Data Streaming and Real-Time Database”, highlighting a strategy for scaling data streaming and real-time databases at the edge.
5. Cloud Service and Networking: This category deals with the management of cloud services and networking in containerized environments. “Multi-Tenant Service with Docker and Cloudant” from the “Cloud Service Management” subcategory illustrates how Docker and Cloudant can be utilized to manage multi-tenant services. From the “Networking” subcategory, the strategy “Edge Scheduling with Location-Awareness” proposes a way to schedule network services at the edge based on location.
6. Data Storage and Persistence: This category discusses strategies related to data management in containerized environments. For instance, the strategy “Firebase Cloud for Real-Time Database” presents an approach that uses Firebase Cloud to manage real-time databases in container-based applications.
[…]
Table 8. Container implementation strategies in multi-cloud environment
Category Subcategory Implementation Strategies ID
Container Deployment and Management Deployment Strategies Automated Orchestration for Web Applications S02
Kubernetes for Containerized Applications S05
Multi-cloud App Development Framework S07
Modular Approach to Container Implementation S08
Deployment with Docker Images S09
Docker Image Creation and Management S31
Container Management Container Image Optimization S32
Efficient Image Reorganization S36
Orchestrating with Docker and Kubernetes S37
Executable Images on Multi-cloud Platforms S79
Container Orchestration Cloud Orchestration SLA-Based Container Strategies S04
Task Division and Scaling in Container Strategy S12
Location-Aware Service Brokering S14
Docker Containers for Multi-tenant Services S17
Resource Acquisition for Aneka Containers S51
Cross-Cloud Container and VM Management S60
Container Orchestration Technologies Overview S33
Microservices Orchestration Microservices Architecture in IoT Cloud Apps S16
ADAPT Deployment of Microservices S27
Microservices Benchmark with Docker and Kubernetes S62
Container Security Container Placement Secure Placement with Deep Learning S06
Container Placement Strategies S39
Security Policies SLA-Oriented Performance Isolation S23
Access Control and Defense for Container Security S25
HIP Integration for Docker Security S77
Secure Connectivity in Multi-tenant Environment S72
Signed URLs for Access Control in Multi-cloud S71
Performance and Scaling Performance Optimization LXC Virtual Cluster for Performance S21
Scalable Multi-layer Architecture with Containers S28
Reactive Auto-scaling and Resource Provisioning S48
Istio for Control Plane Impact and Latency Reduction S59
Scaling Strategies Edge Data Streaming and Real-Time Database S35
MicroCloud Architecture for Efficient Scaling S42
Distributed Workloads with Kubernetes S70
Hybrid Abstractions for Container and Host S40
Cloud Service and Networking Cloud Service Management Multi-Tenant Service with Docker and Cloudant S17
Multi-Cloud Runtime Environment S76
Abstraction Layer for Cloud Service Standardization S63
Networking Edge Scheduling with Location-Awareness S22
Openstack Networking with Kuryr S55
Federated Framework for Cloud Storage S64
Host Identity Protocol for Docker Networking S77
Data Storage and Persistence Data Management Firebase Cloud for Real-Time Database S35
High Availability with SpyStorage MCSS S75
Data Storage with Object Store Service S76
3.3. Pattern and Strategies (RQ2)
Table 9 presents various patterns and strategies employed in containerized-based applications in multi-cloud environment. These patterns and strategies are divided into various categories and subcategories, each of which highlights different aspects of multi-cloud architecture, management, security, resilience, and migration patterns and strategies. Each category is briefly reported below.
1. Cloud Architectural Patterns and Models: This category reports different architecture patterns, communication and networking patterns, deployment patterns, and service models for containerized-based applications in multi-cloud environment. Architecture Patterns involve several design models for structuring the components of a system, providing a template that guides system designers in building complex systems. The leading patterns identified are microservice architecture and service-oriented architecture (SOA), both of which focus on building applications as suites of independently deployable services. Communication and Networking Patterns relate to how different services within the architecture communicate and network with each other. Service mesh and Service Chaining (SC) are the main patterns observed. Deployment Patterns pertain to different strategies for deploying software systems in multi-cloud environment. The two leading patterns are Blue-Green Deployment and Object Store Service. Service Models involve different ways of providing cloud services, with Platform-as-a-Service (PaaS) and Software-as-a-Servic (SaaS) leading the pack.
2. Cloud Management and Resource Allocation Strategies: This category presents the patterns and strategies related to the management of multi-cloud environment, container management, edge computing, IoT strategies, and resource management. For example, Multi-Cloud Management Strategies include various approaches for managing multiple cloud environments. Multi-Cloud Computing Strategy and Hybrid Cloud Architecture are found leading strategies in this category. Container Management Strategies involve the management of containers in multi-cloud environment, with the Linux Container (LXC) project and Container Engine as the primary patterns identified. Edge Computing and IoT Strategies pertain to the use of edge computing and IoT in multi-cloud environment. Mobile Edge Computing (MEC) and Fog Computing are leading strategies for container management in multi-cloud environment. Resource Management Strategies involve methods for managing resources in the multi-cloud with the help of Kubernetes and Minikube.
3. Cloud Security and Resilience Strategies: In this category, the emphasis is on strategies for ensuring the security and resilience of software systems in multi-cloud environment. Security and Resiliency Patterns are strategies to ensure the system is secure and can recover from failures. The two most prominent patterns are the security-by-design approach and encryption of files. Fault-tolerance Strategies are methods used to ensure that systems can continue to function even when there are failures. Fault-tolerance with Redundant Engines and Multi-cloud Systems Fault-Tolerant Workflow are the leading patterns.
4. Cloud Migration Strategies: This category involves patterns and strategies for moving applications and data between cloud environments and optimizing the migration process. Image-based Migration, for example, is an efficient way to migrate applications because it packages the applications along with their dependencies, configurations, and runtime environment. Migration to a virtualized container, such as Docker, is increasingly popular because it guarantees application portability across various environments and scalability on different cloud platforms.
Table 9. Patterns and strategies for container-based applications a multi-cloud environment
Categories Subcategories
Patterns
ID
Cloud Architectural Patterns and Models Architecture Patterns
Microservice Architecture
S02, S09, S16, S77, S80, S24, S31, S33, S68
Service-Oriented Architecture (SOA)
S51, S52, S86
Multitier Architecture, Layered Architecture
S02, S04, S54
Client-Server Architecture
S03, S54
Component-based Architecture
S22
Event-driven Architecture
S44
Multi-agent Architecture
S74, S48
Master-worker Nodes Architecture
S33
Netflix Zuul-based Seeker Component
S17
Multi-tenant Cloud Service Architecture
S17
Stateful Engine Architecture
S17
Communication and Networking Patterns
Service Mesh
S05
Service Chaining (SC)
S50
Multi-Cloud APIs
S71
Third-Party APIs
S71
Service Request Broker
S07
Publish and Subscribe Communication Protocol
S45
Synchronous Network Communication Protocols
S01
Asynchronous Network Communication Protocols
S01
Network Function Virtualization (NFV)
S50
Service Discovery
S76
Deployment Pattern
Blue-Green deployment
S27, S30
Object store service
S76
Multi-cloud Deployment Model
S10
Distributed Deployment Model
S10
Accelerated Particle Swarm Optimization for Deployment
S16
Service Models
Platform-as-a-Service (PaaS)
S24, S41
Software as a Service
S79
Infrastructure-as-a-Service (IaaS)
S24
Function-as-a-Service (FaaS)
S41
Cloud Management and Resource Allocation Strategies Multi-Cloud Management Strategies
Multi-cloud computing strategy
S78, S81
Hybrid cloud architecture
S78
Hybrid/Multi-cloud Approach
S05, S83
Introducing multi-cloud Middleware
S52, S53
MAPE-K control loop
S46, S48
Connecting to multiple cloud service providers
S76
Multi-cloud load balancing
S23, S76
Multi-Cloud BP provisioning
S84
Cross-level orchestration of cloud services
S84
Cross-level monitoring and adaptation of BPs
S84
Container Management Strategies
Linux Container (LXC) project
S24
Container engine
S24
Docker Container Images
S24
Lightweight virtualization
S24
Portable application packaging
S24
One-container-per-app approach
S24
Data volumes and data volume containers
S24
Edge Computing and IoT Strategies
Mobile Edge Computing (MEC)
S22, S70, S43, S54
Fog computing
S32
Edge Services
S35
Connecting IoT Edge Devices to the Cluster
S05
Cloud of Things (CoT)
S05
Resource Management Strategies
Resource allocation based on different SLA levels
S23
Container Orchestration, Kubernetes, Minikube
S24, S29, S33
Cloud Security and Resilience Strategies Security and Resiliency Patterns
Security-by-Design approach
S51
Encryption of files
S71
Standardized APIs for file transfer
S71
Centralized/External ACM Service
S71
Distributed ACM Service
S71
Dynamic Switching between Authentication Methods
S71
ACM based on Signed URLs
S71
Attribute-Based Encryption and Signature
S75
Secure data sharing
S75
Data access control
S75
Local encryption and signing
S75
Authorized Tokens
S75
Byzantine quorum protocol
S75
Fault-tolerance Strategies
Fault-tolerance with Redundant Engines
S17
Multi-cloud systems fault-tolerant workflow
S67
RAFT Consensus Algorithm
S13
Cloud Migration Strategies
Service-oriented Migration
S15
Application-centric Migration
S15
Image-based Migration
S15
Migration to a Virtualized Container
S15
[…]
3.4. Quality Attributes and Tactics (RQ3)
Table 10 provides a comprehensive overview of the quality attributes and related tactics for containerized applications in multi-cloud environment. It is based on data extracted from 86 selected studies, which are organized following the ISO 25010 standard. This standard defines high-level QAs along with their sub-characteristics. Table 10 lists nine QAs, namely Performance (Efficiency), Security, Compatibility, Scalability, Reliability, Portability, Flexibility, Maintainability, and Usability. Each QA has its related terms, associated study numbers, and corresponding tactics to achieve them.
In Table 10, five tactics are provided for each quality attribute, due to the limited space. The complete set of 195 identified tactics is available in the provided dataset (Waseem et al., 2024), presenting a comprehensive map of tactics with QAs. However, it is important to note that these tactics are scattered across different studies, and only a few studies explicitly mentioned these tactics for specific QA.
Below we briefly discuss how each QA can be achieved by applying the mentioned tactics.
Performance (Efficiency): The quality attribute of performance (or efficiency) in containerized applications can be achieved by employing tactics such as efficient resource optimization and allocation, use of ML techniques for performance optimization, utilization of container technology, location-aware service brokering, and performance-oriented Service Level Agreements (SLA). For instance, optimizing resources ensures that the system uses the minimum possible resources while delivering the required output. Using ML techniques for performance optimization can enhance system performance by learning and adjusting the operational parameters. Similarly, utilizing container technology can help maintain optimal performance by encapsulating the application and its environment.
Security: Security can be enhanced using tactics like encryption and strong authentication mechanisms, implementation of firewalls, VPNs, or SDNs for network security, utilization of ML techniques to detect and prevent security threats and attacks, deployment of applications on diverse cloud providers, and consideration of users’ security specifications. For example, encryption provides a secure way of transmitting data, and ML can detect unusual behavior that could signal a threat or attack.
Compatibility: The tactics that can be employed to achieve compatibility include implementation of lightweight communication protocols and modes, standardization of interfaces for seamless integration, componentization and modular design for interoperability, prototyping and exploration of interoperability approaches for multi-cloud deployment, and interoperability standardization and federation between clouds. Standardizing interfaces can ensure that different software components can interact with each other seamlessly, enhancing compatibility.
Scalability: Scalability can be achieved through tactics like combining containerization and microservices for enhanced scalability, implementation of elastic resource allocation, enabling migration between multi-cloud services, combining container-based deployment with runtime monitoring and optimization, and leveraging orchestration, federated networks, and geographic placement. Elastic resource allocation ensures that the system can seamlessly scale up or down in response to changing demands.
Reliability: Tactics such as deploying redundant engines for fault tolerance, distributing resources and replicating applications for improved response time, enforcing redundancy and distributed services for availability, deploying parallel search and multi-cloud distribution, and building fault-tolerant systems can enhance system reliability.
Portability: Portability can be achieved using tactics such as virtual machine-based packaging, Docker image-based packaging, cloud-portable containerization, dynamic cross-level adaptation and provisioning, and adaptive rule-based system modification. These tactics ensure that an application can be easily transferred from one computing environment to another.
Flexibility: Flexibility can be improved through tactics like service-oriented architecture, on-demand dynamic allocation of resources across different cloud platforms, containers, container orchestration, model-driven development, risk analysis, and cross-cloud service orchestration. Using multi-cloud environment, for instance, offers the flexibility to choose services from different providers as per specific needs.
Maintainability: Tactics like microservices for easier maintenance, Infrastructure Provisioning as Code (IaC), containerization and image management, configuration management and templating, version control and change management, and continuous integration and deployment automation can enhance maintainability. For example, IaC allows developers to manage infrastructure more efficiently and minimize human errors, thus increasing maintainability.
Usability: Usability can be enhanced by employing tactics like a feedback-loop controller for multi-cloud infrastructure, container orchestration (e.g., Kubernetes), user-centric interface design, responsive and adaptive user experience, accessible design and compliance, and error handling and feedback mechanisms. For instance, a feedback-loop controller is effective in a multi-cloud context as it dynamically adjusts resources across different cloud platforms based on user interaction, ensuring a seamless and responsive user experience tailored to the multi-cloud setup.
[…]
Table 10. Quality attributes and tactics for containerized applications in multi-cloud environment
Quality Attribute Related Terms ID Tactics
Performance (Efficiency) Time behavior, Resource utilization, Capacity, Throughput, Response time S01, S02, S03, S06, S09, S14, S15, S16, S20, S21, S23, S25, S27, S29, S34, S35, S36, S37, S38, S39, S41, S43, S44, S47, S49, S52, S54, S57, S58, S59, S60, S62, S64, S65, S66 Efficient resource optimization and allocation
Utilizing machine learning techniques for performance optimization
Container technology utilization
Location-aware service brokering
Performance-oriented Service Level Agreements (SLA)
Security Confidentiality, Authenticity, Access control, Authorization, Privacy, Trustworthiness, Integrity S01,S05, S06, S09, S15, S23, S24, S25, S29, S34, S37, S38, S40, S44, S47, S52, S54, S55, S57, S59, S61, S64, S65, S66, S69, S71, S72, S73, S77, S86 Encryption and strong authentication mechanisms
Implementation of firewalls, VPNs, or SDNs for network security
Utilizing machine learning techniques to detect and
prevent security threats and attacks
Deployment of cloud applications on diverse cloud providers
Consideration of users’ security specifications and
addressing CSP incompatibilities
Compatibility Interoperability, Co-existence S15, S16, S21, S24, S27, S42,S44, S45, S49, S63, S64, S71, S77, S80, S82 Implementing lightweight communication protocols and modes
Standardization of interfaces for seamless integration
Componentization and modular design for interoperability
Prototyping and exploring interoperability approaches for
multi-cloud deployment
Interoperability standardization and federation between clouds
Scalability Capacity, Extensibility, Elasticity, Throughput, Responsiveness S01, S08, S09, S14, S15, S16, S17, S18,S19, S21, S22, S24, S25, S27, S29, S31, S35, S41, S42, S44, S45, S54, S57, S62, S66, S69, S72, S74, S76, S77, S82, S84
Combining Containerization and Microservices for enhanced scalability
Implementing Elastic Resource Allocation
Enabling Migration between Multi-Cloud Services
Combining Container-based Deployment with Run-time Monitoring
and Optimization
Leveraging Orchestration, Federated Network,and Geographic Placement
Reliability Maturity, Availability, Fault tolerance, Recoverability S01, S02, S04, S10, S13, S17,S19, S20, S22, S25, S27, S32, S33, S35, S37, S44, S47, S51,S53, S54, S55, S56, S57, S60, S64, S65, S66, S67, S71, S73, S74, S75, S82 Deploying Redundant Engines for Fault-Tolerance
Distributing Resources and Replicating Applications for Improved
Response Time
Enforcing Redundancy and Distributed Service for Availability
Deploying Parallel Search and Multi-Cloud Distribution
Building Fault-Tolerant Systems
Portability Installability, Replaceability, Adaptability S18, S24, S25, S31, S62, S77 Virtual Machine-Based Packaging
Docker Image-Based Packaging
Cloud-Portable Containerization
Dynamic Cross-Level Adaptation and Provisioning
Adaptive Rule-Based System Modification
Flexibility Functional completeness, Functional correctness, Functional appropriateness S16, S21, S27, S31, S41, S61, S60, S70, S76, S84 Service-oriented Architecture
Dynamic allocation of resources across different cloud platforms based on demand
Container orchestration
Model-Driven Development, Risk Analysis
Cross-cloud service orchestration,
Maintainability Modularity, Reusability, Analyzability, Modfiabality, Testability S08, S17, S25, S27, S62, S74, S82 Microservices, easier maintenance
Infrastructure Provisioning as Code (IaC)
Containerization and Image Management
Configuration Management and Templating
Version Control and Change Management
Continuous Integration and Deployment Automation
Usability Appropriateness recognizability, Learnability, Operability , User interface aesthetics, Accessibility S09, S17, S27, S27, S37, S74 Feedback-loop controller for multi-cloud Infrastructure
Container Orchestration (e.g., Kubernetes)
User-Centric Interface Design
Responsive and Adaptive User Experience
Accessible Design and Compliance
Error Handling and Feedback Mechanisms
3.5. Security Challenges and Solution Framework (RQ4 and RQ5)
In an era of rapid technological advancement, ensuring data security is of paramount importance. In this section, we propose a framework designed to address the security challenges in containerized and multi-cloud environment. This proposed framework, composed of seven categories of security challenges and solutions illustrated in Table 11, is derived through a comprehensive review of existing literature. It covers a broad spectrum of categories, each capturing specific security challenges and their corresponding solutions. Each category is briefly described below:
Data Security in container-based applications category covers a wide range of challenges and solutions, including data protection, database security, data compliance, and secure data transfer. Our identified solutions revolve around the use of encryption and secure container orchestration, complying with regulatory requirements such as GDPR, HIPAA, and secure data management strategies. Furthermore, identified solutions emphasizes the role of security measures such as access control and secure transmission protocols in ensuring data privacy, service continuity, fault tolerance, and the secure transfer of data between different cloud environments.
Access and Communication Control category emphasizes maintaining secure multi-cloud connectivity and managing user access control. The identified solutions involve implementing encryption and authentication mechanisms, using secure communication protocols, and executing operations in a provider-agnostic manner to maintain security.
Container Security category addresses issues like co-resident security, container isolation, and providing host security. Implementing secure placement strategies for containers across multi-cloud platforms, utilizing container isolation techniques, and enhancing manual defenses to mitigate risks of compromised containers are some of the identified solutions in this category.
Infrastructure and Deployment Management Security category emphasizes the significance of infrastructure and deployment management security. It encompasses aspects such as secure data deployment, infrastructure orchestration, storage security, and migration security. Essential solutions encompass the application of encryption techniques, implementation of access control mechanisms, adoption of standardized and secure infrastructure orchestration practices, and ensuring the secure migration of containerized applications across multi-cloud environment.
Security Monitoring and Breach Prevention category addresses a range of challenges and solutions related to enhancing security. This includes the implementation of automated security assessments, the use of vulnerability scanners, and the establishment of automated bootstrapping mechanisms for Virtual Machines (VMs) that incorporate security configurations and policies. Additionally, it involves deploying Security Information and Event Management (SIEM) systems, as well as Intrusion Detection System/Intrusion Prevention System (IDS/IPS) tools. The category also covers the application of log analysis solutions to proactively detect and prevent security breaches.
Trust and Compliance Management category identified solutions encompass on implementation of robust security controls, adherence to industry best practices and standards, and establishment of trust-based coalitions to foster secure community formation.
Placement Strategy: It recommends devising a security model for services hosted on volunteering fogs within multi-cloud environment, including formulation of secure communication protocols and implementation of access control mechanisms.
[…]
Table 11. Security challenge and solution framework for containerized applications in multi-cloud environment
ID Challenge Challenge Description Solution Solution Description
Category 1: Data Security
S17 Data Protection Protecting data against threats in containerized multi-cloud setups. Multi-Cloud Data Protection through Secure Containers Deploys sensitive data in encrypted, isolated containers, using secure orchestration tools like Docker and Kubernetes.
S29 Database Security Securing databases within containerized multi-cloud environments. Enhanced multi-cloud Database Security via Containerization Secures databases by implementing encryption (e.g., Transparent Data Encryption) and robust access controls (e.g., RBAC, ABAC).
S37 Data Compliance Ensuring data processing and storage compliance across multi-cloud platforms. Multi-cloud Legal Compliance and Data Protection Maintains privacy and service quality by adhering to regulations (e.g., GDPR, HIPAA) and employing data protection strategies like encryption and audits.
S38 Data Security Adhering to data security standards in multi-cloud containerized systems. Secure CSP Interoperability and Selection in multi-cloud Tackles CSP compatibility issues and enhances security by implementing data handling practices and selecting compliant CSPs.
S53 Data Ownership and Privacy Tackles data ownership, privacy, and availability issues in multi-cloud setups. Multi-cloud Security via NoMISHAP Service Abstraction Uses NoMISHAP for service abstraction, enhancing privacy with encryption and pseudonymization, and ensuring availability with fault tolerance strategies like backups and failovers.
S58 Data Transfer Security Secures data during transfers in multi-cloud environments. Secure multi-cloud Integration via OpenStack Standardization Standardizes OpenStack APIs to prevent vendor lock-in, ensuring safe data transfers with encryption and secure protocols.
S85 Cloud Data Management Manages cloud data security, privacy, and compliance. Multi-cloud Data Security via DRA Framework Implements the DRA framework for comprehensive data management, ensuring encryption, access control, and compliance with standards.
Category 2: Access and Communication Control
S05 Connectivity and Access Control Ensures secure multi-cloud connectivity and robust user access control. Multi-cloud Containerized Application Security Implements TLS, JWT, OAuth 2.0 for secure authentication and authorization in multi-cloud setups.
S43 Secure Communication Secures communications across multi-cloud and MEC platforms. Middleware Secure Deployment for multi-cloud Uses middleware for secure cloud app deployment, with SSL/TLS and access control for secure communication.
S51 Provider-Agnostic Operations Facilitates secure, provider-agnostic operations in multi-cloud environments. Aneka-Based Security and Performance for Containerized multi-cloud Applications Employs Aneka for security and performance, offering transparent access and dynamic provider selection, while ensuring security via secure channels and monitoring.
Category 3: Container Security
S06 Co-resident Security Addresses security risks from container virtualization in multi-cloud environments. Secure Placement Strategy via Deep RL Uses deep reinforcement learning, like Kubernetes Cluster Autoscaler with RL, for secure container placement across multi-cloud platforms.
S25 Container Isolation Aims to prevent information leaks and power attacks through effective container isolation. Power and Leakage Management for multi-cloud Security Manages security via container isolation (e.g., Docker isolation) and Linux kernel-level resource isolation (e.g., SELinux).
S40 Container and Host Security Enhances security against compromised containers and strengthens defenses. Host-Container Cooperation for multi-cloud Defense Develops container-host defense strategies, utilizing tools like Falco, Sysdig, and practices for host hardening.
Category 4: Infrastructure and Deployment Management Security
S20 Data Deployment Secures private data deployment in multi-cloud environments. Secure Private Data Deployment Strategy Utilizes encryption, RBAC, and secure protocols (e.g., AES-256, SFTP, HTTPS) for safeguarding data deployment.
S44 Infrastructure Orchestration Orchestrates secure infrastructure for containerized applications. Standard Interface Implementation for Serverless Security Uses standard interfaces (e.g., OpenStack APIs) for secure, interoperable serverless application deployment across clouds.
S47 Storage Security Enhances security for multi-cloud storage in containerized setups. Multi-cloud Storage Security with MSSF Employs the MSSF for secure storage management, implementing RBAC, data encryption, and regular audits.
S80 Migration Security Ensures the secure migration of applications in multi-cloud environments. Infrastructure-Aware and Agnostic Secure Migration Leverages strategies and tools for safe migration, including image signing and data encryption, adhering to best practices like least privilege.
Category 5: Security Monitoring and Breach Prevention
S30 System Vulnerabilities Addresses vulnerabilities and patching delays in multi-cloud environments. Automated Security Evaluation and Bootstrapping Utilizes tools like OpenVAS and Nessus for security assessments, and automates VM setup with Terraform, Chef, Ansible, or Puppet.
S45 Breach Detection and Prevention Boosts breach detection and prevention in multi-cloud infrastructures. Metric Filtering and Fault Recovery for multi-cloud Security Enhances security with metric filtering, fault recovery, SIEM, IDS/IPS, and log analysis for proactive breach management.
Category 6: Trust and Compliance Management
S52 Application Security and Compliance Ensures security and compliance for applications across multi-clouds. Risk Analysis and Security SLAs with MUSA Framework Uses MUSA for risk analysis and security SLA management, incorporating secure coding, vulnerability management, and compliance audits.
S49 Data Security and Standardization Standardizes data security and cloud interfaces in multi-cloud environments. OpenStack API-based multi-cloud Security and Interoperability Utilizes OpenStack API for secure, standardized cloud interoperability, applying encryption, access control, and CIS benchmarks.
S57 Network Security Secures networks and trust in multi-cloud environments. Inter-cloud Communication Protection with SFC Employs Service Function Chaining (SFC) for secure inter-cloud communications, using encryption, integrity checks, and protocols like IPSec, WireGuard.
S83 Trust and Community Formation Builds secure communities and trust in multi-cloud settings. Trust-based Hedonic Coalitions for multi-cloud Security Forms trust-based coalitions for community security, leveraging IAM and trust frameworks for collaboration in multi-clouds.
S84 Orchestration and Adaptation Manages orchestration and adaptation for security and performance. Multi-cloud Service Orchestration and Cross-Level Adaptation Implements multi-cloud orchestration with tools like Kubernetes, Istio, and Prometheus for secure, adaptive application management.
Category 7: Placement Strategy
S33 Placement Strategy Secures service deployment on volunteer fogs in multi-cloud setups. Security Model for Volunteering Fog Services Creates a security model for fog-deployed services, incorporating SSL/TLS for communication and certificate or token-based access control.
3.6. Automation Challenges and Solution Framework (RQ4 and RQ5)
Table 12 presents an overview of the automation challenges and solution framework for containerized applications in multi-cloud environment. The figure consists of eight categories, each focusing on different aspects of automation challenges and their solutions in a multi-cloud environment.
Multi-Cloud Automation: This category reports challenges and solutions related to automating processes for or combining and administering various cloud computing environments. It addresses challenges such as multi-cloud integration and provisioning, and provides solutions such as proxies for environment interface and model-driven engineering techniques.
Automation in Deployment and Scaling: This category reports challenges and solutions related to automating the deployment and scaling processes of containerized applications across multi-cloud platforms. It tackles challenges like container placement automation and application runtime switch automation. Solutions include intelligent container placement algorithms and service drivers for automated runtime switch.
Resource Management Automation: This category reports challenges and solutions related to automating the management of resources in multi-cloud environment. It addresses challenges such as resource overbooking management and containerized applications management. Solutions include ML-based resource overbooking detection and custom Kubernetes schedulers.
Data and Application Migration Automation: This category reports challenges and solutions related to automating the migration of data and applications in multi-cloud environment. It covers challenges such as data integrity during migration and legacy web applications migration. Solutions encompass the implementation of self-sufficient management cycles for streamlined deployment, alongside the utilization of adaptable architectural patterns designed for the modernization of legacy applications within a containerized multi-cloud environment.
Testing and Benchmarking Automation: This category reports challenges and solutions related to the automation of testing and benchmarking processes for containerized applications in multi-cloud environment. It addresses challenges like user-oriented testing automation and benchmarking automation. Solutions include container-based lightweight virtualization for testing and smart Docker benchmarking orchestrators.
Standardization and Interoperability Challenges: This category reports challenges and solutions related to standardization and interoperability in the automation of containerized applications across multiple cloud environments. It covers challenges like standardization in automation processes and cloud provider selection. Solutions include service mesh networks and AI-driven cloud provider selection tools.
Application and Service Management: This category reports challenges and solutions related to the automation of application and service management processes in containerized multi-cloud environment. It addresses challenges such as 5G ecosystem component selection and end-to-end tail latency management. Solutions include open-source software components for customizable orchestration and modeling approaches for performance evaluation.
Runtime and Service Discovery: This category deals with challenges and solutions related to runtime management and service discovery in multi-cloud environment, which involves managing the complexity of network control and administration within containerized applications across multiple cloud environments.
Table 12. Automation challenge and solution framework for containerized applications in multi-cloud environment
ID Challenge Challenge Description Solution Solution Description
Category 1: Multi-Cloud Automation
S10 Multi-Cloud Integration Automating processes to manage the complexities of integrating multiple cloud environments Proxies for Environment Interface Proxies for individual environment interface, unified view of resources, shielding from underlying complexity in multi-cloud environment
S20 Multi-Cloud Provisioning and Adaptation Automating provisioning, deployment, and dynamic adaptation of multi-cloud applications for efficiency and timely response Model-driven Engineering Techniques Model-driven engineering techniques for dynamic provisioning, deployment, and adaptation in multi-cloud environment
S33 Fog Creation in Multi-Cloud Automating the dynamic creation of nearby fog in multi-cloud environment Dynamic On-demand Fog Computing Dynamic on-demand fog computing architecture in multi-cloud environment
S47 Intelligent User-Friendly Automation Addressing the lack of intelligent, user-friendly automation in delivery platforms MSSF MSSF (Multi-cloud Storage Selection Framework) for automatic selection of multi-cloud redundancy storage based on user needs and security rules
S44 Multi-Stack Orchestration Automating the specification, coordination, and orchestration of different infrastructure stacks and cloud services TOSCA-based Serverless Deployment Topology and Orchestration Specification for Cloud Applications (TOSCA) for automated deployment of multi-cloud serverless applications
S34 Secure Multi-Cloud App Creation Providing tools and mechanisms for non-security expert developers to automate secure multi-cloud app creation MUSA Framework MUSA framework for application modeling, risk analysis, cloud service selection, composite SLA generation in multi-cloud environment
S37 Unified Abstraction for Multi-Cloud Deployment Implementing a common abstraction for products offered in automated multi-cloud deployments Application-level Resource Managers Application-level resource managers, native multi-cloud support for enterprise applications in multi-cloud environment
S52 Multi-Cloud Deployment Integration Addressing integration complexity in multi-cloud deployment automation MUSA Deployer Tool MUSA Deployer tool for automation of deployment process, deployment plan, and security compliance in multi-cloud environment
S57 Deployment Across Federated Clouds Automating deployment across federated cloud providers Orchestrated Deployment Subsystem Orchestrated deployment subsystem, Beacon Broker, and federated environments for coordinated application deployment in multi-cloud environment
S58 Hybrid/Multi-Cloud Setup Automating the setup of hybrid/multi-cloud environment to address security, vendor lock-in, and interoperability issues Interoperability Standardization Interoperability standardization and federation of clouds in multi-cloud environment
S80 Multi-cloud Interoperability Facilitating multi-cloud interoperability, application topology definition/composition, and elastic runtime adaptation for containerized services Multi-cloud Interoperability Multi-cloud interoperability, application topology definition/composition, and elastic runtime adaption in multi-cloud environment
S81 Multi-cloud Brokerage System Overcoming API heterogeneity and promoting semantic interoperability for a brokerage system in multi-cloud container management. Managing Large Number of Cloud Services Interoperability and vendor lock-in caused by the diversity of cloud solutions, complexity of multi-cloud systems administration
S84 Multi-cloud Orchestration Implementing multi-cloud business process provisioning and cross-level orchestration for cloud services in a containerized setting Manual Migration to the Cloud Manual migration of business processes to the Cloud, limited multi-cloud and cross-cloud orchestration
S85 DevOps for Multi-cloud Building a DevOps reference architecture with a CI broker for automation within a containerized multi-cloud environment Agile Software Development Automation Automation of agile software development, application deployment using CI-broker, automated faster application delivery, automated and integrated testing of the application in multi-cloud environment
Category 2: Automation in Deployment and Scaling
S27 Multi-Cloud Deployment and Monitoring Tackling challenges in automating deployment and monitoring of applications across multi-cloud platforms Multi-cloud Strategy Multi-cloud strategy for risk distribution and provider dependency reduction in multi-cloud environment
S06 Container Placement Automation Addressing the complexity of automating container placement in multi-cloud environment Intelligent Container Placement Develop a machine learning-based optimization algorithm for intelligent container placement across multiple cloud environments, focusing on load balancing and minimizing network latency.
S07 Application Runtime Switch Automation Enabling automated application runtime switch among various cloud services Service Driver Service driver, multi-cloud application service access model, runtime service update, development framework in multi-cloud environment
S15 Fine-Grained Automation at Component Level Implementing autonomic fine-grained solutions at the service component level for better automation Mobility Service and Replication Mobility service, migration and replication of critical software components, autonomic fine-grained service migration/replication in multi-cloud environment
S28 DevOps Automation Ensuring effective automation of software development and operations management DevOps Development Model DevOps development model, continuous development, automated tools and management systems, autonomic management system, performance models, systematic decision-making in multi-cloud environment
S29 Service Deployment Automation Addressing manual deployment issues by automating service deployment Automation Scripts Automation scripts, automated deployment, and error reduction in multi-cloud environment
S30 Deployment Time Security Automation Implementing security automation during deployment time Automated Security Measures Automated security measures, security evaluation and bootstrap, vulnerability scanning, firewall deployment, Chef integration in multi-cloud environment
S31 Distributed Simulation Automation Automating distributed simulation environments in cloud infrastructures DevOps Methodologies DevOps methodologies, automated developer operations, rapid microservices deployment, containerization paradigm, DEVS Simulation Farm, Docker technology, SOA Microservices paradigm, DEVSML Stack in multi-cloud environment
S32 Software Deployment Speed Improvement Addressing slow software deployment issues by implementing effective automation, enhancing Quality of Experience (QoE) Reorganized Container Images Reorganized container images, modified Docker deployment process, asynchronous image layer retrieval, profiled container execution, overlay file system in multi-cloud environment
S36 Container Startup Automation Identifying essential files for container startup automation Profiled Container Execution Profiled container execution, essential file determination, special image layer creation in multi-cloud environment
S38 Dynamic Scaling of Resources Automating frequent deployment and dynamic scaling of resources in multi-cloud environment for improved efficiency and performance CSP Automation Capability CSP automation capability for user agility in containerization of multi-cloud environment
S24 Container Deployment and Scaling Implementing automated container deployment and scaling based on resource demand across multiple cloud platforms Kubernetes Kubernetes for automated container deployment and multi-cloud scaling in multi-cloud environment
S48 Elastic Resource Provisioning Implementing elastic resource provisioning for fluctuating business workloads in a containerized multi-cloud environment through automation Elastic Resource Provisioning Design a predictive analytics system coupled with automation tools for real-time, elastic resource provisioning, addressing fluctuating business workloads in a containerized multi-cloud environment
S50 VF Placement and Service Deployment Automating optimal Virtual Function (VF) placement and dynamic service deployment in multi-cloud environment VF Placement Optimization Optimization model for virtual function (VF) placement and dynamic VF allocation in multi-cloud environment
S68 Container and VM Deployment Orchestrating and automating the deployment of containers and VMs in multi-cloud environment Multi-Cloud Orchestration Develop a comprehensive multi-cloud management platform for streamlining orchestration and automation of containers and virtual machine deployment across different cloud environments
Category 3: Resource Management Automation
S04 Resource Overbooking Management Automating the management of resource overbooking in service containers across multi-cloud platforms ML-based Resource Overbooking Detection Machine learning-based resource overbooking detection in multi-cloud environment
S21 VCluster Management Automating cluster creation, deployment, and management for multiple Virtual Clusters (Vclusters) across diverse application frameworks ClaaS Service Model ClaaS service model, virtualized cluster environment, easy deployment and management in multi-cloud environment
S39 Containerized Applications Management Automating deployment, scheduling, and management of containerized applications in multi-user and multi-cloud environment Custom Kubernetes Scheduler Custom Kubernetes scheduler (label-affinity-scheduler) for container orchestration in multi-cloud environment
S41 VM Types Selection in Data Centers Addressing the large and complex search space of different VM types in various data centers through automation GA-based Algorithm GA-based algorithm, domain-tailored service clustering method for complexity reduction in multi-cloud environment
S42 Resource Commissioning and De-Commissioning Speeding up resource commissioning and de-commissioning through automation for timely system scalability MicroCloud Architecture MicroCloud architecture, fine-grained resource allocation, TOSCA library, meta-workflow, coordinated adaptation workflow in multi-cloud environment
S43 Cloud Federation Automation Implementing automation support for federation or inter-platform portability in multi-cloud environment Middleware Platform Middleware platform for security, cost, and performance optimization, data locality, privacy needs in multi-cloud and MEC environments
S45 Resource Provisioning and Elasticity Detection Automating resource provisioning for cloud services, monitoring runtime configurations, and detecting elasticity actions JCatascopia JCatascopia, a fully-automated and modular cloud monitoring framework with elasticity actions, metric rule mechanism, and low-level monitoring metrics in multi-cloud environment
S51 Resource Acquisition Automation Automating resource acquisition processes for applications to function without human intervention Aneka Platform Aneka platform for automatic acquisition and release of resources from multiple clouds in multi-cloud environment
S74 SLA Enforcement Automation Automating the testing and corrective actions for SLA enforcement in containerized multi-cloud environment to reduce expense and time consumption SLA Compliance Testing Manual testing and corrective actions for SLA compliance are expensive and time-consuming in multi-cloud environment
Category 4: Data and Application Migration Automation
S01 Data Integrity during Migration Safeguarding data integrity during the automation of migration processes in multi-cloud environment Autonomous Management Cycle Autonomous management cycle for efficient deployment in multi-cloud environment
S17 Legacy Web Applications Migration Facilitating the automated migration of legacy web applications to cloud services Reusable Architectural Pattern Reusable architectural pattern for containerization in multi-cloud environment
S69 Application Porting to Cloud Automating the porting of existing applications to cloud infrastructures CloudSME Simulation Platform CloudSME Simulation Platform for fast development and deployment of complex multi-cloud applications, commercial accounting, and billing
Category 5: Testing and Banchmakring Automation
S09 User-Oriented Testing Automation Accelerating user-oriented software testing in diverse and heterogeneous environments through automation S Container-based Lightweight Virtualization Container-based lightweight virtualization, just-in-time deployment, streamlined multicomponent prototyping in multi-cloud environment
S02 Benchmarking Automation Implementing an automated benchmarking process for cloud hosts from diverse providers in multi-cloud environment Smart Docker Benchmarking Orchestrator Smart Docker Benchmarking Orchestrator for containerization in multi-cloud environment
S13 CI for Dependable Automation Leveraging Computational Intelligence (CI) to enhance dependability in automated multi-cloud containerized environments Dependability Enhancement with CI Leverage Computational Intelligence (CI) to enhance dependability in automated multi-cloud containerized environments
Category 6: Standardization and Interoperability Challenges
S05 Standardization in Automation Implementing standardization in the automation processes for containerized applications across multiple cloud environments Service Mesh and Multi-cloud Federation Service mesh network and multi-cloud cluster federation for containerization in multi-cloud environment
S70 Cloud Provider Selection Developing optimal strategies for cloud provider and node configuration selection to balance cost and runtime in a multi-cloud, containerized environment Optimal Cloud Provider Selection Create an AI-driven decision-support tool that automatically selects the optimal cloud provider and node configuration, considering both cost-effectiveness and runtime efficiency in a multi-cloud, containerized environment.
S16 Unified Template for Microservices Addressing the lack of a unified template and specification for automating microservices deployment across multiple cloud platforms Tool Service Model Tool service model, invocation relationships, service function chains, objective function optimization, accelerated Particle Swarm Optimization, configurable objective parameters in multi-cloud environment
S64 Cloud Interoperability Addressing the interoperability issues between different clouds through automation multi-cloud Harmony Develop an AI-driven multi-cloud management platform to automate cross-cloud interoperability, ensuring seamless data transfer and communication between disparate cloud environments
S63 Cloud Applications Standardization Promoting the adoption of worldwide standards or interfaces for cloud applications and their management through automation Cloud Service Interoperability Cloud service providers and middleware for interoperability, standardized interfaces, and hybrid solutions in multi-cloud environment
S18 Package Creation and Scaling Automating the creation and scaling of packages using VM and Docker based on machine-readable definition files Package Automation with VM Docker Implement a scriptable automation framework that facilitates creation, scaling, and management of packages using VM and Docker based on machine-readable definition files
Category 7: Application and Service Management
S55 5G Ecosystem Component Selection Simplifying the complexity of 5G ecosystem component selection through automation Open-source Software Components Open-source software components for customizable container and virtual machine orchestration in contemporary deployments
S62 End-to-End Tail Latency Management Automating end-to-end tail latency management in a microservice architecture Performance Modeling Approach Modeling approach for performance evaluation of container-level and VM-level data in multi-cloud environment
S79 Distributed Application Complexity Addressing the increasing complexity of managing distributed applications in containerized multi-cloud environment Complexity of Distributed Applications Increasing complexity of distributed applications and operating multi-cloud management topologies involving CMPs
Category 8: Runtime and Service Discovery
S72 Network Complexity Managing the complexity of network control and administration within containerized applications across multiple cloud environments. Network Management Complexity Complexity of network management and control in multi-cloud environment
S83 Social Network Service Discovery Improving service discovery algorithms for tagging in online social networks within containerized multi-cloud environment Automated Discovery and Selection Automating discovery and selection of relevant and appropriate services for composition and QoS management in multi-cloud environment
S08 Complex Application Administration Handling growing application complexity due to new technologies and needs through automation Autonomic Computing Systems Autonomic Computing Systems with knowledge transfer for containerization in multi-cloud environment
[…]
3.7. Deployment Challenges and Solution Framework (RQ4 and RQ5)
Table 13 presents a comprehensive overview of the challenges and corresponding solutions faced during the deployment of containerized applications in multi-cloud environment. As organizations increasingly adopt containerization and multi-cloud strategies to enhance scalability, flexibility, and resource utilization, they encounter various obstacles that require effective solutions. The figure consists of nine categories, each representing a specific deployment challenge, along with the corresponding challenge description and its solution. These categories encompass a wide range of issues, including deployment complexity and Orchestration, access and communication control, security and compliance management, multi-cloud deployment coordination and Integration, monitoring and scalability challenges, microservices architecture and containerization, and network connectivity and hybrid cloud integration challenges.
Deployment Complexity and Orchestration: This category focuses on addressing challenges and providing solutions related to the complexity and orchestration of container deployment in multi-cloud environment. One of the significant challenges faced in this context is Performance Testing Deployment. It involves evaluating testing procedures and assessing performance metrics during container deployment across multiple cloud environments. To overcome this challenge, the corresponding solution is Standardized Benchmarking and Automation. This solution entails adopting Standardized Benchmarking and automatic systems to ensure consistent and reliable performance evaluation. Another challenge is Deployment Validation, which requires conducting thorough validation and artifact checks during container deployment in multi-cloud environment. To effectively address this challenge, the suggested solution is the implementation of a Testing Process Management System. This system enhances and oversees the validation process for deployment, making sure that all essential criteria are met before deployment proceeds. The challenges outlined in this category can simplify and optimise container deployment in complex multi-cloud environment, facilitating efficient and reliable operations across different cloud service providers.
Access and Communication Control: This category emphasizes maintaining secure multi-cloud connectivity and managing user access control. Challenges in this category include ensuring secure access to resources and secure communication between containers and across clouds. The proposed solutions involve implementing encryption and authentication mechanisms, using secure communication protocols, and executing operations in a provider-agnostic manner to maintain security.
Security and Compliance Management in Multi-cloud Deployment: This category focuses on security and compliance challenges faced during multi-cloud container deployment. One challenge, Vendor Lock-in Prevention Deployment, aims to prevent vendor lock-in and ensure modular and loosely-coupled deployment in multi-cloud environment. The solution proposed is the Cloud Modelling Framework, which introduces a framework for specifying provisioning and deployment to enhance compatibility with existing ACSs and cloud solutions. Another challenge in this category is Security Deployment in Multi-container Environment, which addresses security and privacy concerns during multi-container deployment on the same OS kernel in multi-cloud environment. The solution is Cross-Container Isolation, which enforces cross-container isolation to enhance security.
Multi-cloud Deployment Coordination and Integration: This category deals with challenges related to coordination and integration during multi-cloud deployment. One challenge, Multi-cloud PaaS Deployment, aims to overcome high entry barriers for deploying a multi-cloud PaaS infrastructure. The proposed solution is Lightweight Proxies for PaaS Integration, which utilizes lightweight proxies to facilitate individual PaaS services integration. Another challenge is IoT Application Deployment, which addresses challenges related to deploying IoT applications in multi-cloud environment. The solution is based on the DRA Framework and CI Broker for Multi-Cloud Deployment, which leverages the DRA framework and CI broker component for efficient multi-cloud deployment.
Monitoring and Scalability Challenges: This category focuses on challenges related to monitoring, performance, and scalability during multi-cloud container deployment. For instance, Legacy Code Migration Deployment deals with handling legacy code migration, tenant engine separation, and composition modeling during container deployment in multi-cloud environment. The proposed solution is the Reusable Architectural Pattern with Docker and Cloudant, which advocates for adopting a reusable architectural pattern with Docker containers and Cloudant persistence layer.
Microservices Architecture and Containerization: This category addresses challenges associated with microservices architecture and containerization in multi-cloud environment. For example, Autonomic System Deployment aims to implement an autonomic management system for deploying multi-layer, multi-tier web architectures in multi-cloud environment. The solution proposed is the Self-Tuning Performance Model and Autonomic Management, which introduces a self-tuning performance model and autonomic management system.
Network Connectivity and Hybrid Cloud Integration: This category focuses on network connectivity and hybrid cloud integration challenges during multi-cloud deployment. One challenge, Inter-edge Bandwidth Deployment, aims to address challenges related to interconnecting distributed localities and provisioning inter-edge bandwidth.
Cloud Deployment Constraints and Challenges: This category deals with various constraints and challenges encountered during cloud deployment. For instance, ACS Compatibility Deployment aims to ensure compatibility with existing ACSs and cloud solutions during container deployment in multi-cloud environment. The solution proposed is the Scalable Multi-Cloud Deployment Framework, which proposes a scalable deployment framework for multi-cloud applications.
Infrastructure Provisioning and Container Deployment Challenges: This category addresses challenges related to infrastructure provisioning and container deployment in multi-cloud environment. For example, Fog Computing Deployment deals with fog computing hardware and container deployment time challenges. The proposed solution is Reorganized Container Images and Docker Deployment, which involves reorganizing container images and modifying the Docker deployment process.
Application Deployment Workflow and Orchestration Challenges: This category focuses on challenges related to application deployment workflow and orchestration in multi-cloud environment. One challenge is API Management Deployment, which deals with effectively managing different cloud providers’ APIs/interfaces for multi-cloud container deployment. The solution proposed is DevOps Approach for Multi-Cloud Applications, which advocates for adopting a DevOps approach for managing the lifecycle of multi-cloud applications.
Table 13. Deployment challenge and solution framework for containerized applications in multi-cloud environment
ID Challenge Challenge Description Solution Solution Description
Category 1: Deployment Complexity and Orchestration Challenges
S01 Performance Testing Deployment Evaluating testing and performance during container deployment in multi-cloud environment Standardized Benchmarking and Automation Adoption of standardized benchmarking applications and automatic systems
S09 Deployment Validation Conducting rigorous validation and artifact checks during container deployment in multi-cloud environment Testing Process Management System Implementation of a testing process management system
S02 Cost-effective Deployment Ensuring cost-effective and flexible container deployment options in multi-cloud environment Containerization and Microservice Architecture Implementation of containerization and microservice architecture
S38 Resource Management Deployment Managing diverse resources across multiple cloud service providers during container deployment User Preference-Based Multi-Cloud Resource Brokering Adoption of multi-cloud resource brokering dependent on user preferences and application requirements
S50 NFV Site Deployment Managing site selection, dynamic VF allocation, and service chaining during NFV deployment in multi-cloud environment Site Selection and VF Allocation Algorithms Development of algorithms for site selection and dynamic VF allocation
S79 Cloud Management Platform Evaluation Deployment Addressing diverging functionality, lack of standard benchmark, and evaluating CMPs during multi-cloud container deployment Standardized Output Formats and Evaluation Criteria Establishment of standardized output formats and evaluation criteria
S37 Standardized Requirement Deployment Standardizing the expression of application requirements during multi-cloud container deployments Application-Level Resource Managers for Multi-Cloud Emergence of application-level resource managers supporting multi-cloud deployment
S42 Heterogeneous Node Deployment Managing the deployment and control of heterogeneous nodes in multi-cloud environment Technology-Independent Multi-Level Adaptation Utilization of MicroCloud and TOSCA library for technology independence and fine-grained multi-level adaptation
S52 Deployment Plan Generation Identifying suitable cloud services and generating deployment plans for multi-cloud environment Cloud Services Selection and Deployment Planning Utilization of the MUSA Decision Support Tool and Deployer tool for cloud services selection and deployment planning
S57 TOSCA-based Deployment Implementing TOSCA-based container deployment across multiple cloud providers Brokered Multi-Cloud Deployment with TOSCA Adoption of a brokered approach for multi-cloud deployment with TOSCA
S58 Standardization Deployment Addressing compatibility issues and lack of standardization during multi-cloud container deployment Proxy Cloud Virtualization for OpenStack Framework mapping other clouds as sub-clouds to OpenStack with Proxy Cloud Virtualization layer
Category 2: Security and Compliance Management in Multi-Cloud Deployment
S20 Vendor Lock-in Prevention Deployment Developing strategies to prevent vendor lock-in and ensure modular and loosely-coupled deployment in multi-cloud environment Cloud Modelling Framework Introduction of a Cloud Modelling Framework for specifying provisioning and deployment
S25 Security Deployment in Multi-container Environment Addressing security and privacy concerns during multi-container deployment on the same OS kernel in multi-cloud environment Cross-Container Isolation Enforcing cross-container isolation
S34 Security Deployment Ensuring integrated security during the deployment of multi-cloud applications Security Control Specification and Deployment Framework Utilization of the MUSA framework for specifying, deploying, and assessing security controls
S72 Virtual Network Security Deployment Providing secure virtual networks during multi-cloud container deployment Network Virtualization Platform for Multi-Cloud Adoption of Sirius, a network virtualization platform for multi-cloud environment
S83 Malicious Service Management Deployment Managing malicious services and supporting multi-cloud communities formation during container deployment Collusion-Resilient Trust Aggregation Technique Implementation of a trust aggregation technique resilient to collusion attacks
Category 3: Multi-Cloud Deployment Coordination and Integration
S10 Multi-cloud PaaS Deployment Overcoming challenges and high entry barriers for deploying a multi-cloud PaaS infrastructure Lightweight Proxies for PaaS Integration Usage of lightweight proxies for individual PaaS services integration
S85 IoT Application Deployment Addressing challenges related to deploying IoT applications in multi-cloud environment DRA Framework and CI Broker for Multi-Cloud Deployment Utilization of the DRA framework and CI broker component for multi-cloud deployment
Category 4: Monitoring, Performance, and Scalability Challenges
S17 Legacy Code Migration Deployment Handling legacy code migration, tenant engine separation, and composition modeling during container deployment in multi-cloud environment Reusable Architectural Pattern with Docker and Cloudant Adoption of a reusable architectural pattern with Docker containers and Cloudant persistence layer
S23 Tenant Performance Deployment Managing tenant service performance competition during multi-cloud container deployment Multi-Tenant and Multi-Instance Hybrid Deployment Implementation of a multi-tenant and multi-instance hybrid deployment scheme based on container technology
S45 Monitoring Deployment Overcoming the limitations of current monitoring tools during multi-cloud container deployment Platform-Independent Monitoring System Implementation of JCatascopia, a platform-independent and interoperable monitoring system
Category 5: Microservices Architecture and Containerization Challenges
S28 Autonomic System Deployment Implementing an autonomic management system for deploying multi-layer, multi-tier web architectures in multi-cloud environment Self-Tuning Performance Model and Autonomic Management Development of a self-tuning performance model and autonomic management system
S46 Microservice Deployment Managing the complexity of deploying microservice architectures in multi-cloud environment Metrics and Requirements Framework for Microservices Framework for establishing requirements and relevant metrics for the deployment of microservice-based applications
S76 Multi-cloud Application Configuration Deployment Overcoming challenges related to configuring applications during multi-cloud deployment Runtime Environment with Object Store and Artifact Repository Deployment of a runtime environment with an object store service and artifact repository
S80 Elastic Container Platform Dependency Deployment Addressing vendor lock-in and dependency on specific elastic container platforms during multi-cloud deployment Separation of Elastic Platform and Cloud Application Separation of elastic platform definition from cloud application definition
Category 6: Network Connectivity and Hybrid Cloud Integration
S35 Inter-edge Bandwidth Deployment Addressing challenges related to interconnecting distributed localities, provisioning inter-edge bandwidth, and centralized edge nodes during multi-cloud deployment Modular Edge Cloud Computing Architecture Designing a modular edge cloud computing system architecture using containerization techniques
S41 Network Latency Deployment Minimizing the impact of network latency on application performance during multi-cloud deployment VM Selection for Composite Applications Selection of proper VMs from the global cloud marketplace to support multiple composite applications
S43 Latency Reduction Deployment Tackling the issue of high end-to-end latencies during multi-cloud deployment Edge-Cellular Hybrid Infrastructure Provisioning Provisioning infrastructure at the edge of the cellular network combined with traditional cloud infrastructures
S63 Vendor Lock-in Prevention Deployment Preventing negative effects of vendor lock-in during multi-cloud container deployment Interoperability Layer Above Cloud Infrastructure Implementation of a layer above the cloud service provider’s infrastructure for interoperability
Category 7:Cloud Deployment Constraints and Challenges
S08 ACS Compatibility Deployment Ensuring compatibility with existing ACSs and cloud solutions during container deployment in multi-cloud environment Scalable Multi-Cloud Deployment Framework Proposal and development of a scalable deployment framework for multi-cloud applications
S30 Manual Intervention Risk Deployment Managing risks related to manual intervention, vulnerability, and exposed services during multi-cloud deployment Automated Security Measures at Deployment Automated implementation of required security measures at the point of deployment
S44 Multi-provider Deployment Handling the challenges of using multiple providers and technologies during multi-cloud container deployment TOSCA-Based Deployment Modeling Approach Adoption of a TOSCA-based deployment modeling approach
S81 Cloud Provider Transition Deployment Addressing incompatibilities of cloud solution features during multi-cloud deployment Semantic Interoperability in Multi-Clouds Adoption of the FCLOUDS framework for semantic interoperability in multi-clouds
Category 8: Infrastructure Provisioning and Container Deployment Challenges
S32 Fog Computing Deployment Addressing fog computing hardware and container deployment time challenges in multi-cloud environment Reorganized Container Images and Docker Deployment Reorganization of container images and modification of the Docker deployment process
S33 Fog Device Deployment Managing fog device availability and services during multi-cloud container deployment On-Demand Fog and Microservices Deployment Enabling on-demand creation of fogs and deployment of microservices using Docker and Kubeadm
S36 Fog Computing Container Deployment Overcoming low rates of container deployment, especially in fog computing environments FogDocker for Container Deployment Deployment of containers using FogDocker after essential file downloads
S39 Distributed Compute Node Deployment Effectively managing containerized applications across distributed compute nodes in multi-cloud environment Label-Based Scheduling Strategy Development of a label-based scheduling strategy
S40 Kernel Security Deployment Addressing security risks related to kernel sharing, isolation, and OS-level security during multi-cloud container deployment Enhanced Container Security Enhancing container security from the host OS level
Category 9: Application Deployment Workflow and Orchestration Challenges
S27 API Management Deployment Effectively managing different cloud providers’ APIs/interfaces for multi-cloud container deployment DevOps Approach for Multi-Cloud Applications Adoption of a DevOps approach for managing the lifecycle of multi-cloud applications
S29 Web Service Deployment Ensuring correct and efficient deployment of web services across various computing environments Kubernetes-Based Containerized Deployment Deployment of containerized environments using Kubernetes
S51 Complex Application Deployment Overcoming complexities associated with application deployment in multi-cloud environment Aneka Platform for Distributed Applications Adoption of the Aneka platform for programming and deploying distributed applications
S53 High-availability Deployment Ensuring high availability and preventing outages during deployment in multi-cloud environment Middleware Support for High Availability in Multi-Cloud PaaS Introduction of novel Middleware Support for High Availability in multi-cloud PaaS
S69 Complex Workflow Deployment Utilizing multiple cloud providers for deploying complex application workflows CloudSME Simulation Platform Utilization of the CloudSME Simulation Platform for access to various resources and workflow support
S84 Business Process Provisioning Deployment Orchestrating multi-cloud business process provisioning during deployment Multi-Cloud Service Orchestration Frameworks Adoption of multi-cloud service orchestration frameworks
[…]
3.8. Monitoring Challenges and Solution Framework (RQ4 and RQ5)
Table 14 provides an extensive overview of the monitoring challenges and solution framework for containerized applications in multi-cloud environment. It encompasses diverse challenges related to monitoring containerized applications in the complex multi-cloud landscape. Each category in the figure targets specific monitoring aspects and offers solutions to address these challenges.
Performance, Consistency, and Variability Monitoring: This category emphasizes sustaining performance monitoring in dynamically varying cloud environments for containerized applications. It tackles the challenge of ensuring consistent performance monitoring and proposes solutions like performance variation monitoring and benchmarking for optimized performance.
Resource and Infrastructure Monitoring: In this category, the focus is on monitoring resource and infrastructure management for containerized applications in multi-cloud environment. Challenges include network communication, system image and configuration monitoring, and dynamic resource management. Solutions involve service mesh implementation, Docker server deployment, and using Platform-as-a-Service (e.g., Aneka PaaS) for Cloud Computing.
Security, Compliance, and Trust Monitoring: This category addresses the vital aspects of monitoring security, compliance, and trust management for containerized applications in multi-cloud settings. Challenges include accessing provider host information, continuous security control, and multi-cloud security compliance. Solutions include ML-based monitoring and security assurance platforms in the form of a SaaS (e.g.,MUSA), which support multi-cloud application runtime security control and transparency to enhance user trust.
Application Optimization and Adaptability Monitoring: The challenges of this category revolve around monitoring application optimization and adaptability in multi-cloud environment. Challenges include real-time application monitoring, microservices optimization, application behavior, download execution, and design and state monitoring. To address these challenges, solutions encompass the Dynamic Resource Allocation (DRA) framework, Microservices-based Cloud Application-level Dynamic Orchestrator (MiCADO) optimization, Topology and Orchestration Specification for Cloud Applications (TOSCA) event modeling, FogDocker implementation, and adopting a cloud modeling framework.
System Complexity and Standardization Monitoring: This category focuses on addressing system complexity and promoting standardization in monitoring practices across multi-cloud environment. Challenges include proprietary monitoring solutions and resource isolation mechanisms in the Linux kernel. Solutions encompass pervasive monitoring and identifying in-container leakage channels.
Multi-cloud Coordination and Integration Monitoring: In this category, the challenges revolve around coordinating and integrating monitoring data across multiple cloud environments. Challenges include multi-cloud monitoring challenges, cross-level monitoring improvement, and integrated multi-cloud management monitoring. Solutions encompass self-healing techniques, business process monitoring, and integration with tools like Gnocchi.
[…]
Table 14. Monitoring challenge and solution framework for containerized applications in multi-cloud environment
ID Challenge Challenge Description Solution Solution Description
Category 1: Performance Monitoring Consistency and Variability
S01 Performance Monitoring Ensuring sustained performance monitoring in dynamically varying cloud environments Performance Variation Monitoring Monitoring performance fluctuations in containerized applications through benchmarking periods
S02 Benchmarking Performance Variation and Cost Efficiency Monitoring Efficiently tracking performance disparities and cost-effectiveness for refined monitoring in benchmarking SDBO Optimization Maximizing user budget utility with the optimization feature in SDBO
S29 Website Performance and Availability Monitoring Implementing continuous oversight over website availability and performance Kubernetes Monitoring Utilizing Kubernetes for monitoring performance of containerized services in multi-cloud environment
S38 User Experience Variability Monitoring Addressing monitoring fluctuations in user experience based on diverse factors CSP QoS Metrics Collection Collecting Cloud Service Provider (CSP) Quality of Service (QoS) metrics for benchmark data in multi-cloud environment
S43 Independent Performance Metrics Monitoring Establishing application-independent performance metrics for Round Trip Time (RTT) and payload throughput Performance Evaluations Conducting performance evaluations and autonomic resource provisioning for multi-cloud and MEC platforms using containerized applications
S50 Comprehensive Network Performance Measurement Monitoring Ensuring comprehensive performance measurements monitoring and accounting for transmission delays in the network Server Selection Scheme Applying a server selection scheme to minimize latency in multi-cloud environment based on end-to-end performance measurements
S62 Microservice Performance Degradation Monitoring Addressing unpredictable performance degradation in containerized microservices Tail Latency Prediction Implementing a modeling approach to predict end-to-end tail latency in multi-layer cloud data of containerized applications
Category 2: Resource and Infrastructure Management Monitoring
S05 Network Communication Monitoring Managing service-to-service communications over a network Service Mesh Implementation Deploying a network service mesh for service-discovery, secure connectivity, and visibility in multi-cloud platforms
S09 System Image and Configuration Monitoring Monitoring of system images, remote software packages, usage history, and container execution, configuration, and monitoring. Docker Server Deployment Deploying containerized services using Docker Server API version 1.23 in multi-cloud platforms
S21 Vcluster and Framework Management Monitoring Monitoring and managing the performance and resource usage of each Vcluster and application framework can be complex when dealing with multiple tenants and different types of applications. LXC Implementation Using Lightweight Containers (LXC) for efficient resource allocation and scalability in multi-cloud environment
S24 Multi-Cloud Resource Utilization Monitoring Monitoring container performance and resource usage across multiple clouds Prometheus and Grafana Monitoring Implementing Prometheus and Grafana for monitoring performance and availability of containerized applications in multi-cloud environment
S51 Dynamic Resource Management Monitoring Applications should add and release resources dynamically by measuring and monitoring usage Aneka Platform Utilizing the Aneka platform for monitoring resources and managing dynamic resource acquisition and release in multi-cloud environment
Category 3: Security, Compliance, and Trust Management Monitoring
S04 Provider Host Information Access Monitoring Overcoming the restriction in accessing provider host information ML Monitoring Using machine learning to supervise performance of containerized applications and OS usage
S34 Continuous Security Control Monitoring Strengthening continuous monitoring and control over security MUSA Platform Applying the MUSA Security Assurance Platform for continuous monitoring and dynamic security enforcement of containerized applications
S52 Multi-Cloud Security Compliance Monitoring Monitoring security compliance consistently across multiple cloud providers MUSA Security Compliance Using the MUSA Security Assurance Platform to monitor multi-cloud application compliance and ensure security
S53 Service Continuity and Data Privacy Monitoring Ensuring data privacy, uninterrupted service, fault tolerance, and availability NoMISHAP Implementation Implementing the NoMISHAP solution to address monitoring challenges in multi-cloud environment with containerized applications
S83 Trust Relationship and Network Contact Monitoring Building trust relationships, collecting judgments, and maintaining networks of contacts Trust Bootstrapping Implementing a trust bootstrapping mechanism that combines endorsement and decision tree classification for containerized applications in multi-cloud platforms
Category 4: Application Optimization and Adaptability Monitoring
S20 Adaptive System Design and State Monitoring Implementing continuous design with dynamic adaptability; Enhancing system state evaluation through robust monitoring Cloud Modelling Framework Utilizing a Cloud Modelling Framework for continuous design and dynamic adjustment of multi-cloud containerized applications
S36 Partial Download Execution Monitoring Ensuring accurate execution of applications for partially downloaded files FogDocker Implementation Using FogDocker with low runtime overhead to ensure execution correctness of containerized applications
S44 Event-Driven Application Behavior Monitoring Modeling event-triggered behavior of application components TOSCA Event Modeling Applying TOSCA modeling for event sources, events, and functions in containerized applications
S68 Microservice Optimization Monitoring Enhancing surveillance over microservices across multiple cloud environments for optimization MiCADO Optimization Utilizing MiCADO for runtime monitoring and optimization of cloud applications in multi-cloud environment
S85 Real-Time Application Monitoring Real-time monitoring of applications for immediate corrective action DRA Framework Implementing the DRA framework with the Monitoring and Tracking Platform (M4) for containerized applications in multi-cloud platforms
Category 5: System Complexity and Standardization Monitoring
S10 Proprietary Monitoring Solution Standardization Overcoming the constraints of complex, proprietary monitoring solutions and fostering standardization Pervasive Monitoring Implementing pervasive monitoring solutions for containerized applications within multi-cloud PaaS
S25 Kernel Resource Isolation Mechanism Monitoring Addressing the limitations of resource isolation mechanisms in the Linux kernel Leakage Defense Identifying in-container leakage channels and implementing a two-stage defensive approach in multi-cloud environment
S27 Multi-Cloud Application Monitoring Complexity Simplifying monitoring processes in multi-cloud applications DECIDE DevOps Expansion Expanding the DECIDE DevOps framework for native multi-cloud containerized applications
S28 Monitoring Agent Coordination Enhancing the coordination of monitoring agents for improved system behavior Autonomic Management Implementing an autonomic management system with monitoring agents, a performance model, and adaptive actions for containerized applications
S33 Heterogeneous Fog Resource Management Monitoring Efficient management of diverse fog resources Kubernetes Utility Architecture Implementing a master-worker nodes architecture with Kubernetes Utility (Kubeadm) in multi-cloud environment
S42 Application Topology and Dependency Monitoring Enhancing coordinated adaptation through thorough understanding of application topologies and dependencies MicroCloud TOSCA Library Utilizing the MicroCloud TOSCA library for standardized description of topologies in multi-cloud containerized applications
S45 Elastic Cloud Service Monitoring Research Investigating elastic cloud service monitoring as an open research problem JCatascopia Monitoring Utilizing JCatascopia for multi-layer, platform-independent monitoring of cloud containerized applications
S46 Microservice Monitoring Diversity Addressing diverse monitoring requirements and varying metrics quality in microservices QoM Definition Defining Quality of Monitoring (QoM) and implementing an API for cloud provider capabilities in multi-cloud environment
S49 Compute and Network Resource Utilization Monitoring Strengthening surveillance and prediction capabilities of compute and network resource utilization Time Series Resource Utilization Developing algorithms for monitoring and predicting resource utilization of containerized applications using time series data
Category 6: Multi-cloud Coordination and Integration Monitoring
S58 Integrated Multi-Cloud Management Monitoring Implementing integrated monitoring and management across multi-cloud environment Gnocchi Integration Integrating with federated gnocchi for consolidated monitoring and management of multi-cloud environment
S74 Multi-Cloud Monitoring Challenges Addressing monitoring challenges in multi-cloud environment Self-Healing Techniques Implementing a rule-based algorithm for monitoring and self-healing techniques in multi-cloud environment
S84 Cross-Level Monitoring Improvement Ensuring precise and adaptable metrics computation, overcoming inadequate cross-level monitoring approaches Business Process Monitoring Conducting cross-cloud and cross-level business process monitoring and coordinated adaptation of containerized applications
3.9. Tools and Frameworks (RQ6)
In response to our RQ6 concerning the tools and frameworks utilized for developing containerized multi-cloud applications, an in-depth exploration of the selected studies was conducted. This comprehensive analysis led to the identification of a total of 160 distinct tools and frameworks (see Table 15), all of which are systematically classified into 45 overarching categories, further delineated into 6 subcategories. Notably, the three most prominent categories emerged from this investigation are cloud services and management platform, infrastructure management technologies, and networking tools and framework. Each of these categories encapsulates a range of tools aimed at facilitating the seamless development, deployment, and management of multi-cloud applications. In the following we briefly report the each of the category, shedding light on the tools and frameworks that collectively form the foundation for containerized multi-cloud application development.
Cloud Services and Management Platforms: Cloud Services and Management Platforms: This category of tools facilitates the management and orchestration of cloud services across multi-cloud environments. It includes a broad range of platforms covering Infrastructure as a Service (IaaS), Platform as a Service (PaaS), and Software as a Service (SaaS). These tools are essential for provisioning virtual resources, deploying applications, and delivering software across different cloud platforms. Examples include Amazon EC2, Google Cloud App Engine, and Microsoft Office 365.
Infrastructure Management Technologies: This category of tools facilitates the management of cloud infrastructure. It encompasses container platforms, virtualization solutions, Kubernetes services, orchestration platforms, and tools for configuration management and continuous integration. Key technologies like Docker, Kubernetes, Puppet, Chef, and Terraform play a vital role in automating and optimizing the infrastructure underpinning containerized applications.
Monitoring Tools: This category of tools facilitates the monitoring and analysis of cloud and container operations. It includes solutions for cloud monitoring, container monitoring, visualization & alerting, resource monitoring, and logging & analysis. Tools such as Amazon CloudWatch, cAdvisor, Prometheus, Grafana, and the ELK stack are instrumental in ensuring the performance and reliability of cloud services.
Networking Tools: This category of tools facilitates the establishment and management of network functionalities in a multi-cloud environment. It covers load balancers, reverse proxies, network services, controllers, paradigms, types, and identity protocols, along with storage networks and event buses. Technologies like Software-Defined Networking (SDN), Neutron, and Kafka event bus are crucial for maintaining robust and efficient network operations.
Development and Programming: This category of tools facilitates the development and programming aspects in cloud environments. It includes orchestration tools, various programming languages, integrated development environments (IDEs), service meshes, APIs, and cloud modeling tools. Examples are Java, Python, Eclipse IDE, Istio, REST API, and Cloud Modelling Framework (CLOUDMF), which support the creation and deployment of cloud-native applications.
Security and Assurance: This category of tools facilitates the security and compliance of cloud environments. It encompasses vulnerability management, security platforms & tools, access control & policies, Linux security, and cryptography models & tools. Technologies like OpenVAS Vulnerability Scanner, Kubernetes.eBPF, Seccomp-BPF, and Attribute-Based Encryption and Signature are essential for protecting cloud infrastructure and data against various security threats.
Table 15. Tools and frameworks for containerized applications in multi-cloud environment
Subcategory Tools and Frameworks
Category 1: Cloud Services and Management Platforms
Cloud Management Platforms
Cloudyn (S65), IBM Brokerage Services (S65), CloudBroker Platform (S69), CloudcheckR (S79), MistIO (S79), ManageIQ (S67, S79), CMP2 (S79), AWS SDK (S71), CloudSME Simulation Platform (CSSP) (S69), MiCADO (S69), WS-PGRADE/gUSE (S69), Vagrant (S08), AWS OpsWorks (S38), Universal Compute Xchange (UCX) (S65), Open vSwitch (OvS) (S72), Liferay portlets (S69), Cloud Service Providers (CSPs) (S71), AWS Data Pipeline (S78), Cloud computing (S25), Amazon AWS (S42, S43), GENI (S38), Google Cloud Platform (S38, S53), Microsoft Azure (S38, S48, S51, S54, S60), AWS EC2 (S51), IBM (S49), Amazon Web Services (S54), Google Cloud (S54), SalesForce (S53), Multi-Cloud (S60), Cloud computing platforms (S74), Multiple independent clouds (S74), Cloud computing (S84)
Cloud Deployment Model
Public clouds: (S37, S53, S59, S61), Google cloud (S07), Microsoft Azure (S08, S16, S19), AWS (S65, S67, S80, S81), Google’s cloud (S65), Google Compute Engine (S67, S80), Alibaba (S65), Tencent (S65), IBM Watson Health Cloud (S78). Private cloud: OpenStack (S20, S21, S43), VMware vSphere Specialized Cloud Platforms: European Grid Infrastructure Federated Cloud (EGI FC) (S81)
Infrastructure as a Service (IaaS)
Amazon EC2 (S01, S51), AWS (S02, S65, S67, S80, S81), Azure (S02, S08, S16, S19, S67, S80), Google GCE (S03, S67, S80), Aliyun ECS (S03), Baidu BCE (S03), Tencent CVM (S03), JD VM (S03), VMware vSphere (S08), GoGrid (S51), Windows Azure (S51)
Platform as a Service (PaaS)
Google Cloud App Engine (S02, S10, S15), Heroku (S08, S53), IBM Bluemix Workflow Service (S17, s53), Amazon Elastic Beanstalk (S20), Cloud Foundry (S24), OpenShift (S24, S53, S55)
Software as a Service (SaaS)
SaaS (S15), Microsoft Office 365 (S53)
Distributed Processing Frameworks
Hadoop (S21, S37), Spark (S21, S28), Apache Yarn (S37), Storm (S21), Apache Kafka (S34, S35)
Data Storage
Apache Cassandra (S28), Firebase real-time database (S35), Redis (S36, S42), MySQL (S42), NoSQL (S85), SQL (S85)
Messaging Systems
Apache Kafka (S34, S35), Zookeeper (S35)
Category 2: Infrastructure Management Technologies
Container Platforms
Docker (S08, S13, S16, S17, S18, S25, S27, S28, S29, S30, S31, S32, S35, S55, S59, S62, S74, S77, S79, S80), Smart Docker Benchmarking Orchestrator (SDBO) (S01), Docklet (S21), LXC (S24, S25, S29), Docker container engine (S24), Docker container (S23), Docker CE version 18.01 (S36), Docker Hub (S36)
Virtualization
Virtual Machines (S05, S22, S23, S46, S72), Amazon EC2 VMs (S18), Docker container virtualization technology (S74), Xen (S19), VMware (S19), Hyper-V (S19), KVM (S19), OpenVZ (S19), Libvirt API (S19)
Kubernetes Services
Microsoft AKS (S03), Aliyun Kubernetes Engine (S03), Tencent Kubernetes Engine (S03)
Orchestration Platforms
Kubernetes (S18, S22, S29, S55, S56, S59, S62, S63, S80), Docker Swarm (S18, S22, S39, S59, S80), Mesos (S21, S22, S59, S80), Yarn (S21), Kubeadm (S33), Kubernetes for container orchestration (S62), Apache Mesos (S59, S80), MiCADO (S69)
Configuration Management
Puppet (S08), Chef (S08, S30), Terraform (S27, S29), Shell scripts (S44), CloudBees (S08) Continuous Integration CloudBees (S08), Maven (S08), NPM (S08), Github (S30), Github (S30)
Container Isolation
Sandboxie (S24), Warden/Garden (S24), Rocket (S24)
Container Infrastructure
Linux Container (S16), Docker Containers on Linux Host (S13), Kubernetes cluster with a Docker (S04), Amazon AWS ECS (S18), Container Technology (S23)
Category 3: Infrastructure Management Technologies
Load Balancer
Load balancer (S23)
Reverse Proxy
Reverse proxy technology (S23), httpd (S36), nginx (S36)
Ports and Interfaces
TrunkPort (S55)
Network Services
Neutron (S55)
Gateways and Tunnels
Gateways (S64), GRE Tunnel (S72)
Network Controllers
Software-Defined Networking (SDN) controller (S72)
Identity Protocols
Host Identity Protocol (HIP) (S77)
Storage Networks
MCSS (S75), Multi-Cloud storage overlay network (S75)
Network Tools
OpenADN (S50), OpenDayLight (S50), Kuryr (S55), Flannel (S55)
Event Buses
Kafka event bus (S34)
Category 4: Development and Programming
Orchestration
MUSA Deployer tool (S52), MUSA Modeler tool (S52), OSM (S55), ManageIQ (S55)
Programming Languages
Java (S30, S34, S35, S54, S80, S81), Go programming language (S32, S36), Python (S35, S54, S56, S58, S79), C# (S54), Ruby (S54), PHP (S54)
IDEs and Editors
Eclipse IDE with tight integration with OPTIMUS (S27), MCSLA editor (S27)
Service Mesh
Consul (S05), Istio (S05), Istio service mesh (S59)
APIs
REST API (S56), OpenStack API (S64)
Cloud Modeling
Cloud Modelling Framework (CLOUDMF) (S20), Cloud Modelling Language (CLOUDML) (S20, S81), CAMEL modeling language (S52, S80, S84)
Category 5: Development and Programming
Vulnerability Management
OpenVAS Vulnerability Scanner (S30)
Security Platforms and Tools
MUSA Security Assurance Platform (S34), Firewall Deployment (S30)
Access Control and Policies
Access control policies (S25)
Linux Security
Linux top command (S34), Kubernetes.eBPF (S40), Seccomp-BPF (S40), Landlock LSM (S40), Pledge (S40), Security Namespace (S40), Linux Security Modules (LSM) (S40)
Cryptography Models and Tools
Attribute-Based Encryption and Signature (S75), ABEABS cryptographic model (S75)
Category 6: Monitoring Tools
Cloud Monitoring
Amazon CloudWatch (S19)
Container Monitoring
cAdvisor (S22), Prometheus (S22), AlertManager (S22)
Visualization and Alerting
Grafana (S27), Telegraf (S27)
Resource Monitoring
Resource monitoring tool (S60), Ceilometer (S56, S58), Gnocchi (S58)
Logging and Analysis
ELK stack (S74)
[…]
4. Discussion
4.1. Containers in Multi-Cloud: Roles and Strategies (RQ1)
Containers play a crucial role in the development and operation of modern software systems within multi-cloud environment. Understanding the various roles and implementation strategies of containers provides significant insights for both researchers and practitioners.
Multifaceted Role of Containers in Multi-Cloud Environment: Containers fulfill diverse roles within multi-cloud environment, such as application deployment, service hosting, and resource management. They facilitate deployment of applications across a broad range of systems, from web apps to IoT applications. Furthermore, containers contribute to enhancing security in multi-cloud operations. For researchers, this variety of roles signifies a rich area for inquiry, with opportunities to investigate potential security vulnerabilities, management complexities, and performance overheads. For practitioners, understanding these roles can inform efficient and effective deployment strategies, leading to more robust multi-cloud implementations.
Container Technology Features and their Significance: The unique features of container technology, such as lightweight virtualization, compatibility with orchestration tools, and application in edge computing, offer significant advantages. For researchers, these features present opportunities for further exploration, including how these characteristics (e.g., lightweight virtualization) influence container performance and effectiveness in multi-cloud operations. For practitioners, understanding these features can lead to better utilization of container technology and more efficient multi-cloud deployments.
Implementation Strategies and their Real-World Implications: The strategic implementation of containers significantly affects their functionality in multi-cloud environment. Strategies related to deployment, orchestration, security, performance optimization, cloud service management, and data storage can have significant implications. For researchers, these strategies present an area for exploration, including investigation into their real-world implications, benefits, and limitations. For practitioners, a refined understanding of these strategies can lead to more efficient implementations and operations.
Containerization’s Impact on IoT and Application Development: The incorporation of containers into IoT and app development has transform these fields. Containers offer improved software unit encapsulation, portability, and efficiency, which are particularly important in the distributed and resource-constrained nature of IoT. Researchers can further explore potential challenges in detail such as network complexity, security, and the small footprint requirement of IoT devices. For practitioners, understanding the impact of containerization can lead to the development of more robust, efficient, and secure IoT and app solutions.
4.2. Patterns and Strategies (RQ2)
Containerized applications in multi-cloud environment requires an understanding of prevailing patterns and strategies. In this context, we discuss the key findings related identified patterns and strategies presented in Table 9.
Diverse Architectural Patterns and Communication Strategies: Our study found that Microservice Architecture and Service-Oriented Architecture (SOA) are the most common used architectural patterns for designing container-based applications in multi-cloud environment. These results align with the existing literature that emphasizes the use of MSA and SOA because of their support for modular development, scalability, and efficient resource use (Balalaie et al., 2016; Dragoni et al., 2017). The other patterns frequently reported are Service Mesh and Service Chaining, which are employed as communication and networking patterns, likely due to their ability to manage complex, distributed service-to-service communications in a scalable way (Zhamak, 2018). The frequent use of Blue-Green Deployment and Object Store Service deployment patterns has also been identified in this study, which may stem from their ability to minimize downtime and ensure high availability—essential factors in a cloud-based setting (Fowler, 2010). Implications: Practitioners can enhance the design, scalability, and efficiency of their applications in a multi-cloud environment by understanding and applying these architectural and communication patterns. Researchers might explore how these patterns evolve with the emergence of new technologies and standards.
Importance of Container Management and Multi-Cloud Strategies: Our findings report the significance of robust container management and multi-cloud strategies. Techniques like the Linux Container (LXC) project, Docker Container Images, and Lightweight Virtualization are frequently used, emphasizing the industry trend towards containerization due to its benefits in efficiency, scalability, and platform independence (Vaughn, 2020). Furthermore, the popularity of multi-cloud strategies like Hybrid/Multi-cloud Approach and Multi-cloud Load Balancing confirms the growing industry move towards using multiple cloud providers to enhance redundancy, flexibility, and avoid vendor lock-in (Gillam et al., 2013). Implications: Practitioners should focus on gaining expertise in container technologies and multi-cloud strategies to take advantage of their numerous benefits. Meanwhile, researchers might examine the challenges and best practices associated with the deployment and management of container technologies across multiple cloud platforms.
Security, Resilience, and Migration Patterns and Strategies: Our study also identifies patterns and strategies related to security, resilience, and efficient migration strategies in a multi-cloud context. Strategies such as File Encryption, Attribute-Based Encryption, and Secure Data sharing are often employed to ensure data security, an area of growing concern with the increase in cloud-based applications (Svantesson and Clarke, 2010). Moreover, resilience strategies such as Fault-tolerance with Redundant Engines and RAFT Consensus Algorithm are essential to ensure system reliability in distributed, multi-cloud environment (Ongaro and Ousterhout, 2014).
Migration between different cloud environments, facilitated by strategies like Service-oriented Migration and Migration to a Virtualized Container, is another vital aspect, given the growth in cloud services and the need for interoperability (Jamshidi et al., 2013). Migrating applications and data between cloud environments is an important consideration in a multi-cloud scenario (Jamshidi et al., 2013). Strategies such as image-based migration and migration to virtualized containers, like Docker, are commonly used (Vaughn, 2020). These techniques facilitate portability and provide consistent environments across different cloud platforms. Additionally, they have significant implications for scaling and managing applications across different cloud service providers. Implications: These findings suggest that security, resilience, and efficient migration should be key areas of focus for practitioners operating in multi-cloud environment. Researchers may also look into novel techniques for enhancing cloud security, improving system resilience, and easing the process of cloud migration.
4.3. Quality Attributes and Tactics (RQ3)
The exploration and implementation of quality attributes and associated tactics are critical to the successful deployment of containerized applications in multi-cloud environment, shaping performance, security, and compatibility characteristics. In the following, we will discuss the key findings of our study, providing insights into the QAs and tactics specifically tailored for containerized applications operating within multi-cloud environment.
Performance Optimization in Multi-Cloud Environment: Our study identifies the role of machine learning, container technology, and location-aware service brokering in enhancing the performance and efficiency of containerized applications within multi-cloud environment. Hashem et al. previously suggested similar efficiency gains from machine learning and container technology, validating the interpretive value of our results (Hashem et al., 2016). Specifically, the application of machine learning aligns with a broader trend towards intelligent cloud resource management, a domain that Buyya et al. identified as significant in cloud computing research (Buyya et al., 2020). Meanwhile, location-aware service brokering emerges as a novel and impactful strategy, stressing the role of geographical considerations in further optimizing containerized applications’ performance. Implications: For researchers, the results of our study offer opportunities to explore the integration of AI techniques with container-based cloud architectures further, especially the application of location-aware service brokering in multi-cloud environment. For practitioners, the results indicate that a adaptable approach involving AI, container technology, and geographical considerations can lead to significant performance enhancements in multi-cloud environment.
Security Measures in Multi-Cloud Deployments: The results of our study indicates the multi-faceted approach to ensuring security in multi-cloud deployments. It highlights the combination of encryption, network security protocols, machine learning, and user-specified security measures, while also encouraging deployment across different cloud providers. This finding complements Singh et al.’s argument for a comprehensive approach to cloud security, and aligns with the recent concept of distributed cloud security (Singh et al., 2016). Our results also report the importance of user-specified security measures, which are in agreement with the user-centric security model proposed by Shin and Dong-Hee (Shin, 2013). Implications: From a research perspective, the interact between different security measures and their effect on overall system security in multi-cloud environment warrants further investigation. For practitioners, these results point out the importance of employing a comprehensive security approach in multi-cloud deployments, including encryption, network security, machine learning, and user-specific security measures.
Ensuring Compatibility Across Different Cloud Providers: The results of our study suggests strategies to ensure compatibility in multi-cloud environment through standardization of interfaces, componentization, and interoperability approaches. Emphasizing lightweight communication protocols and interoperability between clouds, the results related to compatibility mirrors Mell and Grance’s principles of service-oriented architecture, reinforcing the need for effective communication across different cloud platforms (Mell et al., 2011). Petcu et al. have also previously stressed the importance of standardization and componentization for ensuring compatibility (Petcu et al., 2018). Implications: From a research standpoint, the dynamic between lightweight communication protocols, interface standardization, and overall system compatibility in multi-cloud deployments offers fertile ground for future work. Practitioners, meanwhile, can adopt these approaches to ensure seamless communication and interoperability across different cloud providers.
These findings not only confirm existing literature results but also offer valuable insights into the practical application of machine learning, container technology, and location-aware services in multi-cloud environment. Future research could further explore the intersection of these quality attributes to develop more robust, efficient, and secure multi-cloud applications by using containerise technologies.
4.4. Automation Challenges and Solution Framework (RQ4 - RQ5)
The Automation Challenge-Solution Framework for Containerized Applications in multi-cloud environment (see Table 12) suggests that multi-cloud management for containerized application is a complex task encompassing several dimensions. These include multi-cloud automation, deployment and scaling, resource management, data and application migration, testing and bench marking, standardization and interoperability, application and service management, and runtime and service discovery. This indicates that addressing one challenge may potentially impact the other areas, necessitating a holistic approach to tackle these issues. In the following we discuss the some of the key observations about this framework.
Multi-dimensional Challenges: The Automation Challenge-Solution Framework for Containerized Applications in multi-cloud environment suggests that multi-cloud management is a complex task encompassing several dimensions. These include multi-cloud automation, deployment and scaling, resource management, data and application migration, testing and bench marking, standardization and interoperability, application and service management, and runtime and service discovery. This indicates that addressing one challenge may potentially impact the other areas, necessitating a holistic approach to tackle these issues. The multi-dimensionality of the challenges implies the need for an integrated approach that considers the interdependencies among these dimensions. It emphasizes the need for solutions that cater to this complexity rather than address individual challenges in isolation. Implications: For researchers, these findings underscore the need to develop multi-faceted, holistic solutions that can address the various interconnected challenges in managing multi-cloud environment. For practitioners, this implies the importance of strategic planning and implementing solutions that address the multiple facets of multi-cloud management while considering their potential impact on one another.
Emergence of AI/ML Solutions: The framework highlights the increasing utilization of AI and ML in providing intelligent solutions to manage containerized applications in multi-cloud environment. This suggests the growing maturity and applicability of these technologies in complex scenarios, a significant step towards achieving efficient multi-cloud management. Previous studies (e.g., (de Laat et al., 2020), (Chen et al., 2021), (Gill et al., 2022)) have recognized the potential of AI and ML in resolving complex cloud management issues, which is in line with the findings from the framework. This further validates the significant role AI and ML have started to play in this domain. The increasing adoption of AI/ML in multi-cloud environment signifies an essential evolution in the field. This stresses the need for further research and development in these technologies to exploit their full potential in resolving multi-cloud management complexities. Implications: The growing use of AI/ML solutions implies that researchers need to focus on improving these technologies and adapting them to specific multi-cloud management challenges. For practitioners, this points towards the increasing necessity to invest in AI/ML capabilities and integrate them into their multi-cloud management strategies.
Importance of Standardization and Interoperability: The framework underscores the crucial role of standardization and interoperability in multi-cloud environment, facilitating seamless integration and management across different cloud platforms. Recent literature validates the importance of standardization and interoperability in multi-cloud environment, emphasizing their role in enhancing the efficiency of cloud services (Alonso et al., 2023a). This agrees with our results, further confirming the critical need for these factors in the domain. The need for standardization and interoperability reiterates the necessity for a unified framework or protocol across different cloud environments. It suggests that attention to these aspects can lead to more seamless, efficient, and secure management of multi-cloud services. Implications: For researchers, the results stress the need to investigate methods to achieve better standardization and interoperability in multi-cloud environment. For practitioners, the findings suggest the need to adopt standards and focus on interoperability while selecting and implementing cloud services, ensuring more efficient management across different platforms.
4.5. Security Challenges and Solutions Framework (RQ4 - RQ5)
The framework proposed in our study provides a comprehensive and systematic approach to addressing the complex security issues associated with containerized applications in multi-cloud environment. In the following, we discuss the key takeaway from security challenges and solution framwrork presented in Section 3.5 and Table 11.
Emphasize on Container-Specific Security Mechanisms: Security solutions specifically designed for containerized applications are critical to managing threats in multi-cloud environment. The selected studies suggest a variety of methods, including secure container orchestration, container isolation, and advanced deployment strategies that make use of deep reinforcement learning. This aligns with the existing literature (e.g., (Zhong et al., 2022), (Li et al., 2023) (Combe et al., 2016)) which highlights the unique security considerations brought forth by containerization. Containers, being lighter than traditional virtual machines and offering process-level isolation, have revolutionized the way applications are packaged and run. However, they also introduce new security challenges that need to be addressed by using container-specific security mechanisms. Containerization, while a powerful tool for application deployment, necessitates its own suite of security strategies. Implementing these will be key to leveraging the benefits of containerization in a secure manner. Implications: For researchers, this highlights an area of potential study: the development of advanced container-specific security strategies, including those that use advanced techniques such as machine learning. For practitioners, this suggests the need to be well-versed in the specifics of container security. Leveraging features of containerization platforms, such as Docker’s isolation features, or incorporating tools specifically designed for container security will be crucial.
Implementing Multi-Layered Security and Compliance Measures: Implementing multi-layered security measures that encompass data protection, access control, and communication security, along with compliance with legal and regulatory requirements, is of utmost importance. Solutions range from data encryption, role-based access control, secure container orchestration, adherence to regulations like GDPR (Regulation, 2016). This is consistent with current cybersecurity frameworks, such as the NIST cybersecurity framework (of Standards and Technology, 2018), which emphasize a multi-layered approach to security. Furthermore, several studies point to the criticality of data security, user access control, and secure communication in cloud environments (Austin, 2018) (Singh et al., 2016). In a multi-cloud, containerized environment, it is not enough to secure data; access and communication channels must also be secured, and all actions must comply with legal and regulatory requirements. Implications: For practitioners is to design and implement a comprehensive security strategy that encompasses multiple layers of protection. Compliance with legal and industry standards should be an integral part of this strategy. For researchers, this underlines the need for studies that address the integration of various layers of security and the development of comprehensive security frameworks.
Ensuring Security through Standardization and Interoperability: Ensuring security through standardization and interoperability is highlighted as a critical step. Secure and standardized interfaces, such as OpenStack APIs, can be used to enhance security and performance across multiple cloud providers. This finding is aligned with existing literature, which emphasizes the role of standardization in achieving security and interoperability in multi-cloud environment (Buyya et al., 2010) (Petcu, 2011). By implementing standardized interfaces and practices, organizations can ensure a level of interoperability and security across cloud environments. Implications: For practitioners, this means choosing platforms and tools that support standardized interfaces and practices, or working to implement those standards within their own environments. Researchers, on the other hand, could focus on the development of new standards or the improvement of existing ones to better address security challenges in multi-cloud, containerized environments.
4.6. Deployment Challenges and Solutions Framework (RQ4 - RQ5)
Table 14 offers a comprehensive overview of challenges and their corresponding solutions encountered during the deployment of containerized applications in multi-cloud environment. In the following, we provide a discussion on key findings pertaining to the challenges and solutions related to the deployment of containerized applications in multi-cloud environment.
Standardization and Automation in Deployment: Standardization and automation stand out as foundational pillars for achieving streamlined and uniform deployments in multi-cloud landscapes. As Raj et al. (Raj et al., 2018) highlighted, embracing automated deployment pipelines paired with standardized toolsets can significantly diminish complexity and mitigate human-induced errors within multi-cloud frameworks. The challenges, captured under Performance Testing Deployment, Deployment Validation, and Cloud Management Platform Evaluation Deployment in the table, elucidate the nuanced complexities associated with deployments spanning multiple cloud platforms. Solutions such as Standardized Benchmarking and Automation, Testing Process Management System, and Standardized Output Formats and Evaluation Criteria emphasize the imperative nature of adopting uniform procedures and harnessing automated solutions. Multi-cloud deployments, inherently intricate, benefit profoundly from standardization, bringing forth predictability and coherence, while automation paves the way for diminishing human-driven inconsistencies and bolstering deployment speeds. Implications: For researchers, the quest for optimized frameworks championing standardization and automation within multi-cloud environment presents a promising research trajectory. For practitioners, adeptness with standardized methodologies and tools, combined with proficiency in automation implementation, becomes crucial in navigating the multi-cloud management maze efficiently
Security and Compliance in multi-cloud environment: Security remains paramount in cloud deployments. Fernandes et al. underscore that the labyrinth of diverse platforms and differing standards in multi-cloud configurations amplifies security complexities (Fernandes et al., 2014). Challenges articulated as Security Deployment in Multi-container Environment and Malicious Service Management Deployment in Table 13 accentuate the primacy of security considerations. Propounded solutions like Cross-Container Isolation and Collusion-Resilient Trust Aggregation Technique spotlight the industry’s shift towards advanced and meticulous security protocols within multi-cloud paradigms. The ever-evolving landscape of multi-cloud deployments demands security solutions that are not only rigorous but also malleable and formidable. To attain comprehensive security across diverse platforms, strategies must encapsulate both depth of understanding and breadth of application. Implications: For the academic community, pioneering avant-garde security solutions apt for multi-cloud environment offers a fertile domain of inquiry. Meanwhile, practitioners should enshrine principles of data integrity, trust assurance, and vulnerability mitigation at the heart of their multi-cloud deployment blueprints.
Integration, Monitoring, and Infrastructure Management: As highlighted by Ferrer et al. (Ferrer et al., 2016) the ability to operate seamlessly across disparate cloud providers and maintain flexible infrastructures is paramount in harnessing the full potential of multi-cloud deployments. The presented challenges, including Network Latency Deployment and Vendor Lock-in Prevention Deployment, underscore the indispensability of agility and cross-platform compatibility. Proposed solutions such as VM Selection for Composite Applications and Interoperability Layer Above Cloud Infrastructure accentuate the importance of crafting adaptable infrastructure and formulating strategies to mitigate vendor entrenchment. In the context of multi-cloud environment, flexibility and interoperability transcend mere advantageous attributes; they become pivotal determinants shaping adaptability, scalability, and the sustained success of deployment blueprints. Implications: For the academic community, delving into novel approaches that champion infrastructure pliability and seamless operation across diverse cloud platforms emerges as a prospective research avenue. For industry professionals, anchoring strategies around flexibility and interoperability not only optimizes current deployments but also fortifies them against future technological evolutions and shifting business dynamics.
4.7. Monitoring Challenges and Solutions (RQ4 - RQ5)
Table 14 presents a comprehensive view of the Monitoring Challenge-Solution Framework for Containerized Applications in multi-cloud environment. It covers a wide range of monitoring challenges within the intricate multi-cloud setting. In the following, we are providing discussion on the key findings.
Enhancing Performance Stability through Multi-Cloud Monitoring: In the context of multi-cloud environment, ensuring consistent and reliable performance monitoring for containerized applications can be challenging due to the dynamic nature of cloud resources. Existing studies (e.g., (Zhong et al., 2022) (Mungoli, 2023)) highlighted the impact of varying resource availability on the performance of containerized applications across multiple clouds. They emphasized the need for performance benchmarking to track fluctuations and maintain a consistent user experience. The solutions presented in our framework, such as benchmarking performance variation and utilizing Kubernetes for monitoring, align with existing studies suggestion for benchmarking. Our proposed framework expands on this by incorporating Kubernetes for improved multi-cloud performance monitoring. Our framework also suggest reliable performance monitoring requires a combination of standardized benchmarking techniques and cloud-specific tools like Kubernetes. This approach can effectively address performance variations and ensure a consistent user experience across diverse cloud environments. Implications: For practitioners, adopt benchmarking practices to track performance fluctuations, and leverage Kubernetes for unified multi-cloud performance monitoring. For researchers, explore further refinements in benchmarking methodologies and investigate the impact of dynamic multi-cloud environment on performance consistency.
Streamlining Resource and Network Management across Multi-Clouds: Managing network communications, system images, and resource utilization in a multi-cloud setting poses challenges due to the diversity of cloud platforms. Sedghpour et al. discussed the complexities of managing network communications in multi-cloud environment. They emphasized the need for standardized approaches like service mesh to enhance network visibility and control (Sedghpour and Townend, 2022). Solutions in our proposed framework, such as using Docker Server for containerized services and Prometheus/Grafana for resource monitoring, align with Sedghpour et al.’s (Sedghpour and Townend, 2022) suggestion for standardized approaches. Additionally, our framework extends these solutions to address broader resource management challenges. Employing standardized tools like Docker Server, along with platforms like Prometheus and Grafana, can help streamline resource management and network communication monitoring across diverse cloud platforms. Implications: For practitioners, implement Docker Server and utilize platforms like Prometheus and Grafana for efficient resource monitoring and management across multi-cloud environment. For researchers, investigate further integration possibilities for standardized resource management tools and explore the impact of such tools on overall system performance and reliability.
Strengthening Security, Compliance, and Trust: Ensuring data privacy, continuous security control, and compliance in multi-cloud containerized applications is essential but challenging due to the distributed nature of cloud resources. Existing studies (e.g., (Alonso et al., 2023a) (Raj et al., 2018)) discussed the challenges of ensuring security and compliance across multi-cloud environment. They highlighted the importance of continuous monitoring and enforcement mechanisms to address security and compliance gaps. Solutions presented in our proposed framework, such as using machine learning for performance supervision and employing security assurance platforms, are aligned with existing studies that emphasis on continuous monitoring and enforcement for security and compliance. Implementing advanced techniques like machine learning for monitoring and adopting security assurance platforms can significantly enhance security, compliance, and trust management in multi-cloud containerized applications. Implications: For practitioners, integrate machine learning-based monitoring and adopt security assurance platforms to ensure continuous security control and compliance across diverse cloud environments. For researchers, explore the scalability and effectiveness of machine learning-driven monitoring techniques and investigate novel methods for enforcing security and compliance in multi-cloud setups.
4.8. Tools and Frameworks (RQ6)
This study report 160 distinct tools and frameworks for container based applications in multi-cloud environment, all of which are systematically classified into 6 categories, further delineated into 46 subcategories. In the following we briefly discuss some of the key findings about identified tools and frameworks.
Diverse Cloud Services by Leading Providers: Our study reveals a diverse array of cloud services, including IaaS, PaaS, and SaaS offerings from major providers like AWS, Azure, and Google Cloud, emphasizing the extensive options available to developers (see Table 15). This underscores the significant range of choices available to developers. Noteworthy cloud service providers highlighted in our study include Amazon EC2, Google Cloud App Engine, and Azure, further underscoring this finding. A pivotal insight gleaned from our research underscores the paramount importance of harmonizing cloud service models with specific application prerequisites, thereby facilitating efficient allocation of resources and optimization of costs. Implications: For practitioners, our outcomes underscore the critical necessity of meticulously evaluating and selecting the most fitting service model based on the unique demands of a given project. This strategic approach ensures the best alignment between technology and goals. Moreover, the implications for researchers are promising: an avenue emerges for in-depth exploration of trade-offs and the identification of optimal practices while navigating the intricate array of cloud service models. This exploration holds potential for enhancing both the efficiency and effectiveness of cloud solutions.
Docker and Kubernetes: Pioneers of Cross-Cloud Containerization: Our investigation spotlights the profound significance of Docker and Kubernetes in streamlining the containerization and orchestration of applications across a spectrum of cloud platforms. Docker adeptly encapsulates applications and their dependencies within self-contained containers, while Kubernetes takes center stage as the orchestration juggernaut, automating deployment, scaling, and management of these containers. These technologies receive direct validation through table references, serving as prime exemplars. Our research yields a compelling revelation: Docker and Kubernetes hold pivotal roles in ensuring consistent application deployment and seamless scalability, effectively simplifying the complexities of multi-cloud environment. Implications: Practitioners stand to gain significantly by adopting Docker and Kubernetes, as this potent combination bolsters efficiency and cultivates heightened portability. Meanwhile, researchers have a promising avenue to explore deeper facets, including security, performance optimization, and the intricate domains of advanced use cases for Docker and Kubernetes within multi-cloud landscapes. This exploration promises valuable insights that can finely calibrate the potential of these cutting-edge technologies.
The Rise of DevOps: Embracing Automation Tools: Our exploration spotlights the important role of DevOps practices and tools such as Puppet, Chef, and Terraform within contemporary software development landscapes. This emphasis revolves around the power of automation in facilitating streamlined deployment and management processes. These tools, by enabling consistent provisioning and deployment of infrastructure, effectively curtail errors and ensure the replicability of processes. The table substantiates this notion through references to Puppet, Chef, and Terraform. Our study unearths a pivotal insight: DevOps practices wield the ability to augment the software development life cycle by automating tasks, ushering in a realm of efficient and dependable deployment. Implications: Practitioners are advised to seamlessly integrate DevOps principles into their workflows by harnessing the prowess of these tools. Furthermore, a realm of exploration opens for researchers, inviting deeper analyses into the evolutionary trajectory of DevOps practices and their profound impact on elevating software quality and hastening development velocities.
5. Threats to Validity
This systematic mapping study is susceptible to various threats that could influence its outcomes. To address these potential threats, we adhered to the established guidelines for conducting SMSs and SLRs as outlined in (Keele et al., 2007) and (Petersen et al., 2008b). We further categorized and analyzed the validity threats to our study based on the four distinct types of validity threats mentioned in (Zhou et al., 2016) and (Wohlin et al., 2012). In the subsequent sections, we delve into the specific validity threats that pertain to the different phases of this SMS.
5.1. Internal Validity
Internal validity refers to the factors that could affect the analysis of the data extracted from the selected studies. The threats to internal validity could happen in the following steps of this SMS:
  • Study Search: The potential to overlook studies during the search process requires careful measures. To mitigate this risk, we used a combination of primary and snowballing search methodologies, as detailed in Section 2.2.1. To increase the collection of primary studies during the initial search phase, we also took additional measures to address search-related issues. Specifically, we improved our search strings through pilot searches before applying them to the databases. This process helped to create a more effective search strategy.
  • Study Selection: We have outlined the study screening and selection process in Table 3. To ensure objectivity and eliminate personal bias in study selection, we implemented a two-phase approach: (i) initial study screening, and (ii) qualitative evaluation of the shortlisted studies. During this procedure, the lead two authors conducted the study screening based on the criteria explicitly detailed in Section 3. To assess the objectivity of the screening process and the level of agreement between the two authors, we applied Cohen’s Kappa (Cohen, 1968). In cases where the lead authors did not agree, the second and third authors independently reviewed the disputed studies to reach a consensus. All the researchers involved in this study have extensive knowledge and research experience in containerization applications within a multi-cloud environment.
  • Data Extraction: Bias from researchers during data extraction can pose a significant challenge in both SMSs and SLRs. To address this potential issue, we developed a standardized data extraction form (see Table 5) to ensure consistent data retrieval. Initially, the first and fourth authors extracted the data. If any uncertainties arose regarding the extracted data, comprehensive discussions were convened among all authors. Following the recommendations in (Wohlin et al., 2012), a subset of the extracted data was cross-verified by the second and third authors.
  • Bias on Themes Classification: Misclassification of data and primary studies may result in biases stemming from subjective interpretation. To minimize this risk, we adhered to the thematic analysis guidelines established by Braun et al. (Braun and Clarke, 2006), and implemented a six-step process for thematic classification, as detailed in Section 2.3.
  • Data Synthesis: We employed both qualitative and quantitative approaches to assess the gathered data. Potential biases in data synthesis could influence the interpretation of our findings. To address this concern, we used open coding and constant comparison techniques from Grounded Theory (Glaser and Strauss, 1967) to analyze the qualitative data extracted from the selected studies.
5.2. External Validity
Concerns about external validity relate to the applicability and generalizability of study findings. Our research provides a thorough examination of containerization in multi-cloud environments, with findings, analyses, and conclusions specifically tailored to this domain. To ensure robust external validity, we developed a study protocol that outlines our research methodology. The literature review spanned a decade, from January 2013 to March 2023, and included peer-reviewed articles from eight preeminent databases in the fields of software engineering and computer science, which are listed in Table 2. While the review is anchored in academic research, which may not encapsulate unpublished industry practices, the systematic approach and comprehensive timeframe of our analysis make the insights valuable for both academia and practitioners. To complement and enrich our findings, future work could include an industrial study (e.g., blogs, white papers) that would provide a broader view of the application of containerization in practice
5.3. Construct Validity
Construct validity concerns to the accuracy of the operational measures used for data collection in a study. The primary constructs of this study revolve around two concepts: “containerization” and “multi-cloud environment”. Utilizing imprecise or incomplete search terms, or employing unsuitable search strategies, can lead to potential pitfalls such as overlooking pertinent papers or including numerous irrelevant ones during the search phase, and omitting relevant papers during the selection phase. To counter these risks, we implemented the following measures: (i) we initiated a pilot search to verify the relevance and comprehensiveness of our search terms; and (ii) we searched paper on eight databases renowned for computer science and software engineering research. Additionally, we customized search string according to each database syntax.
5.4. Conclusion Validity
Threats to conclusion validity relate to factors, such as data inaccuracies, that can impede drawing accurate conclusions. To mitigate these threats, we adhered to best practices, including the search protocol, pilot search, and pilot data selection, as recommended by Kitchenham et al. (Keele et al., 2007) and Petersen et al. (Petersen et al., 2008b). Moreover, to further ensure the validity of our conclusions, the authors engaged in multiple brainstorming sessions to collaboratively interpret the results and finalize the conclusions.
6. Related Work
This section reviews the most relevant existing research in terms of secondary studies such as literature surveys, state-of-the-art analysis, systematic reviews, and mapping studies that consolidate published literature on containerization in multi-cloud environments. The review focuses on (i) existing challenges, proposed solutions, and emerging trends detailed in Section 6.1 and (ii) container-based deployment of multi-cloud systems in Section 6.2.
6.1. Challenges, Solutions, and Performance of Multi-cloud Containers
In recent years, several literature surveys have been published to analyze state-of-the-art on the applications (Pahl et al., 2017), tools and technologies (Casalicchio and Iannucci, 2020), potential and limitations (Bentaleb et al., 2022), as well as emerging trends (Watada et al., 2019) of container-based solutions for cloud computing systems, discussed below.
6.1.1. Challenges, Applications, Tools, and Emerging Trends
Containers in multi-cloud environments provide a standardized and portable way to package, deploy, and run applications, ensuring seamless deployment and management across diverse cloud platforms (Alonso et al., 2023b) (Saxena et al., 2021). One of the earliest SLR-based studies by Pahl et al. (Pahl et al., 2017) on containerized clouds reviewed a total of 46 studies to identify, taxonomically classify and systematically compare the existing research on containers and their application in the context of cloud-based systems. The SLR classified and compared the selected research studies using a conceptual framework to highlight existing trends and needs for future research. SLR results indicate container orchestration, microservice delivery, continuous development and deployment as emerging trends of research on containerized clouds. The SLR in (Alonso et al., 2023b) reviewed a total of 88 studies (published from 2011 - 2021) on multi-cloud systems. Similar survey studies such as (Saxena et al., 2021) concepts, challenges, requirements and future directions for multi-cloud environments are discussed. A survey of existing approaches and solutions provided by different multi-cloud architectures is entailed along with analysis of the pros and cons of different architectures while comparing the same (Imran et al., 2020).
6.1.2. Performance and Resource Utilization
In a study by Bentaleb et al. (Bentaleb et al., 2022), the authors follow the taxonomical classification from (Pahl et al., 2017) to categorize container-based technologies for cloud systems. The study argues for the need of performance metrics to objectively define and evaluate quality attributes such as resource virtualization, service elasticity, orchestration, and multi-tenancy in containerized cloud systems. The study also highlights the needs for future research in terms of best practices, i.e., patterns and tactics (enabling reuse) and tools (supporting automation) for container-based cloud deployment. In the context of performance, the studies (Casalicchio and Iannucci, 2020) (Watada et al., 2019) report literature review and experimental analysis of factors that influence performance of container-based cloud systems. Specifically, Casalicchio et al. (Casalicchio and Iannucci, 2020) reviewed a total of 97 research studies investigating performance evaluation and run-time adaptation of container-based solution. The study highlighted several unsolved challenges such as I/O throughput optimization, performance prediction, and multi-layer monitoring performance bottlenecks. Moreover, Watada et al. (Watada et al., 2019) conducted an experimental study to compare the performance of VMs, containers and uni kernels in terms of and technological maturity using standard benchmarks and observed containers to optimise performance of container-based. The performance of containerized clouds is also determined by resource utilization, which includes but is not limited to CPU utilization, memory footprints, energy consumption, network bandwidth, and execution time.
Kapil et al. (Netaji and Bhole, 2021) reviewed 64 research papers to gain insight into resource allocation, management, and scheduling. Furthermore, limitations of existing resource allocation algorithms are discussed, indicating the need to investigate algorithms or techniques of performance optimization of containers for cloud systems. The study by Maenhaut et al. (Maenhaut et al., 2020) provided an overview of the current state of the art regarding resource management within the broad sense of cloud computing, complementary to existing surveys in the literature. The study investigated how research is adapting to the recent evolution within cloud solutions, including container technologies.
6.2. Container-based Cloud Orchestration, Deployment, and Security
This section reviews the most relevant research on the container-based deployment with a focus on orchestration and security issues of cloud-based systems, detailed below.
6.2.1. Orchestration and Deployment
Naweiluo et al. (Zhou et al., 2022) explored containerization in High Performance Computing (HPC) environments, contrasting it with cloud computing. The results of this study indicate that containers enhance application deployment efficiency, but face challenges in HPC due to high security levels and the need for extensive libraries and packages that affect cloud portability, often resulting in vendor lock-ins. The study also provides a survey and taxonomy on containerization and orchestration efforts in HPC, pointing out differences with cloud environments and identifying future research and engineering potentials. A systematic mapping study by Naylor et al.(Bachiega et al., 2018) explores virtualization in container-based cloud deployments. The mapping study, based on a comprehensive review of major databases, identifies a significant research gap in the performance evaluation of containers, underscoring the need for further investigation on cloud deployment. Carlos et al. (Awada, 2018) present an optimization strategy for deploying microservices in multi-cloud environments focusing on minimizing cloud service costs, network latency, and startup time for new microservices. Their approach uses a Non-dominated Sorting Genetic Algorithm II (NSGA-II) compared to a Greedy First-Fit algorithm, demonstrating a 300% improvement with NSGA-II. This highlights its effectiveness in container and VM orchestration, offering a significant enhancement in deployment of multi-cloud solution in containers. Emiliano et al. (Paladi et al., 2018) survey container orchestration, proposing a reference architecture for autonomic orchestrators, and identifying research challenges in the field. Their work emphasizes the importance of container technologies in cloud environments and the need for advanced orchestration solutions to manage complex multi-container applications effectively. Uchechukwu Awada (Guerrero et al., 2018a) reviews container orchestration tools and platforms, comparing the architectures, components, and capabilities of several Container Service Platforms (CSPs) and orchestration tools like Amazon ECS, Kubernetes, Docker Swarm, and Mesos. The study offers insights into the current state of container orchestration and suggests future research directions, serving as a guide for developers and organizations.
Koustabh et al. (Dolui and Kiraly, 2018) explore multi-container deployment on IoT gateways to meet the stringent latency requirements of advanced IoT applications. Through their study within the AGILE project, they highlight containerization’s role in overcoming the diversity and resource constraints of IoT gateways, showcasing containerization’s advantages for IoT gateway performance optimization. Their research underscores the potential of containerized environments in enhancing application compatibility, portability, and efficient deployment across diverse hardware architectures. Matteo et al. (Nardelli et al., 2018) introduce the Adaptive Container Deployment (ACD) model to optimize containerized application deployment in geo-distributed environments, focusing on IoT and fog computing resources. ACD, formulated as an Integer Linear Programming problem, aims to improve application performance by leveraging containers’ horizontal and vertical elasticity. The study evaluates ACD’s effectiveness against greedy heuristics, highlighting the need for advanced orchestration solutions to exploit emerging computing environments’ characteristics efficiently.
6.2.2. Security of Containerized Clouds
Nicolae et al. (Casalicchio, 2019) analyze security challenges in cloud orchestration for multi-cloud deployments, proposing a security-enabled orchestration framework. Their research identifies potential attack scenarios and security enforcement mechanisms, aiming to enhance security guarantees for cloud operators in a multi-cloud setting. Mohammad et al. (Bokhari et al., 2018) offer a detailed analysis of cloud computing, defining its essential characteristics, architecture, service models (SaaS, PaaS, IaaS), and deployment models. They discuss the security requirements for public and private clouds across different service models, aiming to provide researchers with a comprehensive understanding of cloud computing’s potential and security challenges. Sari et al. (Sultan et al., 2019) survey the landscape of container security, addressing challenges and solutions across four generalized use cases within the host-container threat landscape. Their analysis covers both software-based solutions utilizing Linux kernel features and hardware-based solutions for enhanced security. The study aims to clarify container security requirements and encourage further research in addressing potential vulnerabilities and attacks. Hendrik et al. (Graupner et al., 2015a) tackle privacy, security, and trust issues in cloud computing through a multi-cloud architecture perspective. They propose a novel technique for enhancing security and unifying access control mechanisms across cloud providers, addressing the challenges of inadequate cross-provider APIs and non-unified access control. Rohan et al. (Gupta et al., 2016) propose a security-enhancing methodology for cloud data storage using container clustering and Docker instances for on-demand encryption. This approach aims to secure data while optimizing resource usage and cost, highlighting the benefits of container technology in improving cloud computing security and efficiency.
6.3. Conclusive Summary
Current survey-based studies and systematic reviews, discussed above (e.g., (Casalicchio and Iannucci, 2020) (Pahl et al., 2017)), aim to consolidate the latest findings in published research concerning container-based solutions for cloud computing systems. These studies investigated various aspects of multi-cloud, including exploring applications, tools, and technologies, assessing potential and limitations, and identifying emerging trends.
Our study distinguishes itself from prior research in several ways. First, our work systematically categorizes container utilization in multi-cloud environments into coherent categories and themes. Some of the key themes are ‘Scalability and High Availability’, ‘Performance and Optimization’, ‘Security and Privacy’, and ‘Multi-Cloud Container Monitoring and Adaptation’. We identified and classified seventy-four patterns and strategies for containerization, which existing studies have not provided. Furthermore, we explored the QAs and associated tactics for containerization in multi-cloud environment. This contribution offers practitioners actionable insights into tailoring their containerization strategies to meet desired quality standards. Lastly, our study introduces four distinct frameworks consisting of challenges and solutions related to security, automation, deployment, and monitoring within the context of containerization in multi-cloud environment.
7. Conclusions
This SMS presents the current state of research regarding containerization in multi-cloud environment, focusing on aims of selected studies, the roles of containers, containers implementation strategies, architectural patterns and strategies, quality attributes and corresponding tactics. Additionally, this SMS also explores deployment, monitoring, security challenges, accompanied by their respective solutions, tools, and frameworks used to implement containerized applications in multi-cloud environment. From an initial pool of extensive search results, 86 relevant studies are were finally chosen to answer the research questions presented in Table 1. The key findings of this SMS are listed below.
  • The findings of this SMS reveal various insights regarding SMS demographics, publishers, publication types, authors’ affiliations, and research themes. The yearly distribution reached its peak in 2016, signifying a shifting level of interest over time. The field of publishing is largely dominated by IEEE (72.09%), with conferences emerging as the primary avenue for publication. Notably, academia (93.02%) stands out as the leading affiliation among authors. The prominent research themes encompass Orchestration and Management, Scalability and High Availability, Performance and Optimization, and Security and Privacy. These themes underscore crucial aspects such as container management, scalability, performance, and security within the context of multi-cloud environment.
  • This SMS identifies and classifies a total of 74 patterns and strategies across 10 subcategories and 4 categories for container-based applications in multi-cloud environment. These patterns and strategies encompass a wide range of aspects such as cloud architecture models, communication and networking, deployment, service models, cloud management, container management, edge computing, IoT strategies, security, resilience, fault-tolerance, and cloud migration. These findings are significant for the application and dvancement of containerization-based applications in multi-cloud environment, providing valuable insights into optimizing architecture, communication, management, security, and migration strategies.
  • This SMS identifies and classifies QAs and tactics from selected studies regarding containerized applications in multi-cloud environment. A total of ten QAs and their related terms are identified. Along with these QAs, a total of 47 tactics are identified to enhance the QAs. Notable takeaways include the utilization of machine learning and location-aware service brokering to enhance the performance and efficiency of container-based applications in multi-cloud environment.
  • This SMS presents a comprehensive security challenge-solution framework for containerized applications in multi-cloud environment. This framework offers valuable insights and actionable strategies for practitioners and researchers to enhance the security of containerized applications, address complex multi-cloud security concerns, and ensure compliance with various regulatory requirements, thereby enabling the development of more robust and secure containerized applications in multi-cloud environment.
  • This SMS presents a framework for the automation challenge-solution framework for containerized applications in multi-cloud environment. The automation challenge-solution framework offers a comprehensive guide for practitioners and researchers to address the complex challenges associated with automating various aspects of deploying, managing, scaling, testing, migrating, and ensuring interoperability of containerized applications in multi-cloud environment. By offering specific solutions for each challenge, this framework aids in the development of more efficient, secure, and adaptable automation processes, empowering practitioners to streamline their operations and enabling researchers to advance the field of multi-cloud containerization.
  • The third framework presented in this SMS is the deployment challenge-solution framework for containerized applications in multi-cloud environment. This framework provides valuable insights and practical guidance to practitioners and researchers alike, aiding practitioners in navigating the complexities of multi-cloud deployment through specific solutions to challenges, while also serving as a comprehensive resource for researchers to delve into various aspects of containerized application deployment, orchestration, security, scalability, and more, thus contributing to the advancement of knowledge and innovation in the field of multi-cloud deployment.
  • Fourth framework presented in this SMS is monitoring challenge-solution framework for containerized applications in multi-cloud environment. This framework serves as a comprehensive guide for practitioners, offering specific solutions to challenges related to monitoring performance, resource management, security, optimization, system complexity, and multi-cloud coordination, facilitating their understanding of complex monitoring scenarios and providing actionable insights. Furthermore, researchers can benefit from this framework by gaining a deeper insight into the nuances of monitoring containerized applications in multi-cloud environment, and it can serve as a foundation for further research and innovation in the field of multi-cloud monitoring.
The findings of this SMS will benefit researchers who are interested in understanding the state of research on containerized applications in multi-cloud environment and conducting further investigations to address the open research issues highlighted in Section 4. Additionally, the insights gained from this SMS will support knowledge transfer to practitioners by providing insights into the challenges, solutions, and methods for monitoring, securing, and optimizing containerized applications in multi-cloud environment. We emphasize the importance for practitioners to develop targeted solutions that effectively tackle monitoring, security, and performance degradation concerns within multi-cloud environment deployment. As a future endeavor, we intend to enhance our SMS by conducting industrial case studies with companies, thereby obtaining practical insights and perspectives from practitioners on the effectiveness and applicability of our proposed frameworks in real-world scenarios. This approach will allow us to bridge the gap between research and practical implementations concerning containerized applications in multi-cloud environment and contribute to the advancement of both academic and industry understanding in this domain.
Acknowledgments
This work is partially sponsored by the National Natural Science Foundation of China (NSFC) with Grant No. 62172311.
8. Appendices
List of selected studies for containerization in multi-cloud environments
Study ID Title Venue Year Reference
[…]
Study ID Title Venue Year Reference
[…]
A component-based adaptation approach for multi-cloud applications
IEEE Conference on Computer Communications (INFOCOM) Workshops
2014 (Almeida et al., 2014)
S02
A Cost-Efficient Multi-cloud Orchestrator for Benchmarking Containerized Web-Applications
International Conference on Web Information Systems Engineering (WISE)
2019 (Jha et al., 2019)
S03
A framework to support multi-cloud collaboration
IEEE World Congress on Services (SERVICES)
2020 (Hua et al., 2020)
S04
A Machine Learning Model for Detection of Docker-based APP Overbooking on Kubernetes
International Conference on Communications (ICC)
2021 (Ramos et al., 2021)
S05
A multi-cloud service mesh approach applied to Internet of Things
Annual Conference of Industrial Electronics Society (IECON)
2022 (Gattobigio et al., 2022)
S06
A Secure Container Placement Strategy Using Deep Reinforcement Learning in Cloud
International Conference on Computer Supported Cooperative Work in Design (CSCWD)
2022 (Deng et al., 2022)
S07
A Service Driver Based Application Execution and Development Method in Multi-cloud Context
International Conference on Data Science and Computer Application (ICDSCA)
2021 (Zhang et al., 2021)
S08
An adaptable framework to deploy complex applications onto multi-cloud platforms
International Conference on Computing Communication Technologies - Research, Innovation, and Vision for Future (RIVF)
2015 (Pham et al., 2015)
S09
An agile container-based approach to TaaS
International Congress Future Telecommunications: Infrastructure and Sustainability (FITCE)
2016 (Verdugo et al., 2017)
S10
An Efficient and Reliable Multi-Cloud Provider Monitoring Solution
Conference on Global Communications (GLOBECOM)
2022 (Sabbioni et al., 2020)
S11
An Efficient Multi-Keyword Search Scheme over Encrypted Data in Multi-Cloud Environment
International Conference on Smart Cloud (SmartCloud)
2015 (He et al., 2022)
S12
An Empirical Investigation of a Fault Tolerant Containerized Application Deployment
International Conference on Informatics (ICI)
2015 (Bisht and Kaur, 2022)
S13
Applying Computational Intelligence for enhancing the dependability of multi-cloud systems using Docker Swarm
Symposium Series on Computational Intelligence (SSCI)
2017 (Naik, 2016a)
S14
Automatically Design Heuristics for Multi-Objective Location-Aware Service Brokering in Multi-Cloud
International Conference on Services Computing (SCC)
2020 (Chen et al., 2022)
S15
Autonomic fine-grained migration and replication of component-based applications across multi-clouds
National Foundation for Science and Technology Development Conference on Information and Computer Science (NICS)
2022 (Pham and Pham, 2015)
S16
Band-Area Resource Management Platform and Accelerated Particle Swarm Optimization Alg. for Container Deployment
IEEE Access
2016 (Ouyang et al., 2022)
S17
Building a Multi-tenant Cloud Service from Legacy Code with Docker Containers
International Conference on Cloud Engineering (IC2E)
2022 (Slominski et al., 2015)
S18
Building cloud native application - Analysis for multi-component application deployment
International Conference on Computer Communication and Informatics (ICCCI)
2021 (Pai and Kumar, 2021)
S19
CLAMS: Cross-layer Multi-cloud Application Monitoring-as-a-Service Framework
International Conference on Services Computing (SCC)
2014 (Alhamazani et al., 2014)
S20
Cloud MF: Applying MDE to Tame the Complexity of Managing Multi-Cloud Applications
International Conference on Utility and Cloud Computing (UCC)
2014 (Ferry et al., 2014)
S21
Cluster as a Service: A Container Based Cluster Sharing Approach with Multi-user Support
IEEE Symposium on Service-Oriented System Engineering (SOSE)
2016 (Cui et al., 2016)
S22
Container Deployment Strategy for Edge Networking
Workshop on Middleware for Edge Clouds and Cloudlets (MECC)
2019 (Wong et al., 2019)
S23
Container-Based Performance Isolation for Multi-Tenant SaaS Applications in Micro-Service Architecture
Journal of Physics: Conference Series
2020 (Wang et al., 2020)
S24
Containerization and the PaaS Cloud
IEEE Cloud Computing
2015 (Pahl, 2015)
S25
ContainerLeaks: Emerging Security Threats of Information Leakages in Container Clouds
International Conference on Dependable Systems and Networks (DSN)
2017 (Gao et al., 2017)
S26
ContMEC: An Architecture of Multi-access Edge Computing for Offloading Container-Based Mobile Applications
IEEE International Conference on Communications (ICC)
2022 (Watanabe et al., 2022)
S27
DECIDE: An Extended DevOps Framework for Multi-Cloud Applications
International Conference on Cloud and Big Data Computing (ICCBDC)
2022 (Alonso et al., 2019)
S28
Delivering Elastic Containerized Cloud Applications to Enable DevOps
International Workshop on Software Engineering for Adaptive and Self-Managing Systems (SEAMS)
2017 (Barna et al., 2017)
S29
Deployment of Multi-Tier Application on Cloud and Continuous Monitoring using Kubernetes
International Conference on Electrical, Electronics, Communication, Computer Technologies and Optimization Techniques (ICEECCOT)
2021 (Gupta et al., 2021)
S30
Deployment-Time Multi-Cloud Application Security
International Conference on Smart Computing (SMARTCOMP)
2017 (Sheridan et al., 2017)
S31
DEVSML 3.0 Stack: Rapid deployment of DEVS Farm in distributed cloud environment using microservices and containers
Symposium on Theory of Modeling and Simulation (TMS/DEVS)
2017 (Mittal and Risco-Martín, 2017)
S32
Disaster Recovery and Business Continuity for Database Services in Multi-Cloud
International Conference on Computer Applications & Information Security (ICCAIS)
2018 (Al-shammari and Alwan, 2018)
S33
Dynamic On-Demand Fog Formation Offering On-the-Fly IoT Service Deployment
IEEE Transactions on Network and Service Management
2020 (Sami and Mourad, 2020)
S34
Dynamic security assurance in multi-cloud DevOps
IEEE Conference on Communications and Network Security (CNS)
2017 (Rios et al., 2017)
S35
Engineering and Experimentally Benchmarking a Container-based Edge Computing System
International Conference on Communications (ICC)
2020 (Carpio et al., 2020)
S36
FogDocker: Start Container Now, Fetch Image Later
International Conference on Utility and Cloud Computing (UCC)
2019 (Lorenzo et al., 2019)
S37
Importance of Application-Level Resource Management in Multi-Cloud Deployments
International Conference on Cloud Engineering (IC2E)
2019 (Dimitrijevic et al., 2019)
S38
Knowledge-Engineered Multi-Cloud Resource Brokering for Application Workflow Optimization
IEEE Transactions on Network and Service Management
2022 (Pandey et al., 2022)
S39
Label-affinity-Scheduler: Considering Business Requirements in Container Scheduling for Multi-Cloud and Multi-Tenant
Brazilian Symposium on Computing System Engineering (SBESC)
2022 (Altran et al., 2022)
S40
Leveraging kernel security mechanisms to improve container security: A survey
International Conference on Availability, Reliability and Security (ARES)
2019 (Bélair et al., 2019)
S41
Location-Aware and Budget-Constrained Service Deployment for Composite Applications in Multi-Cloud Environment
IEEE Transactions on Parallel and Distributed Systems
2020 (Shi et al., 2020)
S42
MicroCloud: A Container-Based Solution for Efficient Resource Management in the Cloud
International Conference on Smart Cloud (SmartCloud)
2016 (Baresi et al., 2016)
S43
Mobile Edge as Part of the Multi-Cloud Ecosystem: A Performance Study
Euromicro International Conference on Parallel, Distributed and Network-Based Processing (PDP)
2019 (Dreibholz et al., 2019)
S44
Modeling and Automated Deployment of Serverless Applications Using TOSCA
Conference on Service-Oriented Computing and Applications (SOCA)
2018 (Wurster et al., 2018)
S45
Monitoring Elastically Adaptive Multi-Cloud Services
IEEE Transactions on Cloud Computing
2018 (Trihinas et al., 2015)
S46
Monitoring-Aware Optimal Deployment for Applications Based on Microservices
IEEE Transactions on Services Computing
2021 (Fadda et al., 2019)
S47
MSSF: A Step towards User-Friendly Multi-cloud Data Dispersal
International Conference on Cloud Computing (CLOUD)
2014 (Libardi et al., 2014)
S48
Multi-agent QoS-aware autonomic resource provisioning framework for elastic BPM in containerized multi-cloud environment
Journal of Ambient Intelligence and Humanized Computing
2022 (Saif et al., 2023)
S49
Multicloud- auto scale with prediction and delta correction algorithm
International Conference on Intelligent Computing, Instrumentation and Control Technologies (ICICICT)
2019 (Harwalkar et al., 2019b)
50
Multi-cloud Distribution of Virtual Functions and Dynamic Service Deployment: Open ADN Perspective
IEEE International Conference on Cloud Engineering (IC2E)
2015 (Bhamare et al., 2015)
S51
Multi-cloud resource provisioning with Aneka: A unified and integrated utilization of Microsoft Azure and Amazon EC2
International Conference on Computing and Network Communications (CoCoNet)
2015 (Buyya and Barreto, 2015)
S52
MUSA Deployer: Deployment of Multi-cloud Applications
IEEE International Workshop on Enabling Technologies: Infrastructure for Collaborative Enterprises (WETICE)
2017 (Casola et al., 2017)
S53
NoMISHAP: A Novel Middleware Support for High Availability in Multicloud PaaS
IEEE Cloud Computing
2017 (Acquaviva et al., 2017)
S54
Object storage in the cloud and multi-cloud: State of the art and the research challenges
IEEE International Conference on Automation, Quality and Testing, Robotics (AQTR)
2018 (Bucur et al., 2018)
S55
On the Deployment of an Open-Source, 5G-Aware Evaluation Testbed
International Conference on Mobile Cloud Computing, Services, and Engineering (MobileCloud)
2018 (Bolivar et al., 2018)
S56
OpenStack Federation in Experimentation Multi-cloud Testbeds
IEEE International Conference on Cloud Computing Technology and Science (CloudCom)
2013 (del Castillo et al., 2013)
S57
Orchestrated Multi-Cloud Application Deployment in OpenStack with TOSCA
International Conference on Smart Computing (SMARTCOMP)
2017 (Tricomi et al., 2017)
S58
Orchestration Based Hybrid or Multi Clouds and Interoperability Standardization
IEEE International Conference on Cloud Computing in Emerging Markets (CCEM)
2018 (Sitaram et al., 2018)
S59
Performance Analysis of Zero-Trust multi-cloud
IEEE International Conference on Cloud Computing (CLOUD)
2021 (Rodigari et al., 2021)
S60
Policies Based Container Migration Using Cross-Cloud Management Platform
International Conference on Information and Automation for Sustainability (ICIAfS)
2018 (Janarthanan et al., 2018)
S61
Policy-based deployment in a hybrid and multi-cloud environment
International Conference on Cloud Computing and Services Science (CLOSER)
2019 (Di Modica et al., 2019)
S62
Predicting the End-to-End Tail Latency of Containerized Microservices in the Cloud
International Conference on Cloud Engineering (IC2E)
2019 (Rahman and Lama, 2019)
S63
Preventing vendor lock-ins via an interoperable multi-cloud deployment approach
International Conference for Internet Technology and Secured Transactions (ICITST)
2017 (Pellegrini et al., 2017)
S64
Private STaaS with OpenStack Cinder Volumes for Hybrid/Multi-cloud
International Conference on Advances in Computing and Communication Engineering (ICACCE)
2019 (Harwalkar et al., 2019a)
S65
Real-World Experience with a Multicloud Exchange
IEEE Cloud Computing
2017 (Zeck and Bouroudjian, 2017)
S66
Reliability Management for Blockchain-Based Decentralized Multi-Cloud
International Symposium on Cluster Computing and the Grid (CCGRID)
2020 (Aral et al., 2020)
S67
Reliability-Aware Cost-Efficient Scientific Workflows Scheduling Strategy on Multi-Cloud Systems
IEEE Transactions on Cloud Computing
2022 (Tang, 2021)
S68
Resource optimization of container orchestration: a case study in multi-cloud microservices-based applications
The Journal of Supercomputing
2018 (Guerrero et al., 2018b)
S69
Scalable multi-cloud platform to support industry and scientific applications
International Convention on Information and Communication Technology, Electronics and Microelectronics (MIPRO)
2018 (Kiss, 2018)
S70
Search-based Methods for Multi-Cloud Configuration
International Conference on Cloud Computing (CLOUD)
2022 (Lazuka et al., 2022)
S71
Secure access control for multi-cloud resources
IEEE Conference on ALocal Computer Networks Workshops (LCN Workshops)
2015 (Graupner et al., 2015b)
S72
Secure and Dependable Multi-Cloud Network Virtualization
International Workshop on Security and Dependability of Multi-Domain Infrastructures
2017 (Alaluna et al., 2017)
S73
Secure and Parallel Expressive Search over Encrypted Data with Access Control in Multi-CloudIoT
Cloudification of the Internet of Things (CIoT)
2018 (Guechi and Maamri, [n. d.])
S74
Self-healing cloud services in private multi-clouds
International Conference on High Performance Computing & Simulation (HPCS)
2018 (Mfula and Nurminen, 2018)
S75
SpyStorage: A Highly Reliable Multi-Cloud Storage with Secure and Anonymous Data Sharing
International Conference on Networking, Architecture, and Storage (NAS)
2017 (Shen et al., 2017)
S76
Support Services for Applications Execution in Multi-clouds Environments
International Conference on Autonomic Computing (ICAC)
2016 (Pop et al., 2016)
S77
SynAPTIC: Secure and persistent connectivity for containers
International Symposium on Cluster, Cloud and Grid Computing (CCGRID)
2017 (Ranjbar et al., 2017)
S78
System Restore in a Multi-cloud Data Pipeline Platform
International Conference on Dependable Systems and Networks – Industry Track
2019 (Wang et al., 2019)
S79
Systematic and Recomputable Comparison of Multi-cloud Management Platforms
International Conference on Cloud Computing Technology and Science (CloudCom)
2018 (Serhiienko and Spillner, 2018)
S80
Towards a lightweight multi-cloud DSL for elastic and transferable cloud-native applications
International Conference on Cloud Computing and Services Science (CLOSER)
2018 (Quint and Kratzke, 2018)
S81
Towards Formal-Based Semantic Interoperability in Multi-Clouds: The FCLOUDS Framework
IEEE International Conference on Cloud Computing (CLOUD)
2017 (Challita et al., 2017)
S82
Towards Model-Driven Provisioning, Deployment, Monitoring, and Adaptation of Multi-cloud Systems
IEEE International Conference on Cloud Computing (CLOUD)
2013 (Ferry et al., 2013)
S83
Towards Trustworthy Multi-Cloud Services Communities: A Trust-Based Hedonic Coalitional Game
IEEE Transactions on Services Computing
2018 (Wahab et al., 2016)
S84
Multi-cloud provisioning of business processes
Journal of Cloud Computing
2019 (Kritikos et al., 2019)
S85
Evaluating the DevOps Reference Architecture for Multi-cloud IoT-Applications
SN Computer Science
2021 (Bou Ghantous and Gill, 2021)
S86
Hybrid IT and Multi Cloud an Emerging Trend and Improved Performance in Cloud Computing
SN Computer Science
2020 (Gundu et al., 2020)

### v2

@@
 The aim of this research is to systematically identify and categorize the multiple aspects of containerization in multi-cloud environment.
-We conduct the Systematic Mapping Study (SMS) on the literature published between January 2013 and March 2023.
-Eighty-six studies were finally selected and the key results are: (1) Four leading themes on cloud computing and network systems research were identified: ‘Scalability and High Availability’, ‘Performance and Optimization’, ‘Security and Privacy’, and ‘Multi-Cloud Container Monitoring and Adaptation’.
-(2) Seventy-four patterns and strategies for containerization in multi-cloud environment were classified across 10 subcategories and 4 categories.
+We conducted the Systematic Mapping Study (SMS) on the literature published between January 2013 and July 2024.
+One hundred twenty one studies were selected and the key results are: (1) Four leading themes on containerization in multi-cloud environment are identified: ‘Scalability and High Availability’, ‘Performance and Optimization’, ‘Security and Privacy’, and ‘Multi-Cloud Container Monitoring and Adaptation’.
+(2) Ninety-eight patterns and strategies for containerization in multi-cloud environment were classified across 10 subcategories and 4 categories.
 (3) Ten quality attributes considered were identified with 47 associated tactics.
-(4) Four frameworks were introduced from the analysis of identified challenges and solutions: security, automation, deployment, and monitoring challenge-solution frameworks.
-The results of this SMS will assist researchers and practitioners in pursuing further studies on containerization in multi-cloud environment and developing specialized solutions.
-Containerization, Multi-Cloud Environment, Challenges, Solutions, Frameworks, Systematic Mapping Study
-journal: TOSEM
+(4) Four catalogs consisting of challenges and solutions related to security, automation, deployment, and monitoring were introduced.
+The results of this SMS will assist researchers and practitioners in pursuing further studies on containerization in multi-cloud environment and developing specialized solutions for containerization applications in multi-cloud environment.
+Containerization, Multi-Cloud Environment, Systematic Mapping Study, Cloud Computing
+journal: TOSEMccs: Software and its engineering Software development techniques
 1. Introduction
@@
 Containers are standalone and executable packages of software that include everything needed to run an application, such as code, system tools, libraries, and settings (Merkel et al., 2014).
-Containers allow the packaging of an application along with its required dependencies, which makes it easy to move the application across different environments with little or no modification (Docker, 2021).
+Containers allow the packaging of an application and its required dependencies, which makes it easy to move the application across different environments with minimal modification (Docker, 2021).
 On the other hand, multi-cloud environment is the distribution of cloud assets (Petcu, 2013).
-Using containers in multi-cloud environment allows organizations to achieve flexibility, agility, and cost efficiency.
+Using containers in multi-cloud environment allows organizations to achieve flexibility, agility, and cost-efficiency.
 Developers can build applications in a consistent environment and easily move them between multiple cloud platforms, thereby leveraging the unique strengths (e.g., extensive infrastructure, seamless integration, advanced data analytic) of each platform.
 Containerization in multi-cloud is illustrated in Figure 1.
-This figure provides an overview of applications deployed within various cloud configurations including public, private, and hybrid models.
-Each cloud hosts multiple applications encapsulated in containers, and completes different types of tasks with their required binaries and libraries, emphasizing the portability and isolation that containerization offers.
+This figure provides an overview of applications deployed within various cloud configurations, including public, private, and hybrid models.
+Each cloud hosts multiple applications encapsulated in containers, which can support different types of applications, such as those in healthcare or finance, that require flexibility in deployment and scalability across cloud providers.
+These containers complete various tasks with their required binaries and libraries, emphasizing the portability and isolation that containerization provides.
 A container platform layer, which could be exemplified by systems like Kubernetes or Docker, manages and orchestrates these containers across the different cloud environments.
+For instance, a healthcare application can use a private cloud for processing sensitive data for compliance and a public cloud for data analysis to achieve cost-efficiency and scalability.
 This architecture enables a flexible and scalable approach, allowing efficient application deployment and operations in a multi-cloud setting, while maintaining consistency and resilience across platforms.
@@
 Furthermore, the challenges can also present themselves during system testing, the coding process, and the deployment phase.
-For instance, during the architectural design phase, challenges may involve ensuring seamless integration of containerized components with existing systems and selecting appropriate container orchestration patterns, strategies, and tools that can effectively manage containers across multiple clouds (Raj et al., 2018) (Baby and Vysala, 2015), and while establishing an automated development infrastructure, organizations may encounter difficulties in creating efficient Continuous Integration/Continuous Deployment (CI/CD) pipelines that handle multi-cloud deployment scenarios while maintaining consistent performance and security standards (Debroy et al., 2018).
-Similarly, In the system implementation phase, challenges may include dealing with potential compatibility issues between containers and various cloud platforms, as well as optimizing resource utilization to control costs effectively (Helali and Omri, 2021).
+For instance, during the architectural design phase, challenges may include ensuring seamless integration of containerized components with existing systems and selecting appropriate container orchestration patterns, strategies, and tools that can effectively manage containers across multiple clouds (Raj et al., 2018) (Baby and Vysala, 2015), and while establishing an automated development infrastructure, organizations may encounter difficulties in creating efficient Continuous Integration/Continuous Deployment (CI/CD) pipelines that handle multi-cloud deployment scenarios while maintaining consistent performance and security standards (Debroy et al., 2018).
+Similarly, in the system implementation phase, challenges may include resolving potential compatibility issues between containers and various cloud platforms, as well as optimizing resource utilization to control costs effectively (Helali and Omri, 2021).
+Motivation: Our study is part of the QLEAP project (2022–2024) funded by Business Finland, which investigates the use of containers in multi-cloud environment specifically for architecture design (Ahmad et al., 2025).
+The project brings together a consortium of four companies (Bittium [^1] https://www.bittium.com/, M-Files [^2] https://www.m-files.com/, Solita [^3] https://www.solita.fi/en/, and Vaadin [^4] https://vaadin.com/), each with distinct requirements for containerization, and is further supported by industry leaders Nokia [^5] https://www.nokia.com/ and TietoEVRY [^6] https://www.tietoevry.com/.
+This study fulfills an industry demand for optimized container strategies, addressing critical gaps in deployment consistency, security, and scalability that are necessary for practical applications in multi-cloud contexts.
+Despite the growing interest in containerization for multi-cloud environment, there remains a significant gap in systematically organized knowledge to address these challenges comprehensively.
+Our study not only fulfills an immediate industry demand but also aims to provide a structured view of containerization practices, addressing the needs of both practitioners and researchers in navigating this fragmented knowledge space.
 Recent research (e.g., see Table 8) has highlighted the growing significance of container utilization in multi-cloud environment.
@@
 Additionally, these studies have explored the challenges associated with automation, deployment, monitoring, and security of containerization applications in multi-cloud environment.
-Despite the breadth of knowledge available, such valuable information is dispersed across different publications, spanning scientific research papers and the grey literature.
+Despite the breadth of knowledge available, such valuable information is dispersed across different publications, spanning scientific research papers and gray literature.
 This fragmentation of knowledge about containerization applications in multi-cloud environment presents a significant navigational challenge for practitioners tasked with real-world implementations.
 They need to thoroughly scrutinize numerous aspects, such as patterns and strategies for containerization applications in multi-cloud environment, to locate specific information relevant to their use cases, whether it be a pattern, a challenge, or a solution to an existing problem.
-Consequently, this spreading of knowledge restricts the development of a holistic understanding of the containerization applications in multi-cloud environment, leaving practitioners underprepared to devise comprehensive solutions.
+Consequently, this spreading of knowledge restricts the development of a holistic understanding of the containerization applications in multi-cloud environment, leaving practitioners underprepared to devise comprehensive, secure solutions.
 Furthermore, this state of affairs of containerization in multi-cloud also poses challenges for the academic community.
@@
 Recent studies (e.g., (Bordeleau et al., 2020) (Mihai et al., 2022) (Sarker, 2022)) have notably highlighted that challenges in design, development, monitoring, and testing of containerization applications in multi-cloud environment are deeply interconnected with the software development life cycle.
-To address this knowledge gap, we established the objective of this research is to systematically identify and categorize various facets of container utilization in multi-cloud environment.
-These facets encompass the roles of containers, their implementation strategies, architectural patterns, quality attributes, as well as the motivations of the customers.
-Additionally, we aspire to explore the challenges and solutions related to automation, deployment, monitoring, and security of containerization applications in multi-cloud environment.
+To bridge this knowledge gap and support the practical demands of the QLEAP project consortium, our study systematically identifies and categorizes the various facets of container utilization in multi-cloud environment.
+These facets include container roles, implementation strategies, architectural patterns, and quality attributes.
+In addition, we explore the challenges and solutions associated with automation, deployment, monitoring, and security, providing practitioners and researchers with actionable insights to tackle these critical issues effectively.
 To achieve this objective, we employed a systematic approach to conduct a mapping study.
 We finally selected 86 studies that focus on containerized applications in multi-cloud environment for further analysis.
-The empirical findings and frameworks of this Systematic Mapping Study (SMS) offer significant potential benefits for developers of container-based applications.
-By exploring a diverse range of implementation strategies, architectural patterns, and quality attributes, practitioners can gain valuable insights to enhance their practices.
-Moreover, a detailed examination of challenges and solutions related to automation, deployment, monitoring, and security further provides practitioners with the knowledge to address problems effectively and improve dedicated quality attributes.
-Additionally, these findings and frameworks offer valuable support to practitioners in comprehending various patterns, strategies, and tactics specific to containerized applications in multi-cloud environment.
-The proposed frameworks can act as a facilitator of knowledge sharing, training, and ensure a shared understanding among practitioners regarding the complexities and solutions within containerized applications in multi-cloud environment.
-Ultimately, this SMS serves to empower software developers and practitioners, enabling them to optimize their approaches and elevate the quality and efficiency of containerized applications in multi-cloud environment.
-Our SMS makes the following key contributions:
-  • We developed four frameworks to address challenges and solutions in security, automation, deployment, and monitoring for containerization in multi-cloud environment.
-  • We systematically identified and classified an extensive list of patterns and strategies for container-based applications in multi-cloud environment.
-  • We systematically identified and classified quality attributes and tactics for containerized applications in multi-cloud environment.
-  • We identified and classified a wide range of tools and frameworks for containerized applications in multi-cloud environment.
-  • We have publicly released a dataset, which is available online (Waseem et al., 2024), to enable researchers and practitioners to access all collected data, replicate, and validate our SMS.
-This dataset includes detailed hierarchies of the developed frameworks, identified and classified container implementation strategies, and the roles of containers in multi-cloud environment.
+We adopted the following Goal-Question-Metric (GQM) framework (Caldiera and Rombach, 1994):
+  • Goal: To enhance the understanding and efficiency of deploying and managing container-based applications in multi-cloud environment.
+  • Questions: To explore specific aspects such as container roles, implementation strategies, architectural patterns, quality attributes, challenges, and solutions.
+  • Metrics: To evaluate the effectiveness of our findings based on the identified roles, strategies, patterns, attributes, challenges, solutions, and tools.
+The empirical findings and challenge-solution catalogs from this Systematic Mapping Study (SMS) provide significant benefits for developers of container-based applications.
+By examining implementation strategies, architectural patterns, and quality attributes, practitioners gain insights to improve practices.
+Detailed analyses of challenges and solutions in automation, deployment, monitoring, and security equip practitioners to address issues and enhance quality attributes.
+These challenge-solution catalogs facilitate knowledge sharing, training, and a shared understanding of containerized applications in multi-cloud environments, empowering developers to optimize their approaches and improve efficiency and quality.
+In response to the real-world needs identified in the QLEAP project and the broader research gap, our SMS presents the following key contributions:
+  • We systematically identified and categorized challenges and their solutions related to security, automation, deployment, and monitoring for containerization in a multi-cloud environment.
+  • We systematically identified and categorized an extensive list of patterns and strategies for container-based applications in multi-cloud environment.
+  • We systematically identified and categorized quality attributes and tactics for containerized applications in multi-cloud environment.
+  • We systematically identified and classified a wide range of tools and frameworks for containerized applications in multi-cloud environment.
+  • We have publicly released a dataset, which is available online (Waseem et al., 2024), to enable researchers and practitioners to access, replicate, and validate all collected data from our SMS.
+This dataset includes detailed hierarchies of the developed catalogs, identified and classified container implementation strategies, and the roles of containers in multi-cloud environment.
 It also encompasses patterns, strategies, quality attributes, tactics, tools, and frameworks for containerized-based applications in multi-cloud environment.
@@
 Figure 2 illustrates the SMS execution process.
+Reliability and consistency are critical aspects of any empirical study.
+In this study, we followed established statistical approaches, including inter-rater agreement (IRA) computation, to ensure the reliability of our data coding and extraction processes.
+Additionally, we applied systematic quality assessment techniques to evaluate the studies included in our SMS.
 Figure 2. Schematic representation of the research methodology implemented in this study
 2.1. Research Questions
-To conduct this SMS, we formulated six Research Questions (RQs), as outlined in Table 1, aligning with the objective of our SMS.
-The table provides a structured view of the RQs, their corresponding rationale, and categorization.
+To conduct this SMS, we formulated six Research Questions (RQs) based on the GQM approach presented in the Introduction section (see Section 1) aligns with the objective of our SMS, as outlined in Table 1.
+The table provides a structured view of the RQs with their corresponding rationale and categorization.
 Table 1. Research questions and their rationale
@@
 Container Roles and Implementation
-RQ1
-What roles do containers play, and what strategies can be employed for implementing container-based applications in multi-cloud environment?
+RQ1 What roles do containers play, and what strategies can be employed for implementing container-based applications in multi-cloud environment?
 To investigate the role of containers and explore implementation strategies in multi-cloud environment, with the objective of enhancing understanding and facilitating efficient deployment and management of applications across multiple cloud environments.
-RQ2
-What architectural patterns and strategies are utilized in the implementation of container-based applications in multi-cloud environment?
+RQ2 What architectural patterns and strategies are utilized in the implementation of container-based applications in multi-cloud environment?
 To explore and document the various architectural patterns and strategies that are utilized in implementing container-based applications in multi-cloud environment, thus providing a roadmap for efficient and scalable application development and deployment.
-RQ3
-What are the quality attributes and the tactics associated with their implementation in container-based applications in multi-cloud environment?
+RQ3 What are the quality attributes and the tactics associated with their implementation in container-based applications in multi-cloud environment?
 To examine and analyze the quality attributes required for container-based applications in multi-cloud environment and the corresponding tactics employed in their implementation, enabling the identification of best practices and guidelines for developing robust and reliable applications in such environments.
 Challenges and Solutions
-RQ4
-What are the challenges related to security, automation, deployment, and monitoring in container-based applications in multi-cloud environment?
-To comprehensively evaluate and gain insights into the challenges surrounding security, automation, deployment, and monitoring concerning container-based applications in multi-cloud environment, this question aims to provide valuable insights into the obstacles and intricacies that organizations encounter while implementing and managing these applications in such a dynamic setting.
-RQ5
-How to address security, automation, deployment, and monitoring in container-based applications within multi-cloud environment?
-This question seeks to investigate solutions aimed at tackling the challenges of security, automation, deployment, and monitoring within container-based applications operating in multi-cloud environment.
+RQ4 Challenges in Container-Based Applications in Multi-Cloud Environment RQ4.1:What are the challenges related to security in container-based applications in multi-cloud environment?
+RQ4.2: What are the challenges related to automation in container-based applications in a multi-cloud environment?
+RQ4.3: What are the challenges related to deployment in container-based applications in a multi-cloud environment?
+RQ4.4: What are the challenges related to monitoring in container-based applications in a multi-cloud environment?
+To comprehensively evaluate and gain insights into the challenges surrounding security (RQ4.1), automation (RQ4.2), deployment (RQ4.3), and monitoring (RQ4.4) concerning container-based applications in multi-cloud environment.
+These sub-RQs aim to provide valuable insights into the obstacles and intricacies that organizations encounter while implementing and managing container-based applications in a multi-cloud environment.
+RQ5 Solutions to Address Challenges in Container-Based Applications in Multi-Cloud Environment RQ5.1:How to address the challenges related to security in container-based applications within a multi-cloud environment?
+RQ5.2: How to address the challenges related to automation in container-based applications within a multi-cloud environment?
+RQ5.3: How to address the challenges related to deployment in container-based applications within a multi-cloud environment?
+RQ5.4: How to address the challenges related to monitoring in container-based applications within a multi-cloud environment?
+These sub-RQs seek to investigate solutions aimed at tackling the challenges of security (RQ5.1), automation (RQ5.2), deployment (RQ5.3), and monitoring (RQ5.4) within container-based applications operating in multi-cloud environment.
 By providing actionable guidance and recommendations, the goal is to elevate the overall management and performance of these applications in the context of multi-cloud environment.
 Tools and Frameworks
-RQ6
-What tools and frameworks are used to implement container-based multi-cloud applications?
+RQ6 What tools and frameworks are used to implement container-based multi-cloud applications?
 To investigate and identify the specific tools and frameworks utilized in the implementation of container-based multi-cloud applications, providing an overview of the technological landscape and aiding developers and organizations in making informed decisions when selecting appropriate tools for their application deployment and management.
+These RQs are designed to address important gaps in the current understanding of container-based applications in multi-cloud environment.
+RQ1–RQ3 focus on understanding the roles of containers, implementation strategies, architectural patterns, and quality attributes.
+These questions aim to provide a comprehensive perspective, building on and organizing fragmented insights from prior studies.
+RQ4 and RQ5 focus on specific challenges and corresponding solutions related to security, automation, deployment, and monitoring in container-based multi-cloud environment.
+These questions aim to systematically address issues that are often only partially explored in existing works.
+Finally, RQ6 investigates tools and frameworks used for implementation, helping bridge the gap between academic research and practical applications.
+Together, these RQs provide a structured framework that enhances understanding and supports the development and management of container-based applications in multi-cloud environment, advancing the state-of-the-art in this domain
+2.2. Search String Composition
+Initially, we considered the PICO (Population, Intervention, Comparison, Outcome) framework (Schardt et al., 2007) (Ralph et al., 2020) for constructing our search string.
+However, given the complexity and breadth of our topic, applying PICO strictly required an extensive array of terms and logical operators, exceeding the query limits of databases like ScienceDirect.
+Therefore, we developed the search string for this study based on authors’ knowledge and iterative trial searches.
+This approach allowed us to tailor the search strategy to the specific goal of our study, ensuring comprehensive coverage within practical constraints.
 2.3. Study Search and Selection Process
@@
 The primary search involved querying digital databases (see Table 2) using customized search string.
-The search strings were executed concurrently on eight databases, as illustrated in Figure 2, spanning between January 2013 and March 2023 when we started this SMS.
+The search strings were executed concurrently on eight databases, as illustrated in Figure 2, spanning between January 2013 and July 2024 when we started this SMS.
 We chose 2013 as the starting point because it is the year when container technologies like Docker emerged (Docker, 2023), marking a key phase in containerization and cloud computing.
-This period also saw a rise in relevant research to containers in multi-cloud environment, making it ideal to understand the development of containerized applications in multi-cloud environments over the past decade.
+This period also saw a rise in relevant research to containers in multi-cloud environment, making it ideal to understand the development of containerized applications in multi-cloud environment over the past decade.
 We followed the steps mentioned below to select the relevant studies.
@@
 Selection Criteria Inclusion Criteria Exclusion Criteria
-Language
-English
-Non-English
-Study Type
-Primary studies including peer-reviewed journal articles, book chapters, conference papers, workshops, and symposium papers
-Secondary studies or non-peer-reviewed content (e.g., blogs, webpages, videos, white papers, technical reports, grey literature)
-Study Focus
-Studies addressing container roles and strategies, architectural patterns, quality attributes, adoption drivers, automation, deployment, monitoring, security challenges, solutions, and tools/frameworks for multi-cloud environment.
+Language English Non-English
+Study Type Primary studies including peer-reviewed journal articles, book chapters, conference papers, workshops, and symposium papers Secondary studies or non-peer-reviewed content (e.g., blogs, webpages, videos, white papers, technical reports, grey literature)
+Study Focus Studies addressing container roles and strategies, architectural patterns, quality attributes, adoption drivers, automation, deployment, monitoring, security challenges, solutions, and tools/frameworks for multi-cloud environment.
 Studies that focus solely on traditional cloud computing, single-cloud environments, or unrelated topics
-Study Duration
-Studies published between January 2013 and March 2023
-Studies published before January 2013 or after March 2023
+Study Duration Studies published between January 2013 and July 2024 Studies published before January 2013 or after July 2024
   • Step 1: Extraction of Studies: We ran custom search strings in the selected databases (see Table 2) to retrieve study titles, author names, publication years, venues, publication types, and abstracts.
-This initial search yielded 5,676 studies from eight databases.Following initial retrieval, we did not apply Cohen’s Kappa as this step is automated and does not involve subjective decision-making.
+This initial search yielded 6,727 studies from eight databases.
+Following initial retrieval, we did not apply Cohen’s Kappa (Cohen, 1960) (Cohen, 1968) as this step is automated and does not involve subjective decision-making.
   • Step 2: Title and Keyword Screening: This step involves reviewing the titles and keywords of the collected studies to assess their relevance to the research topic.
-Initially, we removed any duplicated studies obtained from different databases (such as IEEE Xplore and Scopus) by sorting them in ascending order.
-This resulted in the elimination of several hundred duplicate studies.
-We divided the remaining studies between two authors (R1, R2), who independently applied the inclusion and exclusion criteria to identify studies relevant to our research objectives, as detailed in Table 3.
-To assess the consistency of study selection and mitigate bias, a Cohen’s Kappa analysis (Cohen, 1968) was performed on a random sample of 40 studies, as illustrated in Formula 1:
+Initially, we removed any duplicate studies obtained from different databases (e.g., IEEE Xplore, Scopus) by sorting them in ascending order.
+This resulted in the elimination of several thousand duplicate studies.
+We divided the remaining studies between two authors (R1, R2), who independently applied the inclusion and exclusion criteria to identify studies relevant to our research objective, as detailed in Table 3.
+    To assess the consistency of study selection and mitigate bias, a Cohen’s Kappa analysis (Cohen, 1960) (Cohen, 1968) was performed on a random sample of 400 studies (10% of the total dataset of 4,000 studies), as illustrated in Formula 1:
     (1) $\kappa=\frac{P_{o}-P_{e}}{1-P_{e}}$
     where $P_{o}$ represents the relative observed agreement among authors, and $P_{e}$ the hypothetical probability of chance agreement.
-The results, recorded in Table 4.a, indicated a Cohen’s Kappa of 0.531.
-This value reflects a moderate level of agreement between R1 and R2.
-Any disagreements were systematically resolved through discussions among all authors, leading to a consensus.
-Following this process, the number of studies was reduced to 809 for further analysis.
+The results, recorded in Table 4.a, indicated a Cohen’s Kappa of $\kappa=0.531$, which reflects moderate agreement between R1 and R2 (Sim and Wright, 2005).
+Specifically, R1 and R2 agreed on 360 studies being relevant (“YES”) and 20 studies being irrelevant (“NO”).
+Disagreements occurred in 20 studies, which were systematically resolved through discussions among all authors, leading to a final consensus.
+Using a random subset is a common practice in systematic reviews to measure inter-rater agreement without overburdening resources, ensuring reliability while maintaining efficiency (Keele et al., 2007; Chandler et al., 2019).
+Following this process, the number of studies was reduced to 1,047 for further analysis.
   • Step 3: Abstract-based Screening: During this step, we conducted a thorough review of the abstracts of the collected studies to determine their relevance to our research topic.
-Two researchers (R1 and R2) independently examined each abstract, assigning a status of ”relevant,” ”irrelevant,” or ”doubtful” with respect to our research goals.
-To reduce subjective bias and to quantify the consensus between the researchers, we conducted a Cohen’s Kappa analysis on a subset of 30 randomly selected studies.
-The outcomes of this assessment are depicted in Table 4.b, which show substantial agreement between R1 and R2, with 25 studies being unanimously considered relevant and only 4 studies disagree (3 with one researcher favoring relevance and the other not, and 1 where both disagreed) The Kappa value, calculated based on the observed agreements and disagreements was approximately K= 0.268 which suggests a fair agreement between the two researchers.
-Consequently, upon the compilation of step, we got 303 studies for more detailed scrutiny.
-  • Step 4: Full-Text Screening: In this step, we reviewed the full texts of 303 studies that had been shortlisted through abstract-based screening.
-To ensure consistency in our review process, we randomly chose 20 studies (see Table 4.c) for a Cohen’s Kappa analysis to measure the agreement between two researchers.
-The resulting Kappa value was K=0.318, which represents a fair level of agreement.
-This measure confirms that our selection criteria were applied consistently during the full-text screening studies.
+Two researchers (R1 and R2) independently examined each abstract, assigning a status of “relevant”, “irrelevant”, or “doubtful” with respect to our research goals.
+To reduce subjective bias and to quantify the consensus between the researchers, we performed a Cohen’s Kappa analysis on a random subset of 105 studies (approximately 10% of the 1,047 studies remaining after the Title and Keyword Screening phase).
+    The outcomes of this assessment are depicted in Table 4.b.
+The analysis showed substantial agreement between R1 and R2, with 88 studies being unanimously considered relevant and 5 studies unanimously considered irrelevant.
+Disagreements occurred in 12 cases (7 studies where one researcher marked “relevant” while the other marked “irrelevant”, and 5 studies initially classified as “doubtful”).
+The kappa value, calculated based on observed agreements and disagreements, was $\kappa=0.521$, calculated based on observed agreements and disagreements, indicates moderate agreement (Sim and Wright, 2005; Keele et al., 2007).
+    To compute Cohen’s Kappa during the abstract-based screening phase, the original classifications of “relevant”, “irrelevant”, and “doubtful” were simplified into a binary scale (“YES” or “NO”) due to the binary nature of Cohen’s Kappa computation.
+Specifically:
+    • Studies classified as “relevant” were mapped to “YES”.
+    • Studies classified as “irrelevant” were mapped to “NO”.
+    • Studies classified as “doubtful” were resolved through discussions between R1 and R2. If consensus was reached on “relevant”, the study was marked as “YES”; otherwise, it was marked as “NO”.
+    Using subsets for kappa computation ensures that inter-rater reliability can be rigorously evaluated while optimizing the use of resources, as recommended by systematic review guidelines (Keele et al., 2007; Chandler et al., 2019).
+After resolving all disagreements through systematic discussions among the authors, a final consensus was reached for the subset.
+Consequently, upon completing this step, the number of studies was reduced to 346 for more detailed scrutiny.
+  • Step 4: Full-Text Screening: In this step, we reviewed the full texts of 346 studies that had been shortlisted through abstract-based screening.
+To ensure consistency in our review process, we randomly selected a subset of 35 studies (10% of the total 346 studies) for a Cohen’s Kappa analysis to measure the agreement between two researchers (R1 and R2).
+The results, recorded in Table 4.c, indicated a Cohen’s Kappa value of $\kappa=0.578$, which represents moderate agreement between the two researchers (Sim and Wright, 2005).
+Specifically, R1 and R2 agreed on 30 studies as relevant (“YES”) and 2 studies as irrelevant (“NO”).
+Disagreements occurred in 3 studies (2 marked as “YES” by one researcher and “NO” by the other, and 1 marked as “NO” by both but flagged as uncertain).
+All disagreements were systematically resolved through discussions among the authors, leading to a final consensus.
+This subset-based approach, coupled with systematic resolution of disagreements, provides confidence in the reliability of the full-text screening process (Chandler et al., 2019).
+At the end of this step, we obtained a total of 110 studies, which were deemed relevant for further analysis.
 Table 4. Kappa Agreement across different screening steps.
-(a) Title and Keyword Screening on 40 Randomly Selected Studies
+(a) Title & Keyword Screening on 400 Randomly Selected Studies
 R1 Yes R1 No
-R2 Yes 35 2
-R2 No 1 2
-Total 36 4
-(b) Abstract-based Screening on 30 Randomly Selected Studies
+R2 Yes 360 20
+R2 No 10 10
+Total 370 30
+(b) Abstract-based Screening on 105 Randomly Selected Studies
 R1 Yes R1 No
-R2 Yes 25 3
-R2 No 1 1
-Total 26 4
-(c) Full-Text Screening on 20 Randomly Selected Studies
+R2 Yes 88 7
+R2 No 5 5
+Total 93 12
+(c) Full-Text Screening on 35 Randomly Selected Studies
 R1 Yes R1 No
-R2 Yes 16 2
-R2 No 1 1
-Total 17 3
+R2 Yes 30 3
+R2 No 2 0
+Total 32 3
 2.3.2. Snowballing
-In Phase 2, we utilized the snowballing technique, as described in (Wohlin, 2014), to examine references within primary studies for identifying additional relevant studies.
+In Phase 2, we utilized the snowballing technique, as described in (Wohlin, 2014), to examine references within 110 primary studies to identify additional relevant studies.
 This strategy was augmented by forward snowballing, where we gathered studies that cited the selected studies, and backward snowballing, which involved using references within the selected studies.
 Notably, we encountered several dozen studies during the forward and backward snowballing that had been excluded in primary search.
-This phase resulted in the addition of 8 more studies, bringing the final count to 86.
+This phase resulted in the addition of 11 more studies, bringing the final count to 121.
+2.3.3. Quality Assessment of Selected Studies
+We assessed the quality of the selected studies based on a set of predefined criteria, including relevance to multi-cloud containerization, clarity of research design, empirical validation (e.g., experiments, case studies), reporting quality (e.g., discussion of limitations and challenges), publication in peer-reviewed venues, and the generalizability of findings to multi-cloud environment.
+Each study was evaluated using a scoring system where criteria were rated as 0 (not satisfied), 0.5 (partially satisfied), or 1 (fully satisfied), with a maximum total score of 6.
+Studies scoring below 3 were flagged for further review and discussion but retained to ensure alignment with the study’s objectives.
+This quality assessment was conducted after the study selection process to ensure consistency and thoroughness across all 121 studies.
+Detailed results are available in the replication package (Waseem et al., 2024), specifically in the sheet titled “QA of Selected Studies”.
 2.4. Data Extraction and Analysis
@@
 To mitigate the personal bias and ambiguity, all authors engaged in discussions regarding the extracted data.
-The data items labeled as D1 to D11 present a summary of the demographics of the primary studies selected, while D12 to D26 are specifically employed to address RQ1 to RQ6.
-A concise description of each data item is also presented in Table 5.
-Finally, Google Sheets were used to record and further analysis of the extracted data.
-Data Analysis: We employed a descriptive statistics to analyze the quantitative data from data items D1, D5, D7, D8, and D9.
+The data items labeled as D1 to D8 present a summary of the demographics of the primary studies selected, while D9 to D22 are specifically employed to address RQ1 to RQ6.
+A concise description of each data item is presented in Table 5.
+Finally, Google Sheets were used to record and further analyze the extracted data.
+Data Analysis: We employed descriptive statistics to analyze the quantitative data from data items D1, D5, D7, D8, and D9.
 For the remainder of the data, which primarily consist of qualitative free-text descriptions (e.g., study aim, roles of containers, challenges, and solutions), we conducted a thematic analysis in accordance with the guidelines outlined in (Braun and Clarke, 2006).
@@
 D13 Tactics Tactics used in the study
-D14 Automation Challenges Challenges in automation discussed in the study RQ4
-D15 Automation Solutions Solutions for automation challenges discussed in the study RQ5
+D14 Automation Challenges Challenges in automation discussed in the study RQ4.2
+D15 Automation Solutions Solutions for automation challenges discussed in the study RQ5.2
 D16 Deployment Challenges Challenges in deployment discussed in the study RQ4
@@
 D19 Monitoring Solutions Solutions for monitoring challenges discussed in the study RQ6
-D20 Security Challenges Challenges in security discussed in the study RQ4
-D21 Security Solutions Solutions for security challenges discussed in the study RQ5
+D20 Security Challenges Challenges in security discussed in the study RQ4,1
+D21 Security Solutions Solutions for security challenges discussed in the study RQ5.1
 D22 Tools, Languages, and Frameworks Tools, programming languages, and frameworks RQ6
-The process of obtaining research themes executed in this SMS is shown in Figure 2.
-Two researchers participated in the process in order to reduce personal bias.
+Two researchers participated in the process to reduce personal bias.
 The most important activity of this process is brainstorming sessions that were mainly conducted while reviewing, defining, and naming the research themes.
 In these sessions, both researchers discussed and validated the research themes found.
-Concerning data item D9 (testing approaches) and D10 (testing challenges), we used the open coding and constant comparison techniques from Grounded Theory (Glaser and Strauss, 1967) to analyze the qualitative data extracted from the selected studies.
-Finally, we provided a replication package (Waseem et al., 2024) containing the results of each phase of the study selection process (e.g., Phase 1, Phase 2) and detailed results (e.g., Contributions, Patterns, QAs, Challenges, Solutions) for verification and validation purposes of this SMS.
+Concerning data item D9 (Roles of Containers) and D10 (Container Implementation strategy), we used the open coding and constant comparison techniques from Grounded Theory (Glaser and Strauss, 1967) to analyze the qualitative data extracted from the selected studies.
+Finally, we provided a replication package (Waseem et al., 2024) with the results of each phase of the study selection process (e.g., Phase 1, Phase 2) and detailed results (e.g., Contributions, Patterns, QAs, Challenges, Solutions) for verification and validation purposes of this SMS.
 3. Results
 This section presents the results of the SMS, derived from analyzing data from chosen studies.
-Section 3.1 provide a summary of the demographic data, classifications, and the mapping of research themes identified in the studies.
+Section 3.1 provides a summary of the demographic data, classifications, and the mapping of research themes identified in the studies.
 Section 3.2 presents the results concerning container roles and implementation strategies.
-Section 3.3 reports on the patterns and strategies employed in containerized-based applications in multi-cloud environment.
+Section 3.3 reports on the patterns and strategies employed in container-based applications in multi-cloud environment.
 Section 3.4 presents the QAs and related tactics for containerized applications in multi-cloud environment.
-Section 3.5 presents the security challenges and solution framework, while Section 3.6 presents the automation challenges and solution framework.
-Section 3.7 presents the deployment challenges and solution framework, and Section 3.8 presents the monitoring challenges and solution framework.
-Lastly, the tools and frameworks are reported in Section 3.9.
+Section 3.5 presents the security challenges and solutions framework, while Section 3.6 presents the automation challenges and solutions framework.
+Section 3.7 presents the deployment challenges and solutions framework, and Section 3.8 presents the monitoring challenges and solutions framework.
+Finally, the tools and frameworks are reported in Section 15.
 3.1. Demographics, Research Aims, and Contributions
-Studies Yearly Distribution: Figure 3-a illustrates the annual distribution of studies found in eight databases from January 2013 to March 2023.
-After applying our selection criteria, the final dataset includes studies published from 2014 to 2022.
-We found no primary studies on containerization in multi-cloud environments meeting our criteria in the first three months of 2023.
-The bars displays yearly study counts, with a trend showing the overall direction.
-A peak in 2017, followed by a decrease and occasional fluctuations, suggests an early spike in interest then more targeted research.
+Yearly Distribution of Studies: Figure 3-a illustrates the annual distribution of studies retrieved from eight databases between January 2013 and July 2024.
+We found only two primary studies on containerization in multi-cloud environment that meet our criteria in 2013.
+The bars display yearly study counts and trends.
+A peak in 2017, followed by a decrease and occasional fluctuations, suggests an early spike in interest followed by more targeted research.
+Notably, the reduction in publication numbers in 2020 and 2021 may be partially attributed to the global disruptions caused by the COVID-19 pandemic, which affected research outputs across various fields.
+However, there is a notable increase in studies in 2023 and continuing into 2024.
+This could be due to delayed research projects getting back on track and more funding in tech sectors after the pandemic, showing a renewed focus on improving containerization in multi-cloud environment
 Figure 3. Demographics of the Selected Studies, Including: a) Frequency and Categories of Publications; b) Publisher of the Research; c) Affiliations of the Authors
-Publishers Distribution: The pie chart presented in Figure 3-b illustrates the distribution of identified studies among different publishers.
-The largest portion of the studies, accounting for 72.09%, is published by IEEE, indicating its significant contribution to the field.
-Scopus follows with 11.62%, while Engineering Village, ACM, and Springer Link represent smaller portions at 6.97%, 5.81%, and 3.48%, respectively.
+Publishers Distribution: Figure 3-b shows the distribution of studies by publisher.
+IEEE accounts for 65.29%, highlighting its major contribution.
+Scopus follows with 11.57%, while Engineering Village, ACM, and Springer Link represent smaller portions at 12.40%, 7.44%, and 2.48%, respectively.
 The chart emphasizes the dominance of IEEE in the identified studies, suggesting its importance as a primary source of relevant literature.
-Study Types: Figure 3-a provides a breakdown of the types of the selected studies.
-It is apparent that conferences dominate as the most common venue, comprising a substantial 72.09% of the total publications.
-Journals, which are often considered as venues for more mature and rigorously peer-reviewed research, make up 19.76%.
-This indicates a strong preference for disseminating information through conferences, which may be due to the faster publication process and opportunities for networking and collaboration.
-Symposiums and workshops account for 6.97% and 1.16% respectively, suggesting that these are less common avenues, possibly due to their more specialized or focused nature.
-Authors’ Affiliations: Figure 3-c illustrates the distribution of authors’ affiliations in the collected publications.
-A staggering 93.02% of the affiliations are associated with academia.
-This overwhelmingly high percentage emphasizes the central role that academic institutions play in the production and dissemination of research in the field.
-Industry affiliations represent a mere 4.65%, which could imply that industry professionals may be more focused on practical applications and might not publish frequently in the venues considered.
-Mixed affiliations, at 2.32%, indicate a small proportion of collaborations between academia and industry.
+This distribution also suggests to researchers where to primarily search and publish on containerization in multi-cloud environment.
+Study Types: Figure 3-a provides a breakdown of the types of selected studies.
+It is apparent that conferences dominate as the most common venue, comprising 62.5% of the total publications.
+Journals, often considered venues for mature and rigorously peer-reviewed research, make up 32.5%.
+This indicates a preference for conferences, likely due to faster publication and networking opportunities Symposiums and workshops represent for 5.0% and 0.83% respectively, suggesting that these are less common avenues, possibly due to their more specialized or focused nature.
+Authors’ Affiliations: Figure 3-c illustrates the distribution of authors’ affiliations in collected publications.
+90.90% of the affiliations are associated with academia.
+This highlights the central role of academic institutions in producing and disseminating research.
+implying that industry professionals focus more on practical applications.
+The low industry affiliation may reflect the SMS’s focus on academic databases, where industry participation is limited, indicating typical academic database trends and suggesting broader industry inclusion in future studies.
+Mixed affiliations, at 4.95%, indicate a small proportion of collaborations between academia and industry.
 This low rate of cross-affiliation collaboration may suggest that there is potential for increased synergy between academic and industrial entities in future research endeavors.
-Mapping of Research Themes: The thematic analysis of selected studies aims on containerization in multi-cloud environment, as illustrated in Figure 4, indicating various significant insights.The prominent research themes identified include “Scalability and High Availability” with 20 studies, and three other themes: ’Security and Privacy”m “Multi-Cloud Container Monitoring and Adaptation”, and “Performance and Optimization”, each attracting attention in 15 studies.
+Classification of Research Theme: The identified research themes and their notable trends are illustrated in Figure 4 and detailed below.
+Figure 4-(a) presents the taxonomical classification of existing research.
+This taxonomy systematically identifies, names, and represents various topics based on their similarities or distinctions.
+Figure 4-(b) depicts the trends and temporal distribution of the themes over the years (2013–2024) based on the studies identified in our review.
+  • Taxonomy of Research Themes: The prominent research themes identified include “Scalability and High Availability” with 20 studies, and three other themes: “Security and Privacy” in “Multi-Cloud Container Monitoring and Adaptation”, and “Performance and Optimization”, each attracting attention in 15 studies.
 This highlights that the core areas of interest in multi-cloud environment are proficient container management, scalability, and performance optimization.
-“Security and Privacy’ also stands as a considerable theme, as demonstrated by the 15 studies focusing on areas such as secure container handling, data security, and blockchain applications.
+“Security and Privacy” also stands as a considerable theme, as demonstrated by the 15 studies focusing on areas such as secure container handling, data security, and blockchain applications.
 This emphasizes the necessity of protecting data and applications in the context of containers spreading across multi-cloud environment.
@@
 For instance, study S19 is catalogued under both ‘Multi-cloud Orchestration and Management’ and ‘Application Deployment in Multi-cloud Container’, signifying that it encompasses aspects of both Kubernetes and microservices architecture.
-In essence, the exploration into containerization in multi-cloud environment is substantially inclined towards, scalability, performance optimization security, and orchestration, with several studies showcasing inter-thematic relevance.
-This is suggestive of the diverse challenges and opportunities present in deploying containers within multi-cloud contexts.
-Major Contributions of the Selected Studies: Table 6 provides an overview of the contributions made by the selected studies.
+  • Trends and Temporal Distribution of Themes: To highlight prominent trends and their progression over the years (2013–2024), we adopted the thematic and sub-thematic representation from Figure 4(a) and illustrated them via a timeline and relative distribution of each theme in Figure 4-(b).
+Specifically, Figure 4-(b) demonstrates how certain research themes have progressed and matured (i.e., evolved as topics) over the years.
+For instance, the theme multi-cloud orchestration and management is addressed in a total of 18 studies, accounting for approximately 15% of the total reviewed studies.
+Our analysis indicates that one of the earliest research efforts on multi-cloud orchestration, published during 2014–2015, focused on the component-based and model-driven adaptation of applications for multi-cloud containers [S1, S20, S8].
+Following this initial phase of application adaptation for multi-cloud orchestration, subsequent research explored techniques for benchmarking and enabling resource-efficient application orchestration [S2, S41, S39].
+More recent studies, published between 2022 and 2024, have investigated a variety of topics, including but not limited to DevOps-based development [S27] and monitoring orchestrated applications in multi-cloud environments [S39, S87].
+Our review revealed no published studies on the theme of multi-cloud orchestration and management during 2013, 2017, or 2021.
+Two possible reasons for this are: (i) our literature search did not identify any relevant studies (see Section 2.3.1 - primary search), or (ii) the studies from those years did not meet the qualitative evaluation criteria (see Section 2.3.3 - quality assessment) for inclusion in the review.
+Figure 4. Selected Studies: a) Thematic Classification and b) Trends Over the Years
+Major Contributions of the Selected Studies: Table 6 provides an overview of the contributions of the selected studies.
 The contributions are primarily categorized into five main categories and 22 subcategories (see the Contributions sheet in (Waseem et al., 2024)), through thematic analysis.
-The selected studies primarily focused on proposing and validating reference architectures, security approaches, cloud management and interoperability frameworks, orchestration and deployment approaches, as well as approaches and tools for optimization and deployment of containerized applications in multi-cloud environment.
-The studies related to the reference architecture category mainly contributed to the understanding and design of architectures specifically tailored for multi-cloud containerization.
-They aimed to address the complexities associated with containerization in multi-cloud environment.
-On the other hand, the studies in the security approaches category discuss concerns related to cloud analysis, security and performance, cybersecurity, and access control within multi-cloud contexts.
-Figure 4. Selected studies research aims in containerization for multi-cloud environment.
-The studies on cloud management and interoperability frameworks primarily aim to improve resource and service management, monitoring, and deployment in multi-cloud environment.
-Their goal is to enhance the performance and efficiency of managing resources and services across multiple clouds.
-Meanwhile, the research in the area of orchestration and deployment approaches introduces methods for bench-marking, optimizing, and simulating container orchestration in multi-cloud settings.
-The objective of these studies is to streamline the process of deploying and orchestrating containers within multi-cloud environment.
-Lastly, the research on optimization and deployment approaches and tools focuses on improving container deployment and creating efficient management solutions for multi-cloud environment.
-These studies underscore the importance of optimizing resource allocation and developing practical management tools for use in multi-cloud settings.
-Overall, these studies contribute to proposing reference architectures, security approaches, cloud management, and interoperability frameworks.
-These studies also offer strategies and tools for orchestration, deployment operations, and optimization in multi-cloud containerization.
+The selected studies mainly focus on proposing and validating security solutions, reference architectures, cloud management frameworks, interoperability, orchestration, and tools for optimizing and deploying containerized applications in multi-cloud environment.
+Notably, the largest contributions come from the Security Approaches category, representing 27.25% of the total studies, which highlights the significant attention to cybersecurity & access control and cloud analysis approaches.
+Within this category, the Cybersecurity and Access Control subcategory alone accounts for 9.09%, emphasizing the importance of securing multi-cloud environment.
+The studies related to the Reference Architectures and Models for Cloud category, comprising 21.42% of the studies, focus on understanding and designing architectures specifically tailored for multi-cloud containerization.
+Within this category, General Cloud Architectures (4.92%) and Fault Tolerance and Self-Healing (4.13%) are key subcategories.
+These subcategories address critical challenges such as ensuring resilience and performance in cloud-native architectures.
+The studies on Optimization and Deployment Approaches and Tools category accounts for 18.16% of the studies, indicating a strong research focus on optimizing container deployment and resource scheduling.
+Within this category, Optimization & Scheduling Approaches (8.26%) represent the second-largest subcategory across all studies, highlighting the practical need for effective resource allocation in multi-cloud systems.
+Finally, Cloud Management and Interoperability Frameworks, accounting for 12.38%, primarily aim to improve resource utilization, service management, monitoring, and deployment in multi-cloud environments.
+This includes contributions from Resource and Service Management Frameworks, which represent 4.95% of the studies.
 Table 6. Selected studies’ contributions on containerization in multi-cloud environment
 Category Subcategory Study #
-Reference Architectures and Models for Cloud General Cloud Architectures S26, S33, S35, S42, S76
-Fault Tolerance & Self-Healing S12, S55, S67, S74
-Migration & Testing Architecture S17, S79
-Domain-Specific Language S80, S82
-Performance Modeling S28, S62
-Optimization & Reliability Model S50, S66
-IoT & Service Mesh Architecture S05, S59
-Security Approaches Cloud Analysis Approach S18, S24, S37, S65, S69, S86
-Security and Performance Solution S10, S23, S38, S75, S77
-Cybersecurity & Access Control S11, S30, S34, S71, S73
-Security Approach S06, S25, S40, S83
-Cloud Management and Interoperability Framework Resource & Service Management Framework S48, S49, S47, S15, S58
-Monitoring & Deployment Framework S19, S45, S46, S08
-Interoperability Framework S03, S81
-Evaluation & Modeling Framework S20, S85
-Orchestration and Deployment Approach Benchmarking & Optimization Approach S01, S02, S41, S51, S68
-Service Management Approach S07, S14, S29
-Simulation & Integration Approach S27, S31, S56
-Storage & Services Solution S64, S75
-Optimization and Deployment Approaches and Tool Optimization & Scheduling Approach S22, S32, S36, S39
-Multi-Cloud Platform S09, S16, S43, S60, S72
-Deployment Approach and Tool S21, S52, S57, S61
+Reference Architectures and Models for Cloud General Cloud Architectures (6, 4.92%) (Watanabe et al., 2022; Sami and Mourad, 2020; Carpio et al., 2020; Baresi et al., 2016; Pop et al., 2016; Alonso et al., 2023a)
+IoT & Service Mesh Architecture (5, 4.13%) (Gattobigio et al., 2022; Rodigari et al., 2021; Merlino et al., 2024; Hossain et al., 2023; Solayman and Qasha, 2023)
+Fault Tolerance & Self-Healing (5, 4.13%) (Bisht and Kaur, 2022; Bolivar et al., 2018; Tang, 2021; Mfula and Nurminen, 2018; Nagarajan et al., 2023)
+Migration & Testing Architecture (3, 2.47%) (Slominski et al., 2015; Serhiienko and Spillner, 2018; Roy and Kalafatis, 2023)
+Optimization & Reliability Model (3, 2.47%) (Bhamare et al., 2015; Aral et al., 2020; Huang and Wang, 2023)
+Domain-Specific Language (2, 1.65%) (Quint and Kratzke, 2018; Ferry et al., 2013)
+Performance Modeling (2, 1.65%) (Barna et al., 2017; Rahman and Lama, 2019)
+Security Approaches Cybersecurity & Access Control (11, 9.09%) (He et al., 2022; Sheridan et al., 2017; Rios et al., 2017; Graupner et al., 2015a; Guechi and Maamri, [n. d.])
+Cloud Analysis Approaches (7, 5.78%) (Pai and Kumar, 2021; Pahl, 2015; Dimitrijevic et al., 2019; Zeck and Bouroudjian, 2017; Kiss, 2018; Gundu et al., 2020)
+Security and Performance Solution (6, 4.95%) (Sabbioni et al., 2020; Wang et al., 2020; Pandey et al., 2022; Ranjbar et al., 2017)
+Security Approaches (5, 4.13%) (Deng et al., 2022; Gao et al., 2017; Bélair et al., 2019; Wahab et al., 2016)
+Provisioning & Restoration Approach (4, 3.30%) (Pellegrini et al., 2017; Lazuka et al., 2022; Wang et al., 2019; Kritikos et al., 2019)
+Cloud Management and Interoperability Framework Resource & Service Management Framework (6, 4.95%) (Saif et al., 2023; Harwalkar et al., 2019b; Libardi et al., 2014; Pham and Pham, 2015; Sitaram et al., 2018; Barletta et al., 2024)
+Monitoring & Deployment Framework (5, 4.13%) (Alhamazani et al., 2014; Trihinas et al., 2015; Fadda et al., 2019; Pham et al., 2015; Horchulhack et al., 2024; Marques et al., 2024)
+Evaluation & Modeling Framework (4, 3.30%) (Ferry et al., 2014; Bou Ghantous and Gill, 2021; Tundo et al., 2024; Ramesh et al., 2023)
+Interoperability Framework (2, 1.65%) (Hua et al., 2020; Challita et al., 2017)
+Orchestration and Deployment Approach Benchmarking & Optimization Approach (6, 4.13%) (Almeida et al., 2014; Jha et al., 2019; Shi et al., 2020; Buyya and Barreto, 2015; Guerrero et al., 2018a; Petrovski and Gusev, 2024)
+Service Management Approach (4, 3.30%) (Zhang et al., 2021; Chen et al., 2022; Gupta et al., 2021; Benmerar et al., 2023)
+Simulation & Integration Approach (3, 2.47%) (Alonso et al., 2019; Mittal and Risco-Martín, 2017; del Castillo et al., 2013)
+Storage & Services Solution (3, 2.47%) (Harwalkar et al., 2019a; Shen et al., 2017; Orzechowski et al., 2023)
+Optimization and Deployment Approaches and Tool Optimization & Scheduling Approach (10, 8.26%) (Wong et al., 2019; Al-shammari and Alwan, 2018; Lorenzo et al., 2019; Altran et al., 2022; da Silva et al., 2023; Dogani et al., 2023; Pandi et al., 2023; Sofia et al., 2023; Nguyen and Chien, 2023)
+Multi-Cloud Platform (6, 4.95%) (Verdugo et al., 2017; Ouyang et al., 2022; Dreibholz et al., 2019; Janarthanan et al., 2018; Alaluna et al., 2017; Wong et al., 2019; Al-shammari and Alwan, 2018; Lorenzo et al., 2019; Altran et al., 2022; Santos et al., 2023)
+Deployment Approach and Tool (6, 4.95%) (Cui et al., 2016; Casola et al., 2017; Tricomi et al., 2017; Di Modica et al., 2019; Liu and Qin, 2024; Tiwari and Sharma, 2023)
 […]
 3.2. Container Roles and Implementation Strategies (RQ1)
-The next two subsections describe the roles and implementation strategiesS related to containers in multi-cloud environment.
+We report the results related to roles and implementation strategies in the following two subsections.
 3.2.1. Container Roles
-Table 7 presents a systematic breakdown of the roles of containers in various aspects of multi-cloud environment.
-It is organized by the “Category Name” and the “Role of Containers” which detail the specific context or usage of containers in that category.
+We classify the roles of containers in a multi-cloud environment into six categories, as shown in Table 7. Each category is briefly reported below.
 1.
-Container-based Application Deployment: This category discusses the use of containers for the deployment of various types of applications.
-This includes web apps, Platform as a Service (PaaS) solutions, data-intensive applications, and Internet of Things (IoT) apps, among others.
-Containers provide a means of packaging and distributing applications, including their dependencies, across various environments.
+Container-Based Resource and Service Management: This category includes studies discussing the use of containers for resource management and hosting.
+The studies underline how the use of containers enables an efficient multi-tenancy solution, sharing of resources, and reduction in complexity due to efficient management of computational resources.
+It also supports OS-level virtualization and multi-tenant service management, enhancing general resource utilization.
+In the context of hosting services, containers provide scalable and isolated environments for application components and services.
+They support microcloud scaling, ensure efficient service delivery, and are vastly used in edge computing for hosting applications closer to end users.
+Each of these capabilities together shows why containers are very important for optimizing resource management and offering flexible and reliable hosting of services in cloud environments.
 2.
-Container-based Service Hosting: This category refers to the use of containers as a hosting environment for different services.
-Container-based service hosting allow for scaling and isolation of these services, creating an efficient multi-cloud environment.
+Container-based Application Deployment: Containers provide a way to package and distribute applications, along with their dependencies, across different environments.
+This category reports on the use of containers for deploying various types of applications, such as web apps, Platform as a Service (PaaS) solutions, data-intensive applications, and IoT apps.
 3.
-Container-Based Resource Management: This category discusses the use of containers to manage computational resources.
-Containers allow for better multi-tenancy solutions, resource sharing, and reduction in complexity, among other benefits.
+Containerization in App Development and IoT: This category focuses on application development and IoT integration using containers in multi-cloud environment.
+Key areas include improving deployment speed, enhancing DevOps integration for better portability and efficiency, and enabling standardized software units.
+Containers are also use to address issues like privacy, compliance, and vendor lock-in.
+Notable highlights include their role in orchestrating container placement in fog computing and integration with tools like OpenVAS and Chef for enhanced security and automation.
 4.
-Container Technology Features: This category discusses the specific features and advantages of container technology, such as lightweight virtualization, use in edge computing, and compatibility with orchestration tools like Kubernetes.
+Performance and Efficiency through Containers: In this category, we classified studies that discuss how containers improve performance and efficiency.
+The selected studies highlights their role in enabling scalable application development and enhancing performance isolation for SaaS.
+Additionally, studies reported that containers contribute to cloud platform optimization, boosting overall cloud performance, and improving cost efficiency specifically while using microservices.
+Further discussions include evaluating the performance impact of containers across various systems.
 5.
-Containerization in App Development and IoT: The focus of this category is on the use of containers for app development, particularly in the context of IoT.
-Containers provide a means of encapsulating software units, increasing portability, improving efficiency, and integration with DevOps practices, among other benefits.
+Container Technology Features: This category also gathers the studies that report the special characteristics and advantages of container technology, to be specific: lightweight virtualization, where the containers improve resource efficiency in edge computing, and applications are brought closer to end users.
+Additionally, their compatibility with orchestration tools like Kubernetes simplifies container management and deployment across diverse environments.
 6.
-Performance and Efficiency through Containers: This category discusses how containers can enhance performance and efficiency in a cloud environment.
-Containers can improve scalability, isolation, and cost-effectiveness, and also aid in platform optimization.
-7.
-Workload Management via Containers: This category is about the use of containers to manage computational workloads across multiple cloud environments.
-Containers facilitate the migration and execution of these workloads across different platforms.
-8. Security and Compliance via Containers: In this category, containers are discussed as a tool for enhancing security and compliance in multi-cloud environment.
-9.
-Other Container Uses: This category includes various other uses of containers, such as data storage, support in architecture, blockchain reliability, scheduling algorithm evaluation, and virtual network composition.
-Each of these categories demonstrates a dedicated aspect of how containers can contribute to efficient and effective multi-cloud management.
-The studies listed in each row (e.g., S01, S02) provide a more in-depth discussion and evidence for the roles and applications of containers.
+Other Container Uses: This category report several roles of container technology beyond traditional use cases.
+Key roles include container-based data storage, support for system architecture, and enhancing blockchain reliability in multi-cloud environment.
+Additionally, containers are used for evaluating scheduling algorithms and enabling virtual network composition, demonstrating their versatility in addressing a wide range of technical challenges.
 Table 7. Roles of containers in multi-cloud environment
 Category Role of Containers ID
-Container-based Application Deployment Containerized Web-App Deployment S01
-Container-Driven App Deployment S05, S08
-Containerized PaaS Prototyping S09
-Data-Intensive App Deployment via Containers S28
-Containerized App Management S29
-Container-Based Aneka Service Deployment S51
-Container-Driven Microservices Deployment & Scaling S62
-Container-Based Multi-Cloud App Deployment & Scaling S74
-Containerized IoT App Deployment S85
-Container-based Service Hosting Container-Based App Component Hosting S02
-Containerized Service Hosting S15
-Container-Driven MicroCloud Scaling & Isolation S42
-Container-Based Resource Management Container-Based Multi-Tenancy Solution S04
-Container-Driven Resource Sharing & Multi-Tenancy S06
-Multi-Tenant Service Management via Containers S17
-Container-Enabled OS-Level Virtualization S40
-Container-Based Complexity Reduction S41
-Container Technology Features Container Usage in Edge Computing S22
-Container-Enabled Lightweight Virtualization S24
-Key Features of Container Technology S25
-Container-Based Edge Computing Virtualization S35
-Container Leverage in Kubernetes S70
-Containerization in App Development & IoT App Development & IoT Integration via Containers S16
-Container-Based OpenVAS & Integration of Chef S30
-Containers as Standardized Software Units S32
-Container Placement & Orchestration in Fog Computing S33
-Container Deployment Speed & DevOps Integration S36
-Containerizing Apps for Privacy & Compliance S37
-Container-Driven Portability & Efficiency Improvement S39
-Preventing Vendor Lock-In via Containers S63
-Container-Based Extensible App Development S76
-Performance & Efficiency via Containers Scalable App Building via Containers S18
-Container-Driven SaaS Performance Isolation S23
-Cloud Performance Enhancement via Containers S48
-Evaluating Performance Impact of Containers S59
-Cost Efficiency of Microservices via Containers S68
-Container-Based Cloud Platform Optimization S69
-Workload Management via Containers Container-Based Virtual Cluster Implementation S21
-Multi-Cloud Workload Migration via Containers S60, S65
-Seamless Workload Migration via Containers S65
-Container-Based Multi-Cloud Platform Execution S79
-Other Container Uses Container-Based Azure Data Storage S54
-Container & VM Support in Architecture S55
-Blockchain-Based Multi-Cloud Reliability via Containers S66
-Evaluating Scheduling Algorithm via Containers S67
-Container-Based Virtual Network Composition S72
+Container-Based Resource and Service Management Container-Based Multi-Tenancy Solution (Ramos et al., 2021)
+Container-Driven Resource Sharing & Multi-Tenancy (Deng et al., 2022)
+Multi-Tenant Service Management via Containers (Slominski et al., 2015)
+Container-Enabled OS-Level Virtualization (Bélair et al., 2019)
+Container-Based Complexity Reduction (Shi et al., 2020)
+Containerized Service Hosting (Pham and Pham, 2015)
+Multi-Cloud Workload Migration via Containers (Janarthanan et al., 2018; Zeck and Bouroudjian, 2017)
+Seamless Workload Migration via Containers (Zeck and Bouroudjian, 2017)
+Container-Based Multi-Cloud Platform Execution (Serhiienko and Spillner, 2018)
+Container-Driven MicroCloud Scaling & Isolation (Baresi et al., 2016)
+Container-based Application Deployment Containerized Web-App Deployment (Almeida et al., 2014)
+Container-Driven App Deployment (Gattobigio et al., 2022; Pham et al., 2015)
+Containerized PaaS Prototyping (Verdugo et al., 2017)
+Data-Intensive App Deployment via Containers (Barna et al., 2017)
+Containerized App Management (Gupta et al., 2021)
+Container-Based Aneka Service Deployment (Buyya and Barreto, 2015)
+Container-Driven Microservices Deployment & Scaling (Rahman and Lama, 2019)
+Container-Based Multi-Cloud App Deployment & Scaling (Mfula and Nurminen, 2018)
+Containerized IoT App Deployment (Bou Ghantous and Gill, 2021)
+Containerization in App Development & IoT App Development & IoT Integration via Containers (Ouyang et al., 2022)
+Container-Based OpenVAS & Integration of Chef (Sheridan et al., 2017)
+Containers as Standardized Software Units (Al-shammari and Alwan, 2018)
+Container Placement & Orchestration in Fog Computing (Sami and Mourad, 2020)
+Container Deployment Speed & DevOps Integration (Lorenzo et al., 2019)
+Containerizing Apps for Privacy & Compliance (Dimitrijevic et al., 2019)
+Container-Driven Portability & Efficiency Improvement (Altran et al., 2022)
+Preventing Vendor Lock-In via Containers (Pellegrini et al., 2017)
+Container-Based Extensible App Development (Pop et al., 2016)
+Performance & Efficiency via Containers Scalable App Building via Containers (Pai and Kumar, 2021)
+Container-Driven SaaS Performance Isolation (Wang et al., 2020)
+Cloud Performance Enhancement via Containers (Saif et al., 2023)
+Evaluating Performance Impact of Containers (Rodigari et al., 2021)
+Cost Efficiency of Microservices via Containers (Guerrero et al., 2018a)
+Container-Based Cloud Platform Optimization (Kiss, 2018)
+Container Technology Features Container Usage in Edge Computing (Wong et al., 2019)
+Container-Enabled Lightweight Virtualization (Pahl, 2015)
+Key Features of Container Technology (Gao et al., 2017)
+Container-Based Edge Computing Virtualization (Carpio et al., 2020)
+Container Leverage in Kubernetes (Lazuka et al., 2022)
+Other Container Uses Container-Based Azure Data Storage (Bucur et al., 2018)
+Container & VM Support in Architecture (Bolivar et al., 2018)
+Blockchain-Based Multi-Cloud Reliability via Containers (Aral et al., 2020)
+Evaluating Scheduling Algorithm via Containers (Tang, 2021)
+Container-Based Virtual Network Composition (Alaluna et al., 2017)
 […]
 3.2.2. Implementation Strategies
-Table 8 provides an overview of various implementation strategies for container technology in multi-cloud environment.
-The strategies are grouped into several broad categories, each with a unique focus and varying subcategories.
-Within these subcategories, each implementation strategy is listed with a concise and revised name for reference.
+Container implementation strategies in multi-cloud environment refer to the various approaches and techniques used to deploy, manage, and orchestrate containerized applications across multiple cloud platforms.
+These strategies ensure efficient, secure, and scalable management of containerized workloads.
+Table 8 provides an overview of these strategies, categorized into subcategories.
+Overall, we identified 76 strategies classified into 11 subcategories and 5 categories.
+Notably, these strategies were identified from only 76 out of 121 selected studies (62.81%).
+The remaining studies do not explicitly discuss the strategies, or it is unclear what strategies they employed.
+One possible reason for this could be that many studies focus on high-level discussions of container technology without detailing specific implementation strategies.
+Additionally, in some cases, the strategies might have been implicitly integrated into broader architectural frameworks, making them difficult to extract or clearly identify.
 1.
 Container Deployment and Management: This category focuses on how containers are deployed and managed in multi-cloud environment.
-An example from the “Deployment Strategies” subcategory is “Automated Orchestration for Web Applications”, where container deployment is automated for web applications.
-From the “Container Management” subcategory, the strategy “Container Image Optimization” highlights how the size and efficiency of a container can be optimized.
+We identified 23 strategies within this category, grouped into three subcategories: Deployment Strategies, Container Lifecycle Management, and Data Management.
+Deployment strategies, with 10 strategies identified, focus on optimizing the deployment of containers in multi-cloud environment, including approaches such as Kubernetes for containerized applications and Docker image management.
+Nine strategies related to container lifecycle management address resource monitoring, image optimization, and privacy, with examples like multi-cloud container resource usage monitoring and container image reorganization.
+The data management subcategory, comprising 4 strategies, focuses on the storage and management of container data in multi-cloud environment, featuring solutions such as Firebase Cloud for real-time databases and blockchain-based container log management.
+While deployment strategies are well-represented, fewer strategies focus on data management, indicating potential areas for further research in managing data across multi-cloud platforms.
 2.
-Container Orchestration: In this category, the focus is on coordinating and managing the life cycle of containers in a multi-cloud setup.
-An instance from the “Cloud Orchestration” subcategory is “SLA-Based Container Strategies”, illustrating how Service Level Agreements (SLAs) can guide the orchestration of containers.
-In the “Microservices Orchestration” subcategory, “Microservices Architecture in IoT Cloud Apps” represents a way to orchestrate microservices in IoT applications using containers.
+Container Orchestration: We identified a total of 20 strategies in this category, classified into two subcategories: Cloud Orchestration and Microservices Orchestration.
+The Cloud Orchestration subcategory, comprising 13 strategies, focuses on managing containers and resources across multi-cloud environment.
+These strategies include AI/ML-driven container orchestration, cross-cloud orchestration tools, and Kubernetes CI/CD for multi-cloud setups.
+The Microservices Orchestration subcategory, consisting of 7 strategies, emphasizes the orchestration of microservices using technologies like Kubernetes and Docker.
+Key strategies include Kubernetes orchestration for multi-cloud containers and edge computing with Docker and Kubernetes.
+These approaches are essential for enabling scalable and efficient management of containerized applications, with Cloud Orchestration strategies focusing on adaptive resource management, while Microservices Orchestration supports effective multi-cloud microservices deployment.
 3.
-Container Security: This category focus on the security of containers.
-An example from the “Container Placement” subcategory is “Secure Placement with Deep Learning”, which employs deep learning techniques to secure the placement of containers.
-From the “Security Policies” subcategory, the strategy “Access Control and Defense for Container Security” suggests ways to control access and defend containers from security threats.
+Container Security: We identified a total of 11 strategies within the Container Security category, classified into two subcategories: Security Policies and Container Placement.
+The Security Policies subcategory consists of eight strategies and mainly focuses on enhancing container security through mechanisms such as access control, secure connectivity, and deep learning-based policy generation.
+Key strategies include implementing defense mechanisms, integrating HIP for Docker security, and managing access in multi-cloud environments using signed URLs.
+These approaches help address security challenges, safeguard containerized applications, and ensure compliance in multi-cloud infrastructures.
+The Container Placement subcategory consists of three strategies and mainly focuses on optimizing container placement to balance security and efficiency.
+These strategies include secure placement using deep learning and the dynamic relocation of microservices across clouds.
+By enabling secure and adaptive placement, these strategies enhance resource utilization, reduce latency, and maintain consistent application performance across cloud environments.
 4.
-Performance and Scaling: This category addresses the optimization of container performance and the scaling of container-based applications.
-For example, from the “Performance Optimization” subcategory, “LXC Virtual Cluster for Performance” shows how Linux Containers (LXC) can be used to enhance the performance of virtual clusters.
-An example from the “Scaling Strategies” subcategory is “Edge Data Streaming and Real-Time Database”, highlighting a strategy for scaling data streaming and real-time databases at the edge.
+Performance Optimization and Scaling Strategies: We identified a total of 12 strategies within the Performance Optimization and Scaling Strategies category, divided into two subcategories: Container Performance Optimization and Scaling Strategies.
+The Container Performance Optimization subcategory, consisting of seven strategies, focuses on improving container performance in multi-cloud environments.
+Key approaches include reactive auto-scaling and resource provisioning, AI-driven resource allocation, and multi-cloud resource monitoring.
+These strategies aim to enhance performance by optimizing resource usage and reducing latency across cloud platforms.
+The Scaling Strategies subcategory, with five strategies, centers on improving the scalability of containerized applications.
+Notable approaches include managing distributed workloads with Kubernetes, utilizing hybrid abstractions for container scaling, and implementing MicroCloud architectures for efficient scaling.
+Together, these strategies support the dynamic adjustment of containerized applications to fluctuating resource demands in multi-cloud setups while maintaining performance and efficiency.
 5.
-Cloud Service and Networking: This category deals with the management of cloud services and networking in containerized environments.
-“Multi-Tenant Service with Docker and Cloudant” from the “Cloud Service Management” subcategory illustrates how Docker and Cloudant can be utilized to manage multi-tenant services.
-From the “Networking” subcategory, the strategy “Edge Scheduling with Location-Awareness” proposes a way to schedule network services at the edge based on location.
-6.
-Data Storage and Persistence: This category discusses strategies related to data management in containerized environments.
-For instance, the strategy “Firebase Cloud for Real-Time Database” presents an approach that uses Firebase Cloud to manage real-time databases in container-based applications.
+Cloud Service and Networking: We identified a total of 10 strategies within the Cloud Service and Networking category, classified into two subcategories: The Cloud Service Integration subcategory, with five strategies, focuses on integrating services across multi-cloud platforms.
+Key approaches include multi-cloud runtime environments, abstraction layers for cloud service standardization, and the integration of OpenStack and IoT in multi-cloud setups.
+These strategies aim to simplify and streamline the seamless integration of services across different cloud platforms.
+The Multi-Cloud Networking Strategies subcategory, also comprising five strategies, addresses networking challenges in multi-cloud environments.
+Notable strategies include edge scheduling with location awareness, federated frameworks for cloud storage, and TOSCA-based deployment models for multi-cloud setups.
+These approaches ensure efficient and secure communication between clouds, supporting smooth service operation and data transfer across diverse cloud infrastructures.
 […]
@@
 Category Subcategory Implementation Strategies ID
-Container Deployment and Management Deployment Strategies Automated Orchestration for Web Applications S02
-Kubernetes for Containerized Applications S05
-Multi-cloud App Development Framework S07
-Modular Approach to Container Implementation S08
-Deployment with Docker Images S09
-Docker Image Creation and Management S31
-Container Management Container Image Optimization S32
-Efficient Image Reorganization S36
-Orchestrating with Docker and Kubernetes S37
-Executable Images on Multi-cloud Platforms S79
-Container Orchestration Cloud Orchestration SLA-Based Container Strategies S04
-Task Division and Scaling in Container Strategy S12
-Location-Aware Service Brokering S14
-Docker Containers for Multi-tenant Services S17
-Resource Acquisition for Aneka Containers S51
-Cross-Cloud Container and VM Management S60
-Container Orchestration Technologies Overview S33
-Microservices Orchestration Microservices Architecture in IoT Cloud Apps S16
-ADAPT Deployment of Microservices S27
-Microservices Benchmark with Docker and Kubernetes S62
-Container Security Container Placement Secure Placement with Deep Learning S06
-Container Placement Strategies S39
-Security Policies SLA-Oriented Performance Isolation S23
-Access Control and Defense for Container Security S25
-HIP Integration for Docker Security S77
-Secure Connectivity in Multi-tenant Environment S72
-Signed URLs for Access Control in Multi-cloud S71
-Performance and Scaling Performance Optimization LXC Virtual Cluster for Performance S21
-Scalable Multi-layer Architecture with Containers S28
-Reactive Auto-scaling and Resource Provisioning S48
-Istio for Control Plane Impact and Latency Reduction S59
-Scaling Strategies Edge Data Streaming and Real-Time Database S35
-MicroCloud Architecture for Efficient Scaling S42
-Distributed Workloads with Kubernetes S70
-Hybrid Abstractions for Container and Host S40
-Cloud Service and Networking Cloud Service Management Multi-Tenant Service with Docker and Cloudant S17
-Multi-Cloud Runtime Environment S76
-Abstraction Layer for Cloud Service Standardization S63
-Networking Edge Scheduling with Location-Awareness S22
-Openstack Networking with Kuryr S55
-Federated Framework for Cloud Storage S64
-Host Identity Protocol for Docker Networking S77
-Data Storage and Persistence Data Management Firebase Cloud for Real-Time Database S35
-High Availability with SpyStorage MCSS S75
-Data Storage with Object Store Service S76
+Container Deployment and Management Deployment Strategies Automated Orchestration for Web Applications (Jha et al., 2019)
+Kubernetes for Containerized APPs (Gattobigio et al., 2022)
+Multi-cloud App Development Framework (Zhang et al., 2021)
+Modular Approach to Container Implementation (Pham et al., 2015)
+Deployment with Docker Images (Verdugo et al., 2017)
+Docker Image Creation and Management (Mittal and Risco-Martín, 2017)
+Real-time Kubernetes workloads in multi-cloud (Barletta et al., 2024)
+AWS multi-cloud container app deployment (Kataru et al., 2023)
+Multi-cloud container policy enforcement (Liu and Qin, 2024)
+Probes and orchestration for cloud containers (Tundo et al., 2024)
+Container Lifecycle Management Container Image Optimization (Al-shammari and Alwan, 2018)
+Efficient Image Reorganization (Lorenzo et al., 2019)
+Orchestrating with Docker and Kubernetes (Dimitrijevic et al., 2019)
+Executable Images on Multi-cloud Platforms (Serhiienko and Spillner, 2018)
+Privacy monitoring for containerized clouds (Barati et al., 2023)
+Multi-cloud container resource usage monitoring (Horchulhack et al., 2024)
+Container ecosystem for multi-cloud deployment (Wong et al., 2023)
+Containers in AWS ECS for cloud services (Petrovski and Gusev, 2024)
+Container management for multi-cloud MEC (Santos et al., 2023)
+IoT container deployment in multi-cloud (Solayman and Qasha, 2023)
+Data Management Firebase Cloud for Real-Time Database (Carpio et al., 2020)
+High Availability with SpyStorage MCSS (Shen et al., 2017)
+Data Storage with Object Store Service (Pop et al., 2016)
+Container Orchestration Cloud Orchestration SLA-Based Container Strategies (Ramos et al., 2021)
+Task Division and Scaling in Container Strategy (Bisht and Kaur, 2022)
+Location-Aware Service Brokering (Chen et al., 2022)
+Docker Containers for Multi-tenant Services (Slominski et al., 2015)
+Resource Acquisition for Aneka Containers (Buyya and Barreto, 2015)
+Cross-Cloud Container and VM Management (Janarthanan et al., 2018)
+Container Orchestration Technologies Overview (Sami and Mourad, 2020)
+Multi-cloud dynamic resource adaptation (Struhár et al., 2024)
+AI/ML orchestration in multi-cloud containers (Sofia et al., 2024)
+AI-driven container orchestration across clouds (Benmerar et al., 2023)
+Cross-cloud container orchestration tools (Alonso et al., 2023a)
+Kubernetes CI/CD for multi-cloud setups (Pandi et al., 2023)
+Microservices Orchestration Microservices Architecture in IoT Cloud Apps (Ouyang et al., 2022)
+ADAPT Deployment of Microservices (Alonso et al., 2019)
+MSA Benchmark with Docker and Kubernetes (Rahman and Lama, 2019)
+Kubernetes for power-efficient multi-cloud (Centofanti et al., 2024)
+Edge computing with Docker and Kubernetes (Hossain et al., 2023)
+Kubernetes for edge-cloud orchestration (Roy and Kalafatis, 2023)
+Kubernetes orchestration for multi-cloud containers (Sofia et al., 2023)
+Container Security Container Placement Secure Placement with Deep Learning (Deng et al., 2022)
+Container Placement Strategies (Altran et al., 2022)
+Dynamic microservice relocation across clouds (da Silva et al., 2023)
+Security Policies SLA-Oriented Performance Isolation (Wang et al., 2020)
+Access Control and Defense for Container Security (Gao et al., 2017)
+HIP Integration for Docker Security (Ranjbar et al., 2017)
+Secure Connectivity in Multi-tenant Environment (Alaluna et al., 2017)
+Signed URLs for Access Control in Multi-cloud (Graupner et al., 2015a)
+Performance and Scaling Strategies Performance Optimization LXC Virtual Cluster for Performance (Cui et al., 2016)
+Scalable Multi-layer Architecture with Containers (Barna et al., 2017)
+Reactive Auto-scaling and Resource Provisioning (Saif et al., 2023)
+Istio for Control Plane Impact and Latency Reduction (Rodigari et al., 2021)
+Auto-scaling containers across clouds (Dogani et al., 2023)
+Multi-cloud resource monitoring and provisioning (Marques et al., 2024)
+AI resource allocation in multi-cloud containers (Nagarajan et al., 2023)
+Scaling Strategies Edge Data Streaming and Real-Time Database (Carpio et al., 2020)
+MicroCloud Architecture for Efficient Scaling (Baresi et al., 2016)
+Distributed Workloads with Kubernetes (Lazuka et al., 2022)
+Hybrid Abstractions for Container and Host (Bélair et al., 2019)
+Cloud Service and Networking Cloud Service Management Multi-Tenant Service with Docker and Cloudant (Slominski et al., 2015)
+Abstraction Layer for Cloud Service Standardization (Pellegrini et al., 2017)
+OpenStack and IoT in multi-cloud (Merlino et al., 2024)
+Data indexing for multi-cloud exposure (Orzechowski et al., 2023)
+Networking Edge Scheduling with Location-Awareness (Wong et al., 2019)
+Openstack Networking with Kuryr (Bolivar et al., 2018)
+Federated Framework for Cloud Storage (Harwalkar et al., 2019a)
+Host Identity Protocol for Docker Networking (Ranjbar et al., 2017)
 3.3. Pattern and Strategies (RQ2)
-Table 9 presents various patterns and strategies employed in containerized-based applications in multi-cloud environment.
+Table 9 presents various patterns and strategies employed in container-based applications in multi-cloud environment.
 These patterns and strategies are divided into various categories and subcategories, each of which highlights different aspects of multi-cloud architecture, management, security, resilience, and migration patterns and strategies.
+Notably, several studies (e.g., (Pahl, 2015; Graupner et al., 2015a; Shen et al., 2017; Pham and Pham, 2015)) appear across multiple subcategories because they report different patterns that address distinct aspects of cloud architecture and strategies.
+Every pattern is classified based on its specific focus, even when originating from the same study.
+Therefore, we categorized them into separate subcategories to reflect their unique contributions.
 Each category is briefly reported below.
 1.
-Cloud Architectural Patterns and Models: This category reports different architecture patterns, communication and networking patterns, deployment patterns, and service models for containerized-based applications in multi-cloud environment.
-Architecture Patterns involve several design models for structuring the components of a system, providing a template that guides system designers in building complex systems.
-The leading patterns identified are microservice architecture and service-oriented architecture (SOA), both of which focus on building applications as suites of independently deployable services.
-Communication and Networking Patterns relate to how different services within the architecture communicate and network with each other.
-Service mesh and Service Chaining (SC) are the main patterns observed.
-Deployment Patterns pertain to different strategies for deploying software systems in multi-cloud environment.
-The two leading patterns are Blue-Green Deployment and Object Store Service.
-Service Models involve different ways of providing cloud services, with Platform-as-a-Service (PaaS) and Software-as-a-Servic (SaaS) leading the pack.
+Cloud Architectural Patterns and Models: We identified 48 patterns within this category, grouped into four subcategories: Architecture Patterns, Communication and Networking Patterns, Deployment Patterns, and Service Models.
+The Architecture Patterns subcategory, with 22 patterns, is the largest, focusing on designing scalable and flexible cloud systems.
+Leading patterns include Microservice Architecture and Service-Oriented Architecture (SOA), which emphasize modularity and independence of services in multi-cloud environment.
+Emerging patterns such as AI-Driven Edge-Cloud Architecture and Blockchain-Based Architecture reflect the incorporation of advanced technologies in cloud design.
+The Communication and Networking Patterns subcategory consists of 11 patterns, with notable strategies such as Service Mesh and Service Chaining (SC) providing essential frameworks for ensuring secure and efficient communication between distributed services across cloud environments.
+These patterns are critical for enabling seamless integration and operation of cloud services in multi-cloud setups.
+The Deployment Patterns subcategory, containing 8 patterns, focuses on strategies for deploying applications across multiple cloud platforms.
+Key patterns include black-Green Deployment, which supports smooth application updates by alternating between two environments, and Object Store Service, which ensures efficient storage of unstructured data across clouds.
+Advanced orchestration patterns like Decentralized Orchestration Architecture and Policy-Driven Orchestration Architecture further enhance deployment efficiency by automating resource management across clouds.
+Lastly, the Service Models subcategory comprises 7 models that define various cloud service delivery mechanisms.
+The leading models, Platform-as-a-Service (PaaS) and Software-as-a-Service (SaaS), provide essential platforms for building, deploying, and managing applications in the cloud without the need for managing underlying infrastructure.
+Other models, like Function-as-a-Service (FaaS) and Serverless Container-Oriented Architecture, emphasize serverless computing, allowing developers to focus on code rather than infrastructure management.
 2.
-Cloud Management and Resource Allocation Strategies: This category presents the patterns and strategies related to the management of multi-cloud environment, container management, edge computing, IoT strategies, and resource management.
-For example, Multi-Cloud Management Strategies include various approaches for managing multiple cloud environments.
-Multi-Cloud Computing Strategy and Hybrid Cloud Architecture are found leading strategies in this category.
-Container Management Strategies involve the management of containers in multi-cloud environment, with the Linux Container (LXC) project and Container Engine as the primary patterns identified.
-Edge Computing and IoT Strategies pertain to the use of edge computing and IoT in multi-cloud environment.
-Mobile Edge Computing (MEC) and Fog Computing are leading strategies for container management in multi-cloud environment.
-Resource Management Strategies involve methods for managing resources in the multi-cloud with the help of Kubernetes and Minikube.
+Cloud Management and Resource Allocation Strategies: We identified 30 strategies within this category, grouped into four subcategories: Multi-Cloud Management Strategies, Container Management Strategies, Edge Computing and IoT Strategies, and Resource Management Strategies.
+The Multi-Cloud Management Strategies subcategory, with 8 strategies, focuses on managing and optimizing resources across multiple cloud environments.
+Leading strategies include the Multi-cloud Computing Strategy, Hybrid/Multi-cloud Approach, and MAPE-K control loop, all of which emphasize seamless integration and resource optimization across cloud platforms.
+The Container Management Strategies subcategory contains 7 strategies that address the management of containers in cloud environments.
+Key strategies include Docker Container Images and Lightweight Virtualization, both of which streamline container management and enhance resource efficiency in multi-cloud setups.
+The Edge Computing and IoT Strategies subcategory, with 8 strategies, highlights the integration of edge computing and IoT devices with cloud systems.
+Notable strategies include Mobile Edge Computing (MEC) and Fog Computing, which focus on distributing computation and storage closer to IoT devices for improved performance.
+Finally, the Resource Management Strategies subcategory includes 7 strategies that focus on efficient resource allocation in cloud environments.
+Leading approaches like Resource Allocation based on SLA levels, Container Orchestration, Kubernetes, Minikube, and AI-Driven Edge-Cloud Architecture emphasize intelligent resource distribution and orchestration across cloud and edge environments.
 3.
-Cloud Security and Resilience Strategies: In this category, the emphasis is on strategies for ensuring the security and resilience of software systems in multi-cloud environment.
-Security and Resiliency Patterns are strategies to ensure the system is secure and can recover from failures.
-The two most prominent patterns are the security-by-design approach and encryption of files.
-Fault-tolerance Strategies are methods used to ensure that systems can continue to function even when there are failures.
-Fault-tolerance with Redundant Engines and Multi-cloud Systems Fault-Tolerant Workflow are the leading patterns.
+Cloud Security and Resilience Strategies: We identified 20 strategies within this category, divided into two subcategories: Security and Resiliency Patterns and Fault-tolerance Strategies.
+The Security and Resiliency Patterns subcategory, comprising 14 strategies, focuses on enhancing security and data resilience in cloud environments.
+Prominent strategies include Security-by-Design approach, Encryption of files, and Centralized/External ACM Service, which aim to secure data sharing and access control.
+Emerging patterns like Blockchain-Driven SLA Management Architecture and Hybrid Malware Detection Architecture highlight the use of advanced technologies to ensure resilient cloud infrastructures.
+The Fault-tolerance Strategies subcategory contains 6 strategies that focus on ensuring system reliability in case of failures.
+Key strategies include Fault-tolerance with Redundant Engines and Multi-cloud Systems Fault-tolerant Workflow, both of which emphasize maintaining uninterrupted service in multi-cloud environment by building in redundancy and fault-tolerant processes.
+The RAFT Consensus Algorithm ensures consistency across distributed systems, further improving the resilience of containerized applications in a multi-cloud environment.
+.
 4.
-Cloud Migration Strategies: This category involves patterns and strategies for moving applications and data between cloud environments and optimizing the migration process.
-Image-based Migration, for example, is an efficient way to migrate applications because it packages the applications along with their dependencies, configurations, and runtime environment.
-Migration to a virtualized container, such as Docker, is increasingly popular because it guarantees application portability across various environments and scalability on different cloud platforms.
+Cloud Migration Strategies: We identified 4 strategies within this category, which focus on various approaches to migrating applications and services to the cloud.
+These strategies include Service-oriented Migration, Application-centric Migration, Image-based Migration, and Migration to a Virtualized Container.
+Each of these strategies provides a unique method for transferring workloads and data from on-premises or legacy systems to cloud environments.
+The emphasis is on ensuring a smooth transition with minimal disruption to services, whether through the migration of individual services, entire applications, or the use of virtualized containers.
+These strategies are essential for organizations looking to leverage cloud technologies while preserving the integrity of their existing systems during the migration process.
 Table 9. Patterns and strategies for container-based applications a multi-cloud environment
-Categories Subcategories
-Patterns
-ID
-Cloud Architectural Patterns and Models Architecture Patterns
-Microservice Architecture
-S02, S09, S16, S77, S80, S24, S31, S33, S68
-Service-Oriented Architecture (SOA)
-S51, S52, S86
-Multitier Architecture, Layered Architecture
-S02, S04, S54
-Client-Server Architecture
-S03, S54
-Component-based Architecture
-S22
-Event-driven Architecture
-S44
-Multi-agent Architecture
-S74, S48
-Master-worker Nodes Architecture
-S33
-Netflix Zuul-based Seeker Component
-S17
-Multi-tenant Cloud Service Architecture
-S17
-Stateful Engine Architecture
-S17
-Communication and Networking Patterns
-Service Mesh
-S05
-Service Chaining (SC)
-S50
-Multi-Cloud APIs
-S71
-Third-Party APIs
-S71
-Service Request Broker
-S07
-Publish and Subscribe Communication Protocol
-S45
-Synchronous Network Communication Protocols
-S01
-Asynchronous Network Communication Protocols
-S01
-Network Function Virtualization (NFV)
-S50
-Service Discovery
-S76
-Deployment Pattern
-Blue-Green deployment
-S27, S30
-Object store service
-S76
-Multi-cloud Deployment Model
-S10
-Distributed Deployment Model
-S10
-Accelerated Particle Swarm Optimization for Deployment
-S16
-Service Models
-Platform-as-a-Service (PaaS)
-S24, S41
-Software as a Service
-S79
-Infrastructure-as-a-Service (IaaS)
-S24
-Function-as-a-Service (FaaS)
-S41
-Cloud Management and Resource Allocation Strategies Multi-Cloud Management Strategies
-Multi-cloud computing strategy
-S78, S81
-Hybrid cloud architecture
-S78
-Hybrid/Multi-cloud Approach
-S05, S83
-Introducing multi-cloud Middleware
-S52, S53
-MAPE-K control loop
-S46, S48
-Connecting to multiple cloud service providers
-S76
-Multi-cloud load balancing
-S23, S76
-Multi-Cloud BP provisioning
-S84
-Cross-level orchestration of cloud services
-S84
-Cross-level monitoring and adaptation of BPs
-S84
-Container Management Strategies
-Linux Container (LXC) project
-S24
-Container engine
-S24
-Docker Container Images
-S24
-Lightweight virtualization
-S24
-Portable application packaging
-S24
-One-container-per-app approach
-S24
-Data volumes and data volume containers
-S24
-Edge Computing and IoT Strategies
-Mobile Edge Computing (MEC)
-S22, S70, S43, S54
-Fog computing
-S32
-Edge Services
-S35
-Connecting IoT Edge Devices to the Cluster
-S05
-Cloud of Things (CoT)
-S05
-Resource Management Strategies
-Resource allocation based on different SLA levels
-S23
-Container Orchestration, Kubernetes, Minikube
-S24, S29, S33
-Cloud Security and Resilience Strategies Security and Resiliency Patterns
-Security-by-Design approach
-S51
-Encryption of files
-S71
-Standardized APIs for file transfer
-S71
-Centralized/External ACM Service
-S71
-Distributed ACM Service
-S71
-Dynamic Switching between Authentication Methods
-S71
-ACM based on Signed URLs
-S71
-Attribute-Based Encryption and Signature
-S75
-Secure data sharing
-S75
-Data access control
-S75
-Local encryption and signing
-S75
-Authorized Tokens
-S75
-Byzantine quorum protocol
-S75
-Fault-tolerance Strategies
-Fault-tolerance with Redundant Engines
-S17
-Multi-cloud systems fault-tolerant workflow
-S67
-RAFT Consensus Algorithm
-S13
-Cloud Migration Strategies
-Service-oriented Migration
-S15
-Application-centric Migration
-S15
-Image-based Migration
-S15
-Migration to a Virtualized Container
-S15
+Categories Subcategories Patterns ID
+Cloud Architectural Patterns and Models Architecture Patterns Microservice Architecture (Jha et al., 2019; Verdugo et al., 2017; Ouyang et al., 2022; Ranjbar et al., 2017; Quint and Kratzke, 2018; Pahl, 2015; Mittal and Risco-Martín, 2017; Sami and Mourad, 2020; Guerrero et al., 2018a; Centofanti et al., 2024; Hossain et al., 2023; Tiwari and Sharma, 2023)
+Service-Oriented Architecture (SOA) (Buyya and Barreto, 2015; Casola et al., 2017; Gundu et al., 2020; Merlino et al., 2024)
+Multitier Architecture, Layered Architecture (Jha et al., 2019; Ramos et al., 2021; Bucur et al., 2018; Barati et al., 2023; Horchulhack et al., 2024; Liu and Qin, 2024; Tundo et al., 2024)
+Client-Server Architecture (Hua et al., 2020; Bucur et al., 2018)
+Component-based Architecture (Wong et al., 2019)
+Event-driven Architecture (Wurster et al., 2018)
+Multi-agent Architecture (Mfula and Nurminen, 2018; Saif et al., 2023)
+Master-worker Nodes Architecture (Sami and Mourad, 2020)
+Netflix Zuul-based Seeker Component (Slominski et al., 2015)
+Multi-tenant Cloud Service Architecture (Slominski et al., 2015)
+Stateful Engine Architecture (Slominski et al., 2015)
+Distributed Multi-Cloud Native Architecture (Alonso et al., 2023a)
+Plugin-Based Deployment Automation Architecture (Harzenetter et al., 2023)
+Cloud-Edge Integrated Robotics Architecture (Huang and Wang, 2023)
+Hybrid Malware Detection Architecture (Wang et al., 2023)
+Cloudfront-Enhanced Container Architecture (Kataru et al., 2023)
+AI-Driven Edge-Cloud Architecture (Benmerar et al., 2023)
+Metadata-Driven Architecture (Orzechowski et al., 2023)
+Blockchain-Based Architecture (Ahmad and Aujla, 2023)
+Cross-Layered Edge-Cloud Architecture (Sofia et al., 2023; Santos et al., 2023)
+Communication and Networking Patterns Service Mesh (Gattobigio et al., 2022)
+Service Chaining (SC) (Bhamare et al., 2015)
+Multi-Cloud APIs (Graupner et al., 2015a)
+Third-Party APIs (Graupner et al., 2015a)
+Service Request Broker (Zhang et al., 2021)
+Publish and Subscribe Communication Protocol (Trihinas et al., 2015)
+Synchronous Network Communication Protocols (Almeida et al., 2014)
+Asynchronous Network Communication Protocols (Almeida et al., 2014)
+Network Function Virtualization (NFV) (Bhamare et al., 2015)
+Service Discovery (Pop et al., 2016)
+Blockchain-Based Identity and Access Management (IAM) Architecture (Fugkeaw, 2023)
+Federated Cloud-Edge Architecture (Roy and Kalafatis, 2023)
+Deployment Pattern black-Green deployment (Alonso et al., 2019; Sheridan et al., 2017)
+Object store service (Pop et al., 2016)
+Multi-cloud Deployment Model (Sabbioni et al., 2020)
+Distributed Deployment Model (Sabbioni et al., 2020)
+Multi-Cloud Deployment Architecture (Petrovski and Gusev, 2024; Pandi et al., 2023)
+Decentralized Orchestration Architecture (Sofia et al., 2024)
+Policy-Driven Orchestration Architecture (Lee and Nam, 2023)
+DevOps-Oriented Architecture (Solayman and Qasha, 2023)
+Service Models Platform-as-a-Service (PaaS) (Pahl, 2015; Shi et al., 2020)
+Software-as-a-Service (Serhiienko and Spillner, 2018)
+Infrastructure-as-a-Service (IaaS) (Pahl, 2015)
+Function-as-a-Service (FaaS) (Pahl, 2015; Ramesh et al., 2023)
+Multi-Cloud Orchestration Architecture (Pandi et al., 2023)
+Serverless Container-Oriented Architecture (Shi et al., 2020)
+Cloud Management and Resource Allocation Strategies Multi-Cloud Management Strategies Multi-cloud computing strategy S78, S81
+Hybrid cloud architecture (Wang et al., 2019)
+Hybrid/Multi-cloud Approach (Gattobigio et al., 2022; Wahab et al., 2016)
+Introducing multi-cloud Middleware (Casola et al., 2017; Acquaviva et al., 2017)
+MAPE-K control loop (Fadda et al., 2019; Saif et al., 2023)
+Connecting to multiple cloud service providers (Pop et al., 2016)
+Multi-cloud load balancing (Wang et al., 2020; Pop et al., 2016)
+Multi-Cloud BP provisioning (Kritikos et al., 2019)
+Cross-level orchestration of cloud services (Kritikos et al., 2019)
+Cross-level monitoring and adaptation of BPs (Kritikos et al., 2019)
+Container Management Strategies Linux Container (LXC) project (Pahl, 2015)
+Container engine (Pahl, 2015)
+Docker Container Images (Pahl, 2015)
+Lightweight virtualization (Pahl, 2015)
+Portable application packaging (Pahl, 2015)
+One-container-per-app approach (Pahl, 2015)
+Data volumes and data volume containers (Pahl, 2015)
+Edge Computing and IoT Strategies Mobile Edge Computing (MEC) (Wong et al., 2019; Lazuka et al., 2022; Dreibholz et al., 2019; Bucur et al., 2018)
+Fog computing (Al-shammari and Alwan, 2018)
+Edge Services (Carpio et al., 2020)
+Connecting IoT Edge Devices to the Cluster (Gattobigio et al., 2022)
+Cloud of Things (CoT) (Gattobigio et al., 2022)
+Multi-Access Edge Computing (MEC) Architecture (Santos et al., 2023)
+Cross-Layered Edge-Cloud Architecture (Sofia et al., 2023)
+Cloud-Edge Integrated Robotics Architecture (Huang and Wang, 2023)
+AI-Driven Edge-Cloud Architecture (Benmerar et al., 2023)
+Resource Management Strategies Resource allocation based on different SLA levels (Wang et al., 2020)
+Container Orchestration, Kubernetes, Minikube (Pahl, 2015; Gupta et al., 2021; Sami and Mourad, 2020)
+Rate-Based Stream Processing Architecture (Nguyen and Chien, 2023)
+Multi-Agent Resource Optimization Architecture (Nagarajan et al., 2023)
+Cloud Security and Resilience Strategies Security and Resiliency Patterns Security-by-Design approach (Buyya and Barreto, 2015)
+Encryption of files (Graupner et al., 2015a)
+Standardized APIs for file transfer (Graupner et al., 2015a)
+Centralized/External ACM Service (Graupner et al., 2015a)
+Distributed ACM Service (Graupner et al., 2015a)
+Dynamic Switching between Authentication Methods (Graupner et al., 2015a)
+ACM based on Signed URLs (Graupner et al., 2015a)
+Attribute-Based Encryption and Signature (Shen et al., 2017)
+Secure data sharing (Shen et al., 2017)
+Data access control (Shen et al., 2017)
+Local encryption and signing (Shen et al., 2017)
+Authorized Tokens (Shen et al., 2017)
+Byzantine quorum protocol (Shen et al., 2017)
+Blockchain-Driven SLA Management Architecture (Neeraj et al., 2023)
+Hybrid Malware Detection Architecture (Wang et al., 2023)
+Fault-tolerance Strategies Fault-tolerance with Redundant Engines (Slominski et al., 2015)
+Multi-cloud systems fault-tolerant workflow (Tang, 2021)
+RAFT Consensus Algorithm (Naik, 2016a)
+Cloud Migration Strategies Service-oriented Migration (Pham and Pham, 2015)
+Application-centric Migration (Pham and Pham, 2015)
+Image-based Migration (Pham and Pham, 2015)
+Migration to a Virtualized Container (Pham and Pham, 2015)
 […]
@@
 Table 10 provides a comprehensive overview of the quality attributes and related tactics for containerized applications in multi-cloud environment.
-It is based on data extracted from 86 selected studies, which are organized following the ISO 25010 standard.
+It is based on data extracted from 121 selected studies, which are organized following the ISO 25010 standard.
 This standard defines high-level QAs along with their sub-characteristics.
@@
 Each QA has its related terms, associated study numbers, and corresponding tactics to achieve them.
-In Table 10, five tactics are provided for each quality attribute, due to the limited space.
-The complete set of 195 identified tactics is available in the provided dataset (Waseem et al., 2024), presenting a comprehensive map of tactics with QAs.
-However, it is important to note that these tactics are scattered across different studies, and only a few studies explicitly mentioned these tactics for specific QA.
+In Table 10, due to limited space, we provide only five tactics for each QA.
+The complete set of 70 tactics is listed in the provided dataset (Waseem et al., 2024).
+However, it is important to note that these tactics are scattered across different studies.
 Below we briefly discuss how each QA can be achieved by applying the mentioned tactics.
-Performance (Efficiency): The quality attribute of performance (or efficiency) in containerized applications can be achieved by employing tactics such as efficient resource optimization and allocation, use of ML techniques for performance optimization, utilization of container technology, location-aware service brokering, and performance-oriented Service Level Agreements (SLA).
+Performance (Efficiency): Performance or efficiency is the most dominant QA, frequently discussed in 60 (49.58%) of the selected studies, along with related terms or characteristics such as time behavior, resource utilization, and capacity.
+We also listed several tactics that can be used to achieve performance (or efficiency) in containerized applications, such as efficient resource optimization and allocation, use of machine learning techniques for performance optimization, utilization of container technology, location-aware service brokering, and performance-oriented Service Level Agreements (SLA).
 For instance, optimizing resources ensures that the system uses the minimum possible resources while delivering the required output.
@@
 Similarly, utilizing container technology can help maintain optimal performance by encapsulating the application and its environment.
-Security: Security can be enhanced using tactics like encryption and strong authentication mechanisms, implementation of firewalls, VPNs, or SDNs for network security, utilization of ML techniques to detect and prevent security threats and attacks, deployment of applications on diverse cloud providers, and consideration of users’ security specifications.
+These results also indicate that performance is a critical QA, with a strong focus in containerized applications, and the majority of studies advocate for advanced tactics to achieve optimal performance.
+Security: This QA has been discussed in 38 studies along with various tactics.
+We also identified several characteristics of security that have been highlighted in the selected studies, such as confidentiality, authenticity, access control, authorization, privacy, trustworthiness, and integrity.
+Some of the identified tactics to improve the security of containerized applications from the selected studies include encryption and strong authentication mechanisms, implementation of firewalls, VPNs, or SDNs for network security, utilization of ML techniques to detect and prevent security threats and attacks, deployment of applications on diverse cloud providers, and consideration of users’ security specifications.
 For example, encryption provides a secure way of transmitting data, and ML can detect unusual behavior that could signal a threat or attack.
-Compatibility: The tactics that can be employed to achieve compatibility include implementation of lightweight communication protocols and modes, standardization of interfaces for seamless integration, componentization and modular design for interoperability, prototyping and exploration of interoperability approaches for multi-cloud deployment, and interoperability standardization and federation between clouds.
+Compatibility: This QA has been reported in 15 studies, along with characteristics such as interoperability and co-existence.
+Several tactics that can be employed to achieve compatibility include the implementation of lightweight communication protocols and modes, standardization of interfaces for seamless integration, componentization and modular design for interoperability, prototyping and exploring interoperability approaches for multi-cloud deployment, and interoperability standardization and federation between clouds.
 Standardizing interfaces can ensure that different software components can interact with each other seamlessly, enhancing compatibility.
-Scalability: Scalability can be achieved through tactics like combining containerization and microservices for enhanced scalability, implementation of elastic resource allocation, enabling migration between multi-cloud services, combining container-based deployment with runtime monitoring and optimization, and leveraging orchestration, federated networks, and geographic placement.
+The results suggest that compatibility is essential for ensuring smooth integration and operation in multi-cloud environment.
+Scalability: Scalability refers to the ability to expand capacity or performance without losing the efficiency or functionality of a software system.
+We identified 40 studies that report scalability along with other characteristics like capacity, extensibility, elasticity, and throughput.
+We also identified ten tactics, five of which are listed in Table 10, to achieve scalability.
+For example, scalability can be achieved through tactics like combining containerization and microservices for enhanced scalability, implementing elastic resource allocation, enabling migration between multi-cloud services, combining container-based deployment with runtime monitoring and optimization, and leveraging orchestration, federated networks, and geographic placement.
 Elastic resource allocation ensures that the system can seamlessly scale up or down in response to changing demands.
-Reliability: Tactics such as deploying redundant engines for fault tolerance, distributing resources and replicating applications for improved response time, enforcing redundancy and distributed services for availability, deploying parallel search and multi-cloud distribution, and building fault-tolerant systems can enhance system reliability.
-Portability: Portability can be achieved using tactics such as virtual machine-based packaging, Docker image-based packaging, cloud-portable containerization, dynamic cross-level adaptation and provisioning, and adaptive rule-based system modification.
+Reliability: This QA is about the system’s correct operation and consistent performance without failure over a specified period, while ensuring the availability of the system.
+We identified XX studies that report on reliability, along with characteristics such as maturity, availability, and fault tolerance.
+We identified nine tactics overall (see Table 10 and the QA and Tactics Sheet in (Waseem et al., 2024)) to achieve reliability from the selected studies.
+These include deploying redundant engines for fault tolerance, distributing resources and replicating applications for improved response time, enforcing redundancy and distributed services for availability, deploying parallel search and multi-cloud distribution, and building fault-tolerant systems to enhance system reliability.
+Portability: Portability refers to a system’s ability to function correctly across different platforms (e.g., operating systems, hardware configurations).
+We identified 6 studies that report on portability, along with characteristics such as installability, replaceability, and adaptability, as well as 5 tactics that can help achieve portability.
+For example, portability can be achieved using tactics such as virtual machine-based packaging, Docker image-based packaging, cloud-portable containerization, dynamic cross-level adaptation and provisioning, and adaptive rule-based system modification.
 These tactics ensure that an application can be easily transferred from one computing environment to another.
-Flexibility: Flexibility can be improved through tactics like service-oriented architecture, on-demand dynamic allocation of resources across different cloud platforms, containers, container orchestration, model-driven development, risk analysis, and cross-cloud service orchestration.
+Flexibility: This QA represents the system’s capacity to accommodate new features based on user requirements with minimal effort.
+We identified 10 studies (see Table 10 and the QA and Tactics Sheet in (Waseem et al., 2024)) that report this QA, along with five tactics that can help achieve it.
+For example, flexibility can be improved through tactics such as service-oriented architecture, on-demand dynamic allocation of resources across different cloud platforms, containers, container orchestration, model-driven development, risk analysis, and cross-cloud service orchestration.
 Using multi-cloud environment, for instance, offers the flexibility to choose services from different providers as per specific needs.
-Maintainability: Tactics like microservices for easier maintenance, Infrastructure Provisioning as Code (IaC), containerization and image management, configuration management and templating, version control and change management, and continuous integration and deployment automation can enhance maintainability.
+Maintainability: This QA has been reported in 7 studies, along with several characteristics such as modularity, reusability, and analyzability.
+We also identified 6 tactics from these studies that can help achieve maintainability.
+For example, tactics like microservices for easier maintenance, Infrastructure Provisioning as Code (IaC), containerization and image management, configuration management and templating, version control and change management, and continuous integration and deployment automation can enhance maintainability.
 For example, IaC allows developers to manage infrastructure more efficiently and minimize human errors, thus increasing maintainability.
-Usability: Usability can be enhanced by employing tactics like a feedback-loop controller for multi-cloud infrastructure, container orchestration (e.g., Kubernetes), user-centric interface design, responsive and adaptive user experience, accessible design and compliance, and error handling and feedback mechanisms.
-For instance, a feedback-loop controller is effective in a multi-cloud context as it dynamically adjusts resources across different cloud platforms based on user interaction, ensuring a seamless and responsive user experience tailored to the multi-cloud setup.
+Usability: This QA refers to the ease with which users can interact with a system to perform various operations.
+We identified 6 studies that report on usability, along with several characteristics such as learnability, and operability, as well as 6 other tactics.
+According to the selected studies, usability can be enhanced by employing tactics such as a feedback-loop controller for multi-cloud infrastructure, container orchestration (e.g., Kubernetes), user-centric interface design, responsive and adaptive user experience, accessible design and compliance, and error handling and feedback mechanisms.
+For instance, a feedback-loop controller is effective in a multi-cloud context as it dynamically adjusts resources across different cloud platforms based on user interactions, ensuring a seamless and responsive user experience tailored to the multi-cloud setup.
 […]
@@
 Quality Attribute Related Terms ID Tactics
-Performance (Efficiency) Time behavior, Resource utilization, Capacity, Throughput, Response time S01, S02, S03, S06, S09, S14, S15, S16, S20, S21, S23, S25, S27, S29, S34, S35, S36, S37, S38, S39, S41, S43, S44, S47, S49, S52, S54, S57, S58, S59, S60, S62, S64, S65, S66 Efficient resource optimization and allocation
+Performance (Efficiency) Time behavior, Resource utilization, Capacity, Throughput, Response time (Almeida et al., 2014; Jha et al., 2019; Hua et al., 2020; Deng et al., 2022; Verdugo et al., 2017; Chen et al., 2022; Pham and Pham, 2015), (Ouyang et al., 2022; Ferry et al., 2014; Cui et al., 2016; Wang et al., 2020; Gao et al., 2017; Alonso et al., 2019; Gupta et al., 2021), (Rios et al., 2017; Carpio et al., 2020; Lorenzo et al., 2019; Dimitrijevic et al., 2019; Pandey et al., 2022; Altran et al., 2022; Shi et al., 2020), (Dreibholz et al., 2019; Wurster et al., 2018; Libardi et al., 2014; Harwalkar et al., 2019b; Casola et al., 2017; Bucur et al., 2018; Tricomi et al., 2017), (Sitaram et al., 2018; Rodigari et al., 2021; Janarthanan et al., 2018; Rahman and Lama, 2019; Harwalkar et al., 2019a; Zeck and Bouroudjian, 2017; Aral et al., 2020) Efficient resource optimization and allocation
 Utilizing machine learning techniques for performance optimization
@@
 Performance-oriented Service Level Agreements (SLA)
-Security Confidentiality, Authenticity, Access control, Authorization, Privacy, Trustworthiness, Integrity S01,S05, S06, S09, S15, S23, S24, S25, S29, S34, S37, S38, S40, S44, S47, S52, S54, S55, S57, S59, S61, S64, S65, S66, S69, S71, S72, S73, S77, S86 Encryption and strong authentication mechanisms
+Security Confidentiality, Authenticity, Access control, Authorization, Privacy, Trustworthiness, Integrity (Almeida et al., 2014; Gattobigio et al., 2022; Deng et al., 2022; Verdugo et al., 2017; Pham and Pham, 2015; Wang et al., 2020; Pahl, 2015), (Gao et al., 2017; Gupta et al., 2021; Rios et al., 2017; Dimitrijevic et al., 2019; Pandey et al., 2022; Bélair et al., 2019; Wurster et al., 2018), (Libardi et al., 2014; Casola et al., 2017; Bucur et al., 2018; Bolivar et al., 2018; Tricomi et al., 2017; Rodigari et al., 2021; Di Modica et al., 2019), (Harwalkar et al., 2019a; Zeck and Bouroudjian, 2017; Aral et al., 2020; Kiss, 2018; Graupner et al., 2015a; Alaluna et al., 2017; Guechi and Maamri, [n.
+d.]; Ranjbar et al., 2017; Gundu et al., 2020) Encryption and strong authentication mechanisms
 Implementation of firewalls, VPNs, or SDNs for network security
@@
 addressing CSP incompatibilities
-Compatibility Interoperability, Co-existence S15, S16, S21, S24, S27, S42,S44, S45, S49, S63, S64, S71, S77, S80, S82 Implementing lightweight communication protocols and modes
+Compatibility Interoperability, Co-existence (Pham and Pham, 2015; Ouyang et al., 2022; Cui et al., 2016; Pahl, 2015; Alonso et al., 2019; Baresi et al., 2016; Wurster et al., 2018), (Trihinas et al., 2015; Harwalkar et al., 2019b; Pellegrini et al., 2017; Harwalkar et al., 2019a; Graupner et al., 2015a; Ranjbar et al., 2017; Quint and Kratzke, 2018; Ferry et al., 2013) Implementing lightweight communication protocols and modes
 Standardization of interfaces for seamless integration
@@
 Interoperability standardization and federation between clouds
-Scalability Capacity, Extensibility, Elasticity, Throughput, Responsiveness S01, S08, S09, S14, S15, S16, S17, S18,S19, S21, S22, S24, S25, S27, S29, S31, S35, S41, S42, S44, S45, S54, S57, S62, S66, S69, S72, S74, S76, S77, S82, S84
+Scalability Capacity, Extensibility, Elasticity, Throughput, Responsiveness (Almeida et al., 2014; Pham et al., 2015; Verdugo et al., 2017; Chen et al., 2022; Pham and Pham, 2015; Ouyang et al., 2022; Slominski et al., 2015), (Pai and Kumar, 2021; Alhamazani et al., 2014; Cui et al., 2016; Wong et al., 2019; Pahl, 2015; Gao et al., 2017; Alonso et al., 2019), (Gupta et al., 2021; Mittal and Risco-Martín, 2017; Carpio et al., 2020; Shi et al., 2020; Baresi et al., 2016; Wurster et al., 2018; Trihinas et al., 2015), (Bucur et al., 2018; Tricomi et al., 2017; Rahman and Lama, 2019; Aral et al., 2020; Kiss, 2018; Alaluna et al., 2017; Mfula and Nurminen, 2018), (Pop et al., 2016; Ranjbar et al., 2017; Ferry et al., 2013; Kritikos et al., 2019)
 Combining Containerization and Microservices for enhanced scalability
@@
 Leveraging Orchestration, Federated Network,and Geographic Placement
-Reliability Maturity, Availability, Fault tolerance, Recoverability S01, S02, S04, S10, S13, S17,S19, S20, S22, S25, S27, S32, S33, S35, S37, S44, S47, S51,S53, S54, S55, S56, S57, S60, S64, S65, S66, S67, S71, S73, S74, S75, S82 Deploying Redundant Engines for Fault-Tolerance
+Reliability Maturity, Availability, Fault tolerance, Recoverability (Almeida et al., 2014; Jha et al., 2019; Ramos et al., 2021; Sabbioni et al., 2020; Naik, 2016a; Slominski et al., 2015; Alhamazani et al., 2014), (Ferry et al., 2014; Wong et al., 2019; Gao et al., 2017; Alonso et al., 2019; Al-shammari and Alwan, 2018; Sami and Mourad, 2020; Carpio et al., 2020), (Dimitrijevic et al., 2019; Wurster et al., 2018; Libardi et al., 2014; Buyya and Barreto, 2015; Acquaviva et al., 2017; Bucur et al., 2018; Bolivar et al., 2018), (del Castillo et al., 2013; Tricomi et al., 2017; Janarthanan et al., 2018; Harwalkar et al., 2019a; Zeck and Bouroudjian, 2017; Aral et al., 2020; Tang, 2021), (Graupner et al., 2015a; Guechi and Maamri, [n.
+d.]; Mfula and Nurminen, 2018; Shen et al., 2017; Ferry et al., 2013) Deploying Redundant Engines for Fault-Tolerance
 Distributing Resources and Replicating Applications for Improved
@@
 Building Fault-Tolerant Systems
-Portability Installability, Replaceability, Adaptability S18, S24, S25, S31, S62, S77 Virtual Machine-Based Packaging
+Portability Installability, Replaceability, Adaptability (Pai and Kumar, 2021; Pahl, 2015; Gao et al., 2017; Mittal and Risco-Martín, 2017; Rahman and Lama, 2019; Ranjbar et al., 2017) Virtual Machine-Based Packaging
 Docker Image-Based Packaging
@@
 Adaptive Rule-Based System Modification
-Flexibility Functional completeness, Functional correctness, Functional appropriateness S16, S21, S27, S31, S41, S61, S60, S70, S76, S84 Service-oriented Architecture
+Flexibility Functional completeness, Functional correctness, Functional appropriateness (Ouyang et al., 2022; Cui et al., 2016; Alonso et al., 2019; Mittal and Risco-Martín, 2017; Shi et al., 2020; Di Modica et al., 2019; Janarthanan et al., 2018), (Lazuka et al., 2022; Pop et al., 2016; Kritikos et al., 2019) Service-oriented Architecture
 Dynamic allocation of resources across different cloud platforms based on demand
@@
 Cross-cloud service orchestration,
-Maintainability Modularity, Reusability, Analyzability, Modfiabality, Testability S08, S17, S25, S27, S62, S74, S82 Microservices, easier maintenance
+Maintainability Modularity, Reusability, Analyzability, Modfiabality, Testability (Pham et al., 2015; Slominski et al., 2015; Gao et al., 2017; Alonso et al., 2019; Rahman and Lama, 2019; Mfula and Nurminen, 2018; Ferry et al., 2013) Microservices, easier maintenance
 Infrastructure Provisioning as Code (IaC)
@@
 Continuous Integration and Deployment Automation
-Usability Appropriateness recognizability, Learnability, Operability , User interface aesthetics, Accessibility S09, S17, S27, S27, S37, S74 Feedback-loop controller for multi-cloud Infrastructure
+Usability Appropriateness recognizability, Learnability, Operability , User interface aesthetics, Accessibility (Verdugo et al., 2017; Slominski et al., 2015; Alonso et al., 2019, 2019; Dimitrijevic et al., 2019; Mfula and Nurminen, 2018) Feedback-loop controller for multi-cloud Infrastructure
 Container Orchestration (e.g., Kubernetes)
@@
 3.5. Security Challenges and Solution Framework (RQ4 and RQ5)
-In an era of rapid technological advancement, ensuring data security is of paramount importance.
-In this section, we propose a framework designed to address the security challenges in containerized and multi-cloud environment.
-This proposed framework, composed of seven categories of security challenges and solutions illustrated in Table 11, is derived through a comprehensive review of existing literature.
-It covers a broad spectrum of categories, each capturing specific security challenges and their corresponding solutions.
-Each category is briefly described below:
-Data Security in container-based applications category covers a wide range of challenges and solutions, including data protection, database security, data compliance, and secure data transfer.
-Our identified solutions revolve around the use of encryption and secure container orchestration, complying with regulatory requirements such as GDPR, HIPAA, and secure data management strategies.
-Furthermore, identified solutions emphasizes the role of security measures such as access control and secure transmission protocols in ensuring data privacy, service continuity, fault tolerance, and the secure transfer of data between different cloud environments.
-Access and Communication Control category emphasizes maintaining secure multi-cloud connectivity and managing user access control.
-The identified solutions involve implementing encryption and authentication mechanisms, using secure communication protocols, and executing operations in a provider-agnostic manner to maintain security.
-Container Security category addresses issues like co-resident security, container isolation, and providing host security.
-Implementing secure placement strategies for containers across multi-cloud platforms, utilizing container isolation techniques, and enhancing manual defenses to mitigate risks of compromised containers are some of the identified solutions in this category.
-Infrastructure and Deployment Management Security category emphasizes the significance of infrastructure and deployment management security.
-It encompasses aspects such as secure data deployment, infrastructure orchestration, storage security, and migration security.
-Essential solutions encompass the application of encryption techniques, implementation of access control mechanisms, adoption of standardized and secure infrastructure orchestration practices, and ensuring the secure migration of containerized applications across multi-cloud environment.
-Security Monitoring and Breach Prevention category addresses a range of challenges and solutions related to enhancing security.
-This includes the implementation of automated security assessments, the use of vulnerability scanners, and the establishment of automated bootstrapping mechanisms for Virtual Machines (VMs) that incorporate security configurations and policies.
-Additionally, it involves deploying Security Information and Event Management (SIEM) systems, as well as Intrusion Detection System/Intrusion Prevention System (IDS/IPS) tools.
-The category also covers the application of log analysis solutions to proactively detect and prevent security breaches.
-Trust and Compliance Management category identified solutions encompass on implementation of robust security controls, adherence to industry best practices and standards, and establishment of trust-based coalitions to foster secure community formation.
-Placement Strategy: It recommends devising a security model for services hosted on volunteering fogs within multi-cloud environment, including formulation of secure communication protocols and implementation of access control mechanisms.
+This section presents a framework for addressing security challenges in containerized applications in multi-cloud environment.
+Based on a detailed review, it identifies seven key categories of challenges and solutions, outlined in Table LABEL:tab:Security-challenges-solutions.
+Each category is briefly reported below:
+Data Security in Container-based Applications category covers multiple challenges related to data security in container-based applications, including data protection, database security, data compliance, and secure data transfer in a multi-cloud environment.
+Out of the 121 reviewed studies, 11 specifically highlighted these issues and provided corresponding solutions (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)).
+The identified solutions focus on implementing robust encryption techniques, secure container orchestration, and compliance with regulatory standards such as GDPR and HIPAA.
+The solutions also include using Role-Based and Attribute-Based Access Controls (RBAC and ABAC) to manage secure access effectively.
+Standardization efforts like Secure CSP Interoperability and OpenStack API integration aim to mitigate risks of unauthorized access and vendor lock-in.
+Furthermore, comprehensive frameworks like DRA emphasize secure data management, ensuring privacy, fault tolerance, and secure data transfers across different cloud platforms.
+Table 11. Identified security challenges and solutions for containerized applications in multi-cloud environments
+ID Challenge Solution
+Category 1: Data Security
+(Slominski et al., 2015) Data Protection Multi-Cloud Data Protection through Secure Containers
+(Gupta et al., 2021) Database Security Enhanced multi-cloud Database Security via Containerization
+(Dimitrijevic et al., 2019) Data Compliance Multi-cloud Legal Compliance and Data Protection
+(Pandey et al., 2022) Data Security Secure CSP Interoperability and Selection in multi-cloud
+(Acquaviva et al., 2017) Data Ownership and Privacy Multi-cloud Security via NoMISHAP Service Abstraction
+(Sitaram et al., 2018) Data Transfer Security Secure multi-cloud Integration via OpenStack Standardization
+(Bou Ghantous and Gill, 2021) Cloud Data Management Multi-cloud Data Security via DRA Framework
+(Ahmad and Aujla, 2023) Data leaks and access risks Cryptography, smart contracts, encryption
+(Sofia et al., 2024) Decentralized data security challenge Secure data handling with decentralized AI
+(Neeraj et al., 2023) SLA data integrity challenge Blockchain for SLA security and integrity
+(Santos et al., 2023) Multi-cloud data privacy issue Stringent security controls for data protection
+Category 2: Access and Communication Control
+(Gattobigio et al., 2022) Connectivity and Access Control Multi-cloud Containerized Application Security
+(Dreibholz et al., 2019) Secure Communication Middleware Secure Deployment for multi-cloud
+(Buyya and Barreto, 2015) Provider-Agnostic Operations Aneka-Based Security and Performance for Containerized multi-cloud Applications
+(Hossain et al., 2023) Microservices privacy and access control Encryption, authentication, service mesh for security
+(Fugkeaw, 2023) Decentralized IAM security issue Blockchain for IAM security and data integrity
+(Roy and Kalafatis, 2023) Secure communication across edge-cloud environments WireGuard for secure edge-cloud communication
+(Benmerar et al., 2023) Multi-domain edge communication security Secure communication and AI anomaly detection
+(Huang and Wang, 2023) Secure robot-cloud communication Secure communication and encryption for data integrity
+Category 3: Container Security
+(Deng et al., 2022) Co-resident Security Secure Placement Strategy via Deep RL
+(Gao et al., 2017) Container Isolation Power and Leakage Management for multi-cloud Security
+(Bélair et al., 2019) Container and Host Security Host-Container Cooperation for multi-cloud Defense
+(Wang et al., 2023) Container security and malware detection DockerWatch for malicious activity detection
+(Tiwari and Sharma, 2023) Multi-tenant serverless security challenge Security best practices for serverless deployments
+Category 4: Infrastructure and Deployment Management Security
+(Ferry et al., 2014) Data Deployment Secure Private Data Deployment Strategy
+(Wurster et al., 2018) Infrastructure Orchestration Standard Interface Implementation for Serverless Security
+(Libardi et al., 2014) Storage Security Multi-cloud Storage Security with MSSF
+(Quint and Kratzke, 2018) Migration Security Infrastructure-Aware and Agnostic Secure Migration
+(Solayman and Qasha, 2023) Distributed infrastructure security challenge Orchestration security for IoT data protection
+(Kataru et al., 2023) Securing AWS app deployments AWS best practices for CI/CD security
+Category 5: Security Monitoring and Breach Prevention
+(Sheridan et al., 2017) System Vulnerabilities Automated Security Evaluation and Bootstrapping
+(Trihinas et al., 2015) Breach Detection and Prevention Metric Filtering and Fault Recovery for multi-cloud Security
+(Lee and Nam, 2023) Network policy misconfiguration risks Automated network policy discovery and verification
+(Nguyen and Chien, 2023) Stream processing security risks RBAM framework for secure data handling
+Category 6: Trust and Compliance Management
+(Casola et al., 2017) Application Security and Compliance Risk Analysis and Security SLAs with MUSA Framework
+(Harwalkar et al., 2019b) Data Security and Standardization OpenStack API-based multi-cloud Security and Interoperability
+(Tricomi et al., 2017) Network Security Inter-cloud Communication Protection with SFC
+(Wahab et al., 2016) Trust and Community Formation Trust-based Hedonic Coalitions for multi-cloud Security
+(Kritikos et al., 2019) Orchestration and Adaptation Multi-cloud Service Orchestration and Cross-Level Adaptation
+(Petrovski and Gusev, 2024) Multicloud compliance and security issue AWS security measures for multicloud environments
+(Alonso et al., 2023a) Cross-cloud security and compliance Security frameworks for multi-cloud environment
+Category 7: Placement Strategy
+(Sami and Mourad, 2020) Placement Strategy Security Model for Volunteering Fog Services
+(Pandi et al., 2023) Multi-cloud security concerns Security groups, VPCs, and infrastructure security
+(Ramesh et al., 2023) Multicloud environment security issue Multicloud security for risk diversification
+(Sofia et al., 2023) Edge-cloud security concerns Enhanced security protocols for decentralized environments
+(Harzenetter et al., 2023) Cross-technology update security risks Integrated security in workflows for management
+(Nagarajan et al., 2023) Resource allocation security risks Security measures in resource allocation framework
+Access and Communication Control category reports on multiple challenges related to maintaining secure multi-cloud connectivity and managing user access control.
+Out of the 121 reviewed studies, 9 specifically identified issues and proposed solutions addressing secure communication, provider-agnostic operations, and decentralized identity management (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)).
+Advanced encryption and authentication mechanisms, e.g., TLS, JWT, and OAuth 2.0, were identified as solutions to secure communications across cloud and edge environments.
+Secure deployment of applications using middleware solutions was also discussed, which make use of SSL/TLS and access controls to enhance security attributes.
+Provider-agnostic middleware, e.g., the one based on Aneka, enables runtime provider selection and thus facilitates seamless operations across multi-cloud environment.
+Solutions for decentralized IAM security relate to blockchain utilization for improving data integrity and confidentiality.
+Secure communications in multi-domain edge environments are achieved through frameworks like WireGuard, backed by strong service mesh systems that maintain connectivity and protect against unauthorized access.
+Security monitoring for anomaly detection and policy enforcement was also noted as important for sustaining secure operations.
+Container Security category reports on challenges such as co-resident security, container isolation, host security, and multi-tenant serverless security in multi-cloud environment.
+Among the 121 reviewed studies, 5 identified these issues and proposed solutions to address them (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)).
+The solutions identified include the use of secure placement strategies via deep reinforcement learning to manage secure container placement in multi-cloud platforms.
+Security measures for container isolation include the technique of Power and Leakage Management, followed by Linux kernel-level isolation using SELinux.
+In order to enhance security on both container and host sides, solutions such as Host-Container Cooperation and Docker-based detection tools, like DockerWatch, enhance the defenses and malicious activity detection.
+Finally, with a view to further securing multi-tenant serverless environments, the implementation of best practices regarding security in the CLI tool used for managing serverless deployments was underlined.
+Infrastructure and Deployment Management Security category reports on challenges related to securing data deployment, infrastructure orchestration, storage security, and migration security in multi-cloud environment.
+Out of the reviewed 121 studies, 6 specifically addressed these types of challenges and proposed solutions (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)).
+The identified solutions include the use of a Secure Private Data Deployment Strategy with encryption techniques like AES-256 and secure protocols such as SFTP and HTTPS to safeguard data.
+In the case of infrastructure orchestration, the focus is on the implementation of standard interfaces, such as OpenStack APIs, to ensure security and interoperability during deployments.
+For storage security, the solutions introduced the Multi-Cloud Storage Framework (MSSF), providing secure storage management through the implementation of RBAC and data encryption, along with periodic audits.
+The solutions also introduced measures for secure migration through both Infrastructure-Aware and Agnostic Secure Migration strategies, emphasizing image signing and data encryption under the least-privilege principle.
+Additionally, the reviewed studies proposed orchestration security for IoT distributed data and operations, as well as AWS best practices to help protect CI/CD pipelines running in multi-cloud environment.
+Security Monitoring and Breach Prevention category reports on challenges related to system vulnerabilities, breach detection and prevention, and network policy misconfiguration risks in multi-cloud environment.
+Out of the 121 reviewed studies, 4 specifically addressed these issues and proposed corresponding solutions (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)).
+The identified solutions suggest implementing automated security evaluations and bootstrapping mechanisms for Virtual Machines with security configurations using tools like Terraform, Chef, Ansible, or Puppet.
+The studies also highlighted several other tools for vulnerability assessment, such as OpenVAS and Nessus, which can be used to address delays in patching and reduce system vulnerabilities.
+In terms of breach detection and prevention, the reviewed studies emphasized the need to enhance security through metric filtering, fault recovery, and the automated deployment of SIEM, IDS, and IPS tools.
+Log analysis solutions were considered necessary for proactive breach management.
+As a result, automated network policy discovery and verification mechanisms, such as KUNERVA, were proposed, incorporating rigorous verification to prevent most potential breaches in network policy management.
+Trust and Compliance Management category reports on solutions related to ensuring security and compliance in multi-cloud environment.
+Among the 121 reviewed studies, 7 specifically discussed these issues and proposed corresponding solutions (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)).
+The identified solutions emphasize the need for security controls and standardized practices, such as risk analysis and secure coding using MUSA.
+The use of standardized API-based cloud interfaces, like the OpenStack API, was also emphasized to ensure both data security and interoperability between different cloud platforms.
+Establishing trust-based coalitions for community security was another important approach, supported by IAM systems and secure communication frameworks like SFC to enable secure interactions and collaborations within multi-cloud settings.
+Placement Strategy category highlights the need for secure service deployment on volunteering fog nodes in multi-cloud environment.
+Of the 121 reviewed studies, 6 proposed solutions for these challenges (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)).
+The solutions involve developing a security model that integrates SSL/TLS protocols for communication and certificate-based access controls to protect deployed services on fog nodes.
+The strategies include implementing a combination of security groups, VPCs, and infrastructure-level measures to distribute risks and prevent unauthorized access.
+The reviewed studies recommend adaptive methods for handling security in dynamic and decentralized edge-cloud environments by proposing cross-technology security updates and resource allocation frameworks to avoid conflicts and vulnerabilities.
 […]
-Table 11. Security challenge and solution framework for containerized applications in multi-cloud environment
-ID Challenge Challenge Description Solution Solution Description
-Category 1: Data Security
-S17 Data Protection Protecting data against threats in containerized multi-cloud setups.
-Multi-Cloud Data Protection through Secure Containers Deploys sensitive data in encrypted, isolated containers, using secure orchestration tools like Docker and Kubernetes.
-S29 Database Security Securing databases within containerized multi-cloud environments.
-Enhanced multi-cloud Database Security via Containerization Secures databases by implementing encryption (e.g., Transparent Data Encryption) and robust access controls (e.g., RBAC, ABAC).
-S37 Data Compliance Ensuring data processing and storage compliance across multi-cloud platforms.
-Multi-cloud Legal Compliance and Data Protection Maintains privacy and service quality by adhering to regulations (e.g., GDPR, HIPAA) and employing data protection strategies like encryption and audits.
-S38 Data Security Adhering to data security standards in multi-cloud containerized systems.
-Secure CSP Interoperability and Selection in multi-cloud Tackles CSP compatibility issues and enhances security by implementing data handling practices and selecting compliant CSPs.
-S53 Data Ownership and Privacy Tackles data ownership, privacy, and availability issues in multi-cloud setups.
-Multi-cloud Security via NoMISHAP Service Abstraction Uses NoMISHAP for service abstraction, enhancing privacy with encryption and pseudonymization, and ensuring availability with fault tolerance strategies like backups and failovers.
-S58 Data Transfer Security Secures data during transfers in multi-cloud environments.
-Secure multi-cloud Integration via OpenStack Standardization Standardizes OpenStack APIs to prevent vendor lock-in, ensuring safe data transfers with encryption and secure protocols.
-S85 Cloud Data Management Manages cloud data security, privacy, and compliance.
-Multi-cloud Data Security via DRA Framework Implements the DRA framework for comprehensive data management, ensuring encryption, access control, and compliance with standards.
-Category 2: Access and Communication Control
-S05 Connectivity and Access Control Ensures secure multi-cloud connectivity and robust user access control.
-Multi-cloud Containerized Application Security Implements TLS, JWT, OAuth 2.0 for secure authentication and authorization in multi-cloud setups.
-S43 Secure Communication Secures communications across multi-cloud and MEC platforms.
-Middleware Secure Deployment for multi-cloud Uses middleware for secure cloud app deployment, with SSL/TLS and access control for secure communication.
-S51 Provider-Agnostic Operations Facilitates secure, provider-agnostic operations in multi-cloud environments.
-Aneka-Based Security and Performance for Containerized multi-cloud Applications Employs Aneka for security and performance, offering transparent access and dynamic provider selection, while ensuring security via secure channels and monitoring.
-Category 3: Container Security
-S06 Co-resident Security Addresses security risks from container virtualization in multi-cloud environments.
-Secure Placement Strategy via Deep RL Uses deep reinforcement learning, like Kubernetes Cluster Autoscaler with RL, for secure container placement across multi-cloud platforms.
-S25 Container Isolation Aims to prevent information leaks and power attacks through effective container isolation.
-Power and Leakage Management for multi-cloud Security Manages security via container isolation (e.g., Docker isolation) and Linux kernel-level resource isolation (e.g., SELinux).
-S40 Container and Host Security Enhances security against compromised containers and strengthens defenses.
-Host-Container Cooperation for multi-cloud Defense Develops container-host defense strategies, utilizing tools like Falco, Sysdig, and practices for host hardening.
-Category 4: Infrastructure and Deployment Management Security
-S20 Data Deployment Secures private data deployment in multi-cloud environments.
-Secure Private Data Deployment Strategy Utilizes encryption, RBAC, and secure protocols (e.g., AES-256, SFTP, HTTPS) for safeguarding data deployment.
-S44 Infrastructure Orchestration Orchestrates secure infrastructure for containerized applications.
-Standard Interface Implementation for Serverless Security Uses standard interfaces (e.g., OpenStack APIs) for secure, interoperable serverless application deployment across clouds.
-S47 Storage Security Enhances security for multi-cloud storage in containerized setups.
-Multi-cloud Storage Security with MSSF Employs the MSSF for secure storage management, implementing RBAC, data encryption, and regular audits.
-S80 Migration Security Ensures the secure migration of applications in multi-cloud environments.
-Infrastructure-Aware and Agnostic Secure Migration Leverages strategies and tools for safe migration, including image signing and data encryption, adhering to best practices like least privilege.
-Category 5: Security Monitoring and Breach Prevention
-S30 System Vulnerabilities Addresses vulnerabilities and patching delays in multi-cloud environments.
-Automated Security Evaluation and Bootstrapping Utilizes tools like OpenVAS and Nessus for security assessments, and automates VM setup with Terraform, Chef, Ansible, or Puppet.
-S45 Breach Detection and Prevention Boosts breach detection and prevention in multi-cloud infrastructures.
-Metric Filtering and Fault Recovery for multi-cloud Security Enhances security with metric filtering, fault recovery, SIEM, IDS/IPS, and log analysis for proactive breach management.
-Category 6: Trust and Compliance Management
-S52 Application Security and Compliance Ensures security and compliance for applications across multi-clouds.
-Risk Analysis and Security SLAs with MUSA Framework Uses MUSA for risk analysis and security SLA management, incorporating secure coding, vulnerability management, and compliance audits.
-S49 Data Security and Standardization Standardizes data security and cloud interfaces in multi-cloud environments.
-OpenStack API-based multi-cloud Security and Interoperability Utilizes OpenStack API for secure, standardized cloud interoperability, applying encryption, access control, and CIS benchmarks.
-S57 Network Security Secures networks and trust in multi-cloud environments.
-Inter-cloud Communication Protection with SFC Employs Service Function Chaining (SFC) for secure inter-cloud communications, using encryption, integrity checks, and protocols like IPSec, WireGuard.
-S83 Trust and Community Formation Builds secure communities and trust in multi-cloud settings.
-Trust-based Hedonic Coalitions for multi-cloud Security Forms trust-based coalitions for community security, leveraging IAM and trust frameworks for collaboration in multi-clouds.
-S84 Orchestration and Adaptation Manages orchestration and adaptation for security and performance.
-Multi-cloud Service Orchestration and Cross-Level Adaptation Implements multi-cloud orchestration with tools like Kubernetes, Istio, and Prometheus for secure, adaptive application management.
-Category 7: Placement Strategy
-S33 Placement Strategy Secures service deployment on volunteer fogs in multi-cloud setups.
-Security Model for Volunteering Fog Services Creates a security model for fog-deployed services, incorporating SSL/TLS for communication and certificate or token-based access control.
 3.6. Automation Challenges and Solution Framework (RQ4 and RQ5)
-Table 12 presents an overview of the automation challenges and solution framework for containerized applications in multi-cloud environment.
+Table LABEL:tab:Automation-solutions presents an overview of the automation challenges and solution framework for containerized applications in multi-cloud environment.
 The figure consists of eight categories, each focusing on different aspects of automation challenges and their solutions in a multi-cloud environment.
-Multi-Cloud Automation: This category reports challenges and solutions related to automating processes for or combining and administering various cloud computing environments.
-It addresses challenges such as multi-cloud integration and provisioning, and provides solutions such as proxies for environment interface and model-driven engineering techniques.
-Automation in Deployment and Scaling: This category reports challenges and solutions related to automating the deployment and scaling processes of containerized applications across multi-cloud platforms.
-It tackles challenges like container placement automation and application runtime switch automation.
-Solutions include intelligent container placement algorithms and service drivers for automated runtime switch.
-Resource Management Automation: This category reports challenges and solutions related to automating the management of resources in multi-cloud environment.
-It addresses challenges such as resource overbooking management and containerized applications management.
-Solutions include ML-based resource overbooking detection and custom Kubernetes schedulers.
-Data and Application Migration Automation: This category reports challenges and solutions related to automating the migration of data and applications in multi-cloud environment.
-It covers challenges such as data integrity during migration and legacy web applications migration.
-Solutions encompass the implementation of self-sufficient management cycles for streamlined deployment, alongside the utilization of adaptable architectural patterns designed for the modernization of legacy applications within a containerized multi-cloud environment.
-Testing and Benchmarking Automation: This category reports challenges and solutions related to the automation of testing and benchmarking processes for containerized applications in multi-cloud environment.
-It addresses challenges like user-oriented testing automation and benchmarking automation.
-Solutions include container-based lightweight virtualization for testing and smart Docker benchmarking orchestrators.
-Standardization and Interoperability Challenges: This category reports challenges and solutions related to standardization and interoperability in the automation of containerized applications across multiple cloud environments.
-It covers challenges like standardization in automation processes and cloud provider selection.
-Solutions include service mesh networks and AI-driven cloud provider selection tools.
-Application and Service Management: This category reports challenges and solutions related to the automation of application and service management processes in containerized multi-cloud environment.
-It addresses challenges such as 5G ecosystem component selection and end-to-end tail latency management.
-Solutions include open-source software components for customizable orchestration and modeling approaches for performance evaluation.
-Runtime and Service Discovery: This category deals with challenges and solutions related to runtime management and service discovery in multi-cloud environment, which involves managing the complexity of network control and administration within containerized applications across multiple cloud environments.
-Table 12. Automation challenge and solution framework for containerized applications in multi-cloud environment
-ID Challenge Challenge Description Solution Solution Description
+Multi-Cloud Automation category reports on challenges and solutions related to automating processes for combining and administering various cloud computing environments.
+Out of the 121 reviewed studies, 13 identified these challenges and provided associated solutions (see Table LABEL:tab:Automation-solutions and Automation Challenges-Solutions sheet in (Waseem et al., 2024)).
+Key challenges involve the integration, provisioning, orchestration, and adaptation of multiple clouds.
+The solutions focus on utilizing proxies for environment interfaces to ease management across multiple cloud platforms and employing model-driven engineering techniques to manage provisioning and adaptation more efficiently.
+Dynamic on-demand fog computing was identified as an essential approach for handling multi-cloud environment dynamically.
+At the level of cloud redundancy and resource management driven by security policies, smart and user-friendly automation techniques, such as MSSF, were recommended.
+For managing the complexity of multi-cloud orchestration, serverless deployment methodologies were proposed, using models like TOSCA-based orchestration to improve scalability and coordination between different cloud stacks.
+Recommendations also include deploying application-level resource managers to support multi-cloud operations and adopting secure abstraction models to create uniform interfaces for managing a wide range of cloud products.
+MUSA Deployer tools were highlighted for their key role in automating deployment and monitoring activities, ensuring overall security compliance, and simplifying the deployment pipeline across federated clouds.
+Automation in Deployment and Scaling category reports on challenges and solutions related to automating deployment and scaling processes for containerized applications across multi-cloud platforms.
+Out of the 121 reviewed studies, 12 identified these challenges and proposed corresponding solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)).
+Key challenges include intelligent container placement, application runtime switch automation, and service deployment issues.
+The proposed solutions emphasized intelligent algorithms for container placement, using machine learning for multi-cloud workload balancing.
+Runtime switch techniques are also being automated with methods such as service drivers, which are designed to be lightweight and enable seamless application migration and replication.
+DevOps methodologies were frequently mentioned in the literature in relation to managing deployment and scaling through continuous development and error reduction.
+Security during deployment is ensured through automated security measures integrated into the pipeline, aiming for a secure setup with reduced vulnerabilities.
+The automation strategies include elastic resource provisioning using predictive analytics to handle fluctuating workloads in real time.
+Optimization models were involved in solutions for virtual function placement and container deployment to enhance efficiency.
+This emphasis on automating and optimizing deployment strategies led to the proposal of comprehensive multi-cloud orchestration tools that facilitate the deployment of containers and VMs, aimed at improving agility and scalability within diverse cloud environments.
+Table 12. Identified automation challenge and solution for containerized applications in multi-cloud environment
+ID Challenge Solution
 Category 1: Multi-Cloud Automation
-S10 Multi-Cloud Integration Automating processes to manage the complexities of integrating multiple cloud environments Proxies for Environment Interface Proxies for individual environment interface, unified view of resources, shielding from underlying complexity in multi-cloud environment
-S20 Multi-Cloud Provisioning and Adaptation Automating provisioning, deployment, and dynamic adaptation of multi-cloud applications for efficiency and timely response Model-driven Engineering Techniques Model-driven engineering techniques for dynamic provisioning, deployment, and adaptation in multi-cloud environment
-S33 Fog Creation in Multi-Cloud Automating the dynamic creation of nearby fog in multi-cloud environment Dynamic On-demand Fog Computing Dynamic on-demand fog computing architecture in multi-cloud environment
-S47 Intelligent User-Friendly Automation Addressing the lack of intelligent, user-friendly automation in delivery platforms MSSF MSSF (Multi-cloud Storage Selection Framework) for automatic selection of multi-cloud redundancy storage based on user needs and security rules
-S44 Multi-Stack Orchestration Automating the specification, coordination, and orchestration of different infrastructure stacks and cloud services TOSCA-based Serverless Deployment Topology and Orchestration Specification for Cloud Applications (TOSCA) for automated deployment of multi-cloud serverless applications
-S34 Secure Multi-Cloud App Creation Providing tools and mechanisms for non-security expert developers to automate secure multi-cloud app creation MUSA Framework MUSA framework for application modeling, risk analysis, cloud service selection, composite SLA generation in multi-cloud environment
-S37 Unified Abstraction for Multi-Cloud Deployment Implementing a common abstraction for products offered in automated multi-cloud deployments Application-level Resource Managers Application-level resource managers, native multi-cloud support for enterprise applications in multi-cloud environment
-S52 Multi-Cloud Deployment Integration Addressing integration complexity in multi-cloud deployment automation MUSA Deployer Tool MUSA Deployer tool for automation of deployment process, deployment plan, and security compliance in multi-cloud environment
-S57 Deployment Across Federated Clouds Automating deployment across federated cloud providers Orchestrated Deployment Subsystem Orchestrated deployment subsystem, Beacon Broker, and federated environments for coordinated application deployment in multi-cloud environment
-S58 Hybrid/Multi-Cloud Setup Automating the setup of hybrid/multi-cloud environment to address security, vendor lock-in, and interoperability issues Interoperability Standardization Interoperability standardization and federation of clouds in multi-cloud environment
-S80 Multi-cloud Interoperability Facilitating multi-cloud interoperability, application topology definition/composition, and elastic runtime adaptation for containerized services Multi-cloud Interoperability Multi-cloud interoperability, application topology definition/composition, and elastic runtime adaption in multi-cloud environment
-S81 Multi-cloud Brokerage System Overcoming API heterogeneity and promoting semantic interoperability for a brokerage system in multi-cloud container management.
-Managing Large Number of Cloud Services Interoperability and vendor lock-in caused by the diversity of cloud solutions, complexity of multi-cloud systems administration
-S84 Multi-cloud Orchestration Implementing multi-cloud business process provisioning and cross-level orchestration for cloud services in a containerized setting Manual Migration to the Cloud Manual migration of business processes to the Cloud, limited multi-cloud and cross-cloud orchestration
-S85 DevOps for Multi-cloud Building a DevOps reference architecture with a CI broker for automation within a containerized multi-cloud environment Agile Software Development Automation Automation of agile software development, application deployment using CI-broker, automated faster application delivery, automated and integrated testing of the application in multi-cloud environment
+(Sabbioni et al., 2020) Integrating multi-cloud environment Implementing Environment Proxies
+(Ferry et al., 2014) Automating Multi-Cloud Provisioning Adopting Model-Driven Engineering
+(Sami and Mourad, 2020) Creating On-Demand Fog in Multi-Cloud Enabling On-Demand Fog Computing
+(Libardi et al., 2014) Enhancing User-Friendly Automation Using Multi-Cloud Storage Selection
+(Wurster et al., 2018) Orchestrating Multi-Stack Environments Deploying with TOSCA-based Approach
+(Rios et al., 2017) Securing Multi-Cloud Application Creation Applying MUSA Framework
+(Dimitrijevic et al., 2019) Unifying Resource Abstraction Managing with Resource Managers
+(Casola et al., 2017) Integrating Multi-Cloud Deployments Automating MUSA Deployment
+(Tricomi et al., 2017) Coordinating Federated Cloud Deployments Orchestrating Subsystem Deployment
+(Sitaram et al., 2018) Setting Up Hybrid multi-cloud environment Standardizing Interoperability
+(Quint and Kratzke, 2018) Ensuring Multi-Cloud Interoperability Enabling Multi-Cloud Support
+(Challita et al., 2017) Managing Multi-Cloud Brokerage Systems Managing Cloud Services
+(Kritikos et al., 2019) Implementing Multi-Cloud Orchestration Migrating Manually to Cloud
+(Bou Ghantous and Gill, 2021) Establishing DevOps in Multi-Cloud Automating Agile Development
+(da Silva et al., 2023) Addressing Cost and Migration Challenges Managing Complexity
+(Pandi et al., 2023) Overcoming Multi-Cloud Automation Barriers Integrating Jenkins and Kubernetes
+(Ramesh et al., 2023) Improving Multi-Cloud Cost Efficiency Optimizing Deployment Costs
+(Petrovski and Gusev, 2024) Simplifying Multi-Cloud Deployment Complexity Adopting AWS Deployment Strategies
+(Kataru et al., 2023) Maintaining Multi-Cloud Applications Automating Resource Allocation
 Category 2: Automation in Deployment and Scaling
-S27 Multi-Cloud Deployment and Monitoring Tackling challenges in automating deployment and monitoring of applications across multi-cloud platforms Multi-cloud Strategy Multi-cloud strategy for risk distribution and provider dependency reduction in multi-cloud environment
-S06 Container Placement Automation Addressing the complexity of automating container placement in multi-cloud environment Intelligent Container Placement Develop a machine learning-based optimization algorithm for intelligent container placement across multiple cloud environments, focusing on load balancing and minimizing network latency.
-S07 Application Runtime Switch Automation Enabling automated application runtime switch among various cloud services Service Driver Service driver, multi-cloud application service access model, runtime service update, development framework in multi-cloud environment
-S15 Fine-Grained Automation at Component Level Implementing autonomic fine-grained solutions at the service component level for better automation Mobility Service and Replication Mobility service, migration and replication of critical software components, autonomic fine-grained service migration/replication in multi-cloud environment
-S28 DevOps Automation Ensuring effective automation of software development and operations management DevOps Development Model DevOps development model, continuous development, automated tools and management systems, autonomic management system, performance models, systematic decision-making in multi-cloud environment
-S29 Service Deployment Automation Addressing manual deployment issues by automating service deployment Automation Scripts Automation scripts, automated deployment, and error reduction in multi-cloud environment
-S30 Deployment Time Security Automation Implementing security automation during deployment time Automated Security Measures Automated security measures, security evaluation and bootstrap, vulnerability scanning, firewall deployment, Chef integration in multi-cloud environment
-S31 Distributed Simulation Automation Automating distributed simulation environments in cloud infrastructures DevOps Methodologies DevOps methodologies, automated developer operations, rapid microservices deployment, containerization paradigm, DEVS Simulation Farm, Docker technology, SOA Microservices paradigm, DEVSML Stack in multi-cloud environment
-S32 Software Deployment Speed Improvement Addressing slow software deployment issues by implementing effective automation, enhancing Quality of Experience (QoE) Reorganized Container Images Reorganized container images, modified Docker deployment process, asynchronous image layer retrieval, profiled container execution, overlay file system in multi-cloud environment
-S36 Container Startup Automation Identifying essential files for container startup automation Profiled Container Execution Profiled container execution, essential file determination, special image layer creation in multi-cloud environment
-S38 Dynamic Scaling of Resources Automating frequent deployment and dynamic scaling of resources in multi-cloud environment for improved efficiency and performance CSP Automation Capability CSP automation capability for user agility in containerization of multi-cloud environment
-S24 Container Deployment and Scaling Implementing automated container deployment and scaling based on resource demand across multiple cloud platforms Kubernetes Kubernetes for automated container deployment and multi-cloud scaling in multi-cloud environment
-S48 Elastic Resource Provisioning Implementing elastic resource provisioning for fluctuating business workloads in a containerized multi-cloud environment through automation Elastic Resource Provisioning Design a predictive analytics system coupled with automation tools for real-time, elastic resource provisioning, addressing fluctuating business workloads in a containerized multi-cloud environment
-S50 VF Placement and Service Deployment Automating optimal Virtual Function (VF) placement and dynamic service deployment in multi-cloud environment VF Placement Optimization Optimization model for virtual function (VF) placement and dynamic VF allocation in multi-cloud environment
-S68 Container and VM Deployment Orchestrating and automating the deployment of containers and VMs in multi-cloud environment Multi-Cloud Orchestration Develop a comprehensive multi-cloud management platform for streamlining orchestration and automation of containers and virtual machine deployment across different cloud environments
+(Alonso et al., 2019) Automating Multi-Cloud Deployment Monitoring Adopting Multi-Cloud Strategy
+(Deng et al., 2022) Optimizing Container Placement Implementing Intelligent Placement
+(Zhang et al., 2021) Automating Runtime Application Switching Deploying Service Driver Model
+(Pham and Pham, 2015) Enabling Fine-Grained Component Automation Replicating and Migrating Services
+(Barna et al., 2017) Streamlining DevOps Automation Developing with DevOps Approach
+(Gupta et al., 2021) Automating Service Deployment Using Automated Scripts
+(Sheridan et al., 2017) Securing Deployment Time Securing with Automation Tools
+(Mittal and Risco-Martín, 2017) Automating Distributed Simulations Adopting DevOps Methodologies
+(Al-shammari and Alwan, 2018) Improving Software Deployment Speed Optimizing Container Images
+(Lorenzo et al., 2019) Automating Container Startup Profiling Container Execution
+(Pandey et al., 2022) Dynamically Scaling Resources Automating Scaling with CSP Capability
+(Pahl, 2015) Scaling Container Deployment Scaling with Kubernetes
+(Saif et al., 2023) Provisioning Elastic Resources Provisioning Resources Elastically
+(Bhamare et al., 2015) Automating VF Placement Optimizing VF Placement
+(Guerrero et al., 2018a) Managing Container and VM Deployment Orchestrating Multi-Cloud Deployments
 Category 3: Resource Management Automation
-S04 Resource Overbooking Management Automating the management of resource overbooking in service containers across multi-cloud platforms ML-based Resource Overbooking Detection Machine learning-based resource overbooking detection in multi-cloud environment
-S21 VCluster Management Automating cluster creation, deployment, and management for multiple Virtual Clusters (Vclusters) across diverse application frameworks ClaaS Service Model ClaaS service model, virtualized cluster environment, easy deployment and management in multi-cloud environment
-S39 Containerized Applications Management Automating deployment, scheduling, and management of containerized applications in multi-user and multi-cloud environment Custom Kubernetes Scheduler Custom Kubernetes scheduler (label-affinity-scheduler) for container orchestration in multi-cloud environment
-S41 VM Types Selection in Data Centers Addressing the large and complex search space of different VM types in various data centers through automation GA-based Algorithm GA-based algorithm, domain-tailored service clustering method for complexity reduction in multi-cloud environment
-S42 Resource Commissioning and De-Commissioning Speeding up resource commissioning and de-commissioning through automation for timely system scalability MicroCloud Architecture MicroCloud architecture, fine-grained resource allocation, TOSCA library, meta-workflow, coordinated adaptation workflow in multi-cloud environment
-S43 Cloud Federation Automation Implementing automation support for federation or inter-platform portability in multi-cloud environment Middleware Platform Middleware platform for security, cost, and performance optimization, data locality, privacy needs in multi-cloud and MEC environments
-S45 Resource Provisioning and Elasticity Detection Automating resource provisioning for cloud services, monitoring runtime configurations, and detecting elasticity actions JCatascopia JCatascopia, a fully-automated and modular cloud monitoring framework with elasticity actions, metric rule mechanism, and low-level monitoring metrics in multi-cloud environment
-S51 Resource Acquisition Automation Automating resource acquisition processes for applications to function without human intervention Aneka Platform Aneka platform for automatic acquisition and release of resources from multiple clouds in multi-cloud environment
-S74 SLA Enforcement Automation Automating the testing and corrective actions for SLA enforcement in containerized multi-cloud environment to reduce expense and time consumption SLA Compliance Testing Manual testing and corrective actions for SLA compliance are expensive and time-consuming in multi-cloud environment
+(Ramos et al., 2021) Handling Resource Overbooking Detecting Overbooking with ML
+(Cui et al., 2016) Managing Virtual Clusters Managing with ClaaS Model
+(Altran et al., 2022) Scheduling Containerized Applications Scheduling with Kubernetes
+(Shi et al., 2020) Selecting Optimal VM Types Using GA-based Algorithms
+(Baresi et al., 2016) Automating Resource Commissioning Adopting MicroCloud Architecture
+(Dreibholz et al., 2019) Supporting Cloud Federation Automation Optimizing with Middleware Platform
+(Trihinas et al., 2015) Automating Elastic Resource Detection Automating with JCatascopia
+(Buyya and Barreto, 2015) Automating Resource Acquisition Managing Resources with Aneka
+(Mfula and Nurminen, 2018) Enforcing SLA Compliance Testing for Compliance
 Category 4: Data and Application Migration Automation
-S01 Data Integrity during Migration Safeguarding data integrity during the automation of migration processes in multi-cloud environment Autonomous Management Cycle Autonomous management cycle for efficient deployment in multi-cloud environment
-S17 Legacy Web Applications Migration Facilitating the automated migration of legacy web applications to cloud services Reusable Architectural Pattern Reusable architectural pattern for containerization in multi-cloud environment
-S69 Application Porting to Cloud Automating the porting of existing applications to cloud infrastructures CloudSME Simulation Platform CloudSME Simulation Platform for fast development and deployment of complex multi-cloud applications, commercial accounting, and billing
-Category 5: Testing and Banchmakring Automation
-S09 User-Oriented Testing Automation Accelerating user-oriented software testing in diverse and heterogeneous environments through automation S Container-based Lightweight Virtualization Container-based lightweight virtualization, just-in-time deployment, streamlined multicomponent prototyping in multi-cloud environment
-S02 Benchmarking Automation Implementing an automated benchmarking process for cloud hosts from diverse providers in multi-cloud environment Smart Docker Benchmarking Orchestrator Smart Docker Benchmarking Orchestrator for containerization in multi-cloud environment
-S13 CI for Dependable Automation Leveraging Computational Intelligence (CI) to enhance dependability in automated multi-cloud containerized environments Dependability Enhancement with CI Leverage Computational Intelligence (CI) to enhance dependability in automated multi-cloud containerized environments
+(Almeida et al., 2014) Safeguarding Data Integrity in Migration Maintaining Integrity Autonomously
+(Slominski et al., 2015) Migrating Legacy Web Applications Reusing Architectural Patterns
+(Kiss, 2018) Porting Applications to Cloud Using CloudSME Platform
+Category 5: Testing and Benchmarking Automation
+(Verdugo et al., 2017) Automating User-Oriented Testing Virtualizing with Lightweight Containers
+(Jha et al., 2019) Streamlining Benchmarking Orchestrating Docker Benchmarking
+(Naik, 2016a) Enhancing Automation Dependability Enhancing Dependability with CI
 Category 6: Standardization and Interoperability Challenges
-S05 Standardization in Automation Implementing standardization in the automation processes for containerized applications across multiple cloud environments Service Mesh and Multi-cloud Federation Service mesh network and multi-cloud cluster federation for containerization in multi-cloud environment
-S70 Cloud Provider Selection Developing optimal strategies for cloud provider and node configuration selection to balance cost and runtime in a multi-cloud, containerized environment Optimal Cloud Provider Selection Create an AI-driven decision-support tool that automatically selects the optimal cloud provider and node configuration, considering both cost-effectiveness and runtime efficiency in a multi-cloud, containerized environment.
-S16 Unified Template for Microservices Addressing the lack of a unified template and specification for automating microservices deployment across multiple cloud platforms Tool Service Model Tool service model, invocation relationships, service function chains, objective function optimization, accelerated Particle Swarm Optimization, configurable objective parameters in multi-cloud environment
-S64 Cloud Interoperability Addressing the interoperability issues between different clouds through automation multi-cloud Harmony Develop an AI-driven multi-cloud management platform to automate cross-cloud interoperability, ensuring seamless data transfer and communication between disparate cloud environments
-S63 Cloud Applications Standardization Promoting the adoption of worldwide standards or interfaces for cloud applications and their management through automation Cloud Service Interoperability Cloud service providers and middleware for interoperability, standardized interfaces, and hybrid solutions in multi-cloud environment
-S18 Package Creation and Scaling Automating the creation and scaling of packages using VM and Docker based on machine-readable definition files Package Automation with VM Docker Implement a scriptable automation framework that facilitates creation, scaling, and management of packages using VM and Docker based on machine-readable definition files
+(Gattobigio et al., 2022) Standardizing Automation Processes Integrating Service Mesh
+(Lazuka et al., 2022) Optimizing Cloud Provider Selection Selecting Optimal Providers
+(Ouyang et al., 2022) Unifying Microservices Deployment Adopting Service Chain Models
+(Harwalkar et al., 2019a) Addressing Cloud Interoperability Managing with Multi-Cloud Harmony
+(Pellegrini et al., 2017) Standardizing Cloud Applications Standardizing Interfaces
+(Pai and Kumar, 2021) Automating Package Creation and Scaling Automating Package Frameworks
 Category 7: Application and Service Management
-S55 5G Ecosystem Component Selection Simplifying the complexity of 5G ecosystem component selection through automation Open-source Software Components Open-source software components for customizable container and virtual machine orchestration in contemporary deployments
-S62 End-to-End Tail Latency Management Automating end-to-end tail latency management in a microservice architecture Performance Modeling Approach Modeling approach for performance evaluation of container-level and VM-level data in multi-cloud environment
-S79 Distributed Application Complexity Addressing the increasing complexity of managing distributed applications in containerized multi-cloud environment Complexity of Distributed Applications Increasing complexity of distributed applications and operating multi-cloud management topologies involving CMPs
+(Bolivar et al., 2018) Selecting 5G Ecosystem Components Utilizing Open-Source Components
+(Rahman and Lama, 2019) Managing End-to-End Tail Latency Modeling Performance Efficiently
+(Serhiienko and Spillner, 2018) Handling Distributed Application Complexity Managing Distributed Applications
 Category 8: Runtime and Service Discovery
-S72 Network Complexity Managing the complexity of network control and administration within containerized applications across multiple cloud environments.
-Network Management Complexity Complexity of network management and control in multi-cloud environment
-S83 Social Network Service Discovery Improving service discovery algorithms for tagging in online social networks within containerized multi-cloud environment Automated Discovery and Selection Automating discovery and selection of relevant and appropriate services for composition and QoS management in multi-cloud environment
-S08 Complex Application Administration Handling growing application complexity due to new technologies and needs through automation Autonomic Computing Systems Autonomic Computing Systems with knowledge transfer for containerization in multi-cloud environment
+(Alaluna et al., 2017) Addressing Network Complexity Managing Network Complexity
+(Wahab et al., 2016) Improving Social Network Service Discovery Automating Service Discovery
+(Pham et al., 2015) Administering Complex Applications Building Autonomic Systems
+Resource Management Automation category reports on challenges and solutions related to automating the management of resources in multi-cloud environment.
+Out of the 121 studies reviewed, 9 specifically addressed these challenges and provided appropriate solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)).
+These challenges include resource overbooking management, resource management for containerized applications, and resource commissioning and decommissioning.
+The proposed solutions include machine learning-based resource overbooking detection for efficiently managing service containers across multiple cloud platforms.
+To manage containerized applications, custom Kubernetes schedulers, such as label-affinity schedulers, were proposed to optimize orchestration and allocation.
+Additionally, a GA-based algorithm was suggested to reduce the search space of VM types in data centers, simplifying the resource allocation problem within multi-cloud environment.
+For resource commissioning and decommissioning, proposals such as the MicroCloud architecture enabled fine-grained resource allocation and coordinated adaptation workflows.
+The solutions also addressed cloud federation and inter-platform portability challenges, utilizing middleware platforms for security enhancement, cost management, and performance optimization.
+The JCatascopia automated modular monitoring framework was introduced to monitor runtime configurations and detect elasticity actions.
+Finally, SLA compliance testing automation was proposed to manage resource allocation processes and enforce corrective actions, ensuring that service level agreements are met across cloud environments.
+Data and Application Migration Automation: category reports challenges and solutions related to the automation of testing and benchmarking processes for containerized applications in multi-cloud environment.
+Among the 121 studies reviewed, 3 identified these challenges along with their proposed solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)).
+The main challenges include ensuring data integrity during the migration process and enabling the smooth migration of legacy Web applications to cloud services.
+The proposed solutions involve introducing autonomous management cycles to reduce complexity and ensure effective deployment across multi-cloud environment.
+Other solutions include reusable architectural patterns to accelerate the modernization of legacy applications in containerized multi-cloud environment.
+Additionally, the CloudSME Simulation Platform was highlighted for its ability to accelerate the porting of existing applications to cloud infrastructures, with support for more complex tasks such as billing and resource management.
+Testing and Benchmarking Automation category highlights the challenges and solutions related to automating the testing and benchmarking processes for containerized applications in multi-cloud environment.
+Out of the 121 studies reviewed, 3 specifically addressed these challenges and provided relevant solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)).
+The main challenges revolve around automating user-oriented testing and benchmarking across various cloud platforms.
+The solutions range from just-in-time deployment and streamlined multicomponent prototyping to rapid testing, enabled by container-based lightweight virtualization.
+Additionally, Smart Docker Benchmarking Orchestrators were proposed for automating benchmarking in multi-cloud environment, thereby enhancing the overall efficiency of testing procedures.
+The application of Computational Intelligence (CI) was also suggested to improve dependability in automated testing for containerized multi-cloud systems.
+Standardization and Interoperability Challenges category reports on the challenges and solutions related to standardization and interoperability in the automation of containerized applications across multiple cloud environments.
+Out of the 121 reviewed studies, 6 specifically identified these challenges and proposed corresponding solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)).
+The main challenges involve standardizing automation processes, selecting cloud providers, and addressing the lack of unified templates for microservices deployment.
+The proposed solutions include service mesh networks and multi-cloud cluster federation to standardize the automation processes for containerized applications.
+AI-driven tools for optimal cloud provider selection were also introduced, which help in selecting the most suitable provider and node based on factors such as cost-effectiveness and runtime efficiency in multi-cloud environment.
+To address the lack of common templates for microservices, a Service Model approach was proposed to automate the deployment of microservices across multiple cloud platforms, thereby reducing operational complexity.
+For cloud interoperability, an AI-driven platform, such as multi-cloud Harmony, was developed to facilitate seamless data transfer and communication between disparate cloud environments.
+In terms of cloud application standardization, the solutions emphasized promoting the adoption of global standards and interfaces to improve interoperability between cloud service providers and middleware platforms.
+Finally, automation frameworks for package creation and scaling, such as those using machine-readable definition files, were proposed to efficiently manage both virtual machines and Docker containers across multi-cloud environment.
+Application and Service Management category reports challenges and solutions related to the automation of application and service management processes in containerized multi-cloud environment.Out of the 121 reviewed studies, 2 specifically addressed these challenges and proposed appropriate solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)).
+The key challenges include the selection of 5G ecosystem components and end-to-end tail latency management in microservice architectures.
+The proposed solutions involve using open-source software components to customize the orchestration of containers and virtual machines in modern multi-cloud deployments.
+Additionally, performance modeling approaches were suggested to evaluate the performance of container-level and VM-level data in multi-cloud environment, enabling more efficient latency management and optimized overall performance.
+Runtime and Service Discovery category organizes the challenges and solutions related to runtime management and services discovery in multi-cloud environment, particularly in managing the complexity of network control and administration for containerized applications.
+Out of the 121 reviewed studies, 4 specifically tackled these challenges (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)).
+The proposed solutions include automating network management to address the increasing complexity of distributed applications across multi-cloud topologies.
+Additionally, automated discovery and selection algorithms were proposed to tag and manage social network services within containerized multi-cloud systems.
+Autonomic computing systems were introduced to handle the growing complexity of applications by facilitating knowledge transfer, enabling more effective containerization management through automation.
 […]
 3.7. Deployment Challenges and Solution Framework (RQ4 and RQ5)
-Table 13 presents a comprehensive overview of the challenges and corresponding solutions faced during the deployment of containerized applications in multi-cloud environment.
+Table LABEL:tab:Deployment-solutions presents a comprehensive overview of the challenges and corresponding solutions faced during the deployment of containerized applications in multi-cloud environment.
 As organizations increasingly adopt containerization and multi-cloud strategies to enhance scalability, flexibility, and resource utilization, they encounter various obstacles that require effective solutions.
@@
 These categories encompass a wide range of issues, including deployment complexity and Orchestration, access and communication control, security and compliance management, multi-cloud deployment coordination and Integration, monitoring and scalability challenges, microservices architecture and containerization, and network connectivity and hybrid cloud integration challenges.
-Deployment Complexity and Orchestration: This category focuses on addressing challenges and providing solutions related to the complexity and orchestration of container deployment in multi-cloud environment.
-One of the significant challenges faced in this context is Performance Testing Deployment.
-It involves evaluating testing procedures and assessing performance metrics during container deployment across multiple cloud environments.
-To overcome this challenge, the corresponding solution is Standardized Benchmarking and Automation.
-This solution entails adopting Standardized Benchmarking and automatic systems to ensure consistent and reliable performance evaluation.
-Another challenge is Deployment Validation, which requires conducting thorough validation and artifact checks during container deployment in multi-cloud environment.
-To effectively address this challenge, the suggested solution is the implementation of a Testing Process Management System.
-This system enhances and oversees the validation process for deployment, making sure that all essential criteria are met before deployment proceeds.
-The challenges outlined in this category can simplify and optimise container deployment in complex multi-cloud environment, facilitating efficient and reliable operations across different cloud service providers.
-Access and Communication Control: This category emphasizes maintaining secure multi-cloud connectivity and managing user access control.
-Challenges in this category include ensuring secure access to resources and secure communication between containers and across clouds.
-The proposed solutions involve implementing encryption and authentication mechanisms, using secure communication protocols, and executing operations in a provider-agnostic manner to maintain security.
-Security and Compliance Management in Multi-cloud Deployment: This category focuses on security and compliance challenges faced during multi-cloud container deployment.
+Deployment Complexity and Orchestration category outlines various challenges and solutions related to the complexity and orchestration of container deployment in multi-cloud environment.
+Out of the 121 studies reviewed, 18 focused on these specific challenges and their respective solutions.
+Key challenges include Performance Testing Deployment, where standardized benchmarking and automation ensure consistency in performance testing.
+Deployment Validation is addressed by the Testing Process Management System, which enhances validation to ensure all criteria are met before deployment.
+Cost-Effective Deployment is achieved through Containerization and Microservice Architecture, enabling agile and efficient deployment.
+For Resource Management, a User Preference-Based Resource Brokering system optimizes resource allocation across clouds.
+Additional prominent solutions include orchestration using TOSCA, automation tools like Jenkins and Terraform, and Blockchain-based SLA Monitoring for immutable and auditable SLA tracking.
+Dynamic Adaptation is handled through Context-Aware Resource Allocation, while AWS Wavelength for Edge Cloud Integration addresses regional infrastructure consistency issues.
+Access and Communication Control category organizes challenges and solutions related to maintaining secure connectivity in multi-cloud environment and managing user access control.
+Out of the 121 reviewed studies, 8 focused on these challenges.
+These studies report solutions include ensuring secure access to resources, secure communication between containers, and maintaining secure connectivity across multiple clouds.
+The solutions involve the development of encryption and authentication mechanisms to protect data and ensure authorized access.
+Secure communication protocols were recommended for interactions within containers and across cloud environments.
+It was also suggested that security measures should be provider-agnostic, ensuring that they are not tied to any single cloud provider.
+Security and Compliance Management in Multi-cloud Deployment category focuses on security and compliance challenges and their solutions faced during multi-cloud container deployment.
 One challenge, Vendor Lock-in Prevention Deployment, aims to prevent vendor lock-in and ensure modular and loosely-coupled deployment in multi-cloud environment.
@@
 The solution is Cross-Container Isolation, which enforces cross-container isolation to enhance security.
-Multi-cloud Deployment Coordination and Integration: This category deals with challenges related to coordination and integration during multi-cloud deployment.
-One challenge, Multi-cloud PaaS Deployment, aims to overcome high entry barriers for deploying a multi-cloud PaaS infrastructure.
-The proposed solution is Lightweight Proxies for PaaS Integration, which utilizes lightweight proxies to facilitate individual PaaS services integration.
-Another challenge is IoT Application Deployment, which addresses challenges related to deploying IoT applications in multi-cloud environment.
-The solution is based on the DRA Framework and CI Broker for Multi-Cloud Deployment, which leverages the DRA framework and CI broker component for efficient multi-cloud deployment.
-Monitoring and Scalability Challenges: This category focuses on challenges related to monitoring, performance, and scalability during multi-cloud container deployment.
-For instance, Legacy Code Migration Deployment deals with handling legacy code migration, tenant engine separation, and composition modeling during container deployment in multi-cloud environment.
-The proposed solution is the Reusable Architectural Pattern with Docker and Cloudant, which advocates for adopting a reusable architectural pattern with Docker containers and Cloudant persistence layer.
-Microservices Architecture and Containerization: This category addresses challenges associated with microservices architecture and containerization in multi-cloud environment.
-For example, Autonomic System Deployment aims to implement an autonomic management system for deploying multi-layer, multi-tier web architectures in multi-cloud environment.
-The solution proposed is the Self-Tuning Performance Model and Autonomic Management, which introduces a self-tuning performance model and autonomic management system.
-Network Connectivity and Hybrid Cloud Integration: This category focuses on network connectivity and hybrid cloud integration challenges during multi-cloud deployment.
-One challenge, Inter-edge Bandwidth Deployment, aims to address challenges related to interconnecting distributed localities and provisioning inter-edge bandwidth.
-Cloud Deployment Constraints and Challenges: This category deals with various constraints and challenges encountered during cloud deployment.
-For instance, ACS Compatibility Deployment aims to ensure compatibility with existing ACSs and cloud solutions during container deployment in multi-cloud environment.
-The solution proposed is the Scalable Multi-Cloud Deployment Framework, which proposes a scalable deployment framework for multi-cloud applications.
-Infrastructure Provisioning and Container Deployment Challenges: This category addresses challenges related to infrastructure provisioning and container deployment in multi-cloud environment.
-For example, Fog Computing Deployment deals with fog computing hardware and container deployment time challenges.
-The proposed solution is Reorganized Container Images and Docker Deployment, which involves reorganizing container images and modifying the Docker deployment process.
-Application Deployment Workflow and Orchestration Challenges: This category focuses on challenges related to application deployment workflow and orchestration in multi-cloud environment.
-One challenge is API Management Deployment, which deals with effectively managing different cloud providers’ APIs/interfaces for multi-cloud container deployment.
-The solution proposed is DevOps Approach for Multi-Cloud Applications, which advocates for adopting a DevOps approach for managing the lifecycle of multi-cloud applications.
-Table 13. Deployment challenge and solution framework for containerized applications in multi-cloud environment
-ID Challenge Challenge Description Solution Solution Description
+Table 13. Identified deployment challenge and solution for containerized applications in multi-cloud environment
+ID Challenge Solution
 Category 1: Deployment Complexity and Orchestration Challenges
-S01 Performance Testing Deployment Evaluating testing and performance during container deployment in multi-cloud environment Standardized Benchmarking and Automation Adoption of standardized benchmarking applications and automatic systems
-S09 Deployment Validation Conducting rigorous validation and artifact checks during container deployment in multi-cloud environment Testing Process Management System Implementation of a testing process management system
-S02 Cost-effective Deployment Ensuring cost-effective and flexible container deployment options in multi-cloud environment Containerization and Microservice Architecture Implementation of containerization and microservice architecture
-S38 Resource Management Deployment Managing diverse resources across multiple cloud service providers during container deployment User Preference-Based Multi-Cloud Resource Brokering Adoption of multi-cloud resource brokering dependent on user preferences and application requirements
-S50 NFV Site Deployment Managing site selection, dynamic VF allocation, and service chaining during NFV deployment in multi-cloud environment Site Selection and VF Allocation Algorithms Development of algorithms for site selection and dynamic VF allocation
-S79 Cloud Management Platform Evaluation Deployment Addressing diverging functionality, lack of standard benchmark, and evaluating CMPs during multi-cloud container deployment Standardized Output Formats and Evaluation Criteria Establishment of standardized output formats and evaluation criteria
-S37 Standardized Requirement Deployment Standardizing the expression of application requirements during multi-cloud container deployments Application-Level Resource Managers for Multi-Cloud Emergence of application-level resource managers supporting multi-cloud deployment
-S42 Heterogeneous Node Deployment Managing the deployment and control of heterogeneous nodes in multi-cloud environment Technology-Independent Multi-Level Adaptation Utilization of MicroCloud and TOSCA library for technology independence and fine-grained multi-level adaptation
-S52 Deployment Plan Generation Identifying suitable cloud services and generating deployment plans for multi-cloud environment Cloud Services Selection and Deployment Planning Utilization of the MUSA Decision Support Tool and Deployer tool for cloud services selection and deployment planning
-S57 TOSCA-based Deployment Implementing TOSCA-based container deployment across multiple cloud providers Brokered Multi-Cloud Deployment with TOSCA Adoption of a brokered approach for multi-cloud deployment with TOSCA
-S58 Standardization Deployment Addressing compatibility issues and lack of standardization during multi-cloud container deployment Proxy Cloud Virtualization for OpenStack Framework mapping other clouds as sub-clouds to OpenStack with Proxy Cloud Virtualization layer
+(Almeida et al., 2014) Performance Testing Deployment Standardized Benchmarking and Automation
+(Verdugo et al., 2017) Deployment Validation Testing Process Management System
+(Jha et al., 2019) Cost-effective Deployment Containerization and Microservice Architecture
+(Pandey et al., 2022) Resource Management Deployment User Preference-Based Multi-Cloud Resource Brokering
+(Bhamare et al., 2015) NFV Site Deployment Site Selection and VF Allocation Algorithms
+(Serhiienko and Spillner, 2018) Cloud Management Platform Evaluation Standardized Output Formats and Evaluation Criteria
+(Dimitrijevic et al., 2019) Standardized Requirement Deployment Application-Level Resource Managers for Multi-Cloud
+(Baresi et al., 2016) Heterogeneous Node Deployment Technology-Independent Multi-Level Adaptation
+(Casola et al., 2017) Deployment Plan Generation Cloud Services Selection and Deployment Planning
+(Tricomi et al., 2017) TOSCA-based Deployment Brokered Multi-Cloud Deployment with TOSCA
+(Sitaram et al., 2018) Standardization Deployment Proxy Cloud Virtualization for OpenStack
+(Barletta et al., 2024) Criticality Constraints in Kubernetes Deployment Kubernetes for Critical Deployments
+(da Silva et al., 2023) Automation Tools Deployment Issue Accurate Cost and Resource Tracking
+(Pandi et al., 2023) Multi-cloud Deployment Challenge Jenkins for Multi-cloud Automation
+(Sofia et al., 2024) Decentralized App Deployment Issue CODECO for Edge-Cloud Deployment
+(Neeraj et al., 2023) SLA Monitoring Deployment Complexity Blockchain for SLA Data Integrity
+(Petrovski and Gusev, 2024) Container and Serverless Architecture Deployment Issue AWS ECS and Lambda Deployment
+(Solayman and Qasha, 2023) Heterogeneous Environment Deployment Challenge Edge-cloud Orchestration Management
+(Sofia et al., 2023) Dynamic Adaptation Deployment Issue Context-aware Resource Allocation Strategies
+(Santos et al., 2023) Regional Infrastructure Deployment Consistency Issue AWS Wavelength for Edge Cloud Integration
+(Roy and Kalafatis, 2023) Robotic App Deployment Challenge Hybrid Edge-Cloud Architecture
+(Alonso et al., 2023a) Multi-cloud Deployment Consistency Problem Multi-cloud Deployment Efficiency Models
+(Harzenetter et al., 2023) Cross-Technology Deployment Integrity Issue Workflow Automation for Cross-Technology Deployment
+(Huang and Wang, 2023) Industrial Robot Control System Deployment Problem Seamless Control between Cloud and Robots
+(Nagarajan et al., 2023) Efficient Task Deployment across Clouds Hybrid Approach for Resource Allocation
 Category 2: Security and Compliance Management in Multi-Cloud Deployment
-S20 Vendor Lock-in Prevention Deployment Developing strategies to prevent vendor lock-in and ensure modular and loosely-coupled deployment in multi-cloud environment Cloud Modelling Framework Introduction of a Cloud Modelling Framework for specifying provisioning and deployment
-S25 Security Deployment in Multi-container Environment Addressing security and privacy concerns during multi-container deployment on the same OS kernel in multi-cloud environment Cross-Container Isolation Enforcing cross-container isolation
-S34 Security Deployment Ensuring integrated security during the deployment of multi-cloud applications Security Control Specification and Deployment Framework Utilization of the MUSA framework for specifying, deploying, and assessing security controls
-S72 Virtual Network Security Deployment Providing secure virtual networks during multi-cloud container deployment Network Virtualization Platform for Multi-Cloud Adoption of Sirius, a network virtualization platform for multi-cloud environment
-S83 Malicious Service Management Deployment Managing malicious services and supporting multi-cloud communities formation during container deployment Collusion-Resilient Trust Aggregation Technique Implementation of a trust aggregation technique resilient to collusion attacks
+(Ferry et al., 2014) Vendor Lock-in Prevention Deployment Cloud Modelling Framework
+(Gao et al., 2017) Security Deployment in Multi-container Environment Cross-Container Isolation
+(Rios et al., 2017) Security Deployment Security Control Specification and Deployment Framework
+(Alaluna et al., 2017) Virtual Network Security Deployment Network Virtualization Platform for Multi-Cloud
+(Wahab et al., 2016) Malicious Service Management Deployment Collusion-Resilient Trust Aggregation Technique
+(Barati et al., 2023) Secure Personal Data Deployment Challenge Docker, Kubernetes, Blockchain for GDPR Compliance
+(Fugkeaw, 2023) IAM Deployment across Cloud Platforms Blockchain for Decentralized IAM Deployment
+(Wang et al., 2023) Non-intrusive Malware Detection Deployment DockerWatch for Non-intrusive Container Integration
 Category 3: Multi-Cloud Deployment Coordination and Integration
-S10 Multi-cloud PaaS Deployment Overcoming challenges and high entry barriers for deploying a multi-cloud PaaS infrastructure Lightweight Proxies for PaaS Integration Usage of lightweight proxies for individual PaaS services integration
-S85 IoT Application Deployment Addressing challenges related to deploying IoT applications in multi-cloud environment DRA Framework and CI Broker for Multi-Cloud Deployment Utilization of the DRA framework and CI broker component for multi-cloud deployment
+(Sabbioni et al., 2020) Multi-cloud PaaS Deployment Lightweight Proxies for PaaS Integration
+(Bou Ghantous and Gill, 2021) IoT Application Deployment DRA Framework and CI Broker for Multi-Cloud Deployment
+(Ramesh et al., 2023) Data Transfer and Service Compatibility FaaS and Storage for Multi-Cloud Deployment
+(Nguyen and Chien, 2023) Heterogeneous Cloud and Edge Integration RBAM for Cloud-Edge Integration
+(Tiwari and Sharma, 2023) Private Cloud Integration Problem CLI Tool for Serverless Setup
 Category 4: Monitoring, Performance, and Scalability Challenges
-S17 Legacy Code Migration Deployment Handling legacy code migration, tenant engine separation, and composition modeling during container deployment in multi-cloud environment Reusable Architectural Pattern with Docker and Cloudant Adoption of a reusable architectural pattern with Docker containers and Cloudant persistence layer
-S23 Tenant Performance Deployment Managing tenant service performance competition during multi-cloud container deployment Multi-Tenant and Multi-Instance Hybrid Deployment Implementation of a multi-tenant and multi-instance hybrid deployment scheme based on container technology
-S45 Monitoring Deployment Overcoming the limitations of current monitoring tools during multi-cloud container deployment Platform-Independent Monitoring System Implementation of JCatascopia, a platform-independent and interoperable monitoring system
+(Slominski et al., 2015) Legacy Code Migration Deployment Reusable Architectural Pattern with Docker and Cloudant
+(Wang et al., 2020) Tenant Performance Deployment Multi-Tenant and Multi-Instance Hybrid Deployment
+(Trihinas et al., 2015) Monitoring Deployment Platform-Independent Monitoring System
+(Struhár et al., 2024) Resource Contention and Performance Problem Linux Patch for Dynamic CPU Adaptation
+(Horchulhack et al., 2024) QoS Degradation Detection Problem Kubernetes with Performance Monitoring Tools
+(Marques et al., 2024) Monitoring Integration in Hybrid Cloud Issue Custom Scaling with Kubernetes
+(Kataru et al., 2023) AWS Deployment Performance Consistency AWS Model Comparison for Performance
+(Benmerar et al., 2023) Cloud-Edge Deployment Performance Issue AI-driven Orchestration for Dynamic Resource Management
 Category 5: Microservices Architecture and Containerization Challenges
-S28 Autonomic System Deployment Implementing an autonomic management system for deploying multi-layer, multi-tier web architectures in multi-cloud environment Self-Tuning Performance Model and Autonomic Management Development of a self-tuning performance model and autonomic management system
-S46 Microservice Deployment Managing the complexity of deploying microservice architectures in multi-cloud environment Metrics and Requirements Framework for Microservices Framework for establishing requirements and relevant metrics for the deployment of microservice-based applications
-S76 Multi-cloud Application Configuration Deployment Overcoming challenges related to configuring applications during multi-cloud deployment Runtime Environment with Object Store and Artifact Repository Deployment of a runtime environment with an object store service and artifact repository
-S80 Elastic Container Platform Dependency Deployment Addressing vendor lock-in and dependency on specific elastic container platforms during multi-cloud deployment Separation of Elastic Platform and Cloud Application Separation of elastic platform definition from cloud application definition
+(Barna et al., 2017) Autonomic System Deployment Self-Tuning Performance Model and Autonomic Management
+(Fadda et al., 2019) Microservice Deployment Metrics and Requirements Framework for Microservices
+(Pop et al., 2016) Multi-cloud Application Configuration Deployment Runtime Environment with Object Store and Artifact Repository
+(Quint and Kratzke, 2018) Elastic Container Platform Dependency Deployment Separation of Elastic Platform and Cloud Application
+(Hossain et al., 2023) Microservices Deployment in Edge Environments Docker, CI/CD, Service Mesh for Microservices
 Category 6: Network Connectivity and Hybrid Cloud Integration
-S35 Inter-edge Bandwidth Deployment Addressing challenges related to interconnecting distributed localities, provisioning inter-edge bandwidth, and centralized edge nodes during multi-cloud deployment Modular Edge Cloud Computing Architecture Designing a modular edge cloud computing system architecture using containerization techniques
-S41 Network Latency Deployment Minimizing the impact of network latency on application performance during multi-cloud deployment VM Selection for Composite Applications Selection of proper VMs from the global cloud marketplace to support multiple composite applications
-S43 Latency Reduction Deployment Tackling the issue of high end-to-end latencies during multi-cloud deployment Edge-Cellular Hybrid Infrastructure Provisioning Provisioning infrastructure at the edge of the cellular network combined with traditional cloud infrastructures
-S63 Vendor Lock-in Prevention Deployment Preventing negative effects of vendor lock-in during multi-cloud container deployment Interoperability Layer Above Cloud Infrastructure Implementation of a layer above the cloud service provider’s infrastructure for interoperability
+(Carpio et al., 2020) Inter-edge Bandwidth Deployment Modular Edge Cloud Computing Architecture
+(Shi et al., 2020) Network Latency Deployment VM Selection for Composite Applications
+(Dreibholz et al., 2019) Latency Reduction Deployment Edge-Cellular Hybrid Infrastructure Provisioning
+(Pellegrini et al., 2017) Vendor Lock-in Prevention Deployment Interoperability Layer Above Cloud Infrastructure
+(Lee and Nam, 2023) Dynamic Network Policy Deployment Problem KUNERVA for Network Policy Automation
 Category 7: Cloud Deployment Constraints and Challenges
-S08 ACS Compatibility Deployment Ensuring compatibility with existing ACSs and cloud solutions during container deployment in multi-cloud environment Scalable Multi-Cloud Deployment Framework Proposal and development of a scalable deployment framework for multi-cloud applications
-S30 Manual Intervention Risk Deployment Managing risks related to manual intervention, vulnerability, and exposed services during multi-cloud deployment Automated Security Measures at Deployment Automated implementation of required security measures at the point of deployment
-S44 Multi-provider Deployment Handling the challenges of using multiple providers and technologies during multi-cloud container deployment TOSCA-Based Deployment Modeling Approach Adoption of a TOSCA-based deployment modeling approach
-S81 Cloud Provider Transition Deployment Addressing incompatibilities of cloud solution features during multi-cloud deployment Semantic Interoperability in Multi-Clouds Adoption of the FCLOUDS framework for semantic interoperability in multi-clouds
+(Pham et al., 2015) ACS Compatibility Deployment Scalable Multi-Cloud Deployment Framework
+(Sheridan et al., 2017) Manual Intervention Risk Deployment Automated Security Measures at Deployment
+(Wurster et al., 2018) Multi-provider Deployment TOSCA-Based Deployment Modeling Approach
+(Challita et al., 2017) Cloud Provider Transition Deployment Semantic Interoperability in Multi-Clouds
+(da Silva et al., 2023) Automation Tools Deployment Issue Accurate Cost and Resource Tracking
+(Santos et al., 2023) Regional Infrastructure Deployment Consistency Issue AWS Wavelength for Edge Cloud Integration
 Category 8: Infrastructure Provisioning and Container Deployment Challenges
-S32 Fog Computing Deployment Addressing fog computing hardware and container deployment time challenges in multi-cloud environment Reorganized Container Images and Docker Deployment Reorganization of container images and modification of the Docker deployment process
-S33 Fog Device Deployment Managing fog device availability and services during multi-cloud container deployment On-Demand Fog and Microservices Deployment Enabling on-demand creation of fogs and deployment of microservices using Docker and Kubeadm
-S36 Fog Computing Container Deployment Overcoming low rates of container deployment, especially in fog computing environments FogDocker for Container Deployment Deployment of containers using FogDocker after essential file downloads
-S39 Distributed Compute Node Deployment Effectively managing containerized applications across distributed compute nodes in multi-cloud environment Label-Based Scheduling Strategy Development of a label-based scheduling strategy
-S40 Kernel Security Deployment Addressing security risks related to kernel sharing, isolation, and OS-level security during multi-cloud container deployment Enhanced Container Security Enhancing container security from the host OS level
+(Al-shammari and Alwan, 2018) Fog Computing Deployment Reorganized Container Images and Docker Deployment
+(Sami and Mourad, 2020) Fog Device Deployment On-Demand Fog and Microservices Deployment
+(Lorenzo et al., 2019) Fog Computing Container Deployment FogDocker for Container Deployment
+(Altran et al., 2022) Distributed Compute Node Deployment Label-Based Scheduling Strategy
+(Bélair et al., 2019) Kernel Security Deployment Enhanced Container Security
+(Barletta et al., 2024) Criticality Constraints in Kubernetes Deployment Kubernetes for Critical Deployments
+(Centofanti et al., 2024) Power Tool Compatibility Problem Docker, Kubernetes for Power Monitoring Integration
+(Ramesh et al., 2023) Data Transfer and Service Compatibility FaaS and Storage for Multi-Cloud Deployment
 Category 9: Application Deployment Workflow and Orchestration Challenges
-S27 API Management Deployment Effectively managing different cloud providers’ APIs/interfaces for multi-cloud container deployment DevOps Approach for Multi-Cloud Applications Adoption of a DevOps approach for managing the lifecycle of multi-cloud applications
-S29 Web Service Deployment Ensuring correct and efficient deployment of web services across various computing environments Kubernetes-Based Containerized Deployment Deployment of containerized environments using Kubernetes
-S51 Complex Application Deployment Overcoming complexities associated with application deployment in multi-cloud environment Aneka Platform for Distributed Applications Adoption of the Aneka platform for programming and deploying distributed applications
-S53 High-availability Deployment Ensuring high availability and preventing outages during deployment in multi-cloud environment Middleware Support for High Availability in Multi-Cloud PaaS Introduction of novel Middleware Support for High Availability in multi-cloud PaaS
-S69 Complex Workflow Deployment Utilizing multiple cloud providers for deploying complex application workflows CloudSME Simulation Platform Utilization of the CloudSME Simulation Platform for access to various resources and workflow support
-S84 Business Process Provisioning Deployment Orchestrating multi-cloud business process provisioning during deployment Multi-Cloud Service Orchestration Frameworks Adoption of multi-cloud service orchestration frameworks
+(Alonso et al., 2019) API Management Deployment DevOps Approach for Multi-Cloud Applications
+(Gupta et al., 2021) Web Service Deployment Kubernetes-Based Containerized Deployment
+(Buyya and Barreto, 2015) Complex Application Deployment Aneka Platform for Distributed Applications
+(Acquaviva et al., 2017) High-availability Deployment Middleware Support for High Availability in Multi-Cloud PaaS
+(Kiss, 2018) Complex Workflow Deployment CloudSME Simulation Platform
+(Kritikos et al., 2019) Business Process Provisioning Deployment Multi-Cloud Service Orchestration Frameworks
+(Roy and Kalafatis, 2023) Robotic App Deployment Challenge Hybrid Edge-Cloud Architecture
+(Sofia et al., 2023) Dynamic Adaptation Deployment Issue Context-aware Resource Allocation Strategies
+Multi-cloud Deployment Coordination and Integration This category identifies and classifies the challenges and solutions related to coordination and integration during the multi-cloud deployment phase.
+Of the 121 reviewed studies, 5 addressed these challenges.
+One such challenge is Multi-cloud PaaS Deployment, which seeks to overcome high entry barriers for deploying a PaaS infrastructure across multiple clouds.
+The proposed solution is Lightweight Proxies for PaaS Integration, which allows for the seamless integration of different PaaS services through lightweight proxies.
+One other challenge includes IoT Application Deployment, which presents issues when deploying IoT applications on multi-cloud platforms.
+The solution for this challenge is the DRA Framework and CI Broker for Multi-Cloud Deployment, which ensure smooth deployment and resource coordination across multiple clouds.
+Additionally, Data Transfer and Service Compatibility issues are managed using FaaS and storage services, ensuring that data transfer and compatibility between different cloud providers are not problematic.
+RBAM supports heterogeneous cloud and edge integration, offering frictionless management and operational flexibility for cloud-edge configurations.
+Lastly, private cloud integration is facilitated by the CLI tools that enable easy integration and the establishment of serverless environments.
+Monitoring and Scalability Challenges category identifies and classifies various issues and solutions related to monitoring, performance, and scalability in the context of multi-cloud container deployment.
+Of the 121 reviewed studies, 8 specifically addressed these issues.
+One challenge is Legacy Code Migration Deployment, which involves managing the migration of legacy code, tenant engine separation, and composition modeling during container deployment across multiple clouds.
+The proposed solution is a Reusable Architectural Pattern with Docker and Cloudant, which enables the reuse of an architectural pattern using Docker for containerization and Cloudant for the persistence layer.
+Another challenge is Tenant Performance Deployment, which focuses on managing tenant service performance and competition during multi-cloud containerized deployments.
+A proposed solution is the Multi-tenant and Multi-instance Hybrid Deployment scheme, based on container technology, which improves performance efficiency across tenants.
+A critical challenge is the monitoring of deployments, where existing tools have limitations in multi-cloud environment.
+The solution involves implementing JCatascopia, a platform-independent and interoperable monitoring system, which enhances the monitoring of multi-cloud setups and helps mitigate these challenges.
+Additional challenges include Resource Contention and QoS Degradation.
+The solutions include Dynamic CPU Adaptation through Linux patches, along with Kubernetes and performance monitoring tools.
+These measures ensure optimal resource usage and maintain performance consistency across cloud environments.
+Microservices Architecture and Containerization category classifies the challenges and solutions related to microservices architecture and containerization in multi-cloud environment.
+Of the 121 reviewed studies, 5 explicitly addressed challenges in this area.
+One such challenge involves the deployment of multilayer and multitier Web architectures in multi-cloud environment, referred to as Autonomic System Deployment.
+The proposed solution is a Self-Tuning Performance Model and Autonomic Management, which introduces a self-tuning performance model along with an autonomic management system to optimize deployment and management.
+Another challenge focuses on Microservice Deployment, which explores how microservice architectures can be deployed in multi-cloud environment.
+The proposed solution is a Metrics and Requirements Framework for Microservices, emphasizing the need for a structured framework to develop requirements and relevant metrics for deploying microservice-based applications.
+For Multi-cloud Application Configuration, the challenges involve deployment, configuration, and operation across multi-cloud environment.
+The one solution involves deploying a Runtime Environment with an Object Store and Artifact Repository to effectively manage configurations across different clouds.
+The Elastic Container Platform Dependency Deployment solution addresses the challenge of vendor lock-in and dependency on specific container platforms.
+The solution is the Separation of Elastic Platform and Cloud Application Definitions, which separates platform definitions from cloud application definitions to increase flexibility of application deployment.
+Additionally, deploying microservices in an edge environment can help resolve integration and communication issues in heterogeneous edge computing environments.
+The use of Docker, CI/CD Pipelines, and Service Mesh for Microservices enables the automated and secure deployment and communication of microservices.
+Network Connectivity and Hybrid Cloud Integration category classifies the challenges and solutions related to ensuring proper network connectivity in multi-cloud deployments and hybrid cloud integration.
+Out of the 121 studies reviewed, 5 specifically address these issues.
+One challenge, Inter-edge Bandwidth Deployment, deals with interconnecting distributed localities and managing bandwidth at the edge.
+The solution is a Modular Edge Cloud Computing Architecture, using containerization for better bandwidth management.
+A key challenge is Network Latency Deployment, which focuses on minimizing latency during multi-cloud deployments.
+The proposed solution is VM Selection for Composite Applications, optimizing virtual machine selection to reduce latency.
+Latency Reduction Deployment involves improving end-to-end performance through an Edge-Cellular Hybrid Infrastructure, combining edge and cellular networks.
+To address Vendor Lock-in, an Interoperability Layer Above Cloud Infrastructure enables smoother transitions between cloud providers.
+Finally, Dynamic Network Policy Deployment tackles the challenge of adapting network policies to container workloads, with the solution being Automation of Network Policies using tools like KUNERVA.
+Cloud Deployment Constraints and Challenges category identifies constraints and challenges in multi-cloud deployment, with 6 out of 121 studies addressing these issues.
+One challenge, ACS Compatibility Deployment, involves ensuring container compatibility with existing ACS systems, and the solution is a Scalable Multi-Cloud Deployment Framework.
+Other challenges include Manual Intervention Risk Deployment, which addresses risks from manual intervention.
+The solution is Automated Security Measures at Deployment to reduce vulnerabilities.
+The challenge of Multi-Provider Deployment involves managing multiple cloud providers.
+The proposed solution is the TOSCA-Based Deployment Modeling Approach, which provides a unified framework.
+The challenge of Cloud Provider Transition Deployment lies in addressing feature incompatibilities.
+The proposed solution is Semantic Interoperability in Multi-Clouds, which ensures smooth transitions between providers.
+Automation Tools Deployment challenge is solved with tools like Jenkins and Terraform for accurate cost and resource management.
+Infrastructure Provisioning and Container Deployment Challenges category classifies the challenges and solutions associated with infrastructure provisioning and container deployment in multi-cloud settings.
+Of the 121 reviewed studies, 8 focused on these challenges.
+One such challenge is Fog Computing Deployment, which deals with issues related to fog computing hardware and container deployment time.
+The proposed solution is Reorganized Container Images and Docker Deployment, suggesting that container images should be reorganized, and the Docker deployment process is modified to handle fog computing more effectively, enabling better resource management and optimization.
+Another challenge is Fog Device Deployment, which focuses on managing fog device availability during multi-cloud container deployments.
+The proposed solution is On-Demand Fog and Microservices Deployment, allowing for the on-demand creation of fog nodes and the on-demand deployment of microservices via Docker and Kubeadm.
+Distributed Compute Node Deployment involves challenges related to containerized application management in distributed compute nodes.
+Kernel Security Deployment addresses the security risks of kernel sharing and isolation.
+The proposed solution includes a Label-Based Scheduling Strategy and Enhanced Container Security for secure and efficient deployment.
+Application Deployment Workflow and Orchestration Challenges category classifies the challenges and solutions related to application deployment workflows and orchestration in multi-cloud environment.
+Of the 121 reviewed studies, 7 focused on these challenges.
+One such challenge, API Management Deployment, managing the APIs and interfaces provided by different cloud providers in multi-cloud container deployments.
+The proposed solution is a DevOps Approach for Multi-Cloud Applications, which suggests using DevOps to manage application lifecycles in multi-cloud environment.
+Another challenge, Web Service Deployment, involves ensuring proper and efficient deployment in heterogeneous computing environments.
+The solution is Kubernetes-Based Containerized Deployment, which uses Kubernetes to efficiently manage and deploy containerized environments.
+Complex Application Deployment and High-availability Deployment focus on deploying complex applications while ensuring high availability.
+The solutions include the Aneka Platform for Distributed Applications and Multi-cloud PaaS Middleware Support for High Availability, which ensure the availability and distribution of applications across multi-cloud environment.
 […]
 3.8. Monitoring Challenges and Solution Framework (RQ4 and RQ5)
-Table 14 provides an extensive overview of the monitoring challenges and solution framework for containerized applications in multi-cloud environment.
+Table LABEL:tab:monitoring-solutions provides an extensive overview of the monitoring challenges and solution framework for containerized applications in multi-cloud environment.
 It encompasses diverse challenges related to monitoring containerized applications in the complex multi-cloud landscape.
 Each category in the figure targets specific monitoring aspects and offers solutions to address these challenges.
-Performance, Consistency, and Variability Monitoring: This category emphasizes sustaining performance monitoring in dynamically varying cloud environments for containerized applications.
-It tackles the challenge of ensuring consistent performance monitoring and proposes solutions like performance variation monitoring and benchmarking for optimized performance.
-Resource and Infrastructure Monitoring: In this category, the focus is on monitoring resource and infrastructure management for containerized applications in multi-cloud environment.
-Challenges include network communication, system image and configuration monitoring, and dynamic resource management.
-Solutions involve service mesh implementation, Docker server deployment, and using Platform-as-a-Service (e.g., Aneka PaaS) for Cloud Computing.
-Security, Compliance, and Trust Monitoring: This category addresses the vital aspects of monitoring security, compliance, and trust management for containerized applications in multi-cloud settings.
-Challenges include accessing provider host information, continuous security control, and multi-cloud security compliance.
-Solutions include ML-based monitoring and security assurance platforms in the form of a SaaS (e.g.,MUSA), which support multi-cloud application runtime security control and transparency to enhance user trust.
-Application Optimization and Adaptability Monitoring: The challenges of this category revolve around monitoring application optimization and adaptability in multi-cloud environment.
-Challenges include real-time application monitoring, microservices optimization, application behavior, download execution, and design and state monitoring.
-To address these challenges, solutions encompass the Dynamic Resource Allocation (DRA) framework, Microservices-based Cloud Application-level Dynamic Orchestrator (MiCADO) optimization, Topology and Orchestration Specification for Cloud Applications (TOSCA) event modeling, FogDocker implementation, and adopting a cloud modeling framework.
-System Complexity and Standardization Monitoring: This category focuses on addressing system complexity and promoting standardization in monitoring practices across multi-cloud environment.
-Challenges include proprietary monitoring solutions and resource isolation mechanisms in the Linux kernel.
-Solutions encompass pervasive monitoring and identifying in-container leakage channels.
-Multi-cloud Coordination and Integration Monitoring: In this category, the challenges revolve around coordinating and integrating monitoring data across multiple cloud environments.
-Challenges include multi-cloud monitoring challenges, cross-level monitoring improvement, and integrated multi-cloud management monitoring.
-Solutions encompass self-healing techniques, business process monitoring, and integration with tools like Gnocchi.
+Performance, Consistency, and Variability Monitoring category identifies the challenges and solutions that ensure consistent performance monitoring in dynamically changing multi-cloud environment.
+Nine of the reviewed studies addressed these challenges.
+Performance Monitoring focuses on maintaining consistent performance tracking of applications.
+The solution, Performance Variation Monitoring, uses benchmarking systems to track and analyze runtime changes in performance, ensuring optimized performance throughout.
+Another challenge is Benchmarking for Performance and Cost Efficiency, which balances performance against cost.
+The solution, SBDO Optimization, optimizes resource utilization in real time to maximize efficiency and minimize costs.
+A related challenge, Website Performance Monitoring, is managed through Kubernetes Monitoring, which ensures uptime and performance across widely distributed cloud environments.
+Monitoring User Experience is handled by collecting CSP QoS Metrics, allowing cloud service providers to gather Quality of Service data, anticipate variations in performance, and manage the overall user experience effectively.
+Lastly, low-latency performance optimization in multi-cloud systems is addressed by Tail Latency Prediction, which uses predictive models to minimize network delays and ensure smooth application performance.
+Resource and Infrastructure Monitoring catagory classifies the challenges and corresponding solutions related to monitoring resource and infrastructure management for containerized applications in multi-cloud environment.
+Of the reviewed studies, 8 specifically addressed these challenges.
+One key challenge is Network Communication Monitoring, which involves managing service-to-service communication across cloud networks.
+The proposed solution is the implementation of a Service Mesh, which enhances visibility and security for network services.
+Another challenge is System Image and Configuration Monitoring, which tracks software updates, configuration changes, and performance across platforms.
+The solution proposed is Docker Server Deployment, which allows for easy updates to the containerized system.
+For Vcluster and Framework Management Monitoring, the solution is the implementation of LXC (Linux Containers), which manages application isolation and performance using lightweight containers.
+Multi-Cloud Resource Utilization Monitoring involves tracking the performance of resources, such as compute, storage, and network, used across multiple clouds.
+The solution uses Prometheus and Grafana for monitoring resource usage and availability, with an emphasis on optimization.
+Similarly, Dynamic Resource Management Monitoring solutions, such as the Aneka Platform, dynamically add or release resources based on demand.
+The challenge of cost tracking in resource utilization across multiple cloud environments, such as AWS and Azure, is addressed by the COSTA system, which monitors both resource utilization and cost efficiency.
+Finally, Prometheus is used for Real-Time Dynamic Resource Monitoring, allowing real-time data collection and analysis to optimize performance based on resource usage fluctuations.
+Table 14. Identified monitoring challenge and solution for containerized applications in multi-cloud environment
+ID Challenge Solution
+Category 1: Performance Monitoring Consistency and Variability
+(Almeida et al., 2014) Performance Monitoring Performance Variation Monitoring
+(Jha et al., 2019) Benchmarking Performance Variation and Cost Efficiency Monitoring SDBO Optimization
+(Gupta et al., 2021) Website Performance and Availability Monitoring Kubernetes Monitoring
+(Pandey et al., 2022) User Experience Variability Monitoring CSP QoS Metrics Collection
+(Dreibholz et al., 2019) Independent Performance Metrics Monitoring Performance Evaluations
+(Bhamare et al., 2015) Comprehensive Network Performance Measurement Monitoring Server Selection Scheme
+(Rahman and Lama, 2019) Microservice Performance Degradation Monitoring Tail Latency Prediction
+(Struhár et al., 2024) Container Performance Metrics Tracking CLM and OSLM for Resource Adjustments
+(Horchulhack et al., 2024) Container Performance and Resource Usage Monitoring Metric and Container Monitoring Tools
+(Hossain et al., 2023) Microservice Performance Monitoring Challenge Prometheus, Grafana, Kubernetes for Monitoring
+(Ramesh et al., 2023) Cross-Cloud Service Performance Monitoring FaaS and Storage Services for AI Monitoring
+(Nguyen and Chien, 2023) Stream Processing Performance Monitoring RBAM for Real-Time Performance Insights
+(Kataru et al., 2023) AWS Application Performance Monitoring AWS CloudWatch for Performance Monitoring
+(Alonso et al., 2023a) Multi-Cloud Performance Visibility Issue Monitoring Tools for Multi-Cloud Visibility
+(Harzenetter et al., 2023) Cross-Technology Performance Monitoring Issue TOSCA-based Workflows for Monitoring Applications
+(Huang and Wang, 2023) Industrial Robot System Monitoring Challenge Real-Time System Performance Tracking Tools
+Category 2: Resource and Infrastructure Management Monitoring
+(Gattobigio et al., 2022) Network Communication Monitoring Service Mesh Implementation
+(Verdugo et al., 2017) System Image and Configuration Monitoring Docker Server Deployment
+(Cui et al., 2016) Vcluster and Framework Management Monitoring LXC Implementation
+(Pahl, 2015) Multi-Cloud Resource Utilization Monitoring Prometheus and Grafana Monitoring
+(Buyya and Barreto, 2015) Dynamic Resource Management Monitoring Aneka Platform
+(da Silva et al., 2023) Cost and Resource Usage Monitoring COSTA for Cost and Resource Monitoring
+(Dogani et al., 2023) Real-Time Dynamic Resource Monitoring Prometheus for Real-Time Monitoring
+(Wong et al., 2023) Resource Usage and Performance Issue Prometheus for Real-Time Data Collection
+(Marques et al., 2024) Real-Time Container Monitoring Prometheus with Forecasting for Resource Management
+(Santos et al., 2023) Distributed Cloud Resource Monitoring Issue AWS CloudWatch and Azure Monitor for Insights
+(Nagarajan et al., 2023) Resource Allocation and Energy Monitoring Issue Tools for Resource, Energy, and SLA Monitoring
+Category 3: Security, Compliance, and Trust Management Monitoring
+(Ramos et al., 2021) Provider Host Information Access Monitoring ML Monitoring
+(Rios et al., 2017) Continuous Security Control Monitoring MUSA Platform
+(Casola et al., 2017) Multi-Cloud Security Compliance Monitoring MUSA Security Compliance
+(Acquaviva et al., 2017) Service Continuity and Data Privacy Monitoring NoMISHAP Implementation
+(Wahab et al., 2016) Trust Relationship and Network Contact Monitoring Trust Bootstrapping
+(Fugkeaw, 2023) Access and Authentication Monitoring Issue Blockchain for Secure Access Event Logging
+(Wang et al., 2023) Malware Behavior and System Monitoring DockerWatch for Container Behavior Monitoring
+(Neeraj et al., 2023) Log Data Integrity Monitoring Issue Blockchain-based Logs for SLA Monitoring
+Category 4: Application Optimization and Adaptability Monitoring
+(Ferry et al., 2014) Adaptive System Design and State Monitoring Cloud Modelling Framework
+(Lorenzo et al., 2019) Partial Download Execution Monitoring FogDocker Implementation
+(Wurster et al., 2018) Event-Driven Application Behavior Monitoring TOSCA Event Modeling
+(Guerrero et al., 2018a) Microservice Optimization Monitoring MiCADO Optimization
+(Bou Ghantous and Gill, 2021) Real-Time Application Monitoring DRA Framework
+(Sofia et al., 2023) Container Application Adaptability Monitoring Advanced Monitoring with Orchestration Integration
+(Petrovski and Gusev, 2024) Serverless and Containerized Monitoring Challenge AWS Tools for Performance Monitoring
+Category 5: System Complexity and Standardization Monitoring
+(Sabbioni et al., 2020) Proprietary Monitoring Solution Standardization Pervasive Monitoring
+(Gao et al., 2017) Kernel Resource Isolation Mechanism Monitoring Leakage Defense
+(Alonso et al., 2019) Multi-Cloud Application Monitoring Complexity DECIDE DevOps Expansion
+(Barna et al., 2017) Monitoring Agent Coordination Autonomic Management
+(Sami and Mourad, 2020) Heterogeneous Fog Resource Management Monitoring Kubernetes Utility Architecture
+(Baresi et al., 2016) Application Topology and Dependency Monitoring MicroCloud TOSCA Library
+(Trihinas et al., 2015) Elastic Cloud Service Monitoring Research JCatascopia Monitoring
+(Fadda et al., 2019) Microservice Monitoring Diversity QoM Definition
+(Harwalkar et al., 2019b) Compute and Network Resource Utilization Monitoring Time Series Resource Utilization
+(Ahmad and Aujla, 2023) Real-Time Data Operations Monitoring Smart Contracts for GDPR Monitoring
+(Centofanti et al., 2024) Power Consumption Monitoring Issue Prometheus with Power Monitoring Integration
+(Tiwari and Sharma, 2023) Serverless Function Monitoring Complexity CLI Tool for Function Monitoring
+Category 6: Multi-cloud Coordination and Integration Monitoring
+(Sitaram et al., 2018) Integrated Multi-Cloud Management Monitoring Gnocchi Integration
+(Mfula and Nurminen, 2018) Multi-Cloud Monitoring Challenges Self-Healing Techniques
+(Kritikos et al., 2019) Cross-Level Monitoring Improvement Business Process Monitoring
+(Lee and Nam, 2023) Real-Time Network Policy Monitoring Challenge KUNERVA for Real-Time Network Policy Monitoring
+(Sofia et al., 2024) Decentralized System Observability Issue CODECO for Orchestration and Data Observability
+Security, Compliance, and Trust Monitoring: category addresses the challenges and solutions related to monitoring security, compliance, and trust management for containerized applications in multi-cloud environment.
+Out of the 121 reviewed studies, 9 specifically deal with these concerns.
+One challenge is Provider Host Information Access Monitoring, where limitations in accessing host data are overcome using ML Monitoring, which employs machine learning to track system performance and usage patterns.
+Another challenge, Continuous Security Control Monitoring, is addressed by the MUSA platform, which enforces and continuously monitors security in real time for containerized applications.
+Multi-Cloud Security Compliance Monitoring ensures continuous security compliance across multi-cloud environment.
+This is achieved through the MUSA Security Compliance Solution, a platform for continuous compliance management.
+The challenges of maintaining data privacy, fault tolerance, and service continuity in multi-cloud environments are addressed under Service Continuity and Data Privacy Monitoring.
+The NoMISHAP Implementation is proposed to handle these monitoring issues.
+Additionally, Trust Bootstrapping establishes trust through a chain of endorsement and decision tree classification.
+Access and Authentication Monitoring uses blockchain technology to enable secure and transparent logging of authentication activities.
+For Malware Behavior and System Monitoring, DockerWatch monitors container behavior with minimal performance overhead.
+Application Optimization and Adaptability Monitoring: This category identifies and discusses challenges related to optimizing and adapting applications in multi-cloud environment.
+Out of the 121 reviewed studies, 8 specifically addressed these challenges with corresponding solutions.
+Key challenges include Real-Time Application Monitoring, which focuses on taking immediate corrective actions using the DRA Framework, and Microservice Optimization Monitoring, which enhances monitoring across multiple clouds through MiCADO Optimization.
+Other challenges include Partial Download Execution Monitoring, which ensures the correct execution of partially downloaded files.
+The proposed solution for runtime accuracy is the FogDocker Implementation.
+Another challenge, Event-Driven Application Behavior Monitoring, involves monitoring event-driven application behavior.
+The TOSCA Event Modeling solution efficiently handles application events.
+For Adaptive System Design and State Monitoring, continuous design and dynamic system adaptability are managed using the Cloud Modelling Framework.
+Additionally, the challenge of Container Application Adaptability Monitoring is addressed through advanced orchestration techniques, which provide real-time views of distributed applications.
+The solution involves using AWS tools to ensure high performance for different deployment models, addressing challenges in Serverless and Containerized Monitoring.
+AWS tools are recommended to ensure high performance for different deployment models, addressing challenges in Serverless and Containerized Monitoring.
+System Complexity and Standardization Monitoring cateogy identifies and addresses the challenges of managing system complexity and standardization in monitoring solutions across multi-cloud environment.
+Of the 121 studies reviewed, 4 specifically focused on these challenges and proposed corresponding solutions.
+A key challenge is the standardization of proprietary monitoring solutions, for which the proposed solution is Pervasive Monitoring.
+Pervasive Monitoring focuses on implementing pervasive monitoring techniques for containerized applications within multi-cloud PaaS platforms.
+Another significant challenge is Kernel Resource Isolation Mechanism Monitoring, with the proposed solution being Leakage Defense.
+This solution addresses in-container leakage channels by implementing a two-stage isolation mechanism in Linux-based systems.
+Additionally, the challenge of Multi-Cloud Application Monitoring Complexity is addressed by the DECIDE DevOps Expansion, which extends the framework for analyzing multi-cloud containerized applications.
+The Autonomic Management Framework further enhances monitoring through Monitoring Agent Coordination, improving the coordination of monitoring agents using a performance model and adaptive actions for better system behavior within containerized applications.
+Multi-cloud Coordination and Integration Monitoring category identifies and addresses the challenges related to coordinating and integrating monitoring data across multiple cloud environments.
+Out of the 121 studies reviewed, 5 focused on these challenges.
+One key challenge is Multi-cloud Monitoring Improvement, and the proposed solution is the use of Self-Healing Techniques to autonomously detect and correct issues within cloud infrastructures.
+Another challenge is Cross-Level Monitoring, which is addressed through Business Process Monitoring Integration to ensure that business processes are effectively monitored across multiple cloud platforms.
+Lastly, Integrated Multi-cloud Management Monitoring requires advanced solutions, such as Integration with Gnocchi, to efficiently gather and monitor metrics from diverse cloud resources.
 […]
-Table 14. Monitoring challenge and solution framework for containerized applications in multi-cloud environment
-ID Challenge Challenge Description Solution Solution Description
-Category 1: Performance Monitoring Consistency and Variability
-S01 Performance Monitoring Ensuring sustained performance monitoring in dynamically varying cloud environments Performance Variation Monitoring Monitoring performance fluctuations in containerized applications through benchmarking periods
-S02 Benchmarking Performance Variation and Cost Efficiency Monitoring Efficiently tracking performance disparities and cost-effectiveness for refined monitoring in benchmarking SDBO Optimization Maximizing user budget utility with the optimization feature in SDBO
-S29 Website Performance and Availability Monitoring Implementing continuous oversight over website availability and performance Kubernetes Monitoring Utilizing Kubernetes for monitoring performance of containerized services in multi-cloud environment
-S38 User Experience Variability Monitoring Addressing monitoring fluctuations in user experience based on diverse factors CSP QoS Metrics Collection Collecting Cloud Service Provider (CSP) Quality of Service (QoS) metrics for benchmark data in multi-cloud environment
-S43 Independent Performance Metrics Monitoring Establishing application-independent performance metrics for Round Trip Time (RTT) and payload throughput Performance Evaluations Conducting performance evaluations and autonomic resource provisioning for multi-cloud and MEC platforms using containerized applications
-S50 Comprehensive Network Performance Measurement Monitoring Ensuring comprehensive performance measurements monitoring and accounting for transmission delays in the network Server Selection Scheme Applying a server selection scheme to minimize latency in multi-cloud environment based on end-to-end performance measurements
-S62 Microservice Performance Degradation Monitoring Addressing unpredictable performance degradation in containerized microservices Tail Latency Prediction Implementing a modeling approach to predict end-to-end tail latency in multi-layer cloud data of containerized applications
-Category 2: Resource and Infrastructure Management Monitoring
-S05 Network Communication Monitoring Managing service-to-service communications over a network Service Mesh Implementation Deploying a network service mesh for service-discovery, secure connectivity, and visibility in multi-cloud platforms
-S09 System Image and Configuration Monitoring Monitoring of system images, remote software packages, usage history, and container execution, configuration, and monitoring.
-Docker Server Deployment Deploying containerized services using Docker Server API version 1.23 in multi-cloud platforms
-S21 Vcluster and Framework Management Monitoring Monitoring and managing the performance and resource usage of each Vcluster and application framework can be complex when dealing with multiple tenants and different types of applications.
-LXC Implementation Using Lightweight Containers (LXC) for efficient resource allocation and scalability in multi-cloud environment
-S24 Multi-Cloud Resource Utilization Monitoring Monitoring container performance and resource usage across multiple clouds Prometheus and Grafana Monitoring Implementing Prometheus and Grafana for monitoring performance and availability of containerized applications in multi-cloud environment
-S51 Dynamic Resource Management Monitoring Applications should add and release resources dynamically by measuring and monitoring usage Aneka Platform Utilizing the Aneka platform for monitoring resources and managing dynamic resource acquisition and release in multi-cloud environment
-Category 3: Security, Compliance, and Trust Management Monitoring
-S04 Provider Host Information Access Monitoring Overcoming the restriction in accessing provider host information ML Monitoring Using machine learning to supervise performance of containerized applications and OS usage
-S34 Continuous Security Control Monitoring Strengthening continuous monitoring and control over security MUSA Platform Applying the MUSA Security Assurance Platform for continuous monitoring and dynamic security enforcement of containerized applications
-S52 Multi-Cloud Security Compliance Monitoring Monitoring security compliance consistently across multiple cloud providers MUSA Security Compliance Using the MUSA Security Assurance Platform to monitor multi-cloud application compliance and ensure security
-S53 Service Continuity and Data Privacy Monitoring Ensuring data privacy, uninterrupted service, fault tolerance, and availability NoMISHAP Implementation Implementing the NoMISHAP solution to address monitoring challenges in multi-cloud environment with containerized applications
-S83 Trust Relationship and Network Contact Monitoring Building trust relationships, collecting judgments, and maintaining networks of contacts Trust Bootstrapping Implementing a trust bootstrapping mechanism that combines endorsement and decision tree classification for containerized applications in multi-cloud platforms
-Category 4: Application Optimization and Adaptability Monitoring
-S20 Adaptive System Design and State Monitoring Implementing continuous design with dynamic adaptability; Enhancing system state evaluation through robust monitoring Cloud Modelling Framework Utilizing a Cloud Modelling Framework for continuous design and dynamic adjustment of multi-cloud containerized applications
-S36 Partial Download Execution Monitoring Ensuring accurate execution of applications for partially downloaded files FogDocker Implementation Using FogDocker with low runtime overhead to ensure execution correctness of containerized applications
-S44 Event-Driven Application Behavior Monitoring Modeling event-triggered behavior of application components TOSCA Event Modeling Applying TOSCA modeling for event sources, events, and functions in containerized applications
-S68 Microservice Optimization Monitoring Enhancing surveillance over microservices across multiple cloud environments for optimization MiCADO Optimization Utilizing MiCADO for runtime monitoring and optimization of cloud applications in multi-cloud environment
-S85 Real-Time Application Monitoring Real-time monitoring of applications for immediate corrective action DRA Framework Implementing the DRA framework with the Monitoring and Tracking Platform (M4) for containerized applications in multi-cloud platforms
-Category 5: System Complexity and Standardization Monitoring
-S10 Proprietary Monitoring Solution Standardization Overcoming the constraints of complex, proprietary monitoring solutions and fostering standardization Pervasive Monitoring Implementing pervasive monitoring solutions for containerized applications within multi-cloud PaaS
-S25 Kernel Resource Isolation Mechanism Monitoring Addressing the limitations of resource isolation mechanisms in the Linux kernel Leakage Defense Identifying in-container leakage channels and implementing a two-stage defensive approach in multi-cloud environment
-S27 Multi-Cloud Application Monitoring Complexity Simplifying monitoring processes in multi-cloud applications DECIDE DevOps Expansion Expanding the DECIDE DevOps framework for native multi-cloud containerized applications
-S28 Monitoring Agent Coordination Enhancing the coordination of monitoring agents for improved system behavior Autonomic Management Implementing an autonomic management system with monitoring agents, a performance model, and adaptive actions for containerized applications
-S33 Heterogeneous Fog Resource Management Monitoring Efficient management of diverse fog resources Kubernetes Utility Architecture Implementing a master-worker nodes architecture with Kubernetes Utility (Kubeadm) in multi-cloud environment
-S42 Application Topology and Dependency Monitoring Enhancing coordinated adaptation through thorough understanding of application topologies and dependencies MicroCloud TOSCA Library Utilizing the MicroCloud TOSCA library for standardized description of topologies in multi-cloud containerized applications
-S45 Elastic Cloud Service Monitoring Research Investigating elastic cloud service monitoring as an open research problem JCatascopia Monitoring Utilizing JCatascopia for multi-layer, platform-independent monitoring of cloud containerized applications
-S46 Microservice Monitoring Diversity Addressing diverse monitoring requirements and varying metrics quality in microservices QoM Definition Defining Quality of Monitoring (QoM) and implementing an API for cloud provider capabilities in multi-cloud environment
-S49 Compute and Network Resource Utilization Monitoring Strengthening surveillance and prediction capabilities of compute and network resource utilization Time Series Resource Utilization Developing algorithms for monitoring and predicting resource utilization of containerized applications using time series data
-Category 6: Multi-cloud Coordination and Integration Monitoring
-S58 Integrated Multi-Cloud Management Monitoring Implementing integrated monitoring and management across multi-cloud environment Gnocchi Integration Integrating with federated gnocchi for consolidated monitoring and management of multi-cloud environment
-S74 Multi-Cloud Monitoring Challenges Addressing monitoring challenges in multi-cloud environment Self-Healing Techniques Implementing a rule-based algorithm for monitoring and self-healing techniques in multi-cloud environment
-S84 Cross-Level Monitoring Improvement Ensuring precise and adaptable metrics computation, overcoming inadequate cross-level monitoring approaches Business Process Monitoring Conducting cross-cloud and cross-level business process monitoring and coordinated adaptation of containerized applications
 3.9. Tools and Frameworks (RQ6)
-In response to our RQ6 concerning the tools and frameworks utilized for developing containerized multi-cloud applications, an in-depth exploration of the selected studies was conducted.
-This comprehensive analysis led to the identification of a total of 160 distinct tools and frameworks (see Table 15), all of which are systematically classified into 45 overarching categories, further delineated into 6 subcategories.
-Notably, the three most prominent categories emerged from this investigation are cloud services and management platform, infrastructure management technologies, and networking tools and framework.
-Each of these categories encapsulates a range of tools aimed at facilitating the seamless development, deployment, and management of multi-cloud applications.
-In the following we briefly report the each of the category, shedding light on the tools and frameworks that collectively form the foundation for containerized multi-cloud application development.
-Cloud Services and Management Platforms: Cloud Services and Management Platforms: This category of tools facilitates the management and orchestration of cloud services across multi-cloud environments.
-It includes a broad range of platforms covering Infrastructure as a Service (IaaS), Platform as a Service (PaaS), and Software as a Service (SaaS).
-These tools are essential for provisioning virtual resources, deploying applications, and delivering software across different cloud platforms.
-Examples include Amazon EC2, Google Cloud App Engine, and Microsoft Office 365.
-Infrastructure Management Technologies: This category of tools facilitates the management of cloud infrastructure.
-It encompasses container platforms, virtualization solutions, Kubernetes services, orchestration platforms, and tools for configuration management and continuous integration.
-Key technologies like Docker, Kubernetes, Puppet, Chef, and Terraform play a vital role in automating and optimizing the infrastructure underpinning containerized applications.
-Monitoring Tools: This category of tools facilitates the monitoring and analysis of cloud and container operations.
-It includes solutions for cloud monitoring, container monitoring, visualization & alerting, resource monitoring, and logging & analysis.
-Tools such as Amazon CloudWatch, cAdvisor, Prometheus, Grafana, and the ELK stack are instrumental in ensuring the performance and reliability of cloud services.
-Networking Tools: This category of tools facilitates the establishment and management of network functionalities in a multi-cloud environment.
-It covers load balancers, reverse proxies, network services, controllers, paradigms, types, and identity protocols, along with storage networks and event buses.
-Technologies like Software-Defined Networking (SDN), Neutron, and Kafka event bus are crucial for maintaining robust and efficient network operations.
-Development and Programming: This category of tools facilitates the development and programming aspects in cloud environments.
-It includes orchestration tools, various programming languages, integrated development environments (IDEs), service meshes, APIs, and cloud modeling tools.
-Examples are Java, Python, Eclipse IDE, Istio, REST API, and Cloud Modelling Framework (CLOUDMF), which support the creation and deployment of cloud-native applications.
-Security and Assurance: This category of tools facilitates the security and compliance of cloud environments.
-It encompasses vulnerability management, security platforms & tools, access control & policies, Linux security, and cryptography models & tools.
-Technologies like OpenVAS Vulnerability Scanner, Kubernetes.eBPF, Seccomp-BPF, and Attribute-Based Encryption and Signature are essential for protecting cloud infrastructure and data against various security threats.
+In response to RQ6 regarding tools and frameworks for developing containerized multi-cloud applications, we identified 87 distinct tools and frameworks, classified into 5 subcategories (see Table 15 and the Tools and Frameworks sheet in (Waseem et al., 2024)).
+Below, we provide an overview of each category, along with key tools and frameworks, highlighting their functionality, adoption, and distinctive features.
+Container Orchestration and Management Tools: are essential for automating the deployment, scaling, and management of containerized applications across distributed and multi-cloud environment.
+These tools assist developers in managing large-scale container deployments and efficiently allocating resources across different cloud providers.
+Among the most popular orchestration and management tools, Kubernetes is widely adopted for orchestrating containers at scale, providing high availability, and facilitating load balancing across nodes.
+Docker and its variants, such as Docker Swarm, offer more convenient container orchestration but are less scalable than Kubernetes.
+Mesos is another orchestration tool that supports not only containers but also other workloads, such as big data processing tasks.
+Tools like LXC and runC provide the runtime environment to run a single container with fine-grain control.
+Overall, this category covers key orchestration and container management tools that address diverse deployment needs in multi-cloud environment.
+Multi-cloud and Hybrid Cloud Platforms: are designed to support organizations in deploying and managing applications across different cloud providers smoothly.
+In this category, we identify and organize tools like Cloudify and CMP2, which offer end-to-end multi-cloud orchestration capabilities, along with platforms such as Cloudcheckr and MistIO, which provide monitoring and cost management across various cloud services.
+Similarly, AWS SDK and the CloudSME Simulation Platform (CSSP) are widely adopted for interacting with cloud APIs and simulating cloud environments, respectively.
+Moreover, hybrid platforms like OpenStack and VMware vSphere offer a blend of public and private cloud deployment options, which are essential for organizations pursuing a hybrid cloud strategy.
+This category highlights the tools that enable integration, scalability, and cost management for applications running across multiple clouds.
+Networking, Security, and Access Management Tools: This category represents the tools that help to minimize security risks and ensure connectivity across containerized workloads in multi-cloud systems.
+As multi-cloud deployments become increasingly complex, advanced networking tools like Open vSwitch (OvS) enable developers to manage virtual networks across various cloud platforms.
+Security tools such as Seccomp-BPF and Linux Security Modules (LSM) provide critical protection to containers by enforcing strict access controls and sandboxing.
+Identity management solutions, such as Host Identity Protocol (HIP) and access control policies, ensure proper authentication and authorization for users and services in the cloud.
+Monitoring, Management, and Automation Tools: Tools in this category provide the foundation for a scalable and automated approach to managing distributed applications by ensuring that workloads are continuously optimized and monitored on cloud platforms.
+Tools like Prometheus and Grafana offer powerful visualization capabilities for monitoring CPU usage, memory consumption, and network activity, aiding system administrators.
+With Terraform and Puppet, it is possible to automate infrastructure management, enabling developers to define infrastructure as code and automate provisioning across different cloud environments.
+Other tools, such as Ceilometer and Gnocchi, are responsible for gathering telemetry data in large OpenStack deployments.
+Development, APIs, and Cloud Storage Solutions: This category organizes the main tools that support software development, API integration, and data storage in containerized multi-cloud environment.
+The tools include REST API and YAML, which enable services and applications deployed across multiple cloud providers to communicate for data sharing and inter-service interactions.
+General programming languages such as Java, Go, and Python are commonly used to develop applications based on cloud-native architecture, alongside tools for version control, such as GitHub, and project management using Maven.
+On the data storage side, tools like gcloud storage and azblob provide scalable and fault-tolerant systems for managing large datasets in cloud environments.
 Table 15. Tools and frameworks for containerized applications in multi-cloud environment
-Subcategory Tools and Frameworks
-Category 1: Cloud Services and Management Platforms
-Cloud Management Platforms
-Cloudyn (S65), IBM Brokerage Services (S65), CloudBroker Platform (S69), CloudcheckR (S79), MistIO (S79), ManageIQ (S67, S79), CMP2 (S79), AWS SDK (S71), CloudSME Simulation Platform (CSSP) (S69), MiCADO (S69), WS-PGRADE/gUSE (S69), Vagrant (S08), AWS OpsWorks (S38), Universal Compute Xchange (UCX) (S65), Open vSwitch (OvS) (S72), Liferay portlets (S69), Cloud Service Providers (CSPs) (S71), AWS Data Pipeline (S78), Cloud computing (S25), Amazon AWS (S42, S43), GENI (S38), Google Cloud Platform (S38, S53), Microsoft Azure (S38, S48, S51, S54, S60), AWS EC2 (S51), IBM (S49), Amazon Web Services (S54), Google Cloud (S54), SalesForce (S53), Multi-Cloud (S60), Cloud computing platforms (S74), Multiple independent clouds (S74), Cloud computing (S84)
-Cloud Deployment Model
-Public clouds: (S37, S53, S59, S61), Google cloud (S07), Microsoft Azure (S08, S16, S19), AWS (S65, S67, S80, S81), Google’s cloud (S65), Google Compute Engine (S67, S80), Alibaba (S65), Tencent (S65), IBM Watson Health Cloud (S78).
-Private cloud: OpenStack (S20, S21, S43), VMware vSphere Specialized Cloud Platforms: European Grid Infrastructure Federated Cloud (EGI FC) (S81)
-Infrastructure as a Service (IaaS)
-Amazon EC2 (S01, S51), AWS (S02, S65, S67, S80, S81), Azure (S02, S08, S16, S19, S67, S80), Google GCE (S03, S67, S80), Aliyun ECS (S03), Baidu BCE (S03), Tencent CVM (S03), JD VM (S03), VMware vSphere (S08), GoGrid (S51), Windows Azure (S51)
-Platform as a Service (PaaS)
-Google Cloud App Engine (S02, S10, S15), Heroku (S08, S53), IBM Bluemix Workflow Service (S17, s53), Amazon Elastic Beanstalk (S20), Cloud Foundry (S24), OpenShift (S24, S53, S55)
-Software as a Service (SaaS)
-SaaS (S15), Microsoft Office 365 (S53)
-Distributed Processing Frameworks
-Hadoop (S21, S37), Spark (S21, S28), Apache Yarn (S37), Storm (S21), Apache Kafka (S34, S35)
-Data Storage
-Apache Cassandra (S28), Firebase real-time database (S35), Redis (S36, S42), MySQL (S42), NoSQL (S85), SQL (S85)
-Messaging Systems
-Apache Kafka (S34, S35), Zookeeper (S35)
-Category 2: Infrastructure Management Technologies
-Container Platforms
-Docker (S08, S13, S16, S17, S18, S25, S27, S28, S29, S30, S31, S32, S35, S55, S59, S62, S74, S77, S79, S80), Smart Docker Benchmarking Orchestrator (SDBO) (S01), Docklet (S21), LXC (S24, S25, S29), Docker container engine (S24), Docker container (S23), Docker CE version 18.01 (S36), Docker Hub (S36)
-Virtualization
-Virtual Machines (S05, S22, S23, S46, S72), Amazon EC2 VMs (S18), Docker container virtualization technology (S74), Xen (S19), VMware (S19), Hyper-V (S19), KVM (S19), OpenVZ (S19), Libvirt API (S19)
-Kubernetes Services
-Microsoft AKS (S03), Aliyun Kubernetes Engine (S03), Tencent Kubernetes Engine (S03)
-Orchestration Platforms
-Kubernetes (S18, S22, S29, S55, S56, S59, S62, S63, S80), Docker Swarm (S18, S22, S39, S59, S80), Mesos (S21, S22, S59, S80), Yarn (S21), Kubeadm (S33), Kubernetes for container orchestration (S62), Apache Mesos (S59, S80), MiCADO (S69)
-Configuration Management
-Puppet (S08), Chef (S08, S30), Terraform (S27, S29), Shell scripts (S44), CloudBees (S08) Continuous Integration CloudBees (S08), Maven (S08), NPM (S08), Github (S30), Github (S30)
-Container Isolation
-Sandboxie (S24), Warden/Garden (S24), Rocket (S24)
-Container Infrastructure
-Linux Container (S16), Docker Containers on Linux Host (S13), Kubernetes cluster with a Docker (S04), Amazon AWS ECS (S18), Container Technology (S23)
-Category 3: Infrastructure Management Technologies
-Load Balancer
-Load balancer (S23)
-Reverse Proxy
-Reverse proxy technology (S23), httpd (S36), nginx (S36)
-Ports and Interfaces
-TrunkPort (S55)
-Network Services
-Neutron (S55)
-Gateways and Tunnels
-Gateways (S64), GRE Tunnel (S72)
-Network Controllers
-Software-Defined Networking (SDN) controller (S72)
-Identity Protocols
-Host Identity Protocol (HIP) (S77)
-Storage Networks
-MCSS (S75), Multi-Cloud storage overlay network (S75)
-Network Tools
-OpenADN (S50), OpenDayLight (S50), Kuryr (S55), Flannel (S55)
-Event Buses
-Kafka event bus (S34)
-Category 4: Development and Programming
-Orchestration
-MUSA Deployer tool (S52), MUSA Modeler tool (S52), OSM (S55), ManageIQ (S55)
-Programming Languages
-Java (S30, S34, S35, S54, S80, S81), Go programming language (S32, S36), Python (S35, S54, S56, S58, S79), C# (S54), Ruby (S54), PHP (S54)
-IDEs and Editors
-Eclipse IDE with tight integration with OPTIMUS (S27), MCSLA editor (S27)
-Service Mesh
-Consul (S05), Istio (S05), Istio service mesh (S59)
-APIs
-REST API (S56), OpenStack API (S64)
-Cloud Modeling
-Cloud Modelling Framework (CLOUDMF) (S20), Cloud Modelling Language (CLOUDML) (S20, S81), CAMEL modeling language (S52, S80, S84)
-Category 5: Development and Programming
-Vulnerability Management
-OpenVAS Vulnerability Scanner (S30)
-Security Platforms and Tools
-MUSA Security Assurance Platform (S34), Firewall Deployment (S30)
-Access Control and Policies
-Access control policies (S25)
-Linux Security
-Linux top command (S34), Kubernetes.eBPF (S40), Seccomp-BPF (S40), Landlock LSM (S40), Pledge (S40), Security Namespace (S40), Linux Security Modules (LSM) (S40)
-Cryptography Models and Tools
-Attribute-Based Encryption and Signature (S75), ABEABS cryptographic model (S75)
-Category 6: Monitoring Tools
-Cloud Monitoring
-Amazon CloudWatch (S19)
-Container Monitoring
-cAdvisor (S22), Prometheus (S22), AlertManager (S22)
-Visualization and Alerting
-Grafana (S27), Telegraf (S27)
-Resource Monitoring
-Resource monitoring tool (S60), Ceilometer (S56, S58), Gnocchi (S58)
-Logging and Analysis
-ELK stack (S74)
+Category Tool, Framework, and Language
+Container Orchestration and Management Kubernetes (Pai and Kumar, 2021; Wong et al., 2019; Gupta et al., 2021; Bolivar et al., 2018; del Castillo et al., 2013; Rodigari et al., 2021; Rahman and Lama, 2019; Pellegrini et al., 2017; Quint and Kratzke, 2018; Barletta et al., 2024; Struhár et al., 2024; Barati et al., 2023; Horchulhack et al., 2024; Dogani et al., 2023; Wong et al., 2023; Marques et al., 2024; Orzechowski et al., 2023; Merlino et al., 2024; Ahmad and Aujla, 2023; Centofanti et al., 2024; Hossain et al., 2023; Liu and Qin, 2024; Tundo et al., 2024; Pandi et al., 2023; Sofia et al., 2024; Lee and Nam, 2023; Sofia et al., 2023; Nguyen and Chien, 2023; Tiwari and Sharma, 2023; Santos et al., 2023; Roy and Kalafatis, 2023; Kataru et al., 2023; Benmerar et al., 2023; Alonso et al., 2023a; Harzenetter et al., 2023; Huang and Wang, 2023; Wang et al., 2023; Nagarajan et al., 2023), Docker (Pham et al., 2015; Naik, 2016a; Ouyang et al., 2022; Slominski et al., 2015; Pai and Kumar, 2021; Gao et al., 2017; Alonso et al., 2019; Barna et al., 2017; Gupta et al., 2021; Sheridan et al., 2017; Mittal and Risco-Martín, 2017; Al-shammari and Alwan, 2018; Carpio et al., 2020; Bolivar et al., 2018; Rodigari et al., 2021; Rahman and Lama, 2019; Mfula and Nurminen, 2018; Ranjbar et al., 2017; Serhiienko and Spillner, 2018; Quint and Kratzke, 2018), Docker Swarm (Pai and Kumar, 2021; Wong et al., 2019; Altran et al., 2022; Rodigari et al., 2021; Quint and Kratzke, 2018), Mesos (Cui et al., 2016; Wong et al., 2019; Rodigari et al., 2021; Quint and Kratzke, 2018), LXC (Pahl, 2015; Gao et al., 2017; Gupta et al., 2021), OverlayFS (Al-shammari and Alwan, 2018), runC (Bélair et al., 2019), Container Technology (Wang et al., 2020), Docklet (Cui et al., 2016)
+Multi-cloud and Hybrid Cloud Platforms multi-cloud environment (Shi et al., 2020; Fadda et al., 2019; Di Modica et al., 2019; Tang, 2021), Cloudify (Zeck and Bouroudjian, 2017), CloudBroker Platform (Kiss, 2018), Cloudcheckr (Serhiienko and Spillner, 2018), MistIO (Serhiienko and Spillner, 2018), ManageIQ (Tang, 2021; Serhiienko and Spillner, 2018), CMP2 (Serhiienko and Spillner, 2018), AWS SDK (Graupner et al., 2015a), CloudSME Simulation Platform (CSSP) (Kiss, 2018), Vagrant (Pham et al., 2015), AWS OpsWorks (Pham et al., 2015), Universal Compute Xchange (UCX) (Zeck and Bouroudjian, 2017), Liferay portlets (Kiss, 2018), Cloud Service Providers (CSPs) (Graupner et al., 2015a), AWS Data Pipeline (Sitaram et al., 2018), GENI (Pandey et al., 2022), Google Cloud Platform (Pandey et al., 2022; Acquaviva et al., 2017; Tang, 2021; Quint and Kratzke, 2018), AWS EC2 (He et al., 2022; Buyya and Barreto, 2015; Pai and Kumar, 2021), Alibaba cloud (Zeck and Bouroudjian, 2017; Quint and Kratzke, 2018), Tencent (Zeck and Bouroudjian, 2017; Quint and Kratzke, 2018), OpenStack (Cui et al., 2016; Dreibholz et al., 2019), VMware vSphere (Mfula and Nurminen, 2018), Aliyun ECS (Hua et al., 2020), Baidu BCE (Hua et al., 2020), Tencent CW (Hua et al., 2020), JD VW (Hua et al., 2020), GoGrid (Buyya and Barreto, 2015)
+Networking, Security, and Access Management Open vSwitch (OvS) (Alaluna et al., 2017), Cloud networks (Ranjbar et al., 2017), Host Identity Protocol (HIP) (Ranjbar et al., 2017), Gateways (Harwalkar et al., 2019a), Pledge (Bélair et al., 2019), Seccomp-BPF (Bélair et al., 2019), Landlock LSM (Bélair et al., 2019), Linux Security Modules (LSM) (Bélair et al., 2019), MUSA Security Assurance Platform (Rios et al., 2017), Firewall Deployment (Sheridan et al., 2017), Access control policies (Gao et al., 2017), Attribute-Based Encryption and Signature (Shen et al., 2017), ABE/ABS cryptographic model (Shen et al., 2017)
+Monitoring, Management, and Automation Prometheus (Wong et al., 2019; Dogani et al., 2023; Wong et al., 2023; Marques et al., 2024; Orzechowski et al., 2023; Merlino et al., 2024; Ahmad and Aujla, 2023; Centofanti et al., 2024; Hossain et al., 2023; Liu and Qin, 2024; Tundo et al., 2024; Roy and Kalafatis, 2023), Grafana (Alonso et al., 2019; Hossain et al., 2023; Liu and Qin, 2024; Tundo et al., 2024; Roy and Kalafatis, 2023), Telegraf (Alonso et al., 2019), AWS CloudWatch (Alhamazani et al., 2014), cAdvisor (Wong et al., 2019), Ceilometer (del Castillo et al., 2013; Sitaram et al., 2018), Gnocchi (Sitaram et al., 2018), ELK stack (Mfula and Nurminen, 2018), Terraform (Alonso et al., 2019; Gupta et al., 2021; da Silva et al., 2023; Pandi et al., 2023; Harzenetter et al., 2023), CloudBees (Pham et al., 2015), Jenkins (Pandey et al., 2022; Solayman and Qasha, 2023), Puppet (Pham et al., 2015), Chef (Pham et al., 2015; Sheridan et al., 2017), GitHub (Sheridan et al., 2017), Shell scripts (Rios et al., 2017), Maven (Pham et al., 2015), NPM (Pham et al., 2015), Consul (Pham et al., 2015)
+Development, APIs, and Cloud Storage Public cloud storage (Harwalkar et al., 2019a), Private cloud storage (Harwalkar et al., 2019a), Hybrid cloud storage (Harwalkar et al., 2019a), Google Cloud Storage (Ramesh et al., 2023), Azure Blob Storage (Ramesh et al., 2023), REST API (Harwalkar et al., 2019a), YAML (Serhiienko and Spillner, 2018), OpenVAS Vulnerability Scanner (Sheridan et al., 2017), DAX format (Tang, 2021), Java (Sheridan et al., 2017; Rios et al., 2017; Carpio et al., 2020; Bucur et al., 2018; Quint and Kratzke, 2018; Challita et al., 2017), Go programming language (Al-shammari and Alwan, 2018; Lorenzo et al., 2019), Python (Carpio et al., 2020; Bucur et al., 2018; del Castillo et al., 2013; Sitaram et al., 2018; Serhiienko and Spillner, 2018), Ruby (Bucur et al., 2018), PHP (Bucur et al., 2018), C# (Bucur et al., 2018), Cloud Modelling Framework (CLOUMDF) (Ferry et al., 2014), Cloud Modelling Language (CLOUDML) (Ferry et al., 2014; Challita et al., 2017), CAMEL modeling language (Casola et al., 2017; Quint and Kratzke, 2018; Kritikos et al., 2019), Eclipse IDE (Alonso et al., 2019), MCSLA editor (Alonso et al., 2019)
 […]
@@
 Furthermore, containers contribute to enhancing security in multi-cloud operations.
-For researchers, this variety of roles signifies a rich area for inquiry, with opportunities to investigate potential security vulnerabilities, management complexities, and performance overheads.
-For practitioners, understanding these roles can inform efficient and effective deployment strategies, leading to more robust multi-cloud implementations.
+Implications: Based on these findings, researchers have opportunities to investigate further areas, such as potential security vulnerabilities, management complexities, and performance overheads.
+Similarly, practitioners can use this understanding to inform efficient and effective deployment strategies, leading to more robust multi-cloud implementations.
 Container Technology Features and their Significance: The unique features of container technology, such as lightweight virtualization, compatibility with orchestration tools, and application in edge computing, offer significant advantages.
-For researchers, these features present opportunities for further exploration, including how these characteristics (e.g., lightweight virtualization) influence container performance and effectiveness in multi-cloud operations.
-For practitioners, understanding these features can lead to better utilization of container technology and more efficient multi-cloud deployments.
+Implications: Researchers can explore these features further, including how characteristics like lightweight virtualization influence container performance and effectiveness in multi-cloud operations.
+Practitioners can leverage this understanding for better utilization of container technology and more efficient multi-cloud deployments.
 Implementation Strategies and their Real-World Implications: The strategic implementation of containers significantly affects their functionality in multi-cloud environment.
 Strategies related to deployment, orchestration, security, performance optimization, cloud service management, and data storage can have significant implications.
-For researchers, these strategies present an area for exploration, including investigation into their real-world implications, benefits, and limitations.
-For practitioners, a refined understanding of these strategies can lead to more efficient implementations and operations.
+Implications: Researchers may explore these strategies further, including investigations into their real-world implications, benefits, and limitations.
+Practitioners can refine their understanding of these strategies to achieve more efficient implementations and operations.
 Containerization’s Impact on IoT and Application Development: The incorporation of containers into IoT and app development has transform these fields.
 Containers offer improved software unit encapsulation, portability, and efficiency, which are particularly important in the distributed and resource-constrained nature of IoT.
-Researchers can further explore potential challenges in detail such as network complexity, security, and the small footprint requirement of IoT devices.
+Implications: Researchers can further explore potential challenges in detail such as network complexity, security, and the small footprint requirement of IoT devices.
 For practitioners, understanding the impact of containerization can lead to the development of more robust, efficient, and secure IoT and app solutions.
@@
 Containerized applications in multi-cloud environment requires an understanding of prevailing patterns and strategies.
-In this context, we discuss the key findings related identified patterns and strategies presented in Table 9.
+In this context, we discuss the key findings related to the identified patterns and strategies presented in Table 9.
 Diverse Architectural Patterns and Communication Strategies: Our study found that Microservice Architecture and Service-Oriented Architecture (SOA) are the most common used architectural patterns for designing container-based applications in multi-cloud environment.
@@
 The other patterns frequently reported are Service Mesh and Service Chaining, which are employed as communication and networking patterns, likely due to their ability to manage complex, distributed service-to-service communications in a scalable way (Zhamak, 2018).
-The frequent use of Blue-Green Deployment and Object Store Service deployment patterns has also been identified in this study, which may stem from their ability to minimize downtime and ensure high availability—essential factors in a cloud-based setting (Fowler, 2010).
+The frequent use of black-Green Deployment and Object Store Service deployment patterns has also been identified in this study, which may stem from their ability to minimize downtime and ensure high availability—essential factors in a cloud-based setting (fowler2010blackgreen).
 Implications: Practitioners can enhance the design, scalability, and efficiency of their applications in a multi-cloud environment by understanding and applying these architectural and communication patterns.
@@
 4.4. Automation Challenges and Solution Framework (RQ4 - RQ5)
-The Automation Challenge-Solution Framework for Containerized Applications in multi-cloud environment (see Table 12) suggests that multi-cloud management for containerized application is a complex task encompassing several dimensions.
+The Automation Challenge-Solution Framework for Containerized Applications in multi-cloud environment (see Table LABEL:tab:Automation-solutions) suggests that multi-cloud management for containerized application is a complex task encompassing several dimensions.
 These include multi-cloud automation, deployment and scaling, resource management, data and application migration, testing and bench marking, standardization and interoperability, application and service management, and runtime and service discovery.
@@
 Importance of Standardization and Interoperability: The framework underscores the crucial role of standardization and interoperability in multi-cloud environment, facilitating seamless integration and management across different cloud platforms.
-Recent literature validates the importance of standardization and interoperability in multi-cloud environment, emphasizing their role in enhancing the efficiency of cloud services (Alonso et al., 2023a).
+Recent literature validates the importance of standardization and interoperability in multi-cloud environment, emphasizing their role in enhancing the efficiency of cloud services (Alonso et al., 2023b).
 This agrees with our results, further confirming the critical need for these factors in the domain.
@@
 The framework proposed in our study provides a comprehensive and systematic approach to addressing the complex security issues associated with containerized applications in multi-cloud environment.
-In the following, we discuss the key takeaway from security challenges and solution framwrork presented in Section 3.5 and Table 11.
+In the following, we discuss the key takeaway from security challenges and solution framework presented in Section 3.5 and Table LABEL:tab:Security-challenges-solutions.
 Emphasize on Container-Specific Security Mechanisms: Security solutions specifically designed for containerized applications are critical to managing threats in multi-cloud environment.
@@
 4.6. Deployment Challenges and Solutions Framework (RQ4 - RQ5)
-Table 14 offers a comprehensive overview of challenges and their corresponding solutions encountered during the deployment of containerized applications in multi-cloud environment.
+Table LABEL:tab:monitoring-solutions offers a comprehensive overview of challenges and their corresponding solutions encountered during the deployment of containerized applications in multi-cloud environment.
 In the following, we provide a discussion on key findings pertaining to the challenges and solutions related to the deployment of containerized applications in multi-cloud environment.
@@
 underscore that the labyrinth of diverse platforms and differing standards in multi-cloud configurations amplifies security complexities (Fernandes et al., 2014).
-Challenges articulated as Security Deployment in Multi-container Environment and Malicious Service Management Deployment in Table 13 accentuate the primacy of security considerations.
+Challenges articulated as Security Deployment in Multi-container Environment and Malicious Service Management Deployment in Table LABEL:tab:Deployment-solutions accentuate the primacy of security considerations.
 Propounded solutions like Cross-Container Isolation and Collusion-Resilient Trust Aggregation Technique spotlight the industry’s shift towards advanced and meticulous security protocols within multi-cloud paradigms.
@@
 Implications: For the academic community, pioneering avant-garde security solutions apt for multi-cloud environment offers a fertile domain of inquiry.
-Meanwhile, practitioners should enshrine principles of data integrity, trust assurance, and vulnerability mitigation at the heart of their multi-cloud deployment blueprints.
+Meanwhile, practitioners should enshrine principles of data integrity, trust assurance, and vulnerability mitigation at the heart of their multi-cloud deployment blackprints.
 Integration, Monitoring, and Infrastructure Management: As highlighted by Ferrer et al.
@@
 Proposed solutions such as VM Selection for Composite Applications and Interoperability Layer Above Cloud Infrastructure accentuate the importance of crafting adaptable infrastructure and formulating strategies to mitigate vendor entrenchment.
-In the context of multi-cloud environment, flexibility and interoperability transcend mere advantageous attributes; they become pivotal determinants shaping adaptability, scalability, and the sustained success of deployment blueprints.
+In the context of multi-cloud environment, flexibility and interoperability transcend mere advantageous attributes; they become pivotal determinants shaping adaptability, scalability, and the sustained success of deployment blackprints.
 Implications: For the academic community, delving into novel approaches that champion infrastructure pliability and seamless operation across diverse cloud platforms emerges as a prospective research avenue.
@@
 4.7. Monitoring Challenges and Solutions (RQ4 - RQ5)
-Table 14 presents a comprehensive view of the Monitoring Challenge-Solution Framework for Containerized Applications in multi-cloud environment.
+Table LABEL:tab:monitoring-solutions presents a comprehensive view of the Monitoring Challenge-Solution Framework for Containerized Applications in multi-cloud environment.
 It covers a wide range of monitoring challenges within the intricate multi-cloud setting.
@@
 Strengthening Security, Compliance, and Trust: Ensuring data privacy, continuous security control, and compliance in multi-cloud containerized applications is essential but challenging due to the distributed nature of cloud resources.
-Existing studies (e.g., (Alonso et al., 2023a) (Raj et al., 2018)) discussed the challenges of ensuring security and compliance across multi-cloud environment.
+Existing studies (e.g., (Alonso et al., 2023b) (Raj et al., 2018)) discussed the challenges of ensuring security and compliance across multi-cloud environment.
 They highlighted the importance of continuous monitoring and enforcement mechanisms to address security and compliance gaps.
@@
 Concerns about external validity relate to the applicability and generalizability of study findings.
-Our research provides a thorough examination of containerization in multi-cloud environments, with findings, analyses, and conclusions specifically tailored to this domain.
+Our research provides a thorough examination of containerization in multi-cloud environment, with findings, analyses, and conclusions specifically tailored to this domain.
 To ensure robust external validity, we developed a study protocol that outlines our research methodology.
@@
 6. Related Work
-This section reviews the most relevant existing research in terms of secondary studies such as literature surveys, state-of-the-art analysis, systematic reviews, and mapping studies that consolidate published literature on containerization in multi-cloud environments.
+This section reviews the most relevant existing research in terms of secondary studies such as literature surveys, state-of-the-art analysis, systematic reviews, and mapping studies that consolidate published literature on containerization in multi-cloud environment.
 The review focuses on (i) existing challenges, proposed solutions, and emerging trends detailed in Section 6.1 and (ii) container-based deployment of multi-cloud systems in Section 6.2.
@@
 6.1.1. Challenges, Applications, Tools, and Emerging Trends
-Containers in multi-cloud environments provide a standardized and portable way to package, deploy, and run applications, ensuring seamless deployment and management across diverse cloud platforms (Alonso et al., 2023b) (Saxena et al., 2021).
+Containers in multi-cloud environment provide a standardized and portable way to package, deploy, and run applications, ensuring seamless deployment and management across diverse cloud platforms (Alonso et al., 2023c) (Saxena et al., 2021).
 One of the earliest SLR-based studies by Pahl et al.
@@
 SLR results indicate container orchestration, microservice delivery, continuous development and deployment as emerging trends of research on containerized clouds.
-The SLR in (Alonso et al., 2023b) reviewed a total of 88 studies (published from 2011 - 2021) on multi-cloud systems.
-Similar survey studies such as (Saxena et al., 2021) concepts, challenges, requirements and future directions for multi-cloud environments are discussed.
+The SLR in (Alonso et al., 2023c) reviewed a total of 88 studies (published from 2011 - 2021) on multi-cloud systems.
+Similar survey studies such as (Saxena et al., 2021) concepts, challenges, requirements and future directions for multi-cloud environment are discussed.
 A survey of existing approaches and solutions provided by different multi-cloud architectures is entailed along with analysis of the pros and cons of different architectures while comparing the same (Imran et al., 2020).
@@
 Carlos et al.
-(Awada, 2018) present an optimization strategy for deploying microservices in multi-cloud environments focusing on minimizing cloud service costs, network latency, and startup time for new microservices.
+(Awada, 2018) present an optimization strategy for deploying microservices in multi-cloud environment focusing on minimizing cloud service costs, network latency, and startup time for new microservices.
 Their approach uses a Non-dominated Sorting Genetic Algorithm II (NSGA-II) compared to a Greedy First-Fit algorithm, demonstrating a 300% improvement with NSGA-II.
@@
 Their work emphasizes the importance of container technologies in cloud environments and the need for advanced orchestration solutions to manage complex multi-container applications effectively.
-Uchechukwu Awada (Guerrero et al., 2018a) reviews container orchestration tools and platforms, comparing the architectures, components, and capabilities of several Container Service Platforms (CSPs) and orchestration tools like Amazon ECS, Kubernetes, Docker Swarm, and Mesos.
+Uchechukwu Awada (Guerrero et al., 2018b) reviews container orchestration tools and platforms, comparing the architectures, components, and capabilities of several Container Service Platforms (CSPs) and orchestration tools like Amazon ECS, Kubernetes, Docker Swarm, and Mesos.
 The study offers insights into the current state of container orchestration and suggests future research directions, serving as a guide for developers and organizations.
@@
 Hendrik et al.
-(Graupner et al., 2015a) tackle privacy, security, and trust issues in cloud computing through a multi-cloud architecture perspective.
+(Graupner et al., 2015b) tackle privacy, security, and trust issues in cloud computing through a multi-cloud architecture perspective.
 They propose a novel technique for enhancing security and unifying access control mechanisms across cloud providers, addressing the challenges of inadequate cross-provider APIs and non-unified access control.
@@
 These studies investigated various aspects of multi-cloud, including exploring applications, tools, and technologies, assessing potential and limitations, and identifying emerging trends.
+Table 16 presents an overview of the differences between this study and existing secondary studies.
+The symbols used in the table are as follows:  indicates “Yes”,  indicates “No”, and $\triangle$ indicates “Partially”.
+After thoroughly comparing our results with existing studies (e.g., (Pahl et al., 2017), (Casalicchio and Iannucci, 2020), (Netaji and Bhole, 2021)), it is clear that none of the results are fully identical with our findings.
 Our study distinguishes itself from prior research in several ways.
-First, our work systematically categorizes container utilization in multi-cloud environments into coherent categories and themes.
-Some of the key themes are ‘Scalability and High Availability’, ‘Performance and Optimization’, ‘Security and Privacy’, and ‘Multi-Cloud Container Monitoring and Adaptation’.
+First, our work systematically categorizes container utilization in multi-cloud environment into coherent categories and themes.
+Some of the key themes are Scalability and High Availability, Performance and Optimization, Security and Privacy, and Multi-Cloud Container Monitoring and Adaptation.
 We identified and classified seventy-four patterns and strategies for containerization, which existing studies have not provided.
@@
 Lastly, our study introduces four distinct frameworks consisting of challenges and solutions related to security, automation, deployment, and monitoring within the context of containerization in multi-cloud environment.
+Table 16.
+Comparison of This SLR Results with Existing Secondary Studies.
+MS indicate “Mix Study”, LR indicate “Literature Review”  indicates “Yes”,  indicates “No,” and $\triangle$ indicates “Partially.”
+Contributions This Study (Pahl et al., 2017) (Casalicchio and Iannucci, 2020) (Netaji and Bhole, 2021) (Bentaleb et al., 2022) (Watada et al., 2019) (Saxena et al., 2021) (Imran et al., 2020) (Casalicchio, 2019)
+Number of selected studies 121 46 97 64 40 15 11
+Study Types SMS SMS LR LR LR MS LR LR MS
+Container Roles (Section 3.2.1) $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
+Container Implementation Strategies (Section 3.2.2) $\triangle$ $\triangle$ $\triangle$
+Pattern and Strategies (Section 3.3) $\triangle$ $\triangle$
+Quality Attributes and Tactics (Section 3.4) $\triangle$ $\triangle$
+Security Challenges and Solution (Section 3.5) $\triangle$ $\triangle$ $\triangle$
+Automation Challenges and Solution (Section 3.6) $\triangle$ $\triangle$ $\triangle$
+Deployment Challenges and Solution (Section 3.7) $\triangle$
+Monitoring Challenges and Solution (Section 3.8) $\triangle$
+Tools and Frameworks (Section 3.9) $\triangle$ $\triangle$ $\triangle$
 7. Conclusions
@@
 Additionally, this SMS also explores deployment, monitoring, security challenges, accompanied by their respective solutions, tools, and frameworks used to implement containerized applications in multi-cloud environment.
-From an initial pool of extensive search results, 86 relevant studies are were finally chosen to answer the research questions presented in Table 1.
+From an initial pool of extensive search results, 86 relevant studies were finally chosen to answer the research questions presented in Table 1.
 The key findings of this SMS are listed below.
@@
 The yearly distribution reached its peak in 2016, signifying a shifting level of interest over time.
-The field of publishing is largely dominated by IEEE (72.09%), with conferences emerging as the primary avenue for publication.
-Notably, academia (93.02%) stands out as the leading affiliation among authors.
+The field of publishing is largely dominated by IEEE (55.37%), with conferences emerging as the primary avenue for publication.
+Notably, academia (75.02%) stands out as the leading affiliation among authors.
 The prominent research themes encompass Orchestration and Management, Scalability and High Availability, Performance and Optimization, and Security and Privacy.
 These themes underscore crucial aspects such as container management, scalability, performance, and security within the context of multi-cloud environment.
-  • This SMS identifies and classifies a total of 74 patterns and strategies across 10 subcategories and 4 categories for container-based applications in multi-cloud environment.
+  • This SMS identifies and classifies a total of 98 patterns and strategies across 10 subcategories and 4 categories for container-based applications in multi-cloud environment.
 These patterns and strategies encompass a wide range of aspects such as cloud architecture models, communication and networking, deployment, service models, cloud management, container management, edge computing, IoT strategies, security, resilience, fault-tolerance, and cloud migration.
@@
 This framework offers valuable insights and actionable strategies for practitioners and researchers to enhance the security of containerized applications, address complex multi-cloud security concerns, and ensure compliance with various regulatory requirements, thereby enabling the development of more robust and secure containerized applications in multi-cloud environment.
-  • This SMS presents a framework for the automation challenge-solution framework for containerized applications in multi-cloud environment.
-The automation challenge-solution framework offers a comprehensive guide for practitioners and researchers to address the complex challenges associated with automating various aspects of deploying, managing, scaling, testing, migrating, and ensuring interoperability of containerized applications in multi-cloud environment.
-By offering specific solutions for each challenge, this framework aids in the development of more efficient, secure, and adaptable automation processes, empowering practitioners to streamline their operations and enabling researchers to advance the field of multi-cloud containerization.
-  • The third framework presented in this SMS is the deployment challenge-solution framework for containerized applications in multi-cloud environment.
-This framework provides valuable insights and practical guidance to practitioners and researchers alike, aiding practitioners in navigating the complexities of multi-cloud deployment through specific solutions to challenges, while also serving as a comprehensive resource for researchers to delve into various aspects of containerized application deployment, orchestration, security, scalability, and more, thus contributing to the advancement of knowledge and innovation in the field of multi-cloud deployment.
-  • Fourth framework presented in this SMS is monitoring challenge-solution framework for containerized applications in multi-cloud environment.
-This framework serves as a comprehensive guide for practitioners, offering specific solutions to challenges related to monitoring performance, resource management, security, optimization, system complexity, and multi-cloud coordination, facilitating their understanding of complex monitoring scenarios and providing actionable insights.
-Furthermore, researchers can benefit from this framework by gaining a deeper insight into the nuances of monitoring containerized applications in multi-cloud environment, and it can serve as a foundation for further research and innovation in the field of multi-cloud monitoring.
+  • This SMS presents a catalogs for the automation challenge-solution catalogs for containerized applications in multi-cloud environment.
+The automation challenge-solution catalog offers a comprehensive guide for practitioners and researchers to address the complex challenges associated with automating various aspects of deploying, managing, scaling, testing, migrating, and ensuring interoperability of containerized applications in multi-cloud environment.
+By offering specific solutions for each challenge, this catalog aids in the development of more efficient, secure, and adaptable automation processes, empowering practitioners to streamline their operations and enabling researchers to advance the field of multi-cloud containerization.
+  • The third catalog presented in this SMS is the deployment challenge-solution catalog for containerized applications in multi-cloud environment.
+This catalog provides valuable insights and practical guidance to practitioners and researchers alike, aiding practitioners in navigating the complexities of multi-cloud deployment through specific solutions to challenges, while also serving as a comprehensive resource for researchers to delve into various aspects of containerized application deployment, orchestration, security, scalability, and more, thus contributing to the advancement of knowledge and innovation in the field of multi-cloud deployment.
+  • Fourth catalog presented in this SMS is monitoring challenge-solution catalog for containerized applications in multi-cloud environment.
+This catalog serves as a comprehensive guide for practitioners, offering specific solutions to challenges related to monitoring performance, resource management, security, optimization, system complexity, and multi-cloud coordination, facilitating their understanding of complex monitoring scenarios and providing actionable insights.
+Furthermore, researchers can benefit from this catalog by gaining a deeper insight into the nuances of monitoring containerized applications in multi-cloud environment, and it can serve as a foundation for further research and innovation in the field of multi-cloud monitoring.
 The findings of this SMS will benefit researchers who are interested in understanding the state of research on containerized applications in multi-cloud environment and conducting further investigations to address the open research issues highlighted in Section 4.
@@
 We emphasize the importance for practitioners to develop targeted solutions that effectively tackle monitoring, security, and performance degradation concerns within multi-cloud environment deployment.
-As a future endeavor, we intend to enhance our SMS by conducting industrial case studies with companies, thereby obtaining practical insights and perspectives from practitioners on the effectiveness and applicability of our proposed frameworks in real-world scenarios.
+As a future endeavor, we intend to enhance our SMS by conducting industrial case studies with companies, thereby obtaining practical insights and perspectives from practitioners on the effectiveness and applicability of our proposed catalogs.
 This approach will allow us to bridge the gap between research and practical implementations concerning containerized applications in multi-cloud environment and contribute to the advancement of both academic and industry understanding in this domain.
+Data Availability
+The dataset of this study has been provided online: https://zenodo.org/record/10732611.
 Acknowledgments
-This work is partially sponsored by the National Natural Science Foundation of China (NSFC) with Grant No. 62172311.
+This research is funded by Business Finland through QLEAP (2022-24) project, and by the NSFC China under Grant No. 62172311.
 8. Appendices
-List of selected studies for containerization in multi-cloud environments
+{xltabular}
+—c—p4.5cm—p4.5cm—c—c— List of selected studies for containerization in multi-cloud environments
 Study ID Title Venue Year Reference
@@
 […]
+Continued on next page
 […]
-A component-based adaptation approach for multi-cloud applications
-IEEE Conference on Computer Communications (INFOCOM) Workshops
-2014 (Almeida et al., 2014)
-S02
-A Cost-Efficient Multi-cloud Orchestrator for Benchmarking Containerized Web-Applications
-International Conference on Web Information Systems Engineering (WISE)
-2019 (Jha et al., 2019)
-S03
-A framework to support multi-cloud collaboration
-IEEE World Congress on Services (SERVICES)
-2020 (Hua et al., 2020)
-S04
-A Machine Learning Model for Detection of Docker-based APP Overbooking on Kubernetes
-International Conference on Communications (ICC)
-2021 (Ramos et al., 2021)
-S05
-A multi-cloud service mesh approach applied to Internet of Things
-Annual Conference of Industrial Electronics Society (IECON)
-2022 (Gattobigio et al., 2022)
-S06
-A Secure Container Placement Strategy Using Deep Reinforcement Learning in Cloud
-International Conference on Computer Supported Cooperative Work in Design (CSCWD)
-2022 (Deng et al., 2022)
-S07
-A Service Driver Based Application Execution and Development Method in Multi-cloud Context
-International Conference on Data Science and Computer Application (ICDSCA)
-2021 (Zhang et al., 2021)
-S08
-An adaptable framework to deploy complex applications onto multi-cloud platforms
-International Conference on Computing Communication Technologies - Research, Innovation, and Vision for Future (RIVF)
-2015 (Pham et al., 2015)
-S09
-An agile container-based approach to TaaS
-International Congress Future Telecommunications: Infrastructure and Sustainability (FITCE)
-2016 (Verdugo et al., 2017)
-S10
-An Efficient and Reliable Multi-Cloud Provider Monitoring Solution
-Conference on Global Communications (GLOBECOM)
-2022 (Sabbioni et al., 2020)
-S11
-An Efficient Multi-Keyword Search Scheme over Encrypted Data in Multi-Cloud Environment
-International Conference on Smart Cloud (SmartCloud)
-2015 (He et al., 2022)
-S12
-An Empirical Investigation of a Fault Tolerant Containerized Application Deployment
-International Conference on Informatics (ICI)
-2015 (Bisht and Kaur, 2022)
-S13
-Applying Computational Intelligence for enhancing the dependability of multi-cloud systems using Docker Swarm
-Symposium Series on Computational Intelligence (SSCI)
-2017 (Naik, 2016a)
-S14
-Automatically Design Heuristics for Multi-Objective Location-Aware Service Brokering in Multi-Cloud
-International Conference on Services Computing (SCC)
-2020 (Chen et al., 2022)
-S15
-Autonomic fine-grained migration and replication of component-based applications across multi-clouds
-National Foundation for Science and Technology Development Conference on Information and Computer Science (NICS)
-2022 (Pham and Pham, 2015)
-S16
-Band-Area Resource Management Platform and Accelerated Particle Swarm Optimization Alg. for Container Deployment
-IEEE Access
-2016 (Ouyang et al., 2022)
-S17
-Building a Multi-tenant Cloud Service from Legacy Code with Docker Containers
-International Conference on Cloud Engineering (IC2E)
-2022 (Slominski et al., 2015)
-S18
-Building cloud native application - Analysis for multi-component application deployment
-International Conference on Computer Communication and Informatics (ICCCI)
-2021 (Pai and Kumar, 2021)
-S19
-CLAMS: Cross-layer Multi-cloud Application Monitoring-as-a-Service Framework
-International Conference on Services Computing (SCC)
-2014 (Alhamazani et al., 2014)
-S20
-Cloud MF: Applying MDE to Tame the Complexity of Managing Multi-Cloud Applications
-International Conference on Utility and Cloud Computing (UCC)
-2014 (Ferry et al., 2014)
-S21
-Cluster as a Service: A Container Based Cluster Sharing Approach with Multi-user Support
-IEEE Symposium on Service-Oriented System Engineering (SOSE)
-2016 (Cui et al., 2016)
-S22
-Container Deployment Strategy for Edge Networking
-Workshop on Middleware for Edge Clouds and Cloudlets (MECC)
-2019 (Wong et al., 2019)
-S23
-Container-Based Performance Isolation for Multi-Tenant SaaS Applications in Micro-Service Architecture
-Journal of Physics: Conference Series
-2020 (Wang et al., 2020)
-S24
-Containerization and the PaaS Cloud
-IEEE Cloud Computing
-2015 (Pahl, 2015)
-S25
-ContainerLeaks: Emerging Security Threats of Information Leakages in Container Clouds
-International Conference on Dependable Systems and Networks (DSN)
-2017 (Gao et al., 2017)
-S26
-ContMEC: An Architecture of Multi-access Edge Computing for Offloading Container-Based Mobile Applications
-IEEE International Conference on Communications (ICC)
-2022 (Watanabe et al., 2022)
-S27
-DECIDE: An Extended DevOps Framework for Multi-Cloud Applications
-International Conference on Cloud and Big Data Computing (ICCBDC)
-2022 (Alonso et al., 2019)
-S28
-Delivering Elastic Containerized Cloud Applications to Enable DevOps
-International Workshop on Software Engineering for Adaptive and Self-Managing Systems (SEAMS)
-2017 (Barna et al., 2017)
-S29
-Deployment of Multi-Tier Application on Cloud and Continuous Monitoring using Kubernetes
-International Conference on Electrical, Electronics, Communication, Computer Technologies and Optimization Techniques (ICEECCOT)
-2021 (Gupta et al., 2021)
-S30
-Deployment-Time Multi-Cloud Application Security
-International Conference on Smart Computing (SMARTCOMP)
-2017 (Sheridan et al., 2017)
-S31
-DEVSML 3.0 Stack: Rapid deployment of DEVS Farm in distributed cloud environment using microservices and containers
-Symposium on Theory of Modeling and Simulation (TMS/DEVS)
-2017 (Mittal and Risco-Martín, 2017)
-S32
-Disaster Recovery and Business Continuity for Database Services in Multi-Cloud
-International Conference on Computer Applications & Information Security (ICCAIS)
-2018 (Al-shammari and Alwan, 2018)
-S33
-Dynamic On-Demand Fog Formation Offering On-the-Fly IoT Service Deployment
-IEEE Transactions on Network and Service Management
-2020 (Sami and Mourad, 2020)
-S34
-Dynamic security assurance in multi-cloud DevOps
-IEEE Conference on Communications and Network Security (CNS)
-2017 (Rios et al., 2017)
-S35
-Engineering and Experimentally Benchmarking a Container-based Edge Computing System
-International Conference on Communications (ICC)
-2020 (Carpio et al., 2020)
-S36
-FogDocker: Start Container Now, Fetch Image Later
-International Conference on Utility and Cloud Computing (UCC)
-2019 (Lorenzo et al., 2019)
-S37
-Importance of Application-Level Resource Management in Multi-Cloud Deployments
-International Conference on Cloud Engineering (IC2E)
-2019 (Dimitrijevic et al., 2019)
-S38
-Knowledge-Engineered Multi-Cloud Resource Brokering for Application Workflow Optimization
-IEEE Transactions on Network and Service Management
-2022 (Pandey et al., 2022)
-S39
-Label-affinity-Scheduler: Considering Business Requirements in Container Scheduling for Multi-Cloud and Multi-Tenant
-Brazilian Symposium on Computing System Engineering (SBESC)
-2022 (Altran et al., 2022)
-S40
-Leveraging kernel security mechanisms to improve container security: A survey
-International Conference on Availability, Reliability and Security (ARES)
-2019 (Bélair et al., 2019)
-S41
-Location-Aware and Budget-Constrained Service Deployment for Composite Applications in Multi-Cloud Environment
-IEEE Transactions on Parallel and Distributed Systems
-2020 (Shi et al., 2020)
-S42
-MicroCloud: A Container-Based Solution for Efficient Resource Management in the Cloud
-International Conference on Smart Cloud (SmartCloud)
-2016 (Baresi et al., 2016)
-S43
-Mobile Edge as Part of the Multi-Cloud Ecosystem: A Performance Study
-Euromicro International Conference on Parallel, Distributed and Network-Based Processing (PDP)
-2019 (Dreibholz et al., 2019)
-S44
-Modeling and Automated Deployment of Serverless Applications Using TOSCA
-Conference on Service-Oriented Computing and Applications (SOCA)
-2018 (Wurster et al., 2018)
-S45
-Monitoring Elastically Adaptive Multi-Cloud Services
-IEEE Transactions on Cloud Computing
-2018 (Trihinas et al., 2015)
-S46
-Monitoring-Aware Optimal Deployment for Applications Based on Microservices
-IEEE Transactions on Services Computing
-2021 (Fadda et al., 2019)
-S47
-MSSF: A Step towards User-Friendly Multi-cloud Data Dispersal
-International Conference on Cloud Computing (CLOUD)
-2014 (Libardi et al., 2014)
-S48
-Multi-agent QoS-aware autonomic resource provisioning framework for elastic BPM in containerized multi-cloud environment
-Journal of Ambient Intelligence and Humanized Computing
-2022 (Saif et al., 2023)
-S49
-Multicloud- auto scale with prediction and delta correction algorithm
-International Conference on Intelligent Computing, Instrumentation and Control Technologies (ICICICT)
-2019 (Harwalkar et al., 2019b)
-50
-Multi-cloud Distribution of Virtual Functions and Dynamic Service Deployment: Open ADN Perspective
-IEEE International Conference on Cloud Engineering (IC2E)
-2015 (Bhamare et al., 2015)
-S51
-Multi-cloud resource provisioning with Aneka: A unified and integrated utilization of Microsoft Azure and Amazon EC2
-International Conference on Computing and Network Communications (CoCoNet)
-2015 (Buyya and Barreto, 2015)
-S52
-MUSA Deployer: Deployment of Multi-cloud Applications
-IEEE International Workshop on Enabling Technologies: Infrastructure for Collaborative Enterprises (WETICE)
-2017 (Casola et al., 2017)
-S53
-NoMISHAP: A Novel Middleware Support for High Availability in Multicloud PaaS
-IEEE Cloud Computing
-2017 (Acquaviva et al., 2017)
-S54
-Object storage in the cloud and multi-cloud: State of the art and the research challenges
-IEEE International Conference on Automation, Quality and Testing, Robotics (AQTR)
-2018 (Bucur et al., 2018)
-S55
-On the Deployment of an Open-Source, 5G-Aware Evaluation Testbed
-International Conference on Mobile Cloud Computing, Services, and Engineering (MobileCloud)
-2018 (Bolivar et al., 2018)
-S56
-OpenStack Federation in Experimentation Multi-cloud Testbeds
-IEEE International Conference on Cloud Computing Technology and Science (CloudCom)
-2013 (del Castillo et al., 2013)
-S57
-Orchestrated Multi-Cloud Application Deployment in OpenStack with TOSCA
-International Conference on Smart Computing (SMARTCOMP)
-2017 (Tricomi et al., 2017)
-S58
-Orchestration Based Hybrid or Multi Clouds and Interoperability Standardization
-IEEE International Conference on Cloud Computing in Emerging Markets (CCEM)
-2018 (Sitaram et al., 2018)
-S59
-Performance Analysis of Zero-Trust multi-cloud
-IEEE International Conference on Cloud Computing (CLOUD)
-2021 (Rodigari et al., 2021)
-S60
-Policies Based Container Migration Using Cross-Cloud Management Platform
-International Conference on Information and Automation for Sustainability (ICIAfS)
-2018 (Janarthanan et al., 2018)
-S61
-Policy-based deployment in a hybrid and multi-cloud environment
-International Conference on Cloud Computing and Services Science (CLOSER)
-2019 (Di Modica et al., 2019)
-S62
-Predicting the End-to-End Tail Latency of Containerized Microservices in the Cloud
-International Conference on Cloud Engineering (IC2E)
-2019 (Rahman and Lama, 2019)
-S63
-Preventing vendor lock-ins via an interoperable multi-cloud deployment approach
-International Conference for Internet Technology and Secured Transactions (ICITST)
-2017 (Pellegrini et al., 2017)
-S64
-Private STaaS with OpenStack Cinder Volumes for Hybrid/Multi-cloud
-International Conference on Advances in Computing and Communication Engineering (ICACCE)
-2019 (Harwalkar et al., 2019a)
-S65
-Real-World Experience with a Multicloud Exchange
-IEEE Cloud Computing
-2017 (Zeck and Bouroudjian, 2017)
-S66
-Reliability Management for Blockchain-Based Decentralized Multi-Cloud
-International Symposium on Cluster Computing and the Grid (CCGRID)
-2020 (Aral et al., 2020)
-S67
-Reliability-Aware Cost-Efficient Scientific Workflows Scheduling Strategy on Multi-Cloud Systems
-IEEE Transactions on Cloud Computing
-2022 (Tang, 2021)
-S68
-Resource optimization of container orchestration: a case study in multi-cloud microservices-based applications
-The Journal of Supercomputing
-2018 (Guerrero et al., 2018b)
-S69
-Scalable multi-cloud platform to support industry and scientific applications
-International Convention on Information and Communication Technology, Electronics and Microelectronics (MIPRO)
-2018 (Kiss, 2018)
-S70
-Search-based Methods for Multi-Cloud Configuration
-International Conference on Cloud Computing (CLOUD)
-2022 (Lazuka et al., 2022)
-S71
-Secure access control for multi-cloud resources
-IEEE Conference on ALocal Computer Networks Workshops (LCN Workshops)
-2015 (Graupner et al., 2015b)
-S72
-Secure and Dependable Multi-Cloud Network Virtualization
-International Workshop on Security and Dependability of Multi-Domain Infrastructures
-2017 (Alaluna et al., 2017)
-S73
-Secure and Parallel Expressive Search over Encrypted Data with Access Control in Multi-CloudIoT
-Cloudification of the Internet of Things (CIoT)
-2018 (Guechi and Maamri, [n. d.])
-S74
-Self-healing cloud services in private multi-clouds
-International Conference on High Performance Computing & Simulation (HPCS)
-2018 (Mfula and Nurminen, 2018)
-S75
-SpyStorage: A Highly Reliable Multi-Cloud Storage with Secure and Anonymous Data Sharing
-International Conference on Networking, Architecture, and Storage (NAS)
-2017 (Shen et al., 2017)
-S76
-Support Services for Applications Execution in Multi-clouds Environments
-International Conference on Autonomic Computing (ICAC)
-2016 (Pop et al., 2016)
-S77
-SynAPTIC: Secure and persistent connectivity for containers
-International Symposium on Cluster, Cloud and Grid Computing (CCGRID)
-2017 (Ranjbar et al., 2017)
-S78
-System Restore in a Multi-cloud Data Pipeline Platform
-International Conference on Dependable Systems and Networks – Industry Track
-2019 (Wang et al., 2019)
-S79
-Systematic and Recomputable Comparison of Multi-cloud Management Platforms
-International Conference on Cloud Computing Technology and Science (CloudCom)
-2018 (Serhiienko and Spillner, 2018)
-S80
-Towards a lightweight multi-cloud DSL for elastic and transferable cloud-native applications
-International Conference on Cloud Computing and Services Science (CLOSER)
-2018 (Quint and Kratzke, 2018)
-S81
-Towards Formal-Based Semantic Interoperability in Multi-Clouds: The FCLOUDS Framework
-IEEE International Conference on Cloud Computing (CLOUD)
-2017 (Challita et al., 2017)
-S82
-Towards Model-Driven Provisioning, Deployment, Monitoring, and Adaptation of Multi-cloud Systems
-IEEE International Conference on Cloud Computing (CLOUD)
-2013 (Ferry et al., 2013)
-S83
-Towards Trustworthy Multi-Cloud Services Communities: A Trust-Based Hedonic Coalitional Game
-IEEE Transactions on Services Computing
-2018 (Wahab et al., 2016)
-S84
-Multi-cloud provisioning of business processes
-Journal of Cloud Computing
-2019 (Kritikos et al., 2019)
-S85
-Evaluating the DevOps Reference Architecture for Multi-cloud IoT-Applications
-SN Computer Science
-2021 (Bou Ghantous and Gill, 2021)
-S86
-Hybrid IT and Multi Cloud an Emerging Trend and Improved Performance in Cloud Computing
-SN Computer Science
-2020 (Gundu et al., 2020)
+[…]
+S01 A component-based adaptation approach for multi-cloud applications IEEE Conference on Computer Communications (INFOCOM) Workshops 2014 (Almeida et al., 2014)
+S02 A Cost-Efficient Multi-cloud Orchestrator for Benchmarking Containerized Web-Applications International Conference on Web Information Systems Engineering (WISE) 2019 (Jha et al., 2019)
+S03 A framework to support multi-cloud collaboration IEEE World Congress on Services (SERVICES) 2020 (Hua et al., 2020)
+S04 A Machine Learning Model for Detection of Docker-based APP Overbooking on Kubernetes International Conference on Communications (ICC) 2021 (Ramos et al., 2021)
+S05 A multi-cloud service mesh approach applied to Internet of Things Annual Conference of Industrial Electronics Society (IECON) 2022 (Gattobigio et al., 2022)
+S06 A Secure Container Placement Strategy Using Deep Reinforcement Learning in Cloud International Conference on Computer Supported Cooperative Work in Design (CSCWD) 2022 (Deng et al., 2022)
+S07 A Service Driver Based Application Execution and Development Method in Multi-cloud Context International Conference on Data Science and Computer Application (ICDSCA) 2021 (Zhang et al., 2021)
+S08 An adaptable framework to deploy complex applications onto multi-cloud platforms International Conference on Computing Communication Technologies - Research, Innovation, and Vision for Future (RIVF) 2015 (Pham et al., 2015)
+S09 An agile container-based approach to TaaS International Congress Future Telecommunications: Infrastructure and Sustainability (FITCE) 2016 (Verdugo et al., 2017)
+S10 An Efficient and Reliable Multi-Cloud Provider Monitoring Solution Conference on Global Communications (GLOBECOM) 2022 (Sabbioni et al., 2020)
+S11 An Efficient Multi-Keyword Search Scheme over Encrypted Data in Multi-Cloud Environment International Conference on Smart Cloud (SmartCloud) 2015 (He et al., 2022)
+S12 An Empirical Investigation of a Fault Tolerant Containerized Application Deployment International Conference on Informatics (ICI) 2015 (Bisht and Kaur, 2022)
+S13 Applying Computational Intelligence for enhancing the dependability of multi-cloud systems using Docker Swarm Symposium Series on Computational Intelligence (SSCI) 2017 (Naik, 2016a)
+S14 Automatically Design Heuristics for Multi-Objective Location-Aware Service Brokering in Multi-Cloud International Conference on Services Computing (SCC) 2020 (Chen et al., 2022)
+S15 Autonomic fine-grained migration and replication of component-based applications across multi-clouds National Foundation for Science and Technology Development Conference on Information and Computer Science (NICS) 2022 (Pham and Pham, 2015)
+S16 Band-Area Resource Management Platform and Accelerated Particle Swarm Optimization Alg. for Container Deployment IEEE Access 2016 (Ouyang et al., 2022)
+S17 Building a Multi-tenant Cloud Service from Legacy Code with Docker Containers International Conference on Cloud Engineering (IC2E) 2022 (Slominski et al., 2015)
+S18 Building cloud native application - Analysis for multi-component application deployment International Conference on Computer Communication and Informatics (ICCCI) 2021 (Pai and Kumar, 2021)
+S19 CLAMS: Cross-layer Multi-cloud Application Monitoring-as-a-Service Framework International Conference on Services Computing (SCC) 2014 (Alhamazani et al., 2014)
+S20 Cloud MF: Applying MDE to Tame the Complexity of Managing Multi-Cloud Applications International Conference on Utility and Cloud Computing (UCC) 2014 (Ferry et al., 2014)
+S21 Cluster as a Service: A Container Based Cluster Sharing Approach with Multi-user Support IEEE Symposium on Service-Oriented System Engineering (SOSE) 2016 (Cui et al., 2016)
+S22 Container Deployment Strategy for Edge Networking Workshop on Middleware for Edge Clouds and Cloudlets (MECC) 2019 (Wong et al., 2019)
+S23 Container-Based Performance Isolation for Multi-Tenant SaaS Applications in Micro-Service Architecture Journal of Physics: Conference Series 2020 (Wang et al., 2020)
+S24 Containerization and the PaaS Cloud IEEE Cloud Computing 2015 (Pahl, 2015)
+S25 ContainerLeaks: Emerging Security Threats of Information Leakages in Container Clouds International Conference on Dependable Systems and Networks (DSN) 2017 (Gao et al., 2017)
+S26 ContMEC: An Architecture of Multi-access Edge Computing for Offloading Container-Based Mobile Applications IEEE International Conference on Communications (ICC) 2022 (Watanabe et al., 2022)
+S27 DECIDE: An Extended DevOps Framework for Multi-Cloud Applications International Conference on Cloud and Big Data Computing (ICCBDC) 2022 (Alonso et al., 2019)
+S28 Delivering Elastic Containerized Cloud Applications to Enable DevOps International Workshop on Software Engineering for Adaptive and Self-Managing Systems (SEAMS) 2017 (Barna et al., 2017)
+S29 Deployment of Multi-Tier Application on Cloud and Continuous Monitoring using Kubernetes International Conference on Electrical, Electronics, Communication, Computer Technologies and Optimization Techniques (ICEECCOT) 2021 (Gupta et al., 2021)
+S30 Deployment-Time Multi-Cloud Application Security International Conference on Smart Computing (SMARTCOMP) 2017 (Sheridan et al., 2017)
+S31 DEVSML 3.0 Stack: Rapid deployment of DEVS Farm in distributed cloud environment using microservices and containers Symposium on Theory of Modeling and Simulation (TMS/DEVS) 2017 (Mittal and Risco-Martín, 2017)
+S32 Disaster Recovery and Business Continuity for Database Services in Multi-Cloud International Conference on Computer Applications & Information Security (ICCAIS) 2018 (Al-shammari and Alwan, 2018)
+S33 Dynamic On-Demand Fog Formation Offering On-the-Fly IoT Service Deployment IEEE Transactions on Network and Service Management 2020 (Sami and Mourad, 2020)
+S34 Dynamic security assurance in multi-cloud DevOps IEEE Conference on Communications and Network Security (CNS) 2017 (Rios et al., 2017)
+S35 Engineering and Experimentally Benchmarking a Container-based Edge Computing System International Conference on Communications (ICC) 2020 (Carpio et al., 2020)
+S36 FogDocker: Start Container Now, Fetch Image Later International Conference on Utility and Cloud Computing (UCC) 2019 (Lorenzo et al., 2019)
+S37 Importance of Application-Level Resource Management in Multi-Cloud Deployments International Conference on Cloud Engineering (IC2E) 2019 (Dimitrijevic et al., 2019)
+S38 Knowledge-Engineered Multi-Cloud Resource Brokering for Application Workflow Optimization IEEE Transactions on Network and Service Management 2022 (Pandey et al., 2022)
+S39 Label-affinity-Scheduler: Considering Business Requirements in Container Scheduling for Multi-Cloud and Multi-Tenant Brazilian Symposium on Computing System Engineering (SBESC) 2022 (Altran et al., 2022)
+S40 Leveraging kernel security mechanisms to improve container security: A survey International Conference on Availability, Reliability and Security (ARES) 2019 (Bélair et al., 2019)
+S41 Location-Aware and Budget-Constrained Service Deployment for Composite Applications in Multi-Cloud Environment IEEE Transactions on Parallel and Distributed Systems 2020 (Shi et al., 2020)
+S42 MicroCloud: A Container-Based Solution for Efficient Resource Management in the Cloud International Conference on Smart Cloud (SmartCloud) 2016 (Baresi et al., 2016)
+S43 Mobile Edge as Part of the Multi-Cloud Ecosystem: A Performance Study Euromicro International Conference on Parallel, Distributed and Network-Based Processing (PDP) 2019 (Dreibholz et al., 2019)
+S44 Modeling and Automated Deployment of Serverless Applications Using TOSCA Conference on Service-Oriented Computing and Applications (SOCA) 2018 (Wurster et al., 2018)
+S45 Monitoring Elastically Adaptive Multi-Cloud Services IEEE Transactions on Cloud Computing 2018 (Trihinas et al., 2015)
+S46 Monitoring-Aware Optimal Deployment for Applications Based on Microservices IEEE Transactions on Services Computing 2021 (Fadda et al., 2019)
+S47 MSSF: A Step towards User-Friendly Multi-cloud Data Dispersal International Conference on Cloud Computing (CLOUD) 2014 (Libardi et al., 2014)
+S48 Multi-agent QoS-aware autonomic resource provisioning framework for elastic BPM in containerized multi-cloud environment Journal of Ambient Intelligence and Humanized Computing 2022 (Saif et al., 2023)
+S49 Multicloud- auto scale with prediction and delta correction algorithm International Conference on Intelligent Computing, Instrumentation and Control Technologies (ICICICT) 2019 (Harwalkar et al., 2019b)
+50 Multi-cloud Distribution of Virtual Functions and Dynamic Service Deployment: Open ADN Perspective IEEE International Conference on Cloud Engineering (IC2E) 2015 (Bhamare et al., 2015)
+S51 Multi-cloud resource provisioning with Aneka: A unified and integrated utilization of Microsoft Azure and Amazon EC2International Conference on Computing and Network Communications (CoCoNet) 2015 (Buyya and Barreto, 2015)
+S52 MUSA Deployer: Deployment of Multi-cloud Applications IEEE International Workshop on Enabling Technologies: Infrastructure for Collaborative Enterprises (WETICE) 2017 (Casola et al., 2017)
+S53 NoMISHAP: A Novel Middleware Support for High Availability in Multicloud PaaS IEEE Cloud Computing 2017 (Acquaviva et al., 2017)
+S54 Object storage in the cloud and multi-cloud: State of the art and the research challengesIEEE International Conference on Automation, Quality and Testing, Robotics (AQTR) 2018 (Bucur et al., 2018)
+S55 On the Deployment of an Open-Source, 5G-Aware Evaluation Testbed International Conference on Mobile Cloud Computing, Services, and Engineering (MobileCloud) 2018 (Bolivar et al., 2018)
+S56 OpenStack Federation in Experimentation Multi-cloud Testbeds IEEE International Conference on Cloud Computing Technology and Science (CloudCom) 2013 (del Castillo et al., 2013)
+S57 Orchestrated Multi-Cloud Application Deployment in OpenStack with TOSCA International Conference on Smart Computing (SMARTCOMP) 2017 (Tricomi et al., 2017)
+S58 Orchestration Based Hybrid or Multi Clouds and Interoperability Standardization IEEE International Conference on Cloud Computing in Emerging Markets (CCEM) 2018 (Sitaram et al., 2018)
+S59 Performance Analysis of Zero-Trust multi-cloud IEEE International Conference on Cloud Computing (CLOUD) 2021 (Rodigari et al., 2021)
+S60 Policies Based Container Migration Using Cross-Cloud Management Platform International Conference on Information and Automation for Sustainability (ICIAfS) 2018 (Janarthanan et al., 2018)
+S61 Policy-based deployment in a hybrid and multi-cloud environment International Conference on Cloud Computing and Services Science (CLOSER) 2019 (Di Modica et al., 2019)
+S62 Predicting the End-to-End Tail Latency of Containerized Microservices in the Cloud International Conference on Cloud Engineering (IC2E) 2019 (Rahman and Lama, 2019)
+S63 Preventing vendor lock-ins via an interoperable multi-cloud deployment approach International Conference for Internet Technology and Secured Transactions (ICITST) 2017 (Pellegrini et al., 2017)
+S64 Private STaaS with OpenStack Cinder Volumes for Hybrid/Multi-cloud International Conference on Advances in Computing and Communication Engineering (ICACCE) 2019 (Harwalkar et al., 2019a)
+S65 Real-World Experience with a Multicloud Exchange IEEE Cloud Computing 2017 (Zeck and Bouroudjian, 2017)
+S66 Reliability Management for Blockchain-Based Decentralized Multi-Cloud International Symposium on Cluster Computing and the Grid (CCGRID) 2020 (Aral et al., 2020)
+S67 Reliability-Aware Cost-Efficient Scientific Workflows Scheduling Strategy on Multi-Cloud Systems IEEE Transactions on Cloud Computing 2022 (Tang, 2021)
+S68 Resource optimization of container orchestration: a case study in multi-cloud microservices-based applicationsThe Journal of Supercomputing 2018 (Guerrero et al., 2018a)
+S69 Scalable multi-cloud platform to support industry and scientific applications International Convention on Information and Communication Technology, Electronics and Microelectronics (MIPRO) 2018 (Kiss, 2018)
+S70 Search-based Methods for Multi-Cloud Configuration International Conference on Cloud Computing (CLOUD) 2022 (Lazuka et al., 2022)
+S71 Secure access control for multi-cloud resources IEEE Conference on ALocal Computer Networks Workshops (LCN Workshops) 2015 (Graupner et al., 2015a)
+S72 Secure and Dependable Multi-Cloud Network Virtualization International Workshop on Security and Dependability of Multi-Domain Infrastructures 2017 (Alaluna et al., 2017)
+S73 Secure and Parallel Expressive Search over Encrypted Data with Access Control in Multi-CloudIoT Cloudification of the Internet of Things (CIoT) 2018 (Guechi and Maamri, [n. d.])
+S74 Self-healing cloud services in private multi-clouds International Conference on High Performance Computing & Simulation (HPCS) 2018 (Mfula and Nurminen, 2018)
+S75 SpyStorage: A Highly Reliable Multi-Cloud Storage with Secure and Anonymous Data Sharing International Conference on Networking, Architecture, and Storage (NAS) 2017 (Shen et al., 2017)
+S76 Support Services for Applications Execution in Multi-clouds Environments International Conference on Autonomic Computing (ICAC) 2016 (Pop et al., 2016)
+S77 SynAPTIC: Secure and persistent connectivity for containers International Symposium on Cluster, Cloud and Grid Computing (CCGRID) 2017 (Ranjbar et al., 2017)
+S78 System Restore in a Multi-cloud Data Pipeline Platform International Conference on Dependable Systems and Networks – Industry Track 2019 (Wang et al., 2019)
+S79 Systematic and Recomputable Comparison of Multi-cloud Management Platforms International Conference on Cloud Computing Technology and Science (CloudCom) 2018 (Serhiienko and Spillner, 2018)
+S80 Towards a lightweight multi-cloud DSL for elastic and transferable cloud-native applications International Conference on Cloud Computing and Services Science (CLOSER) 2018 (Quint and Kratzke, 2018)
+S81 Towards Formal-Based Semantic Interoperability in Multi-Clouds: The FCLOUDS Framework IEEE International Conference on Cloud Computing (CLOUD) 2017 (Challita et al., 2017)
+S82 Towards Model-Driven Provisioning, Deployment, Monitoring, and Adaptation of Multi-cloud Systems IEEE International Conference on Cloud Computing (CLOUD) 2013 (Ferry et al., 2013)
+S83 Towards Trustworthy Multi-Cloud Services Communities: A Trust-Based Hedonic Coalitional Game IEEE Transactions on Services Computing 2018 (Wahab et al., 2016)
+S84 Multi-cloud provisioning of business processes Journal of Cloud Computing 2019 (Kritikos et al., 2019)
+S85 Evaluating the DevOps Reference Architecture for Multi-cloud IoT-Applications SN Computer Science 2021 (Bou Ghantous and Gill, 2021)
+S86 Hybrid IT and Multi Cloud an Emerging Trend and Improved Performance in Cloud Computing SN Computer Science 2020 (Gundu et al., 2020)

### v3

@@
 Containerization in Multi-Cloud Environment: Roles, Strategies, Challenges, and Solutions for Effective Implementation
-Abstract.
-Containerization in a multi-cloud environment facilitates workload portability and optimized resource utilization.
+Abstract
 Containerization in multi-cloud environments has received significant attention in recent years both from academic research and industrial development perspectives.
@@
 The results of this SMS will assist researchers and practitioners in pursuing further studies on containerization in multi-cloud environment and developing specialized solutions for containerization applications in multi-cloud environment.
-Containerization, Multi-Cloud Environment, Systematic Mapping Study, Cloud Computing
-journal: TOSEMccs: Software and its engineering Software development techniques
-1. Introduction
-The use of containers in multi-cloud environment has been widespread in the industry for many years (Naik, 2016b).
-Containers are standalone and executable packages of software that include everything needed to run an application, such as code, system tools, libraries, and settings (Merkel et al., 2014).
-Containers allow the packaging of an application and its required dependencies, which makes it easy to move the application across different environments with minimal modification (Docker, 2021).
-On the other hand, multi-cloud environment is the distribution of cloud assets (Petcu, 2013).
+keywords:
+[…]
+1 Introduction
+The use of containers in multi-cloud environment has been widespread in the industry for many years [1].
+Containers are standalone and executable packages of software that include everything needed to run an application, such as code, system tools, libraries, and settings [2].
+Containers allow the packaging of an application and its required dependencies, which makes it easy to move the application across different environments with minimal modification [3].
+On the other hand, multi-cloud environment is the distribution of cloud assets [4].
 Using containers in multi-cloud environment allows organizations to achieve flexibility, agility, and cost-efficiency.
@@
 This architecture enables a flexible and scalable approach, allowing efficient application deployment and operations in a multi-cloud setting, while maintaining consistency and resilience across platforms.
-Figure 1. Context: Containerization in multi-cloud environment
-However, utilizing containerization in multi-cloud environment is not without challenges that organizations may face (Saif et al., 2022).
+Figure 1: Context: Containerization in multi-clouds
+However, utilizing containerization in multi-cloud environment is not without challenges that organizations may face [5].
 These challenges may arise during different phases and activities of container-based application development including architectural design phase, system implementation, or when establishing an automated development infrastructure.
 Furthermore, the challenges can also present themselves during system testing, the coding process, and the deployment phase.
-For instance, during the architectural design phase, challenges may include ensuring seamless integration of containerized components with existing systems and selecting appropriate container orchestration patterns, strategies, and tools that can effectively manage containers across multiple clouds (Raj et al., 2018) (Baby and Vysala, 2015), and while establishing an automated development infrastructure, organizations may encounter difficulties in creating efficient Continuous Integration/Continuous Deployment (CI/CD) pipelines that handle multi-cloud deployment scenarios while maintaining consistent performance and security standards (Debroy et al., 2018).
-Similarly, in the system implementation phase, challenges may include resolving potential compatibility issues between containers and various cloud platforms, as well as optimizing resource utilization to control costs effectively (Helali and Omri, 2021).
-Motivation: Our study is part of the QLEAP project (2022–2024) funded by Business Finland, which investigates the use of containers in multi-cloud environment specifically for architecture design (Ahmad et al., 2025).
+For instance, during the architectural design phase, challenges may include ensuring seamless integration of containerized components with existing systems and selecting appropriate container orchestration patterns, strategies, and tools that can effectively manage containers across multiple clouds [6] [7], and while establishing an automated development infrastructure, organizations may encounter difficulties in creating efficient Continuous Integration/Continuous Deployment (CI/CD) pipelines that handle multi-cloud deployment scenarios while maintaining consistent performance and security standards [8].
+Similarly, in the system implementation phase, challenges may include resolving potential compatibility issues between containers and various cloud platforms, as well as optimizing resource utilization to control costs effectively [9].
+Motivation: Our study is part of the QLEAP project (2022–2024) funded by Business Finland, which investigates the use of containers in multi-cloud environment specifically for architecture design [10].
 The project brings together a consortium of four companies (Bittium [^1] https://www.bittium.com/, M-Files [^2] https://www.m-files.com/, Solita [^3] https://www.solita.fi/en/, and Vaadin [^4] https://vaadin.com/), each with distinct requirements for containerization, and is further supported by industry leaders Nokia [^5] https://www.nokia.com/ and TietoEVRY [^6] https://www.tietoevry.com/.
@@
 Our study not only fulfills an immediate industry demand but also aims to provide a structured view of containerization practices, addressing the needs of both practitioners and researchers in navigating this fragmented knowledge space.
-Recent research (e.g., see Table 8) has highlighted the growing significance of container utilization in multi-cloud environment.
+Recent research (e.g., see Selected Studies sheet in [11]) has highlighted the growing significance of container utilization in multi-cloud environment.
 These studies have explored various aspects, including container roles and strategies, architectural patterns, Quality Attributes (QAs), and the tools and frameworks employed.
@@
 Moreover, the distributed nature of this knowledge delays the formation of a cohesive and unified understanding of the subject matter.
-Recent studies (e.g., (Bordeleau et al., 2020) (Mihai et al., 2022) (Sarker, 2022)) have notably highlighted that challenges in design, development, monitoring, and testing of containerization applications in multi-cloud environment are deeply interconnected with the software development life cycle.
+Recent studies (e.g., [12] [13] [14]) have notably highlighted that challenges in design, development, monitoring, and testing of containerization applications in multi-cloud environment are deeply interconnected with the software development life cycle.
 To bridge this knowledge gap and support the practical demands of the QLEAP project consortium, our study systematically identifies and categorizes the various facets of container utilization in multi-cloud environment.
@@
 We finally selected 86 studies that focus on containerized applications in multi-cloud environment for further analysis.
-We adopted the following Goal-Question-Metric (GQM) framework (Caldiera and Rombach, 1994):
+We adopted the following Goal-Question-Metric (GQM) framework [15]:
   • Goal:To enhance the understanding and efficiency of deploying and managing container-based applications in multi-cloud environment.
@@
   • We systematically identified and classified a wide range of tools and frameworks for containerized applications in multi-cloud environment.
-  • We have publicly released a dataset, which is available online (Waseem et al., 2024), to enable researchers and practitioners to access, replicate, and validate all collected data from our SMS.
+  • We have publicly released a dataset, which is available online [11], to enable researchers and practitioners to access, replicate, and validate all collected data from our SMS.
 This dataset includes detailed hierarchies of the developed catalogs, identified and classified container implementation strategies, and the roles of containers in multi-cloud environment.
@@
 Section 7 draws conclusions and outlines future research directions.
-2. Methodology
-We conducted a Systematic Mapping Study (SMS) by following the guidelines in (Petersen et al., 2008a) and augmenting them with SLR strategies (Keele et al., 2007).
+2 Methodology
+We conducted a Systematic Mapping Study (SMS) by following the guidelines in [16] and augmenting them with SLR strategies [17].
 Our SMS consists of three phases: specifying research questions and search string, conducting the literature search, and performing data analysis and documentation.
@@
 Additionally, we applied systematic quality assessment techniques to evaluate the studies included in our SMS.
-Figure 2. Schematic representation of the research methodology implemented in this study
-2.1. Research Questions
+Figure 2: Schematic representation of the research methodology implemented in this study
+2.1 Research Questions
 To conduct this SMS, we formulated six Research Questions (RQs) based on the GQM approach presented in the Introduction section (see Section 1) aligns with the objective of our SMS, as outlined in Table 1.
 The table provides a structured view of the RQs with their corresponding rationale and categorization.
-Table 1. Research questions and their rationale
+Table 1: Research questions and their rationale
 # Research Questions Rationale
@@
 Together, these RQs provide a structured framework that enhances understanding and supports the development and management of container-based applications in multi-cloud environment, advancing the state-of-the-art in this domain
-2.2. Search String Composition
-Initially, we considered the PICO (Population, Intervention, Comparison, Outcome) framework (Schardt et al., 2007) (Ralph et al., 2020) for constructing our search string.
+2.2 Search String Composition
+Initially, we considered the PICO (Population, Intervention, Comparison, Outcome) framework [18] [19] for constructing our search string.
 However, given the complexity and breadth of our topic, applying PICO strictly required an extensive array of terms and logical operators, exceeding the query limits of databases like ScienceDirect.
@@
 This approach allowed us to tailor the search strategy to the specific goal of our study, ensuring comprehensive coverage within practical constraints.
-2.3. Study Search and Selection Process
+2.3 Study Search and Selection Process
 The study search and selection process for this study is divided into two phases.
-In the first phase, we conducted a primary search using a specific search string on various databases (see Table 2 and the Summary sheet in (Waseem et al., 2024)).
+In the first phase, we conducted a primary search using a specific search string on various databases (see Table 2 and the Summary sheet in [11]).
 In the second phase, we applied the snowballing technique on primary studies that were selected from the first phase.
-Table 2. Search string and targeted search area for databases
+Table 2: Search string and targeted search area for databases
 Search String
@@
 Web of Science Paper title, keywords, abstract
-2.3.1. Primary Search
+2.3.1 Primary Search
 The primary search involved querying digital databases (see Table 2) using customized search string.
 The search strings were executed concurrently on eight databases, as illustrated in Figure 2, spanning between January 2013 and July 2024 when we started this SMS.
-We chose 2013 as the starting point because it is the year when container technologies like Docker emerged (Docker, 2023), marking a key phase in containerization and cloud computing.
+We chose 2013 as the starting point because it is the year when container technologies like Docker emerged [20], marking a key phase in containerization and cloud computing.
 This period also saw a rise in relevant research to containers in multi-cloud environment, making it ideal to understand the development of containerized applications in multi-cloud environment over the past decade.
 We followed the steps mentioned below to select the relevant studies.
-Table 3. Inclusion and exclusion criteria for selecting primary studies in this SMS
+Table 3: Inclusion and exclusion criteria for selecting primary studies in this SMS
 Selection Criteria Inclusion Criteria Exclusion Criteria
@@
 Study Type Primary studies including peer-reviewed journal articles, book chapters, conference papers, workshops, and symposium papers Secondary studies or non-peer-reviewed content (e.g., blogs, webpages, videos, white papers, technical reports, grey literature)
-Study Focus Studies addressing container roles and strategies, architectural patterns, quality attributes, adoption drivers, automation, deployment, monitoring, security challenges, solutions, and tools/frameworks for multi-cloud environment.
-Studies that focus solely on traditional cloud computing, single-cloud environments, or unrelated topics
+Study Focus Studies that explicitly address the intersection of containerization and multi-cloud environments.
+This includes studies that investigate the use of container technologies (e.g., Docker, Kubernetes) in the context of multi-cloud deployment, management, or architecture.
+Studies that focus solely on containerization without a multi-cloud context, or vice versa, are excluded.
+Studies that address only traditional cloud computing, single-cloud environments, or containerization without reference to multi-cloud contexts, and vice versa.
+Specifically, studies that examine container technologies without situating them in multi-cloud scenarios, or studies focused on multi-cloud architectures without discussing container-based solutions, are excluded.
 Study Duration Studies published between January 2013 and July 2024 Studies published before January 2013 or after July 2024
@@
 This initial search yielded 6,727 studies from eight databases.
-Following initial retrieval, we did not apply Cohen’s Kappa (Cohen, 1960) (Cohen, 1968) as this step is automated and does not involve subjective decision-making.
+Following initial retrieval, we did not apply Cohen’s Kappa [21] [22] as this step is automated and does not involve subjective decision-making.
   • Step 2: Title and Keyword Screening: This step involves reviewing the titles and keywords of the collected studies to assess their relevance to the research topic.
+To ensure alignment with our inclusion criteria, we manually reviewed study titles and abstracts to include only those that explicitly discuss containerization within multi-cloud contexts.
+Studies that focused solely on single-cloud containerization, general cloud orchestration without containers, or container use cases unrelated to multi-cloud environments were excluded—even if they matched keywords in the search string.
 Initially, we removed any duplicate studies obtained from different databases (e.g., IEEE Xplore, Scopus) by sorting them in ascending order.
@@
 We divided the remaining studies between two authors (R1, R2), who independently applied the inclusion and exclusion criteria to identify studies relevant to our research objective, as detailed in Table 3.
-    To assess the consistency of study selection and mitigate bias, a Cohen’s Kappa analysis (Cohen, 1960) (Cohen, 1968) was performed on a random sample of 400 studies (10% of the total dataset of 4,000 studies), as illustrated in Formula 1:
-    (1) $\kappa=\frac{P_{o}-P_{e}}{1-P_{e}}$
+To assess the consistency of study selection and mitigate bias, a Cohen’s Kappa analysis [21] [22] was performed on a random sample of 400 studies (10% of the total dataset of 4,000 studies), as illustrated in Formula 1:
+    $\kappa=\frac{P_{o}-P_{e}}{1-P_{e}}$ (1)
     where $P_{o}$ represents the relative observed agreement among authors, and $P_{e}$ the hypothetical probability of chance agreement.
-The results, recorded in Table 4.a, indicated a Cohen’s Kappa of $\kappa=0.531$, which reflects moderate agreement between R1 and R2 (Sim and Wright, 2005).
+The results, recorded in Table 4.a, indicated a Cohen’s Kappa of $\kappa=0.531$, which reflects moderate agreement between R1 and R2 [23].
 Specifically, R1 and R2 agreed on 360 studies being relevant (“YES”) and 20 studies being irrelevant (“NO”).
 Disagreements occurred in 20 studies, which were systematically resolved through discussions among all authors, leading to a final consensus.
-Using a random subset is a common practice in systematic reviews to measure inter-rater agreement without overburdening resources, ensuring reliability while maintaining efficiency (Keele et al., 2007; Chandler et al., 2019).
+Using a random subset is a common practice in systematic reviews to measure inter-rater agreement without overburdening resources, ensuring reliability while maintaining efficiency [17, 24].
 Following this process, the number of studies was reduced to 1,047 for further analysis.
@@
 Disagreements occurred in 12 cases (7 studies where one researcher marked “relevant” while the other marked “irrelevant”, and 5 studies initially classified as “doubtful”).
-The kappa value, calculated based on observed agreements and disagreements, was $\kappa=0.521$, calculated based on observed agreements and disagreements, indicates moderate agreement (Sim and Wright, 2005; Keele et al., 2007).
-    To compute Cohen’s Kappa during the abstract-based screening phase, the original classifications of “relevant”, “irrelevant”, and “doubtful” were simplified into a binary scale (“YES” or “NO”) due to the binary nature of Cohen’s Kappa computation.
+The kappa value, calculated based on observed agreements and disagreements, was $\kappa=0.521$, calculated based on observed agreements and disagreements, indicates moderate agreement [23, 17].
+    To compute Cohen’s Kappa during the abstract-based screening phase, the original classifications of “relevant", “irrelevant”, and “doubtful” were simplified into a binary scale (“YES” or “NO”) due to the binary nature of Cohen’s Kappa computation.
 Specifically:
@@
     • Studies classified as “doubtful” were resolved through discussions between R1 and R2. If consensus was reached on “relevant”, the study was marked as “YES”; otherwise, it was marked as “NO”.
-    Using subsets for kappa computation ensures that inter-rater reliability can be rigorously evaluated while optimizing the use of resources, as recommended by systematic review guidelines (Keele et al., 2007; Chandler et al., 2019).
+    Using subsets for kappa computation ensures that inter-rater reliability can be rigorously evaluated while optimizing the use of resources, as recommended by systematic review guidelines [17, 24].
 After resolving all disagreements through systematic discussions among the authors, a final consensus was reached for the subset.
@@
 To ensure consistency in our review process, we randomly selected a subset of 35 studies (10% of the total 346 studies) for a Cohen’s Kappa analysis to measure the agreement between two researchers (R1 and R2).
-The results, recorded in Table 4.c, indicated a Cohen’s Kappa value of $\kappa=0.578$, which represents moderate agreement between the two researchers (Sim and Wright, 2005).
+The results, recorded in Table 4.c, indicated a Cohen’s Kappa value of $\kappa=0.578$, which represents moderate agreement between the two researchers [23].
 Specifically, R1 and R2 agreed on 30 studies as relevant (“YES”) and 2 studies as irrelevant (“NO”).
@@
 All disagreements were systematically resolved through discussions among the authors, leading to a final consensus.
-This subset-based approach, coupled with systematic resolution of disagreements, provides confidence in the reliability of the full-text screening process (Chandler et al., 2019).
+This subset-based approach, coupled with systematic resolution of disagreements, provides confidence in the reliability of the full-text screening process [24].
 At the end of this step, we obtained a total of 110 studies, which were deemed relevant for further analysis.
-Table 4. Kappa Agreement across different screening steps.
-(a) Title & Keyword Screening on 400 Randomly Selected Studies
-R1 Yes R1 No
-R2 Yes 360 20
-R2 No 10 10
-Total 370 30
-(b) Abstract-based Screening on 105 Randomly Selected Studies
-R1 Yes R1 No
-R2 Yes 88 7
-R2 No 5 5
-Total 93 12
-(c) Full-Text Screening on 35 Randomly Selected Studies
-R1 Yes R1 No
-R2 Yes 30 3
-R2 No 2 0
-Total 32 3
-2.3.2. Snowballing
-In Phase 2, we utilized the snowballing technique, as described in (Wohlin, 2014), to examine references within 110 primary studies to identify additional relevant studies.
+Table 4: Kappa agreement across different screening steps
+(a) Title & Keyword
+Screening (400 Studies)
+R1 Yes R1 No R2 Yes 360 20 R2 No 10 10 Total 370 30
+(b) Abstract-Based
+Screening (105 Studies)
+R1 Yes R1 No R2 Yes 88 7 R2 No 5 5 Total 93 12
+(c) Full-Text
+Screening (35 Studies)
+R1 Yes R1 No R2 Yes 30 3 R2 No 2 0 Total 32 3
+2.3.2 Snowballing
+In Phase 2, we utilized the snowballing technique, as described in [25], to examine references within 110 primary studies to identify additional relevant studies.
 This strategy was augmented by forward snowballing, where we gathered studies that cited the selected studies, and backward snowballing, which involved using references within the selected studies.
@@
 This phase resulted in the addition of 11 more studies, bringing the final count to 121.
-2.3.3. Quality Assessment of Selected Studies
+2.3.3 Quality Assessment of Selected Studies
 We assessed the quality of the selected studies based on a set of predefined criteria, including relevance to multi-cloud containerization, clarity of research design, empirical validation (e.g., experiments, case studies), reporting quality (e.g., discussion of limitations and challenges), publication in peer-reviewed venues, and the generalizability of findings to multi-cloud environment.
-Each study was evaluated using a scoring system where criteria were rated as 0 (not satisfied), 0.5 (partially satisfied), or 1 (fully satisfied), with a maximum total score of 6.
+Each study was assessed using six predefined quality criteria: relevance to multi-cloud containerization, clarity of research design, data validity (e.g., empirical evidence), reporting quality (e.g., discussion of limitations), peer-review status, and generalizability of findings.
+Each criterion was scored using a 3-point scale: 1 (fully satisfied), 0.5 (partially satisfied), or 0 (not satisfied), giving a total score out of 6.
+The scoring was applied consistently across all studies based on observable features.
+For example, studies addressing both containerization and multi-cloud contexts were scored 1 for relevance; studies with explicit research methods scored higher on research design; and those lacking empirical evidence were scored lower for data validity.
+Scores for all individual criteria are documented in the “QA of Selected Studies” sheet in our replication package, providing transparency and allowing independent verification of the quality assessment process.
 Studies scoring below 3 were flagged for further review and discussion but retained to ensure alignment with the study’s objectives.
 This quality assessment was conducted after the study selection process to ensure consistency and thoroughness across all 121 studies.
-Detailed results are available in the replication package (Waseem et al., 2024), specifically in the sheet titled “QA of Selected Studies”.
-2.4. Data Extraction and Analysis
+Detailed results are available in the replication package [11], specifically in the sheet titled “QA of Selected Studies”.
+2.4 Data Extraction and Analysis
 Data Extraction: The data extraction form was designed based on predefined data items (see Table 5) that were formulated to address the RQs specified in Table 1.
@@
 Data Analysis: We employed descriptive statistics to analyze the quantitative data from data items D1, D5, D7, D8, and D9.
-For the remainder of the data, which primarily consist of qualitative free-text descriptions (e.g., study aim, roles of containers, challenges, and solutions), we conducted a thematic analysis in accordance with the guidelines outlined in (Braun and Clarke, 2006).
+For the remainder of the data, which primarily consist of qualitative free-text descriptions (e.g., study aim, roles of containers, challenges, and solutions), we conducted a thematic analysis in accordance with the guidelines outlined in [26].
 Our thematic analysis process consists of the following steps:
-  1. (1)
+  1. 1.
     Data Familiarization: We conducted a thorough review of the selected studies by repeatedly reading through them and meticulously noting key points related to study aims (D7), contributions (D8), roles of containers (D9), implementation strategies (D10), architectural patterns (D11), quality attributes (D12), tactics (D13), motivations (D14), and challenges and solutions (D15-D21) in automation, deployment, monitoring, and security challenges and solutions, as well as the tools, languages, and frameworks (D22) employed.
-  2. (2)
+  2. 2.
     Generation of Initial Codes: After developing a thorough understanding of the data, we created an initial set of codes derived from the information extracted concerning the data items identified in the previous step.
-  3. (3)
+  3. 3.
     Identification of Types and Emerging Themes: In this step, we conducted a two-tiered analysis.
@@
 Subsequently, we developed subcategories based on these types, and then formulated overarching categories that encompass the related subcategories.
-  4. (4)
+  4. 4.
     Critical Evaluation of Types, Subcategories, and Categories: All authors actively participated in rigorously reviewing and refining the coded data, including types, subcategories, and categories.
 During this collaborative process, we redefined, merged, or dropped certain themes based on collective input and discussion.
-  5. (5)
+  5. 5.
     Defining and Naming Categories: At this point, we provided explicit definitions for each of the identified themes and further refined them, ensuring that the terminology used for the categories was precise and unambiguous.
-Table 5. Data items to be extracted in this SMS
+Table 5: Data items to be extracted in this SMS
 Code Data Item Description RQ
@@
 In these sessions, both researchers discussed and validated the research themes found.
-Concerning data item D9 (Roles of Containers) and D10 (Container Implementation strategy), we used the open coding and constant comparison techniques from Grounded Theory (Glaser and Strauss, 1967) to analyze the qualitative data extracted from the selected studies.
-Finally, we provided a replication package (Waseem et al., 2024) with the results of each phase of the study selection process (e.g., Phase 1, Phase 2) and detailed results (e.g., Contributions, Patterns, QAs, Challenges, Solutions) for verification and validation purposes of this SMS.
-3. Results
+Concerning data item D9 (Roles of Containers) and D10 (Container Implementation strategy), we used the open coding and constant comparison techniques from Grounded Theory [27] to analyze the qualitative data extracted from the selected studies.
+Finally, we provided a replication package [11] with the results of each phase of the study selection process (e.g., Phase 1, Phase 2) and detailed results (e.g., Contributions, Patterns, QAs, Challenges, Solutions) for verification and validation purposes of this SMS.
+3 Results
 This section presents the results of the SMS, derived from analyzing data from chosen studies.
@@
 Finally, the tools and frameworks are reported in Section 15.
-3.1. Demographics, Research Aims, and Contributions
+3.1 Demographics, Research Aims, and Contributions
 Yearly Distribution of Studies: Figure 3-a illustrates the annual distribution of studies retrieved from eight databases between January 2013 and July 2024.
@@
 However, there is a notable increase in studies in 2023 and continuing into 2024.
-This could be due to delayed research projects getting back on track and more funding in tech sectors after the pandemic, showing a renewed focus on improving containerization in multi-cloud environment
-Figure 3. Demographics of the Selected Studies, Including: a) Frequency and Categories of Publications; b) Publisher of the Research; c) Affiliations of the Authors
+This could be due to delayed research projects getting back on track and more funding in tech sectors after the pandemic, showing a renewed focus on improving containerization in multi-cloud environment.
+Figure 3: Demographics of the selected studies, including: a) Frequency and Types of Publications; b) Research Publisher; c) Author’s Affiliation
 Publishers Distribution: Figure 3-b shows the distribution of studies by publisher.
@@
 This low rate of cross-affiliation collaboration may suggest that there is potential for increased synergy between academic and industrial entities in future research endeavors.
+Takeaway 1: Research on containerization in multi-cloud environments has gained more attention in recent years, especially after 2022.
+Most of the studies are published by IEEE and come from academic authors.
+This shows that the topic is mainly explored in academic settings.
+However, there is little contribution from industry, and most studies are presented at conferences rather than in journals.
+This suggests a need for more collaboration between academia and industry and more detailed research through journal publications.
 Classification of Research Theme: The identified research themes and their notable trends are illustrated in Figure 4 and detailed below.
@@
 Two possible reasons for this are: (i) our literature search did not identify any relevant studies (see Section 2.3.1 - primary search), or (ii) the studies from those years did not meet the qualitative evaluation criteria (see Section 2.3.3 - quality assessment) for inclusion in the review.
-Figure 4. Selected Studies: a) Thematic Classification and b) Trends Over the Years
+Figure 4: Selected studies: a) Thematic Classification and b) Trends and Distribution of Themes Over the Years
+Takeaway 2: The research themes show that when containerization is applied in multi-cloud environments, unique challenges emerge—such as the need for orchestration tools that can manage service consistency across diverse providers, and security mechanisms that adapt to varying cloud configurations.
+Studies have evolved from basic scalability and performance concerns to addressing how to efficiently monitor, deploy, and manage containerized applications across different cloud platforms using DevOps and federation-aware solutions.
+This highlights how multi-cloud settings demand more adaptable and interoperable container strategies than single-cloud systems.
 Major Contributions of the Selected Studies: Table 6 provides an overview of the contributions of the selected studies.
-The contributions are primarily categorized into five main categories and 22 subcategories (see the Contributions sheet in (Waseem et al., 2024)), through thematic analysis.
+The contributions are primarily categorized into five main categories and 22 subcategories (see the Contributions sheet in [11]), through thematic analysis.
 The selected studies mainly focus on proposing and validating security solutions, reference architectures, cloud management frameworks, interoperability, orchestration, and tools for optimizing and deploying containerized applications in multi-cloud environment.
@@
 This includes contributions from Resource and Service Management Frameworks, which represent 4.95% of the studies.
-Table 6. Selected studies’ contributions on containerization in multi-cloud environment
+Table 6: Selected studies’ contributions on containerization in multi-cloud environment
 Category Subcategory Study #
-Reference Architectures and Models for Cloud General Cloud Architectures (6, 4.92%) (Watanabe et al., 2022; Sami and Mourad, 2020; Carpio et al., 2020; Baresi et al., 2016; Pop et al., 2016; Alonso et al., 2023a)
-IoT & Service Mesh Architecture (5, 4.13%) (Gattobigio et al., 2022; Rodigari et al., 2021; Merlino et al., 2024; Hossain et al., 2023; Solayman and Qasha, 2023)
-Fault Tolerance & Self-Healing (5, 4.13%) (Bisht and Kaur, 2022; Bolivar et al., 2018; Tang, 2021; Mfula and Nurminen, 2018; Nagarajan et al., 2023)
-Migration & Testing Architecture (3, 2.47%) (Slominski et al., 2015; Serhiienko and Spillner, 2018; Roy and Kalafatis, 2023)
-Optimization & Reliability Model (3, 2.47%) (Bhamare et al., 2015; Aral et al., 2020; Huang and Wang, 2023)
-Domain-Specific Language (2, 1.65%) (Quint and Kratzke, 2018; Ferry et al., 2013)
-Performance Modeling (2, 1.65%) (Barna et al., 2017; Rahman and Lama, 2019)
-Security Approaches Cybersecurity & Access Control (11, 9.09%) (He et al., 2022; Sheridan et al., 2017; Rios et al., 2017; Graupner et al., 2015a; Guechi and Maamri, [n. d.])
-Cloud Analysis Approaches (7, 5.78%) (Pai and Kumar, 2021; Pahl, 2015; Dimitrijevic et al., 2019; Zeck and Bouroudjian, 2017; Kiss, 2018; Gundu et al., 2020)
-Security and Performance Solution (6, 4.95%) (Sabbioni et al., 2020; Wang et al., 2020; Pandey et al., 2022; Ranjbar et al., 2017)
-Security Approaches (5, 4.13%) (Deng et al., 2022; Gao et al., 2017; Bélair et al., 2019; Wahab et al., 2016)
-Provisioning & Restoration Approach (4, 3.30%) (Pellegrini et al., 2017; Lazuka et al., 2022; Wang et al., 2019; Kritikos et al., 2019)
-Cloud Management and Interoperability Framework Resource & Service Management Framework (6, 4.95%) (Saif et al., 2023; Harwalkar et al., 2019b; Libardi et al., 2014; Pham and Pham, 2015; Sitaram et al., 2018; Barletta et al., 2024)
-Monitoring & Deployment Framework (5, 4.13%) (Alhamazani et al., 2014; Trihinas et al., 2015; Fadda et al., 2019; Pham et al., 2015; Horchulhack et al., 2024; Marques et al., 2024)
-Evaluation & Modeling Framework (4, 3.30%) (Ferry et al., 2014; Bou Ghantous and Gill, 2021; Tundo et al., 2024; Ramesh et al., 2023)
-Interoperability Framework (2, 1.65%) (Hua et al., 2020; Challita et al., 2017)
-Orchestration and Deployment Approach Benchmarking & Optimization Approach (6, 4.13%) (Almeida et al., 2014; Jha et al., 2019; Shi et al., 2020; Buyya and Barreto, 2015; Guerrero et al., 2018a; Petrovski and Gusev, 2024)
-Service Management Approach (4, 3.30%) (Zhang et al., 2021; Chen et al., 2022; Gupta et al., 2021; Benmerar et al., 2023)
-Simulation & Integration Approach (3, 2.47%) (Alonso et al., 2019; Mittal and Risco-Martín, 2017; del Castillo et al., 2013)
-Storage & Services Solution (3, 2.47%) (Harwalkar et al., 2019a; Shen et al., 2017; Orzechowski et al., 2023)
-Optimization and Deployment Approaches and Tool Optimization & Scheduling Approach (10, 8.26%) (Wong et al., 2019; Al-shammari and Alwan, 2018; Lorenzo et al., 2019; Altran et al., 2022; da Silva et al., 2023; Dogani et al., 2023; Pandi et al., 2023; Sofia et al., 2023; Nguyen and Chien, 2023)
-Multi-Cloud Platform (6, 4.95%) (Verdugo et al., 2017; Ouyang et al., 2022; Dreibholz et al., 2019; Janarthanan et al., 2018; Alaluna et al., 2017; Wong et al., 2019; Al-shammari and Alwan, 2018; Lorenzo et al., 2019; Altran et al., 2022; Santos et al., 2023)
-Deployment Approach and Tool (6, 4.95%) (Cui et al., 2016; Casola et al., 2017; Tricomi et al., 2017; Di Modica et al., 2019; Liu and Qin, 2024; Tiwari and Sharma, 2023)
-[…]
-3.2. Container Roles and Implementation Strategies (RQ1)
-We report the results related to roles and implementation strategies in the following two subsections.
-3.2.1. Container Roles
+Reference Architectures and Models for Cloud General Cloud Architectures (6, 4.92%) [28, 29, 30, 31, 32, 33]
+IoT & Service Mesh Architecture (5, 4.13%) [34, 35, 36, 37, 38]
+Fault Tolerance & Self-Healing (5, 4.13%) [39, 40, 41, 42, 43]
+Migration & Testing Architecture (3, 2.47%) [44, 45, 46]
+Optimization & Reliability Model (3, 2.47%) [47, 48, 49]
+Domain-Specific Language (2, 1.65%) [50, 51]
+Performance Modeling (2, 1.65%) [52, 53]
+Security Approaches Cybersecurity & Access Control (11, 9.09%) [54, 55, 56, 57, 58]
+Cloud Analysis Approaches (7, 5.78%) [59, 60, 61, 62, 63, 64]
+Security and Performance Solution (6, 4.95%) [65, 66, 67, 68]
+Security Approaches (5, 4.13%) [69, 70, 71, 72]
+Provisioning & Restoration Approach (4, 3.30%) [73, 74, 75, 76]
+Cloud Management and Interoperability Framework Resource & Service Management Framework (6, 4.95%) [77, 78, 79, 80, 81, 82]
+Monitoring & Deployment Framework (5, 4.13%) [83, 84, 85, 86, 87, 88]
+Evaluation & Modeling Framework (4, 3.30%) [89, 90, 91, 92]
+Interoperability Framework (2, 1.65%) [93, 94]
+Orchestration and Deployment Approach Benchmarking & Optimization Approach (6, 4.13%) [95, 96, 97, 98, 99, 100]
+Service Management Approach (4, 3.30%) [101, 102, 103, 104]
+Simulation & Integration Approach (3, 2.47%) [105, 106, 107]
+Storage & Services Solution (3, 2.47%) [108, 109, 110]
+Optimization and Deployment Approaches and Tool Optimization & Scheduling Approach (10, 8.26%) [111, 112, 113, 114, 115, 116, 117, 118, 119]
+Multi-Cloud Platform (6, 4.95%) [120, 121, 122, 123, 124, 111, 112, 113, 114, 125]
+Deployment Approach and Tool (6, 4.95%) [126, 127, 128, 129, 130, 131]
+3.2 Container Roles and Implementation Strategies (RQ1)
+In this section, we present the roles that containers play in multi-cloud environments and the strategies used to implement them.
+Based on thematic analysis of the selected studies, we classified the findings into six major roles and five strategic categories.
+Each category is supported with examples to help readers relate to real-world use cases.
+The goal is to provide a holistic understanding of how containers are applied and managed across clouds.
+Readers can interpret the tables as a catalog of practical use scenarios and corresponding strategic techniques.
+3.2.1 Container Roles
 We classify the roles of containers in a multi-cloud environment into six categories, as shown in Table 7. Each category is briefly reported below.
+Takeaway 3: The contributions of the reviewed studies highlight how multi-cloud environments introduce specific demands on containerization, particularly in areas such as cross-platform security, cloud-aware architecture design, and adaptive resource scheduling.
+Unlike generic solutions, these studies propose container orchestration frameworks, interoperability tools, and security models tailored to the challenges of managing distributed containerized applications across heterogeneous cloud providers.
 1.
@@
 Additionally, containers are used for evaluating scheduling algorithms and enabling virtual network composition, demonstrating their versatility in addressing a wide range of technical challenges.
-Table 7. Roles of containers in multi-cloud environment
+Table 7: Roles of containers in multi-cloud environment
 Category Role of Containers ID
-Container-Based Resource and Service Management Container-Based Multi-Tenancy Solution (Ramos et al., 2021)
-Container-Driven Resource Sharing & Multi-Tenancy (Deng et al., 2022)
-Multi-Tenant Service Management via Containers (Slominski et al., 2015)
-Container-Enabled OS-Level Virtualization (Bélair et al., 2019)
-Container-Based Complexity Reduction (Shi et al., 2020)
-Containerized Service Hosting (Pham and Pham, 2015)
-Multi-Cloud Workload Migration via Containers (Janarthanan et al., 2018; Zeck and Bouroudjian, 2017)
-Seamless Workload Migration via Containers (Zeck and Bouroudjian, 2017)
-Container-Based Multi-Cloud Platform Execution (Serhiienko and Spillner, 2018)
-Container-Driven MicroCloud Scaling & Isolation (Baresi et al., 2016)
-Container-based Application Deployment Containerized Web-App Deployment (Almeida et al., 2014)
-Container-Driven App Deployment (Gattobigio et al., 2022; Pham et al., 2015)
-Containerized PaaS Prototyping (Verdugo et al., 2017)
-Data-Intensive App Deployment via Containers (Barna et al., 2017)
-Containerized App Management (Gupta et al., 2021)
-Container-Based Aneka Service Deployment (Buyya and Barreto, 2015)
-Container-Driven Microservices Deployment & Scaling (Rahman and Lama, 2019)
-Container-Based Multi-Cloud App Deployment & Scaling (Mfula and Nurminen, 2018)
-Containerized IoT App Deployment (Bou Ghantous and Gill, 2021)
-Containerization in App Development & IoT App Development & IoT Integration via Containers (Ouyang et al., 2022)
-Container-Based OpenVAS & Integration of Chef (Sheridan et al., 2017)
-Containers as Standardized Software Units (Al-shammari and Alwan, 2018)
-Container Placement & Orchestration in Fog Computing (Sami and Mourad, 2020)
-Container Deployment Speed & DevOps Integration (Lorenzo et al., 2019)
-Containerizing Apps for Privacy & Compliance (Dimitrijevic et al., 2019)
-Container-Driven Portability & Efficiency Improvement (Altran et al., 2022)
-Preventing Vendor Lock-In via Containers (Pellegrini et al., 2017)
-Container-Based Extensible App Development (Pop et al., 2016)
-Performance & Efficiency via Containers Scalable App Building via Containers (Pai and Kumar, 2021)
-Container-Driven SaaS Performance Isolation (Wang et al., 2020)
-Cloud Performance Enhancement via Containers (Saif et al., 2023)
-Evaluating Performance Impact of Containers (Rodigari et al., 2021)
-Cost Efficiency of Microservices via Containers (Guerrero et al., 2018a)
-Container-Based Cloud Platform Optimization (Kiss, 2018)
-Container Technology Features Container Usage in Edge Computing (Wong et al., 2019)
-Container-Enabled Lightweight Virtualization (Pahl, 2015)
-Key Features of Container Technology (Gao et al., 2017)
-Container-Based Edge Computing Virtualization (Carpio et al., 2020)
-Container Leverage in Kubernetes (Lazuka et al., 2022)
-Other Container Uses Container-Based Azure Data Storage (Bucur et al., 2018)
-Container & VM Support in Architecture (Bolivar et al., 2018)
-Blockchain-Based Multi-Cloud Reliability via Containers (Aral et al., 2020)
-Evaluating Scheduling Algorithm via Containers (Tang, 2021)
-Container-Based Virtual Network Composition (Alaluna et al., 2017)
-[…]
-3.2.2. Implementation Strategies
+Container-Based Resource and Service Management Container-Based Multi-Tenancy Solution [132]
+Container-Driven Resource Sharing & Multi-Tenancy [69]
+Multi-Tenant Service Management via Containers [44]
+Container-Enabled OS-Level Virtualization [71]
+Container-Based Complexity Reduction [97]
+Containerized Service Hosting [80]
+Multi-Cloud Workload Migration via Containers [123, 62]
+Seamless Workload Migration via Containers [62]
+Container-Based Multi-Cloud Platform Execution [45]
+Container-Driven MicroCloud Scaling & Isolation [31]
+Container-based Application Deployment Containerized Web-App Deployment [95]
+Container-Driven App Deployment [34, 86]
+Containerized PaaS Prototyping [120]
+Data-Intensive App Deployment via Containers [52]
+Containerized App Management [103]
+Container-Based Aneka Service Deployment [98]
+Container-Driven Microservices Deployment & Scaling [53]
+Container-Based Multi-Cloud App Deployment & Scaling [42]
+Containerized IoT App Deployment [90]
+Containerization in App Development & IoT App Development & IoT Integration via Containers [121]
+Container-Based OpenVAS & Integration of Chef [55]
+Containers as Standardized Software Units [112]
+Container Placement & Orchestration in Fog Computing [29]
+Container Deployment Speed & DevOps Integration [113]
+Containerizing Apps for Privacy & Compliance [61]
+Container-Driven Portability & Efficiency Improvement [114]
+Preventing Vendor Lock-In via Containers [73]
+Container-Based Extensible App Development [32]
+Performance & Efficiency via Containers Scalable App Building via Containers [59]
+Container-Driven SaaS Performance Isolation [66]
+Cloud Performance Enhancement via Containers [77]
+Evaluating Performance Impact of Containers [35]
+Cost Efficiency of Microservices via Containers [99]
+Container-Based Cloud Platform Optimization [63]
+Container Technology Features Container Usage in Edge Computing [111]
+Container-Enabled Lightweight Virtualization [60]
+Key Features of Container Technology [70]
+Container-Based Edge Computing Virtualization [30]
+Container Leverage in Kubernetes [74]
+Other Container Uses Container-Based Azure Data Storage [133]
+Container & VM Support in Architecture [40]
+Blockchain-Based Multi-Cloud Reliability via Containers [48]
+Evaluating Scheduling Algorithm via Containers [41]
+Container-Based Virtual Network Composition [124]
+Takeaway 4: The reviewed studies show that containers take on distinct roles in multi-cloud environments, particularly in addressing challenges such as provider heterogeneity, cross-cloud orchestration, and decentralized application deployment.
+Containers are used for multi-tenant resource management, cross-cloud service hosting, and workload migration.
+They also support edge and fog computing, compliance, and vendor lock-in prevention—roles less prominent in single-cloud contexts.
+This shows that multi-cloud scenarios introduce unique architectural and operational needs that shape how containers are implemented.
+3.2.2 Implementation Strategies
 Container implementation strategies in multi-cloud environment refer to the various approaches and techniques used to deploy, manage, and orchestrate containerized applications across multiple cloud platforms.
@@
 These approaches ensure efficient and secure communication between clouds, supporting smooth service operation and data transfer across diverse cloud infrastructures.
-[…]
-Table 8. Container implementation strategies in multi-cloud environment
+Table 8: Container implementation strategies in multi-cloud environment
 Category Subcategory Implementation Strategies ID
-Container Deployment and Management Deployment Strategies Automated Orchestration for Web Applications (Jha et al., 2019)
-Kubernetes for Containerized APPs (Gattobigio et al., 2022)
-Multi-cloud App Development Framework (Zhang et al., 2021)
-Modular Approach to Container Implementation (Pham et al., 2015)
-Deployment with Docker Images (Verdugo et al., 2017)
-Docker Image Creation and Management (Mittal and Risco-Martín, 2017)
-Real-time Kubernetes workloads in multi-cloud (Barletta et al., 2024)
-AWS multi-cloud container app deployment (Kataru et al., 2023)
-Multi-cloud container policy enforcement (Liu and Qin, 2024)
-Probes and orchestration for cloud containers (Tundo et al., 2024)
-Container Lifecycle Management Container Image Optimization (Al-shammari and Alwan, 2018)
-Efficient Image Reorganization (Lorenzo et al., 2019)
-Orchestrating with Docker and Kubernetes (Dimitrijevic et al., 2019)
-Executable Images on Multi-cloud Platforms (Serhiienko and Spillner, 2018)
-Privacy monitoring for containerized clouds (Barati et al., 2023)
-Multi-cloud container resource usage monitoring (Horchulhack et al., 2024)
-Container ecosystem for multi-cloud deployment (Wong et al., 2023)
-Containers in AWS ECS for cloud services (Petrovski and Gusev, 2024)
-Container management for multi-cloud MEC (Santos et al., 2023)
-IoT container deployment in multi-cloud (Solayman and Qasha, 2023)
-Data Management Firebase Cloud for Real-Time Database (Carpio et al., 2020)
-High Availability with SpyStorage MCSS (Shen et al., 2017)
-Data Storage with Object Store Service (Pop et al., 2016)
-Container Orchestration Cloud Orchestration SLA-Based Container Strategies (Ramos et al., 2021)
-Task Division and Scaling in Container Strategy (Bisht and Kaur, 2022)
-Location-Aware Service Brokering (Chen et al., 2022)
-Docker Containers for Multi-tenant Services (Slominski et al., 2015)
-Resource Acquisition for Aneka Containers (Buyya and Barreto, 2015)
-Cross-Cloud Container and VM Management (Janarthanan et al., 2018)
-Container Orchestration Technologies Overview (Sami and Mourad, 2020)
-Multi-cloud dynamic resource adaptation (Struhár et al., 2024)
-AI/ML orchestration in multi-cloud containers (Sofia et al., 2024)
-AI-driven container orchestration across clouds (Benmerar et al., 2023)
-Cross-cloud container orchestration tools (Alonso et al., 2023a)
-Kubernetes CI/CD for multi-cloud setups (Pandi et al., 2023)
-Microservices Orchestration Microservices Architecture in IoT Cloud Apps (Ouyang et al., 2022)
-ADAPT Deployment of Microservices (Alonso et al., 2019)
-MSA Benchmark with Docker and Kubernetes (Rahman and Lama, 2019)
-Kubernetes for power-efficient multi-cloud (Centofanti et al., 2024)
-Edge computing with Docker and Kubernetes (Hossain et al., 2023)
-Kubernetes for edge-cloud orchestration (Roy and Kalafatis, 2023)
-Kubernetes orchestration for multi-cloud containers (Sofia et al., 2023)
-Container Security Container Placement Secure Placement with Deep Learning (Deng et al., 2022)
-Container Placement Strategies (Altran et al., 2022)
-Dynamic microservice relocation across clouds (da Silva et al., 2023)
-Security Policies SLA-Oriented Performance Isolation (Wang et al., 2020)
-Access Control and Defense for Container Security (Gao et al., 2017)
-HIP Integration for Docker Security (Ranjbar et al., 2017)
-Secure Connectivity in Multi-tenant Environment (Alaluna et al., 2017)
-Signed URLs for Access Control in Multi-cloud (Graupner et al., 2015a)
-Performance and Scaling Strategies Performance Optimization LXC Virtual Cluster for Performance (Cui et al., 2016)
-Scalable Multi-layer Architecture with Containers (Barna et al., 2017)
-Reactive Auto-scaling and Resource Provisioning (Saif et al., 2023)
-Istio for Control Plane Impact and Latency Reduction (Rodigari et al., 2021)
-Auto-scaling containers across clouds (Dogani et al., 2023)
-Multi-cloud resource monitoring and provisioning (Marques et al., 2024)
-AI resource allocation in multi-cloud containers (Nagarajan et al., 2023)
-Scaling Strategies Edge Data Streaming and Real-Time Database (Carpio et al., 2020)
-MicroCloud Architecture for Efficient Scaling (Baresi et al., 2016)
-Distributed Workloads with Kubernetes (Lazuka et al., 2022)
-Hybrid Abstractions for Container and Host (Bélair et al., 2019)
-Cloud Service and Networking Cloud Service Management Multi-Tenant Service with Docker and Cloudant (Slominski et al., 2015)
-Abstraction Layer for Cloud Service Standardization (Pellegrini et al., 2017)
-OpenStack and IoT in multi-cloud (Merlino et al., 2024)
-Data indexing for multi-cloud exposure (Orzechowski et al., 2023)
-Networking Edge Scheduling with Location-Awareness (Wong et al., 2019)
-Openstack Networking with Kuryr (Bolivar et al., 2018)
-Federated Framework for Cloud Storage (Harwalkar et al., 2019a)
-Host Identity Protocol for Docker Networking (Ranjbar et al., 2017)
-3.3. Pattern and Strategies (RQ2)
-Table 9 presents various patterns and strategies employed in container-based applications in multi-cloud environment.
-These patterns and strategies are divided into various categories and subcategories, each of which highlights different aspects of multi-cloud architecture, management, security, resilience, and migration patterns and strategies.
-Notably, several studies (e.g., (Pahl, 2015; Graupner et al., 2015a; Shen et al., 2017; Pham and Pham, 2015)) appear across multiple subcategories because they report different patterns that address distinct aspects of cloud architecture and strategies.
-Every pattern is classified based on its specific focus, even when originating from the same study.
-Therefore, we categorized them into separate subcategories to reflect their unique contributions.
-Each category is briefly reported below.
+Container Deployment and Management Deployment Strategies Automated Orchestration for Web Applications [96]
+Kubernetes for Containerized APPs [34]
+Multi-cloud App Development Framework [101]
+Modular Approach to Container Implementation [86]
+Deployment with Docker Images [120]
+Docker Image Creation and Management [106]
+Real-time Kubernetes workloads in multi-cloud [82]
+AWS multi-cloud container app deployment [134]
+Multi-cloud container policy enforcement [130]
+Probes and orchestration for cloud containers [91]
+Container Lifecycle Management Container Image Optimization [112]
+Efficient Image Reorganization [113]
+Orchestrating with Docker and Kubernetes [61]
+Executable Images on Multi-cloud Platforms [45]
+Privacy monitoring for containerized clouds [135]
+Multi-cloud container resource usage monitoring [87]
+Container ecosystem for multi-cloud deployment [136]
+Containers in AWS ECS for cloud services [100]
+Container management for multi-cloud MEC [125]
+IoT container deployment in multi-cloud [38]
+Data Management Firebase Cloud for Real-Time Database [30]
+High Availability with SpyStorage MCSS [109]
+Data Storage with Object Store Service [32]
+Container Orchestration Cloud Orchestration SLA-Based Container Strategies [132]
+Task Division and Scaling in Container Strategy [39]
+Location-Aware Service Brokering [102]
+Docker Containers for Multi-tenant Services [44]
+Resource Acquisition for Aneka Containers [98]
+Cross-Cloud Container and VM Management [123]
+Container Orchestration Technologies Overview [29]
+Multi-cloud dynamic resource adaptation [137]
+AI/ML orchestration in multi-cloud containers [138]
+AI-driven container orchestration across clouds [104]
+Cross-cloud container orchestration tools [33]
+Kubernetes CI/CD for multi-cloud setups [117]
+Microservices Orchestration Microservices Architecture in IoT Cloud Apps [121]
+ADAPT Deployment of Microservices [105]
+MSA Benchmark with Docker and Kubernetes [53]
+Kubernetes for power-efficient multi-cloud [139]
+Edge computing with Docker and Kubernetes [37]
+Kubernetes for edge-cloud orchestration [46]
+Kubernetes orchestration for multi-cloud containers [118]
+Container Security Container Placement Secure Placement with Deep Learning [69]
+Container Placement Strategies [114]
+Dynamic microservice relocation across clouds [115]
+Security Policies SLA-Oriented Performance Isolation [66]
+Access Control and Defense for Container Security [70]
+HIP Integration for Docker Security [68]
+Secure Connectivity in Multi-tenant Environment [124]
+Signed URLs for Access Control in Multi-cloud [57]
+Performance and Scaling Strategies Performance Optimization LXC Virtual Cluster for Performance [126]
+Scalable Multi-layer Architecture with Containers [52]
+Reactive Auto-scaling and Resource Provisioning [77]
+Istio for Control Plane Impact and Latency Reduction [35]
+Auto-scaling containers across clouds [116]
+Multi-cloud resource monitoring and provisioning [88]
+AI resource allocation in multi-cloud containers [43]
+Scaling Strategies Edge Data Streaming and Real-Time Database [30]
+MicroCloud Architecture for Efficient Scaling [31]
+Distributed Workloads with Kubernetes [74]
+Hybrid Abstractions for Container and Host [71]
+Cloud Service and Networking Cloud Service Management Multi-Tenant Service with Docker and Cloudant [44]
+Abstraction Layer for Cloud Service Standardization [73]
+OpenStack and IoT in multi-cloud [36]
+Data indexing for multi-cloud exposure [110]
+Networking Edge Scheduling with Location-Awareness [111]
+Openstack Networking with Kuryr [40]
+Federated Framework for Cloud Storage [108]
+Host Identity Protocol for Docker Networking [68]
+Takeaway 5: The reviewed studies identify a wide range of implementation strategies that reflect the specific challenges of containerizing applications in multi-cloud environments.
+These include orchestrating containers across heterogeneous platforms, ensuring secure placement and access control under varying provider policies, and optimizing performance and scaling through AI-driven resource provisioning and hybrid abstractions.
+Unique to multi-cloud settings are strategies for managing distributed services, networking across clouds, and integrating edge and IoT services.
+3.3 Pattern and Strategies (RQ2)
+Table 9 presents a thematic classification of patterns and strategies employed in container-based applications within multi-cloud environments.
+These patterns were identified through qualitative coding and grouped into multiple categories and subcategories based on their architectural, management, security, resilience, and migration focus.
+Each row corresponds to a pattern extracted from the literature, and studies that reported multiple distinct patterns appear in more than one subcategory (e.g., [60, 57, 109, 80]).
+Table 9 is structured to allow readers to quickly identify and compare solution strategies across different technical domains.
+To improve interpretability, we provide a brief overview of each category below.
 1.
@@
 Leading strategies include the Multi-cloud Computing Strategy, Hybrid/Multi-cloud Approach, and MAPE-K control loop, all of which emphasize seamless integration and resource optimization across cloud platforms.
-The Container Management Strategies subcategory contains 7 strategies that address the management of containers in cloud environments.
-Key strategies include Docker Container Images and Lightweight Virtualization, both of which streamline container management and enhance resource efficiency in multi-cloud setups.
-The Edge Computing and IoT Strategies subcategory, with 8 strategies, highlights the integration of edge computing and IoT devices with cloud systems.
-Notable strategies include Mobile Edge Computing (MEC) and Fog Computing, which focus on distributing computation and storage closer to IoT devices for improved performance.
-Finally, the Resource Management Strategies subcategory includes 7 strategies that focus on efficient resource allocation in cloud environments.
-Leading approaches like Resource Allocation based on SLA levels, Container Orchestration, Kubernetes, Minikube, and AI-Driven Edge-Cloud Architecture emphasize intelligent resource distribution and orchestration across cloud and edge environments.
+The Container Management Strategies subcategory includes 7 approaches that focus on effectively managing containers in cloud environments.
+Key examples are Docker Container Images and Lightweight Virtualization, both of which help simplify container operations and improve resource efficiency in multi-cloud settings.
+The Edge Computing and IoT Strategies subcategory consists of 8 strategies that explore how edge computing and IoT devices can be integrated with cloud systems.
+Noteworthy strategies such as Mobile Edge Computing (MEC) and Fog Computing aim to bring computation and storage closer to IoT devices, enhancing system responsiveness and performance.
+Lastly, the Resource Management Strategies subcategory comprises 7 strategies dedicated to optimizing resource usage in cloud environments.
+Prominent approaches like Resource Allocation based on SLA levels, Container Orchestration, Kubernetes, Minikube, and AI-Driven Edge-Cloud Architecture focus on smart resource distribution and coordination across cloud and edge platforms.
 3.
@@
 These strategies are essential for organizations looking to leverage cloud technologies while preserving the integrity of their existing systems during the migration process.
-Table 9. Patterns and strategies for container-based applications a multi-cloud environment
+Takeaway 6: The studies reveal that when containerized applications are deployed in multi-cloud environments, architecture and management patterns must be adapted to support distributed, heterogeneous infrastructure.
+Patterns like Microservices and SOA are extended with multi-cloud orchestration and federation capabilities, while service mesh and chaining are tailored for secure, cross-cloud communication.
+Deployment and fault-tolerant patterns—such as black-Green deployment, decentralized orchestration, and blockchain-driven resilience.
+Table 9: Patterns and strategies for container-based applications in multi-cloud environment
 Categories Subcategories Patterns ID
-Cloud Architectural Patterns and Models Architecture Patterns Microservice Architecture (Jha et al., 2019; Verdugo et al., 2017; Ouyang et al., 2022; Ranjbar et al., 2017; Quint and Kratzke, 2018; Pahl, 2015; Mittal and Risco-Martín, 2017; Sami and Mourad, 2020; Guerrero et al., 2018a; Centofanti et al., 2024; Hossain et al., 2023; Tiwari and Sharma, 2023)
-Service-Oriented Architecture (SOA) (Buyya and Barreto, 2015; Casola et al., 2017; Gundu et al., 2020; Merlino et al., 2024)
-Multitier Architecture, Layered Architecture (Jha et al., 2019; Ramos et al., 2021; Bucur et al., 2018; Barati et al., 2023; Horchulhack et al., 2024; Liu and Qin, 2024; Tundo et al., 2024)
-Client-Server Architecture (Hua et al., 2020; Bucur et al., 2018)
-Component-based Architecture (Wong et al., 2019)
-Event-driven Architecture (Wurster et al., 2018)
-Multi-agent Architecture (Mfula and Nurminen, 2018; Saif et al., 2023)
-Master-worker Nodes Architecture (Sami and Mourad, 2020)
-Netflix Zuul-based Seeker Component (Slominski et al., 2015)
-Multi-tenant Cloud Service Architecture (Slominski et al., 2015)
-Stateful Engine Architecture (Slominski et al., 2015)
-Distributed Multi-Cloud Native Architecture (Alonso et al., 2023a)
-Plugin-Based Deployment Automation Architecture (Harzenetter et al., 2023)
-Cloud-Edge Integrated Robotics Architecture (Huang and Wang, 2023)
-Hybrid Malware Detection Architecture (Wang et al., 2023)
-Cloudfront-Enhanced Container Architecture (Kataru et al., 2023)
-AI-Driven Edge-Cloud Architecture (Benmerar et al., 2023)
-Metadata-Driven Architecture (Orzechowski et al., 2023)
-Blockchain-Based Architecture (Ahmad and Aujla, 2023)
-Cross-Layered Edge-Cloud Architecture (Sofia et al., 2023; Santos et al., 2023)
-Communication and Networking Patterns Service Mesh (Gattobigio et al., 2022)
-Service Chaining (SC) (Bhamare et al., 2015)
-Multi-Cloud APIs (Graupner et al., 2015a)
-Third-Party APIs (Graupner et al., 2015a)
-Service Request Broker (Zhang et al., 2021)
-Publish and Subscribe Communication Protocol (Trihinas et al., 2015)
-Synchronous Network Communication Protocols (Almeida et al., 2014)
-Asynchronous Network Communication Protocols (Almeida et al., 2014)
-Network Function Virtualization (NFV) (Bhamare et al., 2015)
-Service Discovery (Pop et al., 2016)
-Blockchain-Based Identity and Access Management (IAM) Architecture (Fugkeaw, 2023)
-Federated Cloud-Edge Architecture (Roy and Kalafatis, 2023)
-Deployment Pattern black-Green deployment (Alonso et al., 2019; Sheridan et al., 2017)
-Object store service (Pop et al., 2016)
-Multi-cloud Deployment Model (Sabbioni et al., 2020)
-Distributed Deployment Model (Sabbioni et al., 2020)
-Multi-Cloud Deployment Architecture (Petrovski and Gusev, 2024; Pandi et al., 2023)
-Decentralized Orchestration Architecture (Sofia et al., 2024)
-Policy-Driven Orchestration Architecture (Lee and Nam, 2023)
-DevOps-Oriented Architecture (Solayman and Qasha, 2023)
-Service Models Platform-as-a-Service (PaaS) (Pahl, 2015; Shi et al., 2020)
-Software-as-a-Service (Serhiienko and Spillner, 2018)
-Infrastructure-as-a-Service (IaaS) (Pahl, 2015)
-Function-as-a-Service (FaaS) (Pahl, 2015; Ramesh et al., 2023)
-Multi-Cloud Orchestration Architecture (Pandi et al., 2023)
-Serverless Container-Oriented Architecture (Shi et al., 2020)
+Cloud Architectural Patterns and Models Architecture Patterns Microservice Architecture [96, 120, 121, 68, 50, 60, 106, 29, 99, 139, 37, 131]
+Service-Oriented Architecture (SOA) [98, 127, 64, 36]
+Multitier Architecture, Layered Architecture [96, 132, 133, 135, 87, 130, 91]
+Client-Server Architecture [93, 133]
+Component-based Architecture [111]
+Event-driven Architecture [140]
+Multi-agent Architecture [42, 77]
+Master-worker Nodes Architecture [29]
+Netflix Zuul-based Seeker Component [44]
+Multi-tenant Cloud Service Architecture [44]
+Stateful Engine Architecture [44]
+Distributed Multi-Cloud Native Architecture [33]
+Plugin-Based Deployment Automation Architecture [141]
+Cloud-Edge Integrated Robotics Architecture [49]
+Hybrid Malware Detection Architecture [142]
+Cloudfront-Enhanced Container Architecture [134]
+AI-Driven Edge-Cloud Architecture [104]
+Metadata-Driven Architecture [110]
+Blockchain-Based Architecture [143]
+Cross-Layered Edge-Cloud Architecture [118, 125]
+Communication and Networking Patterns Service Mesh [34]
+Service Chaining (SC) [47]
+Multi-Cloud APIs [57]
+Third-Party APIs [57]
+Service Request Broker [101]
+Publish and Subscribe Communication Protocol [84]
+Synchronous Network Communication Protocols [95]
+Asynchronous Network Communication Protocols [95]
+Network Function Virtualization (NFV) [47]
+Service Discovery [32]
+Blockchain-Based Identity and Access Management (IAM) Architecture [144]
+Federated Cloud-Edge Architecture [46]
+Deployment Pattern black-Green deployment [105, 55]
+Object store service [32]
+Multi-cloud Deployment Model [65]
+Distributed Deployment Model [65]
+Multi-Cloud Deployment Architecture [100, 117]
+Decentralized Orchestration Architecture [138]
+Policy-Driven Orchestration Architecture [145]
+DevOps-Oriented Architecture [38]
+Service Models Platform-as-a-Service (PaaS) [60, 97]
+Software-as-a-Service [45]
+Infrastructure-as-a-Service (IaaS) [60]
+Function-as-a-Service (FaaS) [60, 92]
+Multi-Cloud Orchestration Architecture [117]
+Serverless Container-Oriented Architecture [97]
 Cloud Management and Resource Allocation Strategies Multi-Cloud Management Strategies Multi-cloud computing strategy S78, S81
-Hybrid cloud architecture (Wang et al., 2019)
-Hybrid/Multi-cloud Approach (Gattobigio et al., 2022; Wahab et al., 2016)
-Introducing multi-cloud Middleware (Casola et al., 2017; Acquaviva et al., 2017)
-MAPE-K control loop (Fadda et al., 2019; Saif et al., 2023)
-Connecting to multiple cloud service providers (Pop et al., 2016)
-Multi-cloud load balancing (Wang et al., 2020; Pop et al., 2016)
-Multi-Cloud BP provisioning (Kritikos et al., 2019)
-Cross-level orchestration of cloud services (Kritikos et al., 2019)
-Cross-level monitoring and adaptation of BPs (Kritikos et al., 2019)
-Container Management Strategies Linux Container (LXC) project (Pahl, 2015)
-Container engine (Pahl, 2015)
-Docker Container Images (Pahl, 2015)
-Lightweight virtualization (Pahl, 2015)
-Portable application packaging (Pahl, 2015)
-One-container-per-app approach (Pahl, 2015)
-Data volumes and data volume containers (Pahl, 2015)
-Edge Computing and IoT Strategies Mobile Edge Computing (MEC) (Wong et al., 2019; Lazuka et al., 2022; Dreibholz et al., 2019; Bucur et al., 2018)
-Fog computing (Al-shammari and Alwan, 2018)
-Edge Services (Carpio et al., 2020)
-Connecting IoT Edge Devices to the Cluster (Gattobigio et al., 2022)
-Cloud of Things (CoT) (Gattobigio et al., 2022)
-Multi-Access Edge Computing (MEC) Architecture (Santos et al., 2023)
-Cross-Layered Edge-Cloud Architecture (Sofia et al., 2023)
-Cloud-Edge Integrated Robotics Architecture (Huang and Wang, 2023)
-AI-Driven Edge-Cloud Architecture (Benmerar et al., 2023)
-Resource Management Strategies Resource allocation based on different SLA levels (Wang et al., 2020)
-Container Orchestration, Kubernetes, Minikube (Pahl, 2015; Gupta et al., 2021; Sami and Mourad, 2020)
-Rate-Based Stream Processing Architecture (Nguyen and Chien, 2023)
-Multi-Agent Resource Optimization Architecture (Nagarajan et al., 2023)
-Cloud Security and Resilience Strategies Security and Resiliency Patterns Security-by-Design approach (Buyya and Barreto, 2015)
-Encryption of files (Graupner et al., 2015a)
-Standardized APIs for file transfer (Graupner et al., 2015a)
-Centralized/External ACM Service (Graupner et al., 2015a)
-Distributed ACM Service (Graupner et al., 2015a)
-Dynamic Switching between Authentication Methods (Graupner et al., 2015a)
-ACM based on Signed URLs (Graupner et al., 2015a)
-Attribute-Based Encryption and Signature (Shen et al., 2017)
-Secure data sharing (Shen et al., 2017)
-Data access control (Shen et al., 2017)
-Local encryption and signing (Shen et al., 2017)
-Authorized Tokens (Shen et al., 2017)
-Byzantine quorum protocol (Shen et al., 2017)
-Blockchain-Driven SLA Management Architecture (Neeraj et al., 2023)
-Hybrid Malware Detection Architecture (Wang et al., 2023)
-Fault-tolerance Strategies Fault-tolerance with Redundant Engines (Slominski et al., 2015)
-Multi-cloud systems fault-tolerant workflow (Tang, 2021)
-RAFT Consensus Algorithm (Naik, 2016a)
-Cloud Migration Strategies Service-oriented Migration (Pham and Pham, 2015)
-Application-centric Migration (Pham and Pham, 2015)
-Image-based Migration (Pham and Pham, 2015)
-Migration to a Virtualized Container (Pham and Pham, 2015)
-[…]
-3.4. Quality Attributes and Tactics (RQ3)
-Table 10 provides a comprehensive overview of the quality attributes and related tactics for containerized applications in multi-cloud environment.
-It is based on data extracted from 121 selected studies, which are organized following the ISO 25010 standard.
-This standard defines high-level QAs along with their sub-characteristics.
-Table 10 lists nine QAs, namely Performance (Efficiency), Security, Compatibility, Scalability, Reliability, Portability, Flexibility, Maintainability, and Usability.
-Each QA has its related terms, associated study numbers, and corresponding tactics to achieve them.
-In Table 10, due to limited space, we provide only five tactics for each QA.
-The complete set of 70 tactics is listed in the provided dataset (Waseem et al., 2024).
-However, it is important to note that these tactics are scattered across different studies.
-Below we briefly discuss how each QA can be achieved by applying the mentioned tactics.
+Hybrid cloud architecture [75]
+Hybrid/Multi-cloud Approach [34, 72]
+Introducing multi-cloud Middleware [127, 146]
+MAPE-K control loop [85, 77]
+Connecting to multiple cloud service providers [32]
+Multi-cloud load balancing [66, 32]
+Multi-Cloud BP provisioning [76]
+Cross-level orchestration of cloud services [76]
+Cross-level monitoring and adaptation of BPs [76]
+Container Management Strategies Linux Container (LXC) project [60]
+Container engine [60]
+Docker Container Images [60]
+Lightweight virtualization [60]
+Portable application packaging [60]
+One-container-per-app approach [60]
+Data volumes and data volume containers [60]
+Edge Computing and IoT Strategies Mobile Edge Computing (MEC) [111, 74, 122, 133]
+Fog computing [112]
+Edge Services [30]
+Connecting IoT Edge Devices to the Cluster [34]
+Cloud of Things (CoT) [34]
+Multi-Access Edge Computing (MEC) Architecture [125]
+Cross-Layered Edge-Cloud Architecture [118]
+Cloud-Edge Integrated Robotics Architecture [49]
+AI-Driven Edge-Cloud Architecture [104]
+Resource Management Strategies Resource allocation based on different SLA levels [66]
+Container Orchestration, Kubernetes, Minikube [60, 103, 29]
+Rate-Based Stream Processing Architecture [119]
+Multi-Agent Resource Optimization Architecture [43]
+Cloud Security and Resilience Strategies Security and Resiliency Patterns Security-by-Design approach [98]
+Encryption of files [57]
+Standardized APIs for file transfer [57]
+Centralized/External ACM Service [57]
+Distributed ACM Service [57]
+Dynamic Switching between Authentication Methods [57]
+ACM based on Signed URLs [57]
+Attribute-Based Encryption and Signature [109]
+Secure data sharing [109]
+Data access control [109]
+Local encryption and signing [109]
+Authorized Tokens [109]
+Byzantine quorum protocol [109]
+Blockchain-Driven SLA Management Architecture [147]
+Hybrid Malware Detection Architecture [142]
+Fault-tolerance Strategies Fault-tolerance with Redundant Engines [44]
+Multi-cloud systems fault-tolerant workflow [41]
+RAFT Consensus Algorithm [148]
+Cloud Migration Strategies Service-oriented Migration [80]
+Application-centric Migration [80]
+Image-based Migration [80]
+Migration to a Virtualized Container [80]
+3.4 Quality Attributes and Tactics (RQ3)
+Table 10 presents a thematic classification of QAs and their associated implementation tactics for containerized applications in multi-cloud environments.
+These QAs and tactics were derived through thematic analysis of data extracted from 121 primary studies and are organized according to the ISO 25010 standard.
+This standard defines high-level software quality characteristics and their sub-attributes, which served as a guiding framework for our classification.
+The table includes nine core QAs, namely Performance (Efficiency), Security, Compatibility, Scalability, Reliability, Portability, Flexibility, Maintainability, and Usability.
+For each QA, the table lists representative related terms, example study references, and commonly applied tactics.
+Due to space constraints, only five example tactics per QA are shown in Table 10, while the complete list of 70 tactics is available in the replication package [11].
+These tactics are distributed across multiple studies and highlight recurring practices for achieving specific quality goals.
+To help readers navigate this classification, we briefly explain below how each QA is addressed in the literature and what patterns emerge in terms of implementation strategies.
 Performance (Efficiency): Performance or efficiency is the most dominant QA, frequently discussed in 60 (49.58%) of the selected studies, along with related terms or characteristics such as time behavior, resource utilization, and capacity.
@@
 Reliability: This QA is about the system’s correct operation and consistent performance without failure over a specified period, while ensuring the availability of the system.
-We identified XX studies that report on reliability, along with characteristics such as maturity, availability, and fault tolerance.
-We identified nine tactics overall (see Table 10 and the QA and Tactics Sheet in (Waseem et al., 2024)) to achieve reliability from the selected studies.
+We identified 33 studies that report on reliability, along with characteristics such as maturity, availability, and fault tolerance.
+We identified nine tactics overall (see Table 10 and the QA and Tactics Sheet in [11]) to achieve reliability from the selected studies.
 These include deploying redundant engines for fault tolerance, distributing resources and replicating applications for improved response time, enforcing redundancy and distributed services for availability, deploying parallel search and multi-cloud distribution, and building fault-tolerant systems to enhance system reliability.
@@
 Flexibility: This QA represents the system’s capacity to accommodate new features based on user requirements with minimal effort.
-We identified 10 studies (see Table 10 and the QA and Tactics Sheet in (Waseem et al., 2024)) that report this QA, along with five tactics that can help achieve it.
+We identified 10 studies (see Table 10 and the QA and Tactics Sheet in [11]) that report this QA, along with five tactics that can help achieve it.
 For example, flexibility can be improved through tactics such as service-oriented architecture, on-demand dynamic allocation of resources across different cloud platforms, containers, container orchestration, model-driven development, risk analysis, and cross-cloud service orchestration.
@@
 For instance, a feedback-loop controller is effective in a multi-cloud context as it dynamically adjusts resources across different cloud platforms based on user interactions, ensuring a seamless and responsive user experience tailored to the multi-cloud setup.
-[…]
-Table 10. Quality attributes and tactics for containerized applications in multi-cloud environment
+Takeaway 7: QAs and their associated tactics from the reviewed studies reveal that ensuring performance, security, reliability, and other QAs in containerized applications requires context-specific adaptations when operating in multi-cloud environments.
+Unlike single-cloud setups, these environments introduce heterogeneity in infrastructure, distributed control, and varying security models across providers.
+As a result, tactics such as SLA-aware resource optimization, federated access control, decentralized fault-tolerance mechanisms, and cross-platform provisioning are essential to achieving QA goals.
+Table 10: Quality attributes and tactics for containerized applications in multi-cloud environment
 Quality Attribute Related Terms ID Tactics
-Performance (Efficiency) Time behavior, Resource utilization, Capacity, Throughput, Response time (Almeida et al., 2014; Jha et al., 2019; Hua et al., 2020; Deng et al., 2022; Verdugo et al., 2017; Chen et al., 2022; Pham and Pham, 2015), (Ouyang et al., 2022; Ferry et al., 2014; Cui et al., 2016; Wang et al., 2020; Gao et al., 2017; Alonso et al., 2019; Gupta et al., 2021), (Rios et al., 2017; Carpio et al., 2020; Lorenzo et al., 2019; Dimitrijevic et al., 2019; Pandey et al., 2022; Altran et al., 2022; Shi et al., 2020), (Dreibholz et al., 2019; Wurster et al., 2018; Libardi et al., 2014; Harwalkar et al., 2019b; Casola et al., 2017; Bucur et al., 2018; Tricomi et al., 2017), (Sitaram et al., 2018; Rodigari et al., 2021; Janarthanan et al., 2018; Rahman and Lama, 2019; Harwalkar et al., 2019a; Zeck and Bouroudjian, 2017; Aral et al., 2020) Efficient resource optimization and allocation
+Performance (Efficiency) Time behavior, Resource utilization, Capacity, Throughput, Response time [95, 96, 93, 69, 120, 102, 80], [121, 89, 126, 66, 70, 105, 103], [56, 30, 113, 61, 67, 114, 97], [122, 140, 79, 78, 127, 133, 128], [81, 35, 123, 53, 108, 62, 48] Efficient resource optimization and allocation
 Utilizing machine learning techniques for performance optimization
@@
 Performance-oriented Service Level Agreements (SLA)
-Security Confidentiality, Authenticity, Access control, Authorization, Privacy, Trustworthiness, Integrity (Almeida et al., 2014; Gattobigio et al., 2022; Deng et al., 2022; Verdugo et al., 2017; Pham and Pham, 2015; Wang et al., 2020; Pahl, 2015), (Gao et al., 2017; Gupta et al., 2021; Rios et al., 2017; Dimitrijevic et al., 2019; Pandey et al., 2022; Bélair et al., 2019; Wurster et al., 2018), (Libardi et al., 2014; Casola et al., 2017; Bucur et al., 2018; Bolivar et al., 2018; Tricomi et al., 2017; Rodigari et al., 2021; Di Modica et al., 2019), (Harwalkar et al., 2019a; Zeck and Bouroudjian, 2017; Aral et al., 2020; Kiss, 2018; Graupner et al., 2015a; Alaluna et al., 2017; Guechi and Maamri, [n.
-d.]; Ranjbar et al., 2017; Gundu et al., 2020) Encryption and strong authentication mechanisms
+Security Confidentiality, Authenticity, Access control, Authorization, Privacy, Trustworthiness, Integrity [95, 34, 69, 120, 80, 66, 60], [70, 103, 56, 61, 67, 71, 140], [79, 127, 133, 40, 128, 35, 129], [108, 62, 48, 63, 57, 124, 58, 68, 64] Encryption and strong authentication mechanisms
 Implementation of firewalls, VPNs, or SDNs for network security
@@
 addressing CSP incompatibilities
-Compatibility Interoperability, Co-existence (Pham and Pham, 2015; Ouyang et al., 2022; Cui et al., 2016; Pahl, 2015; Alonso et al., 2019; Baresi et al., 2016; Wurster et al., 2018), (Trihinas et al., 2015; Harwalkar et al., 2019b; Pellegrini et al., 2017; Harwalkar et al., 2019a; Graupner et al., 2015a; Ranjbar et al., 2017; Quint and Kratzke, 2018; Ferry et al., 2013) Implementing lightweight communication protocols and modes
+Compatibility Interoperability, Co-existence [80, 121, 126, 60, 105, 31, 140], [84, 78, 73, 108, 57, 68, 50, 51] Implementing lightweight communication protocols and modes
 Standardization of interfaces for seamless integration
@@
 Interoperability standardization and federation between clouds
-Scalability Capacity, Extensibility, Elasticity, Throughput, Responsiveness (Almeida et al., 2014; Pham et al., 2015; Verdugo et al., 2017; Chen et al., 2022; Pham and Pham, 2015; Ouyang et al., 2022; Slominski et al., 2015), (Pai and Kumar, 2021; Alhamazani et al., 2014; Cui et al., 2016; Wong et al., 2019; Pahl, 2015; Gao et al., 2017; Alonso et al., 2019), (Gupta et al., 2021; Mittal and Risco-Martín, 2017; Carpio et al., 2020; Shi et al., 2020; Baresi et al., 2016; Wurster et al., 2018; Trihinas et al., 2015), (Bucur et al., 2018; Tricomi et al., 2017; Rahman and Lama, 2019; Aral et al., 2020; Kiss, 2018; Alaluna et al., 2017; Mfula and Nurminen, 2018), (Pop et al., 2016; Ranjbar et al., 2017; Ferry et al., 2013; Kritikos et al., 2019)
+Scalability Capacity, Extensibility, Elasticity, Throughput, Responsiveness [95, 86, 120, 102, 80, 121, 44], [59, 83, 126, 111, 60, 70, 105], [103, 106, 30, 97, 31, 140, 84], [133, 128, 53, 48, 63, 124, 42], [32, 68, 51, 76]
 Combining Containerization and Microservices for enhanced scalability
@@
 Leveraging Orchestration, Federated Network,and Geographic Placement
-Reliability Maturity, Availability, Fault tolerance, Recoverability (Almeida et al., 2014; Jha et al., 2019; Ramos et al., 2021; Sabbioni et al., 2020; Naik, 2016a; Slominski et al., 2015; Alhamazani et al., 2014), (Ferry et al., 2014; Wong et al., 2019; Gao et al., 2017; Alonso et al., 2019; Al-shammari and Alwan, 2018; Sami and Mourad, 2020; Carpio et al., 2020), (Dimitrijevic et al., 2019; Wurster et al., 2018; Libardi et al., 2014; Buyya and Barreto, 2015; Acquaviva et al., 2017; Bucur et al., 2018; Bolivar et al., 2018), (del Castillo et al., 2013; Tricomi et al., 2017; Janarthanan et al., 2018; Harwalkar et al., 2019a; Zeck and Bouroudjian, 2017; Aral et al., 2020; Tang, 2021), (Graupner et al., 2015a; Guechi and Maamri, [n.
-d.]; Mfula and Nurminen, 2018; Shen et al., 2017; Ferry et al., 2013) Deploying Redundant Engines for Fault-Tolerance
+Reliability Maturity, Availability, Fault tolerance, Recoverability [95, 96, 132, 65, 148, 44, 83], [89, 111, 70, 105, 112, 29, 30], [61, 140, 79, 98, 146, 133, 40], [107, 128, 123, 108, 62, 48, 41], [57, 58, 42, 109, 51] Deploying Redundant Engines for Fault-Tolerance
 Distributing Resources and Replicating Applications for Improved
@@
 Building Fault-Tolerant Systems
-Portability Installability, Replaceability, Adaptability (Pai and Kumar, 2021; Pahl, 2015; Gao et al., 2017; Mittal and Risco-Martín, 2017; Rahman and Lama, 2019; Ranjbar et al., 2017) Virtual Machine-Based Packaging
+Portability Installability, Replaceability, Adaptability [59, 60, 70, 106, 53, 68] Virtual Machine-Based Packaging
 Docker Image-Based Packaging
@@
 Adaptive Rule-Based System Modification
-Flexibility Functional completeness, Functional correctness, Functional appropriateness (Ouyang et al., 2022; Cui et al., 2016; Alonso et al., 2019; Mittal and Risco-Martín, 2017; Shi et al., 2020; Di Modica et al., 2019; Janarthanan et al., 2018), (Lazuka et al., 2022; Pop et al., 2016; Kritikos et al., 2019) Service-oriented Architecture
+Flexibility Functional completeness, Functional correctness, Functional appropriateness [121, 126, 105, 106, 97, 129, 123], [74, 32, 76] Service-oriented Architecture
 Dynamic allocation of resources across different cloud platforms based on demand
@@
 Cross-cloud service orchestration,
-Maintainability Modularity, Reusability, Analyzability, Modfiabality, Testability (Pham et al., 2015; Slominski et al., 2015; Gao et al., 2017; Alonso et al., 2019; Rahman and Lama, 2019; Mfula and Nurminen, 2018; Ferry et al., 2013) Microservices, easier maintenance
+Maintainability Modularity, Reusability, Analyzability, Modfiabality, Testability [86, 44, 70, 105, 53, 42, 51] Microservices, easier maintenance
 Infrastructure Provisioning as Code (IaC)
@@
 Continuous Integration and Deployment Automation
-Usability Appropriateness recognizability, Learnability, Operability , User interface aesthetics, Accessibility (Verdugo et al., 2017; Slominski et al., 2015; Alonso et al., 2019, 2019; Dimitrijevic et al., 2019; Mfula and Nurminen, 2018) Feedback-loop controller for multi-cloud Infrastructure
+Usability Appropriateness recognizability, Learnability, Operability , User interface aesthetics, Accessibility [120, 44, 105, 105, 61, 42] Feedback-loop controller for multi-cloud Infrastructure
 Container Orchestration (e.g., Kubernetes)
@@
 Error Handling and Feedback Mechanisms
-3.5. Security Challenges and Solution Framework (RQ4 and RQ5)
+3.5 Security Challenges and Solution Framework (RQ4 and RQ5)
 This section presents a framework for addressing security challenges in containerized applications in multi-cloud environment.
@@
 Data Security in Container-based Applications category covers multiple challenges related to data security in container-based applications, including data protection, database security, data compliance, and secure data transfer in a multi-cloud environment.
-Out of the 121 reviewed studies, 11 specifically highlighted these issues and provided corresponding solutions (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)).
+Out of the 121 reviewed studies, 11 specifically highlighted these issues and provided corresponding solutions (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in [11]).
 The identified solutions focus on implementing robust encryption techniques, secure container orchestration, and compliance with regulatory standards such as GDPR and HIPAA.
+Table 11: Identified security challenges and solutions for containerized applications in multi-cloud environment
+ID Challenge Solution
+Category 1: Data Security
+[44] Data Protection Multi-Cloud Data Protection through Secure Containers
+[103] Database Security Enhanced multi-cloud Database Security via Containerization
+[61] Data Compliance Multi-cloud Legal Compliance and Data Protection
+[67] Data Security Secure CSP Interoperability and Selection in multi-cloud
+[146] Data Ownership and Privacy Multi-cloud Security via NoMISHAP Service Abstraction
+[81] Data Transfer Security Secure multi-cloud Integration via OpenStack Standardization
+[90] Cloud Data Management Multi-cloud Data Security via DRA Framework
+[143] Data leaks and access risks Cryptography, smart contracts, encryption
+[138] Decentralized data security challenge Secure data handling with decentralized AI
+[147] SLA data integrity challenge Blockchain for SLA security and integrity
+[125] Multi-cloud data privacy issue Stringent security controls for data protection
+Category 2: Access and Communication Control
+[34] Connectivity and Access Control Multi-cloud Containerized Application Security
+[122] Secure Communication Middleware Secure Deployment for multi-cloud
+[98] Provider-Agnostic Operations Aneka-Based Security and Performance for Containerized multi-cloud Applications
+[37] Microservices privacy and access control Encryption, authentication, service mesh for security
+[144] Decentralized IAM security issue Blockchain for IAM security and data integrity
+[46] Secure communication across edge-cloud environments WireGuard for secure edge-cloud communication
+[104] Multi-domain edge communication security Secure communication and AI anomaly detection
+[49] Secure robot-cloud communication Secure communication and encryption for data integrity
+Category 3: Container Security
+[69] Co-resident Security Secure Placement Strategy via Deep RL
+[70] Container Isolation Power and Leakage Management for multi-cloud Security
+[71] Container and Host Security Host-Container Cooperation for multi-cloud Defense
+[142] Container security and malware detection DockerWatch for malicious activity detection
+[131] Multi-tenant serverless security challenge Security best practices for serverless deployments
+Category 4: Infrastructure and Deployment Management Security
+[89] Data Deployment Secure Private Data Deployment Strategy
+[140] Infrastructure Orchestration Standard Interface Implementation for Serverless Security
+[79] Storage Security Multi-cloud Storage Security with MSSF
+[50] Migration Security Infrastructure-Aware and Agnostic Secure Migration
+[38] Distributed infrastructure security challenge Orchestration security for IoT data protection
+[134] Securing AWS app deployments AWS best practices for CI/CD security
+Category 5: Security Monitoring and Breach Prevention
+[55] System Vulnerabilities Automated Security Evaluation and Bootstrapping
+[84] Breach Detection and Prevention Metric Filtering and Fault Recovery for multi-cloud Security
+[145] Network policy misconfiguration risks Automated network policy discovery and verification
+[119] Stream processing security risks RBAM framework for secure data handling
+Category 6: Trust and Compliance Management
+[127] Application Security and Compliance Risk Analysis and Security SLAs with MUSA Framework
+[78] Data Security and Standardization OpenStack API-based multi-cloud Security and Interoperability
+[128] Network Security Inter-cloud Communication Protection with SFC
+[72] Trust and Community Formation Trust-based Hedonic Coalitions for multi-cloud Security
+[76] Orchestration and Adaptation Multi-cloud Service Orchestration and Cross-Level Adaptation
+[100] Multicloud compliance and security issue AWS security measures for multicloud environments
+[33] Cross-cloud security and compliance Security frameworks for multi-cloud environment
+Category 7: Placement Strategy
+[29] Placement Strategy Security Model for Volunteering Fog Services
+[117] Multi-cloud security concerns Security groups, VPCs, and infrastructure security
+[92] Multicloud environment security issue Multicloud security for risk diversification
+[118] Edge-cloud security concerns Enhanced security protocols for decentralized environments
+[141] Cross-technology update security risks Integrated security in workflows for management
+[43] Resource allocation security risks Security measures in resource allocation framework
 The solutions also include using Role-Based and Attribute-Based Access Controls (RBAC and ABAC) to manage secure access effectively.
@@
 Furthermore, comprehensive frameworks like DRA emphasize secure data management, ensuring privacy, fault tolerance, and secure data transfers across different cloud platforms.
-Table 11. Identified security challenges and solutions for containerized applications in multi-cloud environments
-ID Challenge Solution
-Category 1: Data Security
-(Slominski et al., 2015) Data Protection Multi-Cloud Data Protection through Secure Containers
-(Gupta et al., 2021) Database Security Enhanced multi-cloud Database Security via Containerization
-(Dimitrijevic et al., 2019) Data Compliance Multi-cloud Legal Compliance and Data Protection
-(Pandey et al., 2022) Data Security Secure CSP Interoperability and Selection in multi-cloud
-(Acquaviva et al., 2017) Data Ownership and Privacy Multi-cloud Security via NoMISHAP Service Abstraction
-(Sitaram et al., 2018) Data Transfer Security Secure multi-cloud Integration via OpenStack Standardization
-(Bou Ghantous and Gill, 2021) Cloud Data Management Multi-cloud Data Security via DRA Framework
-(Ahmad and Aujla, 2023) Data leaks and access risks Cryptography, smart contracts, encryption
-(Sofia et al., 2024) Decentralized data security challenge Secure data handling with decentralized AI
-(Neeraj et al., 2023) SLA data integrity challenge Blockchain for SLA security and integrity
-(Santos et al., 2023) Multi-cloud data privacy issue Stringent security controls for data protection
-Category 2: Access and Communication Control
-(Gattobigio et al., 2022) Connectivity and Access Control Multi-cloud Containerized Application Security
-(Dreibholz et al., 2019) Secure Communication Middleware Secure Deployment for multi-cloud
-(Buyya and Barreto, 2015) Provider-Agnostic Operations Aneka-Based Security and Performance for Containerized multi-cloud Applications
-(Hossain et al., 2023) Microservices privacy and access control Encryption, authentication, service mesh for security
-(Fugkeaw, 2023) Decentralized IAM security issue Blockchain for IAM security and data integrity
-(Roy and Kalafatis, 2023) Secure communication across edge-cloud environments WireGuard for secure edge-cloud communication
-(Benmerar et al., 2023) Multi-domain edge communication security Secure communication and AI anomaly detection
-(Huang and Wang, 2023) Secure robot-cloud communication Secure communication and encryption for data integrity
-Category 3: Container Security
-(Deng et al., 2022) Co-resident Security Secure Placement Strategy via Deep RL
-(Gao et al., 2017) Container Isolation Power and Leakage Management for multi-cloud Security
-(Bélair et al., 2019) Container and Host Security Host-Container Cooperation for multi-cloud Defense
-(Wang et al., 2023) Container security and malware detection DockerWatch for malicious activity detection
-(Tiwari and Sharma, 2023) Multi-tenant serverless security challenge Security best practices for serverless deployments
-Category 4: Infrastructure and Deployment Management Security
-(Ferry et al., 2014) Data Deployment Secure Private Data Deployment Strategy
-(Wurster et al., 2018) Infrastructure Orchestration Standard Interface Implementation for Serverless Security
-(Libardi et al., 2014) Storage Security Multi-cloud Storage Security with MSSF
-(Quint and Kratzke, 2018) Migration Security Infrastructure-Aware and Agnostic Secure Migration
-(Solayman and Qasha, 2023) Distributed infrastructure security challenge Orchestration security for IoT data protection
-(Kataru et al., 2023) Securing AWS app deployments AWS best practices for CI/CD security
-Category 5: Security Monitoring and Breach Prevention
-(Sheridan et al., 2017) System Vulnerabilities Automated Security Evaluation and Bootstrapping
-(Trihinas et al., 2015) Breach Detection and Prevention Metric Filtering and Fault Recovery for multi-cloud Security
-(Lee and Nam, 2023) Network policy misconfiguration risks Automated network policy discovery and verification
-(Nguyen and Chien, 2023) Stream processing security risks RBAM framework for secure data handling
-Category 6: Trust and Compliance Management
-(Casola et al., 2017) Application Security and Compliance Risk Analysis and Security SLAs with MUSA Framework
-(Harwalkar et al., 2019b) Data Security and Standardization OpenStack API-based multi-cloud Security and Interoperability
-(Tricomi et al., 2017) Network Security Inter-cloud Communication Protection with SFC
-(Wahab et al., 2016) Trust and Community Formation Trust-based Hedonic Coalitions for multi-cloud Security
-(Kritikos et al., 2019) Orchestration and Adaptation Multi-cloud Service Orchestration and Cross-Level Adaptation
-(Petrovski and Gusev, 2024) Multicloud compliance and security issue AWS security measures for multicloud environments
-(Alonso et al., 2023a) Cross-cloud security and compliance Security frameworks for multi-cloud environment
-Category 7: Placement Strategy
-(Sami and Mourad, 2020) Placement Strategy Security Model for Volunteering Fog Services
-(Pandi et al., 2023) Multi-cloud security concerns Security groups, VPCs, and infrastructure security
-(Ramesh et al., 2023) Multicloud environment security issue Multicloud security for risk diversification
-(Sofia et al., 2023) Edge-cloud security concerns Enhanced security protocols for decentralized environments
-(Harzenetter et al., 2023) Cross-technology update security risks Integrated security in workflows for management
-(Nagarajan et al., 2023) Resource allocation security risks Security measures in resource allocation framework
 Access and Communication Control category reports on multiple challenges related to maintaining secure multi-cloud connectivity and managing user access control.
-Out of the 121 reviewed studies, 9 specifically identified issues and proposed solutions addressing secure communication, provider-agnostic operations, and decentralized identity management (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)).
+Out of the 121 reviewed studies, 9 specifically identified issues and proposed solutions addressing secure communication, provider-agnostic operations, and decentralized identity management (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in [11]).
 Advanced encryption and authentication mechanisms, e.g., TLS, JWT, and OAuth 2.0, were identified as solutions to secure communications across cloud and edge environments.
@@
 Container Security category reports on challenges such as co-resident security, container isolation, host security, and multi-tenant serverless security in multi-cloud environment.
-Among the 121 reviewed studies, 5 identified these issues and proposed solutions to address them (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)).
+Among the 121 reviewed studies, 5 identified these issues and proposed solutions to address them (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in [11]).
 The solutions identified include the use of secure placement strategies via deep reinforcement learning to manage secure container placement in multi-cloud platforms.
@@
 Infrastructure and Deployment Management Security category reports on challenges related to securing data deployment, infrastructure orchestration, storage security, and migration security in multi-cloud environment.
-Out of the reviewed 121 studies, 6 specifically addressed these types of challenges and proposed solutions (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)).
+Out of the reviewed 121 studies, 6 specifically addressed these types of challenges and proposed solutions (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in [11]).
 The identified solutions include the use of a Secure Private Data Deployment Strategy with encryption techniques like AES-256 and secure protocols such as SFTP and HTTPS to safeguard data.
@@
 Security Monitoring and Breach Prevention category reports on challenges related to system vulnerabilities, breach detection and prevention, and network policy misconfiguration risks in multi-cloud environment.
-Out of the 121 reviewed studies, 4 specifically addressed these issues and proposed corresponding solutions (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)).
+Out of the 121 reviewed studies, 4 specifically addressed these issues and proposed corresponding solutions (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in [11]).
 The identified solutions suggest implementing automated security evaluations and bootstrapping mechanisms for Virtual Machines with security configurations using tools like Terraform, Chef, Ansible, or Puppet.
@@
 Trust and Compliance Management category reports on solutions related to ensuring security and compliance in multi-cloud environment.
-Among the 121 reviewed studies, 7 specifically discussed these issues and proposed corresponding solutions (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)).
+Among the 121 reviewed studies, 7 specifically discussed these issues and proposed corresponding solutions (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in [11]).
 The identified solutions emphasize the need for security controls and standardized practices, such as risk analysis and secure coding using MUSA.
@@
 Placement Strategy category highlights the need for secure service deployment on volunteering fog nodes in multi-cloud environment.
-Of the 121 reviewed studies, 6 proposed solutions for these challenges (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in (Waseem et al., 2024)).
+Of the 121 reviewed studies, 6 proposed solutions for these challenges (see Table LABEL:tab:Security-challenges-solutions and the Security Challenges-Solution sheet in [11]).
 The solutions involve developing a security model that integrates SSL/TLS protocols for communication and certificate-based access controls to protect deployed services on fog nodes.
@@
 The reviewed studies recommend adaptive methods for handling security in dynamic and decentralized edge-cloud environments by proposing cross-technology security updates and resource allocation frameworks to avoid conflicts and vulnerabilities.
-[…]
-3.6. Automation Challenges and Solution Framework (RQ4 and RQ5)
-Table LABEL:tab:Automation-solutions presents an overview of the automation challenges and solution framework for containerized applications in multi-cloud environment.
-The figure consists of eight categories, each focusing on different aspects of automation challenges and their solutions in a multi-cloud environment.
+Takeaway 8: Security in containerized multi-cloud environments poses unique challenges due to the combined complexity of container orchestration and distributed cloud platforms.
+The reviewed studies show that addressing these issues requires specific strategies—such as secure container placement, decentralized identity management, and container-aware compliance frameworks—not typically needed in single-cloud or non-containerized setups.
+These solutions highlight the distinct security needs that arise only when containerization is applied across multiple clouds.
+3.6 Automation Challenges and Solution Framework (RQ4 and RQ5)
+Table LABEL:tab:Automation-solutions presents an overview of automation challenges and corresponding solutions for containerized applications in multi-cloud environments.
+The framework is organized into eight categories, each addressing a distinct aspect of automation—such as orchestration, deployment, resource management, and standardization.
+These categories capture recurring challenges and the solution strategies proposed across the selected studies.
 Multi-Cloud Automation category reports on challenges and solutions related to automating processes for combining and administering various cloud computing environments.
-Out of the 121 reviewed studies, 13 identified these challenges and provided associated solutions (see Table LABEL:tab:Automation-solutions and Automation Challenges-Solutions sheet in (Waseem et al., 2024)).
+Out of the 121 reviewed studies, 13 identified these challenges and provided associated solutions (see Table LABEL:tab:Automation-solutions and Automation Challenges-Solutions sheet in [11]).
 Key challenges involve the integration, provisioning, orchestration, and adaptation of multiple clouds.
@@
 Automation in Deployment and Scaling category reports on challenges and solutions related to automating deployment and scaling processes for containerized applications across multi-cloud platforms.
-Out of the 121 reviewed studies, 12 identified these challenges and proposed corresponding solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)).
+Out of the 121 reviewed studies, 12 identified these challenges and proposed corresponding solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in [11]).
 Key challenges include intelligent container placement, application runtime switch automation, and service deployment issues.
@@
 This emphasis on automating and optimizing deployment strategies led to the proposal of comprehensive multi-cloud orchestration tools that facilitate the deployment of containers and VMs, aimed at improving agility and scalability within diverse cloud environments.
-Table 12. Identified automation challenge and solution for containerized applications in multi-cloud environment
+Resource Management Automation category reports on challenges and solutions related to automating the management of resources in multi-cloud environment.
+Out of the 121 studies reviewed, 9 specifically addressed these challenges and provided appropriate solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in [11]).
+These challenges include resource overbooking management, resource management for containerized applications, and resource commissioning and decommissioning.
+Table 12: Identified automation challenges and solutions for containerized applications in multi-cloud environment
 ID Challenge Solution
 Category 1: Multi-Cloud Automation
-(Sabbioni et al., 2020) Integrating multi-cloud environment Implementing Environment Proxies
-(Ferry et al., 2014) Automating Multi-Cloud Provisioning Adopting Model-Driven Engineering
-(Sami and Mourad, 2020) Creating On-Demand Fog in Multi-Cloud Enabling On-Demand Fog Computing
-(Libardi et al., 2014) Enhancing User-Friendly Automation Using Multi-Cloud Storage Selection
-(Wurster et al., 2018) Orchestrating Multi-Stack Environments Deploying with TOSCA-based Approach
-(Rios et al., 2017) Securing Multi-Cloud Application Creation Applying MUSA Framework
-(Dimitrijevic et al., 2019) Unifying Resource Abstraction Managing with Resource Managers
-(Casola et al., 2017) Integrating Multi-Cloud Deployments Automating MUSA Deployment
-(Tricomi et al., 2017) Coordinating Federated Cloud Deployments Orchestrating Subsystem Deployment
-(Sitaram et al., 2018) Setting Up Hybrid multi-cloud environment Standardizing Interoperability
-(Quint and Kratzke, 2018) Ensuring Multi-Cloud Interoperability Enabling Multi-Cloud Support
-(Challita et al., 2017) Managing Multi-Cloud Brokerage Systems Managing Cloud Services
-(Kritikos et al., 2019) Implementing Multi-Cloud Orchestration Migrating Manually to Cloud
-(Bou Ghantous and Gill, 2021) Establishing DevOps in Multi-Cloud Automating Agile Development
-(da Silva et al., 2023) Addressing Cost and Migration Challenges Managing Complexity
-(Pandi et al., 2023) Overcoming Multi-Cloud Automation Barriers Integrating Jenkins and Kubernetes
-(Ramesh et al., 2023) Improving Multi-Cloud Cost Efficiency Optimizing Deployment Costs
-(Petrovski and Gusev, 2024) Simplifying Multi-Cloud Deployment Complexity Adopting AWS Deployment Strategies
-(Kataru et al., 2023) Maintaining Multi-Cloud Applications Automating Resource Allocation
+[65] Integrating multi-cloud environment Implementing Environment Proxies
+[89] Automating Multi-Cloud Provisioning Adopting Model-Driven Engineering
+[29] Creating On-Demand Fog in Multi-Cloud Enabling On-Demand Fog Computing
+[79] Enhancing User-Friendly Automation Using Multi-Cloud Storage Selection
+[140] Orchestrating Multi-Stack Environments Deploying with TOSCA-based Approach
+[56] Securing Multi-Cloud Application Creation Applying MUSA Framework
+[61] Unifying Resource Abstraction Managing with Resource Managers
+[127] Integrating Multi-Cloud Deployments Automating MUSA Deployment
+[128] Coordinating Federated Cloud Deployments Orchestrating Subsystem Deployment
+[81] Setting Up Hybrid multi-cloud environment Standardizing Interoperability
+[50] Ensuring Multi-Cloud Interoperability Enabling Multi-Cloud Support
+[94] Managing Multi-Cloud Brokerage Systems Managing Cloud Services
+[76] Implementing Multi-Cloud Orchestration Migrating Manually to Cloud
+[90] Establishing DevOps in Multi-Cloud Automating Agile Development
+[115] Addressing Cost and Migration Challenges Managing Complexity
+[117] Overcoming Multi-Cloud Automation Barriers Integrating Jenkins and Kubernetes
+[92] Improving Multi-Cloud Cost Efficiency Optimizing Deployment Costs
+[100] Simplifying Multi-Cloud Deployment Complexity Adopting AWS Deployment Strategies
+[134] Maintaining Multi-Cloud Applications Automating Resource Allocation
 Category 2: Automation in Deployment and Scaling
-(Alonso et al., 2019) Automating Multi-Cloud Deployment Monitoring Adopting Multi-Cloud Strategy
-(Deng et al., 2022) Optimizing Container Placement Implementing Intelligent Placement
-(Zhang et al., 2021) Automating Runtime Application Switching Deploying Service Driver Model
-(Pham and Pham, 2015) Enabling Fine-Grained Component Automation Replicating and Migrating Services
-(Barna et al., 2017) Streamlining DevOps Automation Developing with DevOps Approach
-(Gupta et al., 2021) Automating Service Deployment Using Automated Scripts
-(Sheridan et al., 2017) Securing Deployment Time Securing with Automation Tools
-(Mittal and Risco-Martín, 2017) Automating Distributed Simulations Adopting DevOps Methodologies
-(Al-shammari and Alwan, 2018) Improving Software Deployment Speed Optimizing Container Images
-(Lorenzo et al., 2019) Automating Container Startup Profiling Container Execution
-(Pandey et al., 2022) Dynamically Scaling Resources Automating Scaling with CSP Capability
-(Pahl, 2015) Scaling Container Deployment Scaling with Kubernetes
-(Saif et al., 2023) Provisioning Elastic Resources Provisioning Resources Elastically
-(Bhamare et al., 2015) Automating VF Placement Optimizing VF Placement
-(Guerrero et al., 2018a) Managing Container and VM Deployment Orchestrating Multi-Cloud Deployments
+[105] Automating Multi-Cloud Deployment Monitoring Adopting Multi-Cloud Strategy
+[69] Optimizing Container Placement Implementing Intelligent Placement
+[101] Automating Runtime Application Switching Deploying Service Driver Model
+[80] Enabling Fine-Grained Component Automation Replicating and Migrating Services
+[52] Streamlining DevOps Automation Developing with DevOps Approach
+[103] Automating Service Deployment Using Automated Scripts
+[55] Securing Deployment Time Securing with Automation Tools
+[106] Automating Distributed Simulations Adopting DevOps Methodologies
+[112] Improving Software Deployment Speed Optimizing Container Images
+[113] Automating Container Startup Profiling Container Execution
+[67] Dynamically Scaling Resources Automating Scaling with CSP Capability
+[60] Scaling Container Deployment Scaling with Kubernetes
+[77] Provisioning Elastic Resources Provisioning Resources Elastically
+[47] Automating VF Placement Optimizing VF Placement
+[99] Managing Container and VM Deployment Orchestrating Multi-Cloud Deployments
 Category 3: Resource Management Automation
-(Ramos et al., 2021) Handling Resource Overbooking Detecting Overbooking with ML
-(Cui et al., 2016) Managing Virtual Clusters Managing with ClaaS Model
-(Altran et al., 2022) Scheduling Containerized Applications Scheduling with Kubernetes
-(Shi et al., 2020) Selecting Optimal VM Types Using GA-based Algorithms
-(Baresi et al., 2016) Automating Resource Commissioning Adopting MicroCloud Architecture
-(Dreibholz et al., 2019) Supporting Cloud Federation Automation Optimizing with Middleware Platform
-(Trihinas et al., 2015) Automating Elastic Resource Detection Automating with JCatascopia
-(Buyya and Barreto, 2015) Automating Resource Acquisition Managing Resources with Aneka
-(Mfula and Nurminen, 2018) Enforcing SLA Compliance Testing for Compliance
+[132] Handling Resource Overbooking Detecting Overbooking with ML
+[126] Managing Virtual Clusters Managing with ClaaS Model
+[114] Scheduling Containerized Applications Scheduling with Kubernetes
+[97] Selecting Optimal VM Types Using GA-based Algorithms
+[31] Automating Resource Commissioning Adopting MicroCloud Architecture
+[122] Supporting Cloud Federation Automation Optimizing with Middleware Platform
+[84] Automating Elastic Resource Detection Automating with JCatascopia
+[98] Automating Resource Acquisition Managing Resources with Aneka
+[42] Enforcing SLA Compliance Testing for Compliance
 Category 4: Data and Application Migration Automation
-(Almeida et al., 2014) Safeguarding Data Integrity in Migration Maintaining Integrity Autonomously
-(Slominski et al., 2015) Migrating Legacy Web Applications Reusing Architectural Patterns
-(Kiss, 2018) Porting Applications to Cloud Using CloudSME Platform
+[95] Safeguarding Data Integrity in Migration Maintaining Integrity Autonomously
+[44] Migrating Legacy Web Applications Reusing Architectural Patterns
+[63] Porting Applications to Cloud Using CloudSME Platform
 Category 5: Testing and Benchmarking Automation
-(Verdugo et al., 2017) Automating User-Oriented Testing Virtualizing with Lightweight Containers
-(Jha et al., 2019) Streamlining Benchmarking Orchestrating Docker Benchmarking
-(Naik, 2016a) Enhancing Automation Dependability Enhancing Dependability with CI
+[120] Automating User-Oriented Testing Virtualizing with Lightweight Containers
+[96] Streamlining Benchmarking Orchestrating Docker Benchmarking
+[148] Enhancing Automation Dependability Enhancing Dependability with CI
 Category 6: Standardization and Interoperability Challenges
-(Gattobigio et al., 2022) Standardizing Automation Processes Integrating Service Mesh
-(Lazuka et al., 2022) Optimizing Cloud Provider Selection Selecting Optimal Providers
-(Ouyang et al., 2022) Unifying Microservices Deployment Adopting Service Chain Models
-(Harwalkar et al., 2019a) Addressing Cloud Interoperability Managing with Multi-Cloud Harmony
-(Pellegrini et al., 2017) Standardizing Cloud Applications Standardizing Interfaces
-(Pai and Kumar, 2021) Automating Package Creation and Scaling Automating Package Frameworks
+[34] Standardizing Automation Processes Integrating Service Mesh
+[74] Optimizing Cloud Provider Selection Selecting Optimal Providers
+[121] Unifying Microservices Deployment Adopting Service Chain Models
+[108] Addressing Cloud Interoperability Managing with Multi-Cloud Harmony
+[73] Standardizing Cloud Applications Standardizing Interfaces
+[59] Automating Package Creation and Scaling Automating Package Frameworks
 Category 7: Application and Service Management
-(Bolivar et al., 2018) Selecting 5G Ecosystem Components Utilizing Open-Source Components
-(Rahman and Lama, 2019) Managing End-to-End Tail Latency Modeling Performance Efficiently
-(Serhiienko and Spillner, 2018) Handling Distributed Application Complexity Managing Distributed Applications
+[40] Selecting 5G Ecosystem Components Utilizing Open-Source Components
+[53] Managing End-to-End Tail Latency Modeling Performance Efficiently
+[45] Handling Distributed Application Complexity Managing Distributed Applications
 Category 8: Runtime and Service Discovery
-(Alaluna et al., 2017) Addressing Network Complexity Managing Network Complexity
-(Wahab et al., 2016) Improving Social Network Service Discovery Automating Service Discovery
-(Pham et al., 2015) Administering Complex Applications Building Autonomic Systems
-Resource Management Automation category reports on challenges and solutions related to automating the management of resources in multi-cloud environment.
-Out of the 121 studies reviewed, 9 specifically addressed these challenges and provided appropriate solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)).
-These challenges include resource overbooking management, resource management for containerized applications, and resource commissioning and decommissioning.
+[124] Addressing Network Complexity Managing Network Complexity
+[72] Improving Social Network Service Discovery Automating Service Discovery
+[86] Administering Complex Applications Building Autonomic Systems
 The proposed solutions include machine learning-based resource overbooking detection for efficiently managing service containers across multiple cloud platforms.
@@
 Data and Application Migration Automation: category reports challenges and solutions related to the automation of testing and benchmarking processes for containerized applications in multi-cloud environment.
-Among the 121 studies reviewed, 3 identified these challenges along with their proposed solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)).
+Among the 121 studies reviewed, 3 identified these challenges along with their proposed solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in [11]).
 The main challenges include ensuring data integrity during the migration process and enabling the smooth migration of legacy Web applications to cloud services.
@@
 Testing and Benchmarking Automation category highlights the challenges and solutions related to automating the testing and benchmarking processes for containerized applications in multi-cloud environment.
-Out of the 121 studies reviewed, 3 specifically addressed these challenges and provided relevant solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)).
+Out of the 121 studies reviewed, 3 studies specifically addressed these challenges and provided relevant solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in [11]).
 The main challenges revolve around automating user-oriented testing and benchmarking across various cloud platforms.
@@
 Standardization and Interoperability Challenges category reports on the challenges and solutions related to standardization and interoperability in the automation of containerized applications across multiple cloud environments.
-Out of the 121 reviewed studies, 6 specifically identified these challenges and proposed corresponding solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)).
+Out of the 121 reviewed studies, 6 specifically identified these challenges and proposed corresponding solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in [11]).
 The main challenges involve standardizing automation processes, selecting cloud providers, and addressing the lack of unified templates for microservices deployment.
@@
 Finally, automation frameworks for package creation and scaling, such as those using machine-readable definition files, were proposed to efficiently manage both virtual machines and Docker containers across multi-cloud environment.
-Application and Service Management category reports challenges and solutions related to the automation of application and service management processes in containerized multi-cloud environment.Out of the 121 reviewed studies, 2 specifically addressed these challenges and proposed appropriate solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)).
+Application and Service Management category reports challenges and solutions related to the automation of application and service management processes in containerized multi-cloud environment.Out of the 121 reviewed studies, 2 specifically
+addressed these challenges and proposed appropriate solutions (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in [11]).
 The key challenges include the selection of 5G ecosystem components and end-to-end tail latency management in microservice architectures.
@@
 Runtime and Service Discovery category organizes the challenges and solutions related to runtime management and services discovery in multi-cloud environment, particularly in managing the complexity of network control and administration for containerized applications.
-Out of the 121 reviewed studies, 4 specifically tackled these challenges (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in (Waseem et al., 2024)).
+Out of the 121 reviewed studies, 4 specifically tackled these challenges (see Table LABEL:tab:Automation-solutions and the Automation Challenges-Solutions sheet in [11]).
 The proposed solutions include automating network management to address the increasing complexity of distributed applications across multi-cloud topologies.
@@
 Autonomic computing systems were introduced to handle the growing complexity of applications by facilitating knowledge transfer, enabling more effective containerization management through automation.
-[…]
-3.7. Deployment Challenges and Solution Framework (RQ4 and RQ5)
+Takeaway 9: Automation challenges in containerized multi-cloud environments are not merely generic cloud or container issues; they arise from the complex interplay between container portability and cloud heterogeneity.
+Studies reveal that managing orchestration, scaling, and resource allocation across diverse cloud platforms requires specialized solutions, such as TOSCA-based orchestration, intelligent placement algorithms, and DevOps-integrated pipelines, highlighting how containerization uniquely amplifies automation demands in multi-cloud settings.
+3.7 Deployment Challenges and Solution Framework (RQ4 and RQ5)
 Table LABEL:tab:Deployment-solutions presents a comprehensive overview of the challenges and corresponding solutions faced during the deployment of containerized applications in multi-cloud environment.
@@
 The solution is Cross-Container Isolation, which enforces cross-container isolation to enhance security.
-Table 13. Identified deployment challenge and solution for containerized applications in multi-cloud environment
-ID Challenge Solution
-Category 1: Deployment Complexity and Orchestration Challenges
-(Almeida et al., 2014) Performance Testing Deployment Standardized Benchmarking and Automation
-(Verdugo et al., 2017) Deployment Validation Testing Process Management System
-(Jha et al., 2019) Cost-effective Deployment Containerization and Microservice Architecture
-(Pandey et al., 2022) Resource Management Deployment User Preference-Based Multi-Cloud Resource Brokering
-(Bhamare et al., 2015) NFV Site Deployment Site Selection and VF Allocation Algorithms
-(Serhiienko and Spillner, 2018) Cloud Management Platform Evaluation Standardized Output Formats and Evaluation Criteria
-(Dimitrijevic et al., 2019) Standardized Requirement Deployment Application-Level Resource Managers for Multi-Cloud
-(Baresi et al., 2016) Heterogeneous Node Deployment Technology-Independent Multi-Level Adaptation
-(Casola et al., 2017) Deployment Plan Generation Cloud Services Selection and Deployment Planning
-(Tricomi et al., 2017) TOSCA-based Deployment Brokered Multi-Cloud Deployment with TOSCA
-(Sitaram et al., 2018) Standardization Deployment Proxy Cloud Virtualization for OpenStack
-(Barletta et al., 2024) Criticality Constraints in Kubernetes Deployment Kubernetes for Critical Deployments
-(da Silva et al., 2023) Automation Tools Deployment Issue Accurate Cost and Resource Tracking
-(Pandi et al., 2023) Multi-cloud Deployment Challenge Jenkins for Multi-cloud Automation
-(Sofia et al., 2024) Decentralized App Deployment Issue CODECO for Edge-Cloud Deployment
-(Neeraj et al., 2023) SLA Monitoring Deployment Complexity Blockchain for SLA Data Integrity
-(Petrovski and Gusev, 2024) Container and Serverless Architecture Deployment Issue AWS ECS and Lambda Deployment
-(Solayman and Qasha, 2023) Heterogeneous Environment Deployment Challenge Edge-cloud Orchestration Management
-(Sofia et al., 2023) Dynamic Adaptation Deployment Issue Context-aware Resource Allocation Strategies
-(Santos et al., 2023) Regional Infrastructure Deployment Consistency Issue AWS Wavelength for Edge Cloud Integration
-(Roy and Kalafatis, 2023) Robotic App Deployment Challenge Hybrid Edge-Cloud Architecture
-(Alonso et al., 2023a) Multi-cloud Deployment Consistency Problem Multi-cloud Deployment Efficiency Models
-(Harzenetter et al., 2023) Cross-Technology Deployment Integrity Issue Workflow Automation for Cross-Technology Deployment
-(Huang and Wang, 2023) Industrial Robot Control System Deployment Problem Seamless Control between Cloud and Robots
-(Nagarajan et al., 2023) Efficient Task Deployment across Clouds Hybrid Approach for Resource Allocation
-Category 2: Security and Compliance Management in Multi-Cloud Deployment
-(Ferry et al., 2014) Vendor Lock-in Prevention Deployment Cloud Modelling Framework
-(Gao et al., 2017) Security Deployment in Multi-container Environment Cross-Container Isolation
-(Rios et al., 2017) Security Deployment Security Control Specification and Deployment Framework
-(Alaluna et al., 2017) Virtual Network Security Deployment Network Virtualization Platform for Multi-Cloud
-(Wahab et al., 2016) Malicious Service Management Deployment Collusion-Resilient Trust Aggregation Technique
-(Barati et al., 2023) Secure Personal Data Deployment Challenge Docker, Kubernetes, Blockchain for GDPR Compliance
-(Fugkeaw, 2023) IAM Deployment across Cloud Platforms Blockchain for Decentralized IAM Deployment
-(Wang et al., 2023) Non-intrusive Malware Detection Deployment DockerWatch for Non-intrusive Container Integration
-Category 3: Multi-Cloud Deployment Coordination and Integration
-(Sabbioni et al., 2020) Multi-cloud PaaS Deployment Lightweight Proxies for PaaS Integration
-(Bou Ghantous and Gill, 2021) IoT Application Deployment DRA Framework and CI Broker for Multi-Cloud Deployment
-(Ramesh et al., 2023) Data Transfer and Service Compatibility FaaS and Storage for Multi-Cloud Deployment
-(Nguyen and Chien, 2023) Heterogeneous Cloud and Edge Integration RBAM for Cloud-Edge Integration
-(Tiwari and Sharma, 2023) Private Cloud Integration Problem CLI Tool for Serverless Setup
-Category 4: Monitoring, Performance, and Scalability Challenges
-(Slominski et al., 2015) Legacy Code Migration Deployment Reusable Architectural Pattern with Docker and Cloudant
-(Wang et al., 2020) Tenant Performance Deployment Multi-Tenant and Multi-Instance Hybrid Deployment
-(Trihinas et al., 2015) Monitoring Deployment Platform-Independent Monitoring System
-(Struhár et al., 2024) Resource Contention and Performance Problem Linux Patch for Dynamic CPU Adaptation
-(Horchulhack et al., 2024) QoS Degradation Detection Problem Kubernetes with Performance Monitoring Tools
-(Marques et al., 2024) Monitoring Integration in Hybrid Cloud Issue Custom Scaling with Kubernetes
-(Kataru et al., 2023) AWS Deployment Performance Consistency AWS Model Comparison for Performance
-(Benmerar et al., 2023) Cloud-Edge Deployment Performance Issue AI-driven Orchestration for Dynamic Resource Management
-Category 5: Microservices Architecture and Containerization Challenges
-(Barna et al., 2017) Autonomic System Deployment Self-Tuning Performance Model and Autonomic Management
-(Fadda et al., 2019) Microservice Deployment Metrics and Requirements Framework for Microservices
-(Pop et al., 2016) Multi-cloud Application Configuration Deployment Runtime Environment with Object Store and Artifact Repository
-(Quint and Kratzke, 2018) Elastic Container Platform Dependency Deployment Separation of Elastic Platform and Cloud Application
-(Hossain et al., 2023) Microservices Deployment in Edge Environments Docker, CI/CD, Service Mesh for Microservices
-Category 6: Network Connectivity and Hybrid Cloud Integration
-(Carpio et al., 2020) Inter-edge Bandwidth Deployment Modular Edge Cloud Computing Architecture
-(Shi et al., 2020) Network Latency Deployment VM Selection for Composite Applications
-(Dreibholz et al., 2019) Latency Reduction Deployment Edge-Cellular Hybrid Infrastructure Provisioning
-(Pellegrini et al., 2017) Vendor Lock-in Prevention Deployment Interoperability Layer Above Cloud Infrastructure
-(Lee and Nam, 2023) Dynamic Network Policy Deployment Problem KUNERVA for Network Policy Automation
-Category 7: Cloud Deployment Constraints and Challenges
-(Pham et al., 2015) ACS Compatibility Deployment Scalable Multi-Cloud Deployment Framework
-(Sheridan et al., 2017) Manual Intervention Risk Deployment Automated Security Measures at Deployment
-(Wurster et al., 2018) Multi-provider Deployment TOSCA-Based Deployment Modeling Approach
-(Challita et al., 2017) Cloud Provider Transition Deployment Semantic Interoperability in Multi-Clouds
-(da Silva et al., 2023) Automation Tools Deployment Issue Accurate Cost and Resource Tracking
-(Santos et al., 2023) Regional Infrastructure Deployment Consistency Issue AWS Wavelength for Edge Cloud Integration
-Category 8: Infrastructure Provisioning and Container Deployment Challenges
-(Al-shammari and Alwan, 2018) Fog Computing Deployment Reorganized Container Images and Docker Deployment
-(Sami and Mourad, 2020) Fog Device Deployment On-Demand Fog and Microservices Deployment
-(Lorenzo et al., 2019) Fog Computing Container Deployment FogDocker for Container Deployment
-(Altran et al., 2022) Distributed Compute Node Deployment Label-Based Scheduling Strategy
-(Bélair et al., 2019) Kernel Security Deployment Enhanced Container Security
-(Barletta et al., 2024) Criticality Constraints in Kubernetes Deployment Kubernetes for Critical Deployments
-(Centofanti et al., 2024) Power Tool Compatibility Problem Docker, Kubernetes for Power Monitoring Integration
-(Ramesh et al., 2023) Data Transfer and Service Compatibility FaaS and Storage for Multi-Cloud Deployment
-Category 9: Application Deployment Workflow and Orchestration Challenges
-(Alonso et al., 2019) API Management Deployment DevOps Approach for Multi-Cloud Applications
-(Gupta et al., 2021) Web Service Deployment Kubernetes-Based Containerized Deployment
-(Buyya and Barreto, 2015) Complex Application Deployment Aneka Platform for Distributed Applications
-(Acquaviva et al., 2017) High-availability Deployment Middleware Support for High Availability in Multi-Cloud PaaS
-(Kiss, 2018) Complex Workflow Deployment CloudSME Simulation Platform
-(Kritikos et al., 2019) Business Process Provisioning Deployment Multi-Cloud Service Orchestration Frameworks
-(Roy and Kalafatis, 2023) Robotic App Deployment Challenge Hybrid Edge-Cloud Architecture
-(Sofia et al., 2023) Dynamic Adaptation Deployment Issue Context-aware Resource Allocation Strategies
-Multi-cloud Deployment Coordination and Integration This category identifies and classifies the challenges and solutions related to coordination and integration during the multi-cloud deployment phase.
+Multi-cloud Deployment Coordination and Integration: This category identifies and classifies the challenges and solutions related to coordination and integration during the multi-cloud deployment phase.
 Of the 121 reviewed studies, 5 addressed these challenges.
@@
 To address Vendor Lock-in, an Interoperability Layer Above Cloud Infrastructure enables smoother transitions between cloud providers.
+Table 13: Identified deployment challenges and solutions for containerized applications in multi-cloud environment
+      ID        Challenge        Solution
+      Category 1: Deployment Complexity and Orchestration Challenges
+      [95]        Performance Testing Deployment        Standardized Benchmarking and Automation
+      [120]        Deployment Validation        Testing Process Management System
+      [96]        Cost-effective Deployment        Containerization and Microservice Architecture
+      [67]        Resource Management Deployment        User Preference-Based Multi-Cloud Resource Brokering
+      [47]        NFV Site Deployment        Site Selection and VF Allocation Algorithms
+      [45]        Cloud Management Platform Evaluation        Standardized Output Formats and Evaluation Criteria
+      [61]        Standardized Requirement Deployment        Application-Level Resource Managers for Multi-Cloud
+      [31]        Heterogeneous Node Deployment        Technology-Independent Multi-Level Adaptation
+      [127]        Deployment Plan Generation        Cloud Services Selection and Deployment Planning
+      [128]        TOSCA-based Deployment        Brokered Multi-Cloud Deployment with TOSCA
+      [81]        Standardization Deployment        Proxy Cloud Virtualization for OpenStack
+      [82]        Criticality Constraints in Kubernetes Deployment        Kubernetes for Critical Deployments
+      [115]        Automation Tools Deployment Issue        Accurate Cost and Resource Tracking
+      [117]        Multi-cloud Deployment Challenge        Jenkins for Multi-cloud Automation
+      [138]        Decentralized App Deployment Issue        CODECO for Edge-Cloud Deployment
+      [147]        SLA Monitoring Deployment Complexity        Blockchain for SLA Data Integrity
+      [100]        Container and Serverless Architecture Deployment Issue        AWS ECS and Lambda Deployment
+      [38]        Heterogeneous Environment Deployment Challenge        Edge-cloud Orchestration Management
+      [118]        Dynamic Adaptation Deployment Issue        Context-aware Resource Allocation Strategies
+      [125]        Regional Infrastructure Deployment Consistency Issue        AWS Wavelength for Edge Cloud Integration
+      [46]        Robotic App Deployment Challenge        Hybrid Edge-Cloud Architecture
+      [33]        Multi-cloud Deployment Consistency Problem        Multi-cloud Deployment Efficiency Models
+      [141]        Cross-Technology Deployment Integrity Issue        Workflow Automation for Cross-Technology Deployment
+      [49]        Industrial Robot Control System Deployment Problem        Seamless Control between Cloud and Robots
+      Category 2: Security and Compliance Management in Multi-Cloud Deployment
+      [89]        Vendor Lock-in Prevention Deployment        Cloud Modelling Framework
+      [70]        Security Deployment in Multi-container Environment        Cross-Container Isolation
+      [56]        Security Deployment        Security Control Specification and Deployment Framework
+      [124]        Virtual Network Security Deployment        Network Virtualization Platform for Multi-Cloud
+      [72]        Malicious Service Management Deployment        Collusion-Resilient Trust Aggregation Technique
+      [135]        Secure Personal Data Deployment Challenge        Docker, Kubernetes, Blockchain for GDPR Compliance
+      [144]        IAM Deployment across Cloud Platforms        Blockchain for Decentralized IAM Deployment
+      [142]        Non-intrusive Malware Detection Deployment        DockerWatch for Non-intrusive Container Integration
+      Category 3: Multi-Cloud Deployment Coordination and Integration
+      [65]        Multi-cloud PaaS Deployment        Lightweight Proxies for PaaS Integration
+      [90]        IoT Application Deployment        DRA Framework and CI Broker for Multi-Cloud Deployment
+      [92]        Data Transfer and Service Compatibility        FaaS and Storage for Multi-Cloud Deployment
+      [119]        Heterogeneous Cloud and Edge Integration        RBAM for Cloud-Edge Integration
+      [131]        Private Cloud Integration Problem        CLI Tool for Serverless Setup
+      Category 4: Monitoring, Performance, and Scalability Challenges
+      [44]        Legacy Code Migration Deployment        Reusable Architectural Pattern with Docker and Cloudant
+      [66]        Tenant Performance Deployment        Multi-Tenant and Multi-Instance Hybrid Deployment
+      [84]        Monitoring Deployment        Platform-Independent Monitoring System
+      [137]        Resource Contention and Performance Problem        Linux Patch for Dynamic CPU Adaptation
+      [87]        QoS Degradation Detection Problem        Kubernetes with Performance Monitoring Tools
+      [88]        Monitoring Integration in Hybrid Cloud Issue        Custom Scaling with Kubernetes
+      [134]        AWS Deployment Performance Consistency        AWS Model Comparison for Performance
+      [104]        Cloud-Edge Deployment Performance Issue        AI-driven Orchestration for Dynamic Resource Management
+      Category 5: Microservices Architecture and Containerization Challenges
+      [52]        Autonomic System Deployment        Self-Tuning Performance Model and Autonomic Management
+      [85]        Microservice Deployment        Metrics and Requirements Framework for Microservices
+      [32]        Multi-cloud Application Configuration Deployment        Runtime Environment with Object Store and Artifact Repository
+      [50]        Elastic Container Platform Dependency Deployment        Separation of Elastic Platform and Cloud Application
+      [37]        Microservices Deployment in Edge Environments        Docker, CI/CD, Service Mesh for Microservices
+      Category 6: Network Connectivity and Hybrid Cloud Integration
+      [30]        Inter-edge Bandwidth Deployment        Modular Edge Cloud Computing Architecture
+      [97]        Network Latency Deployment        VM Selection for Composite Applications
+      [122]        Latency Reduction Deployment        Edge-Cellular Hybrid Infrastructure Provisioning
+      [73]        Vendor Lock-in Prevention Deployment        Interoperability Layer Above Cloud Infrastructure
+      [145]        Dynamic Network Policy Deployment Problem        KUNERVA for Network Policy Automation
+      Category 7: Cloud Deployment Constraints and Challenges
+      [86]        ACS Compatibility Deployment        Scalable Multi-Cloud Deployment Framework
+      [55]        Manual Intervention Risk Deployment        Automated Security Measures at Deployment
+      [140]        Multi-provider Deployment        TOSCA-Based Deployment Modeling Approach
+      [94]        Cloud Provider Transition Deployment        Semantic Interoperability in Multi-Clouds
+      [115]        Automation Tools Deployment Issue        Accurate Cost and Resource Tracking
+      [125]        Regional Infrastructure Deployment Consistency Issue        AWS Wavelength for Edge Cloud Integration
+      Category 8: Infrastructure Provisioning and Container Deployment Challenges
+      [112]        Fog Computing Deployment        Reorganized Container Images and Docker Deployment
+      [29]        Fog Device Deployment        On-Demand Fog and Microservices Deployment
+      [113]        Fog Computing Container Deployment        FogDocker for Container Deployment
+      [114]        Distributed Compute Node Deployment        Label-Based Scheduling Strategy
+      [71]        Kernel Security Deployment        Enhanced Container Security
+      [82]        Criticality Constraints in Kubernetes Deployment        Kubernetes for Critical Deployments
+      [139]        Power Tool Compatibility Problem        Docker, Kubernetes for Power Monitoring Integration
+      [92]        Data Transfer and Service Compatibility        FaaS and Storage for Multi-Cloud Deployment
+      Category 9: Application Deployment Workflow and Orchestration Challenges
+      [105]        API Management Deployment        DevOps Approach for Multi-Cloud Applications
+      [103]        Web Service Deployment        Kubernetes-Based Containerized Deployment
+      [98]        Complex Application Deployment        Aneka Platform for Distributed Applications
+      [146]        High-availability Deployment        Middleware Support for High Availability in Multi-Cloud PaaS
+      [63]        Complex Workflow Deployment        CloudSME Simulation Platform
+      [76]        Business Process Provisioning Deployment        Multi-Cloud Service Orchestration Frameworks
+      [46]        Robotic App Deployment Challenge        Hybrid Edge-Cloud Architecture
+      [118]        Dynamic Adaptation Deployment Issue        Context-aware Resource Allocation Strategies
 Finally, Dynamic Network Policy Deployment tackles the challenge of adapting network policies to container workloads, with the solution being Automation of Network Policies using tools like KUNERVA.
@@
 The solutions include the Aneka Platform for Distributed Applications and Multi-cloud PaaS Middleware Support for High Availability, which ensure the availability and distribution of applications across multi-cloud environment.
-[…]
-3.8. Monitoring Challenges and Solution Framework (RQ4 and RQ5)
-Table LABEL:tab:monitoring-solutions provides an extensive overview of the monitoring challenges and solution framework for containerized applications in multi-cloud environment.
-It encompasses diverse challenges related to monitoring containerized applications in the complex multi-cloud landscape.
-Each category in the figure targets specific monitoring aspects and offers solutions to address these challenges.
+Takeaway 10: Reviewed studies show that deployment challenges in containerized multi-cloud environments emerge from the tight coupling of container orchestration with heterogeneous cloud infrastructures.
+Solutions such as TOSCA-based modeling, cross-container isolation, and context-aware adaptation specifically address complexities introduced when containers span multiple clouds, highlighting the unique integration, security, and coordination demands identified through this SMS.
+3.8 Monitoring Challenges and Solution Framework (RQ4 and RQ5)
+Table LABEL:tab:monitoring-solutions presents a categorized overview of monitoring challenges and corresponding solutions for containerized applications in multi-cloud environments.
+The framework captures a range of challenges related to performance variability, resource utilization, observability, and integration.
+Each of the six categories highlights a specific aspect of monitoring and summarizes relevant solutions proposed in the literature.
 Performance, Consistency, and Variability Monitoring category identifies the challenges and solutions that ensure consistent performance monitoring in dynamically changing multi-cloud environment.
@@
 Finally, Prometheus is used for Real-Time Dynamic Resource Monitoring, allowing real-time data collection and analysis to optimize performance based on resource usage fluctuations.
-Table 14. Identified monitoring challenge and solution for containerized applications in multi-cloud environment
-ID Challenge Solution
-Category 1: Performance Monitoring Consistency and Variability
-(Almeida et al., 2014) Performance Monitoring Performance Variation Monitoring
-(Jha et al., 2019) Benchmarking Performance Variation and Cost Efficiency Monitoring SDBO Optimization
-(Gupta et al., 2021) Website Performance and Availability Monitoring Kubernetes Monitoring
-(Pandey et al., 2022) User Experience Variability Monitoring CSP QoS Metrics Collection
-(Dreibholz et al., 2019) Independent Performance Metrics Monitoring Performance Evaluations
-(Bhamare et al., 2015) Comprehensive Network Performance Measurement Monitoring Server Selection Scheme
-(Rahman and Lama, 2019) Microservice Performance Degradation Monitoring Tail Latency Prediction
-(Struhár et al., 2024) Container Performance Metrics Tracking CLM and OSLM for Resource Adjustments
-(Horchulhack et al., 2024) Container Performance and Resource Usage Monitoring Metric and Container Monitoring Tools
-(Hossain et al., 2023) Microservice Performance Monitoring Challenge Prometheus, Grafana, Kubernetes for Monitoring
-(Ramesh et al., 2023) Cross-Cloud Service Performance Monitoring FaaS and Storage Services for AI Monitoring
-(Nguyen and Chien, 2023) Stream Processing Performance Monitoring RBAM for Real-Time Performance Insights
-(Kataru et al., 2023) AWS Application Performance Monitoring AWS CloudWatch for Performance Monitoring
-(Alonso et al., 2023a) Multi-Cloud Performance Visibility Issue Monitoring Tools for Multi-Cloud Visibility
-(Harzenetter et al., 2023) Cross-Technology Performance Monitoring Issue TOSCA-based Workflows for Monitoring Applications
-(Huang and Wang, 2023) Industrial Robot System Monitoring Challenge Real-Time System Performance Tracking Tools
-Category 2: Resource and Infrastructure Management Monitoring
-(Gattobigio et al., 2022) Network Communication Monitoring Service Mesh Implementation
-(Verdugo et al., 2017) System Image and Configuration Monitoring Docker Server Deployment
-(Cui et al., 2016) Vcluster and Framework Management Monitoring LXC Implementation
-(Pahl, 2015) Multi-Cloud Resource Utilization Monitoring Prometheus and Grafana Monitoring
-(Buyya and Barreto, 2015) Dynamic Resource Management Monitoring Aneka Platform
-(da Silva et al., 2023) Cost and Resource Usage Monitoring COSTA for Cost and Resource Monitoring
-(Dogani et al., 2023) Real-Time Dynamic Resource Monitoring Prometheus for Real-Time Monitoring
-(Wong et al., 2023) Resource Usage and Performance Issue Prometheus for Real-Time Data Collection
-(Marques et al., 2024) Real-Time Container Monitoring Prometheus with Forecasting for Resource Management
-(Santos et al., 2023) Distributed Cloud Resource Monitoring Issue AWS CloudWatch and Azure Monitor for Insights
-(Nagarajan et al., 2023) Resource Allocation and Energy Monitoring Issue Tools for Resource, Energy, and SLA Monitoring
-Category 3: Security, Compliance, and Trust Management Monitoring
-(Ramos et al., 2021) Provider Host Information Access Monitoring ML Monitoring
-(Rios et al., 2017) Continuous Security Control Monitoring MUSA Platform
-(Casola et al., 2017) Multi-Cloud Security Compliance Monitoring MUSA Security Compliance
-(Acquaviva et al., 2017) Service Continuity and Data Privacy Monitoring NoMISHAP Implementation
-(Wahab et al., 2016) Trust Relationship and Network Contact Monitoring Trust Bootstrapping
-(Fugkeaw, 2023) Access and Authentication Monitoring Issue Blockchain for Secure Access Event Logging
-(Wang et al., 2023) Malware Behavior and System Monitoring DockerWatch for Container Behavior Monitoring
-(Neeraj et al., 2023) Log Data Integrity Monitoring Issue Blockchain-based Logs for SLA Monitoring
-Category 4: Application Optimization and Adaptability Monitoring
-(Ferry et al., 2014) Adaptive System Design and State Monitoring Cloud Modelling Framework
-(Lorenzo et al., 2019) Partial Download Execution Monitoring FogDocker Implementation
-(Wurster et al., 2018) Event-Driven Application Behavior Monitoring TOSCA Event Modeling
-(Guerrero et al., 2018a) Microservice Optimization Monitoring MiCADO Optimization
-(Bou Ghantous and Gill, 2021) Real-Time Application Monitoring DRA Framework
-(Sofia et al., 2023) Container Application Adaptability Monitoring Advanced Monitoring with Orchestration Integration
-(Petrovski and Gusev, 2024) Serverless and Containerized Monitoring Challenge AWS Tools for Performance Monitoring
-Category 5: System Complexity and Standardization Monitoring
-(Sabbioni et al., 2020) Proprietary Monitoring Solution Standardization Pervasive Monitoring
-(Gao et al., 2017) Kernel Resource Isolation Mechanism Monitoring Leakage Defense
-(Alonso et al., 2019) Multi-Cloud Application Monitoring Complexity DECIDE DevOps Expansion
-(Barna et al., 2017) Monitoring Agent Coordination Autonomic Management
-(Sami and Mourad, 2020) Heterogeneous Fog Resource Management Monitoring Kubernetes Utility Architecture
-(Baresi et al., 2016) Application Topology and Dependency Monitoring MicroCloud TOSCA Library
-(Trihinas et al., 2015) Elastic Cloud Service Monitoring Research JCatascopia Monitoring
-(Fadda et al., 2019) Microservice Monitoring Diversity QoM Definition
-(Harwalkar et al., 2019b) Compute and Network Resource Utilization Monitoring Time Series Resource Utilization
-(Ahmad and Aujla, 2023) Real-Time Data Operations Monitoring Smart Contracts for GDPR Monitoring
-(Centofanti et al., 2024) Power Consumption Monitoring Issue Prometheus with Power Monitoring Integration
-(Tiwari and Sharma, 2023) Serverless Function Monitoring Complexity CLI Tool for Function Monitoring
-Category 6: Multi-cloud Coordination and Integration Monitoring
-(Sitaram et al., 2018) Integrated Multi-Cloud Management Monitoring Gnocchi Integration
-(Mfula and Nurminen, 2018) Multi-Cloud Monitoring Challenges Self-Healing Techniques
-(Kritikos et al., 2019) Cross-Level Monitoring Improvement Business Process Monitoring
-(Lee and Nam, 2023) Real-Time Network Policy Monitoring Challenge KUNERVA for Real-Time Network Policy Monitoring
-(Sofia et al., 2024) Decentralized System Observability Issue CODECO for Orchestration and Data Observability
-Security, Compliance, and Trust Monitoring: category addresses the challenges and solutions related to monitoring security, compliance, and trust management for containerized applications in multi-cloud environment.
-Out of the 121 reviewed studies, 9 specifically deal with these concerns.
-One challenge is Provider Host Information Access Monitoring, where limitations in accessing host data are overcome using ML Monitoring, which employs machine learning to track system performance and usage patterns.
-Another challenge, Continuous Security Control Monitoring, is addressed by the MUSA platform, which enforces and continuously monitors security in real time for containerized applications.
-Multi-Cloud Security Compliance Monitoring ensures continuous security compliance across multi-cloud environment.
-This is achieved through the MUSA Security Compliance Solution, a platform for continuous compliance management.
-The challenges of maintaining data privacy, fault tolerance, and service continuity in multi-cloud environments are addressed under Service Continuity and Data Privacy Monitoring.
-The NoMISHAP Implementation is proposed to handle these monitoring issues.
-Additionally, Trust Bootstrapping establishes trust through a chain of endorsement and decision tree classification.
-Access and Authentication Monitoring uses blockchain technology to enable secure and transparent logging of authentication activities.
-For Malware Behavior and System Monitoring, DockerWatch monitors container behavior with minimal performance overhead.
 Application Optimization and Adaptability Monitoring: This category identifies and discusses challenges related to optimizing and adapting applications in multi-cloud environment.
@@
 Lastly, Integrated Multi-cloud Management Monitoring requires advanced solutions, such as Integration with Gnocchi, to efficiently gather and monitor metrics from diverse cloud resources.
-[…]
-3.9. Tools and Frameworks (RQ6)
-In response to RQ6 regarding tools and frameworks for developing containerized multi-cloud applications, we identified 87 distinct tools and frameworks, classified into 5 subcategories (see Table 15 and the Tools and Frameworks sheet in (Waseem et al., 2024)).
+Security, Compliance, and Trust Monitoring: category addresses the challenges and solutions related to monitoring security, compliance, and trust management for containerized applications in multi-cloud environment.
+Out of the 121 reviewed studies, 9 specifically deal with these concerns.
+One challenge is Provider Host Information Access Monitoring, where limitations in accessing host data are overcome using ML Monitoring, which employs machine learning to track system performance and usage patterns.
+Table 14: Identified monitoring challenges and solutions for containerized applications in multi-cloud environment
+      ID        Challenge        Solution
+      Category 1: Performance Monitoring Consistency and Variability
+      [95]        Performance Monitoring        Performance Variation Monitoring
+      [96]        Benchmarking Performance Variation and Cost Efficiency Monitoring        SDBO Optimization
+      [103]        Website Performance and Availability Monitoring        Kubernetes Monitoring
+      [67]        User Experience Variability Monitoring        CSP QoS Metrics Collection
+      [122]        Independent Performance Metrics Monitoring        Performance Evaluations
+      [47]        Comprehensive Network Performance Measurement Monitoring        Server Selection Scheme
+      [53]        Microservice Performance Degradation Monitoring        Tail Latency Prediction
+      [137]        Container Performance Metrics Tracking        CLM and OSLM for Resource Adjustments
+      [87]        Container Performance and Resource Usage Monitoring        Metric and Container Monitoring Tools
+      [37]        Microservice Performance Monitoring Challenge        Prometheus, Grafana, Kubernetes for Monitoring
+      [92]        Cross-Cloud Service Performance Monitoring        FaaS and Storage Services for AI Monitoring
+      [119]        Stream Processing Performance Monitoring        RBAM for Real-Time Performance Insights
+      [134]        AWS Application Performance Monitoring        AWS CloudWatch for Performance Monitoring
+      [33]        Multi-Cloud Performance Visibility Issue        Monitoring Tools for Multi-Cloud Visibility
+      [141]        Cross-Technology Performance Monitoring Issue        TOSCA-based Workflows for Monitoring Applications
+      [49]        Industrial Robot System Monitoring Challenge        Real-Time System Performance Tracking Tools
+      Category 2: Resource and Infrastructure Management Monitoring
+      [34]        Network Communication Monitoring        Service Mesh Implementation
+      [120]        System Image and Configuration Monitoring        Docker Server Deployment
+      [126]        Vcluster and Framework Management Monitoring        LXC Implementation
+      [60]        Multi-Cloud Resource Utilization Monitoring        Prometheus and Grafana Monitoring
+      [98]        Dynamic Resource Management Monitoring        Aneka Platform
+      [115]        Cost and Resource Usage Monitoring        COSTA for Cost and Resource Monitoring
+      [116]        Real-Time Dynamic Resource Monitoring        Prometheus for Real-Time Monitoring
+      [136]        Resource Usage and Performance Issue        Prometheus for Real-Time Data Collection
+      [88]        Real-Time Container Monitoring        Prometheus with Forecasting for Resource Management
+      [125]        Distributed Cloud Resource Monitoring Issue        AWS CloudWatch and Azure Monitor for Insights
+      [43]        Resource Allocation and Energy Monitoring Issue        Tools for Resource, Energy, and SLA Monitoring
+      Category 3: Security, Compliance, and Trust Management Monitoring
+      [132]        Provider Host Information Access Monitoring        ML Monitoring
+      [56]        Continuous Security Control Monitoring        MUSA Platform
+      [127]        Multi-Cloud Security Compliance Monitoring        MUSA Security Compliance
+      [146]        Service Continuity and Data Privacy Monitoring        NoMISHAP Implementation
+      [72]        Trust Relationship and Network Contact Monitoring        Trust Bootstrapping
+      [144]        Access and Authentication Monitoring Issue        Blockchain for Secure Access Event Logging
+      [142]        Malware Behavior and System Monitoring        DockerWatch for Container Behavior Monitoring
+      [147]        Log Data Integrity Monitoring Issue        Blockchain-based Logs for SLA Monitoring
+      Category 4: Application Optimization and Adaptability Monitoring
+      [89]        Adaptive System Design and State Monitoring        Cloud Modelling Framework
+      [113]        Partial Download Execution Monitoring        FogDocker Implementation
+      [140]        Event-Driven Application Behavior Monitoring        TOSCA Event Modeling
+      [99]        Microservice Optimization Monitoring        MiCADO Optimization
+      [90]        Real-Time Application Monitoring        DRA Framework
+      [118]        Container Application Adaptability Monitoring        Advanced Monitoring with Orchestration Integration
+      [100]        Serverless and Containerized Monitoring Challenge        AWS Tools for Performance Monitoring
+      Category 5: System Complexity and Standardization Monitoring
+      [65]        Proprietary Monitoring Solution Standardization        Pervasive Monitoring
+      [70]        Kernel Resource Isolation Mechanism Monitoring        Leakage Defense
+      [105]        Multi-Cloud Application Monitoring Complexity        DECIDE DevOps Expansion
+      [52]        Monitoring Agent Coordination        Autonomic Management
+      [29]        Heterogeneous Fog Resource Management Monitoring        Kubernetes Utility Architecture
+      [31]        Application Topology and Dependency Monitoring        MicroCloud TOSCA Library
+      [84]        Elastic Cloud Service Monitoring Research        JCatascopia Monitoring
+      [85]        Microservice Monitoring Diversity        QoM Definition
+      [78]        Compute and Network Resource Utilization Monitoring        Time Series Resource Utilization
+      [143]        Real-Time Data Operations Monitoring        Smart Contracts for GDPR Monitoring
+      [139]        Power Consumption Monitoring Issue        Prometheus with Power Monitoring Integration
+      [131]        Serverless Function Monitoring Complexity        CLI Tool for Function Monitoring
+      Category 6: Multi-cloud Coordination and Integration Monitoring
+      [81]        Integrated Multi-Cloud Management Monitoring        Gnocchi Integration
+      [42]        Multi-Cloud Monitoring Challenges        Self-Healing Techniques
+      [76]        Cross-Level Monitoring Improvement        Business Process Monitoring
+      [145]        Real-Time Network Policy Monitoring Challenge        KUNERVA for Real-Time Network Policy Monitoring
+      [138]        Decentralized System Observability Issue        CODECO for Orchestration and Data Observability
+Another challenge, Continuous Security Control Monitoring, is addressed by the MUSA platform, which enforces and continuously monitors security in real time for containerized applications.
+Multi-Cloud Security Compliance Monitoring ensures continuous security compliance across multi-cloud environment.
+This is achieved through the MUSA Security Compliance Solution, a platform for continuous compliance management.
+The challenges of maintaining data privacy, fault tolerance, and service continuity in multi-cloud environments are addressed under Service Continuity and Data Privacy Monitoring.
+The NoMISHAP Implementation is proposed to handle these monitoring issues.
+Additionally, Trust Bootstrapping establishes trust through a chain of endorsement and decision tree classification.
+Access and Authentication Monitoring uses blockchain technology to enable secure and transparent logging of authentication activities.
+For Malware Behavior and System Monitoring, DockerWatch monitors container behavior with minimal performance overhead.
+Takeaway 11: Reviewed studies reveal that monitoring containerized applications in multi-cloud environments introduces unique challenges, such as inconsistent performance metrics across cloud providers, real-time trust and security event tracking, and fragmented observability due to container orchestration spanning heterogeneous infrastructures.
+These challenges are addressed through specialized solutions like cross-cloud Prometheus-Grafana stacks, blockchain-based SLA logging, and service mesh integrations, reflecting the need for tightly coupled monitoring mechanisms tailored specifically to the hybrid nature of containerized multi-cloud deployments.
+3.9 Tools and Frameworks (RQ6)
+In response to RQ6 regarding tools and frameworks for developing containerized multi-cloud applications, we identified 87 distinct tools and frameworks, classified into 5 subcategories (see Table 15 and the Tools and Frameworks sheet in [11]).
 Below, we provide an overview of each category, along with key tools and frameworks, highlighting their functionality, adoption, and distinctive features.
@@
 Overall, this category covers key orchestration and container management tools that address diverse deployment needs in multi-cloud environment.
-Multi-cloud and Hybrid Cloud Platforms: are designed to support organizations in deploying and managing applications across different cloud providers smoothly.
+Multi-cloud and Hybrid Cloud Platforms are designed to support organizations in deploying and managing applications across different cloud providers smoothly.
 In this category, we identify and organize tools like Cloudify and CMP2, which offer end-to-end multi-cloud orchestration capabilities, along with platforms such as Cloudcheckr and MistIO, which provide monitoring and cost management across various cloud services.
@@
 Networking, Security, and Access Management Tools: This category represents the tools that help to minimize security risks and ensure connectivity across containerized workloads in multi-cloud systems.
-As multi-cloud deployments become increasingly complex, advanced networking tools like Open vSwitch (OvS) enable developers to manage virtual networks across various cloud platforms.
-Security tools such as Seccomp-BPF and Linux Security Modules (LSM) provide critical protection to containers by enforcing strict access controls and sandboxing.
-Identity management solutions, such as Host Identity Protocol (HIP) and access control policies, ensure proper authentication and authorization for users and services in the cloud.
-Monitoring, Management, and Automation Tools: Tools in this category provide the foundation for a scalable and automated approach to managing distributed applications by ensuring that workloads are continuously optimized and monitored on cloud platforms.
-Tools like Prometheus and Grafana offer powerful visualization capabilities for monitoring CPU usage, memory consumption, and network activity, aiding system administrators.
-With Terraform and Puppet, it is possible to automate infrastructure management, enabling developers to define infrastructure as code and automate provisioning across different cloud environments.
-Other tools, such as Ceilometer and Gnocchi, are responsible for gathering telemetry data in large OpenStack deployments.
+Advanced networking tools like Open vSwitch (OvS) allow developers to manage virtual networks across multiple cloud platforms as multi-cloud deployments grow more complicated.
+By implementing stringent access controls and sandboxing, security tools like Seccomp-BPF and Linux Security Modules (LSM) offer vital container protection.
+Proper authentication and authorization for cloud users and services are ensured by identity management solutions like Host Identity Protocol (HIP) and access control policies.
+Monitoring, Management, and Automation Tools: Tools in this category help manage large-scale cloud applications in an easy and automated way.
+They make sure everything runs smoothly by constantly checking and improving how resources are used.
+For example, Prometheus and Grafana help system administrators by showing important information like CPU use, memory load, and network traffic in clear charts and dashboards.
+Tools like Terraform and Puppet make it possible to write and reuse code to set up servers and other infrastructure automatically across different cloud services.
+This saves time and reduces errors.
+In big systems using OpenStack, tools such as Ceilometer and Gnocchi collect data about how the system is working, which helps with tracking performance and solving issues.
 Development, APIs, and Cloud Storage Solutions: This category organizes the main tools that support software development, API integration, and data storage in containerized multi-cloud environment.
@@
 On the data storage side, tools like gcloud storage and azblob provide scalable and fault-tolerant systems for managing large datasets in cloud environments.
-Table 15. Tools and frameworks for containerized applications in multi-cloud environment
+Table 15: Tools and frameworks for containerized applications in multi-cloud environment
        Category        Tool, Framework, and Language
-Container Orchestration and Management Kubernetes (Pai and Kumar, 2021; Wong et al., 2019; Gupta et al., 2021; Bolivar et al., 2018; del Castillo et al., 2013; Rodigari et al., 2021; Rahman and Lama, 2019; Pellegrini et al., 2017; Quint and Kratzke, 2018; Barletta et al., 2024; Struhár et al., 2024; Barati et al., 2023; Horchulhack et al., 2024; Dogani et al., 2023; Wong et al., 2023; Marques et al., 2024; Orzechowski et al., 2023; Merlino et al., 2024; Ahmad and Aujla, 2023; Centofanti et al., 2024; Hossain et al., 2023; Liu and Qin, 2024; Tundo et al., 2024; Pandi et al., 2023; Sofia et al., 2024; Lee and Nam, 2023; Sofia et al., 2023; Nguyen and Chien, 2023; Tiwari and Sharma, 2023; Santos et al., 2023; Roy and Kalafatis, 2023; Kataru et al., 2023; Benmerar et al., 2023; Alonso et al., 2023a; Harzenetter et al., 2023; Huang and Wang, 2023; Wang et al., 2023; Nagarajan et al., 2023), Docker (Pham et al., 2015; Naik, 2016a; Ouyang et al., 2022; Slominski et al., 2015; Pai and Kumar, 2021; Gao et al., 2017; Alonso et al., 2019; Barna et al., 2017; Gupta et al., 2021; Sheridan et al., 2017; Mittal and Risco-Martín, 2017; Al-shammari and Alwan, 2018; Carpio et al., 2020; Bolivar et al., 2018; Rodigari et al., 2021; Rahman and Lama, 2019; Mfula and Nurminen, 2018; Ranjbar et al., 2017; Serhiienko and Spillner, 2018; Quint and Kratzke, 2018), Docker Swarm (Pai and Kumar, 2021; Wong et al., 2019; Altran et al., 2022; Rodigari et al., 2021; Quint and Kratzke, 2018), Mesos (Cui et al., 2016; Wong et al., 2019; Rodigari et al., 2021; Quint and Kratzke, 2018), LXC (Pahl, 2015; Gao et al., 2017; Gupta et al., 2021), OverlayFS (Al-shammari and Alwan, 2018), runC (Bélair et al., 2019), Container Technology (Wang et al., 2020), Docklet (Cui et al., 2016)
-Multi-cloud and Hybrid Cloud Platforms multi-cloud environment (Shi et al., 2020; Fadda et al., 2019; Di Modica et al., 2019; Tang, 2021), Cloudify (Zeck and Bouroudjian, 2017), CloudBroker Platform (Kiss, 2018), Cloudcheckr (Serhiienko and Spillner, 2018), MistIO (Serhiienko and Spillner, 2018), ManageIQ (Tang, 2021; Serhiienko and Spillner, 2018), CMP2 (Serhiienko and Spillner, 2018), AWS SDK (Graupner et al., 2015a), CloudSME Simulation Platform (CSSP) (Kiss, 2018), Vagrant (Pham et al., 2015), AWS OpsWorks (Pham et al., 2015), Universal Compute Xchange (UCX) (Zeck and Bouroudjian, 2017), Liferay portlets (Kiss, 2018), Cloud Service Providers (CSPs) (Graupner et al., 2015a), AWS Data Pipeline (Sitaram et al., 2018), GENI (Pandey et al., 2022), Google Cloud Platform (Pandey et al., 2022; Acquaviva et al., 2017; Tang, 2021; Quint and Kratzke, 2018), AWS EC2 (He et al., 2022; Buyya and Barreto, 2015; Pai and Kumar, 2021), Alibaba cloud (Zeck and Bouroudjian, 2017; Quint and Kratzke, 2018), Tencent (Zeck and Bouroudjian, 2017; Quint and Kratzke, 2018), OpenStack (Cui et al., 2016; Dreibholz et al., 2019), VMware vSphere (Mfula and Nurminen, 2018), Aliyun ECS (Hua et al., 2020), Baidu BCE (Hua et al., 2020), Tencent CW (Hua et al., 2020), JD VW (Hua et al., 2020), GoGrid (Buyya and Barreto, 2015)
-Networking, Security, and Access Management Open vSwitch (OvS) (Alaluna et al., 2017), Cloud networks (Ranjbar et al., 2017), Host Identity Protocol (HIP) (Ranjbar et al., 2017), Gateways (Harwalkar et al., 2019a), Pledge (Bélair et al., 2019), Seccomp-BPF (Bélair et al., 2019), Landlock LSM (Bélair et al., 2019), Linux Security Modules (LSM) (Bélair et al., 2019), MUSA Security Assurance Platform (Rios et al., 2017), Firewall Deployment (Sheridan et al., 2017), Access control policies (Gao et al., 2017), Attribute-Based Encryption and Signature (Shen et al., 2017), ABE/ABS cryptographic model (Shen et al., 2017)
-Monitoring, Management, and Automation Prometheus (Wong et al., 2019; Dogani et al., 2023; Wong et al., 2023; Marques et al., 2024; Orzechowski et al., 2023; Merlino et al., 2024; Ahmad and Aujla, 2023; Centofanti et al., 2024; Hossain et al., 2023; Liu and Qin, 2024; Tundo et al., 2024; Roy and Kalafatis, 2023), Grafana (Alonso et al., 2019; Hossain et al., 2023; Liu and Qin, 2024; Tundo et al., 2024; Roy and Kalafatis, 2023), Telegraf (Alonso et al., 2019), AWS CloudWatch (Alhamazani et al., 2014), cAdvisor (Wong et al., 2019), Ceilometer (del Castillo et al., 2013; Sitaram et al., 2018), Gnocchi (Sitaram et al., 2018), ELK stack (Mfula and Nurminen, 2018), Terraform (Alonso et al., 2019; Gupta et al., 2021; da Silva et al., 2023; Pandi et al., 2023; Harzenetter et al., 2023), CloudBees (Pham et al., 2015), Jenkins (Pandey et al., 2022; Solayman and Qasha, 2023), Puppet (Pham et al., 2015), Chef (Pham et al., 2015; Sheridan et al., 2017), GitHub (Sheridan et al., 2017), Shell scripts (Rios et al., 2017), Maven (Pham et al., 2015), NPM (Pham et al., 2015), Consul (Pham et al., 2015)
-Development, APIs, and Cloud Storage Public cloud storage (Harwalkar et al., 2019a), Private cloud storage (Harwalkar et al., 2019a), Hybrid cloud storage (Harwalkar et al., 2019a), Google Cloud Storage (Ramesh et al., 2023), Azure Blob Storage (Ramesh et al., 2023), REST API (Harwalkar et al., 2019a), YAML (Serhiienko and Spillner, 2018), OpenVAS Vulnerability Scanner (Sheridan et al., 2017), DAX format (Tang, 2021), Java (Sheridan et al., 2017; Rios et al., 2017; Carpio et al., 2020; Bucur et al., 2018; Quint and Kratzke, 2018; Challita et al., 2017), Go programming language (Al-shammari and Alwan, 2018; Lorenzo et al., 2019), Python (Carpio et al., 2020; Bucur et al., 2018; del Castillo et al., 2013; Sitaram et al., 2018; Serhiienko and Spillner, 2018), Ruby (Bucur et al., 2018), PHP (Bucur et al., 2018), C# (Bucur et al., 2018), Cloud Modelling Framework (CLOUMDF) (Ferry et al., 2014), Cloud Modelling Language (CLOUDML) (Ferry et al., 2014; Challita et al., 2017), CAMEL modeling language (Casola et al., 2017; Quint and Kratzke, 2018; Kritikos et al., 2019), Eclipse IDE (Alonso et al., 2019), MCSLA editor (Alonso et al., 2019)
-[…]
-4. Discussion
-4.1. Containers in Multi-Cloud: Roles and Strategies (RQ1)
+       Container Orchestration and Management        Kubernetes [59, 111, 103, 40, 107, 35, 53, 73, 50, 82, 137, 135, 87, 116, 136, 88, 110, 36, 143, 139, 37, 130, 91, 117, 138, 145, 118, 119, 131, 125, 46, 134, 104, 33, 141, 49, 142, 43], Docker [86, 148, 121, 44, 59, 70, 105, 52, 103, 55, 106, 112, 30, 40, 35, 53, 42, 68, 45, 50], Docker Swarm [59, 111, 114, 35, 50], Mesos [126, 111, 35, 50], LXC [60, 70, 103], OverlayFS [112], runC [71], Container Technology [66], Docklet [126]
+       Multi-cloud and Hybrid Cloud Platforms        multi-cloud environment [97, 85, 129, 41], Cloudify [62], CloudBroker Platform [63], Cloudcheckr [45], MistIO [45], ManageIQ [41, 45], CMP2 [45], AWS SDK [57], CloudSME Simulation Platform (CSSP) [63], Vagrant [86], AWS OpsWorks [86], Universal Compute Xchange (UCX) [62], Liferay portlets [63], Cloud Service Providers (CSPs) [57], AWS Data Pipeline [81], GENI [67], Google Cloud Platform [67, 146, 41, 50], AWS EC2 [54, 98, 59], Alibaba cloud [62, 50], Tencent [62, 50], OpenStack [126, 122], VMware vSphere [42], Aliyun ECS [93], Baidu BCE [93], Tencent CW [93], JD VW [93], GoGrid [98]
+       Networking, Security, and Access Management        Open vSwitch (OvS) [124], Cloud networks [68], Host Identity Protocol (HIP) [68], Gateways [108], Pledge [71], Seccomp-BPF [71], Landlock LSM [71], Linux Security Modules (LSM) [71], MUSA Security Assurance Platform [56], Firewall Deployment [55], Access control policies [70], Attribute-Based Encryption and Signature [109], ABE/ABS cryptographic model [109]
+       Monitoring, Management, and Automation        Prometheus [111, 116, 136, 88, 110, 36, 143, 139, 37, 130, 91, 46], Grafana [105, 37, 130, 91, 46], Telegraf [105], AWS CloudWatch [83], cAdvisor [111], Ceilometer [107, 81], Gnocchi [81], ELK stack [42], Terraform [105, 103, 115, 117, 141], CloudBees [86], Jenkins [67, 38], Puppet [86], Chef [86, 55], GitHub [55], Shell scripts [56], Maven [86], NPM [86], Consul [86]
+       Development, APIs, and Cloud Storage        Public cloud storage [108], Private cloud storage [108], Hybrid cloud storage [108], Google Cloud Storage [92], Azure Blob Storage [92], REST API [108], YAML [45], OpenVAS Vulnerability Scanner [55], DAX format [41], Java [55, 56, 30, 133, 50, 94], Go programming language [112, 113], Python [30, 133, 107, 81, 45], Ruby [133], PHP [133], C# [133], Cloud Modelling Framework (CLOUMDF) [89], Cloud Modelling Language (CLOUDML) [89, 94], CAMEL modeling language [127, 50, 76], Eclipse IDE [105], MCSLA editor [105]
+Takeaway 12: The reviewed studies reveal that tools and frameworks commonly used in general container or cloud settings (e.g., Kubernetes, Docker, Terraform, Prometheus) require specific adaptations or configurations to operate effectively in containerized multi-cloud environments.
+This includes challenges such as consistent orchestration across heterogeneous cloud APIs, maintaining observability across isolated infrastructures, and enabling secure, policy-compliant automation at scale.
+Moreover, hybrid-enabling tools like OpenStack, Cloudify, and MistIO emerged as critical for unifying management across cloud boundaries — highlighting how the combination of containerization and multi-cloud imposes unique operational and integration demands not encountered when using either paradigm alone.
+4 Discussion
+The following discussion synthesizes key findings for each research question and their implications for research and practice.
+These implications also highlight open problems and unresolved challenges that require further research attention, particularly those emerging from the intersection of containerization and multi-cloud environments.
+4.1 Containers in Multi-Cloud: Roles and Strategies (RQ1)
 Containers play a crucial role in the development and operation of modern software systems within multi-cloud environment.
 Understanding the various roles and implementation strategies of containers provides significant insights for both researchers and practitioners.
-Multifaceted Role of Containers in Multi-Cloud Environment: Containers fulfill diverse roles within multi-cloud environment, such as application deployment, service hosting, and resource management.
-They facilitate deployment of applications across a broad range of systems, from web apps to IoT applications.
-Furthermore, containers contribute to enhancing security in multi-cloud operations.
-Implications: Based on these findings, researchers have opportunities to investigate further areas, such as potential security vulnerabilities, management complexities, and performance overheads.
-Similarly, practitioners can use this understanding to inform efficient and effective deployment strategies, leading to more robust multi-cloud implementations.
-Container Technology Features and their Significance: The unique features of container technology, such as lightweight virtualization, compatibility with orchestration tools, and application in edge computing, offer significant advantages.
+Multifaceted Role of Containers in Multi-Cloud Environment: In a multi-cloud environment, containers serve multiple purposes, such as managing resources, hosting services, and deploying applications.
+They simplify the deployment of apps across various platforms, including web applications and Internet of Things (IoT) applications.
+Additionally, containers enhance security in multi-cloud operations.
+These roles are shaped by the inherent heterogeneity of multi-cloud settings, where containers are used to ensure portability, manage provider-specific constraints, and support workload migration across isolated cloud infrastructures.
+Implications: These findings open up opportunities for researchers to explore additional areas, including security vulnerabilities, management challenges, and performance overhead.
+Likewise, practitioners can apply these insights to design more efficient and effective deployment strategies, ultimately contributing to stronger and more reliable multi-cloud implementations.
+Container Technology Features and their Significance: Container technology offers several benefits due to its unique features, such as support for edge computing, compatibility with orchestration tools, and lightweight virtualization.
+In multi-cloud environments, these features enable containers to act as a unifying abstraction layer across diverse infrastructure stacks, allowing consistent deployment behavior despite differing cloud-specific configurations.
 Implications: Researchers can explore these features further, including how characteristics like lightweight virtualization influence container performance and effectiveness in multi-cloud operations.
-Practitioners can leverage this understanding for better utilization of container technology and more efficient multi-cloud deployments.
+Practitioners can use this understanding for better utilization of container technology and more efficient multi-cloud deployments.
 Implementation Strategies and their Real-World Implications: The strategic implementation of containers significantly affects their functionality in multi-cloud environment.
 Strategies related to deployment, orchestration, security, performance optimization, cloud service management, and data storage can have significant implications.
+The reviewed strategies demonstrate how container management must adapt to cloud-specific APIs, variable service guarantees, and cross-cloud policy enforcement, revealing how the interplay between containerization and multi-cloud governance introduces distinctive design trade-offs.
 Implications: Researchers may explore these strategies further, including investigations into their real-world implications, benefits, and limitations.
@@
 Containers offer improved software unit encapsulation, portability, and efficiency, which are particularly important in the distributed and resource-constrained nature of IoT.
+In multi-cloud scenarios, this impact is increase by the need to coordinate deployments across heterogeneous edge-cloud setups and maintain interoperability between varied cloud-native services and orchestration frameworks.
 Implications: Researchers can further explore potential challenges in detail such as network complexity, security, and the small footprint requirement of IoT devices.
 For practitioners, understanding the impact of containerization can lead to the development of more robust, efficient, and secure IoT and app solutions.
-4.2. Patterns and Strategies (RQ2)
+4.2 Patterns and Strategies (RQ2)
 Containerized applications in multi-cloud environment requires an understanding of prevailing patterns and strategies.
@@
 Diverse Architectural Patterns and Communication Strategies: Our study found that Microservice Architecture and Service-Oriented Architecture (SOA) are the most common used architectural patterns for designing container-based applications in multi-cloud environment.
-These results align with the existing literature that emphasizes the use of MSA and SOA because of their support for modular development, scalability, and efficient resource use (Balalaie et al., 2016; Dragoni et al., 2017).
-The other patterns frequently reported are Service Mesh and Service Chaining, which are employed as communication and networking patterns, likely due to their ability to manage complex, distributed service-to-service communications in a scalable way (Zhamak, 2018).
-The frequent use of black-Green Deployment and Object Store Service deployment patterns has also been identified in this study, which may stem from their ability to minimize downtime and ensure high availability—essential factors in a cloud-based setting (fowler2010blackgreen).
+These results align with the existing literature that emphasizes the use of MSA and SOA because of their support for modular development, scalability, and efficient resource use [149, 150].
+The other patterns frequently reported are Service Mesh and Service Chaining, which are employed as communication and networking patterns, likely due to their ability to manage complex, distributed service-to-service communications in a scalable way [151].
+The frequent use of black-Green Deployment and Object Store Service deployment patterns has also been identified in this study, which may stem from their ability to minimize downtime and ensure high availability—essential factors in a cloud-based setting [152].
+Notably, these architectural and communication patterns are adapted in the multi-cloud setting to support distributed orchestration, cross-cloud resilience, and integration across heterogeneous platforms—attributes that are less emphasized in single-cloud deployments.
+For instance, service mesh implementations in multi-clouds enable secure communication across isolated cloud boundaries, and decentralized orchestration patterns accommodate independently governed cloud nodes.
 Implications: Practitioners can enhance the design, scalability, and efficiency of their applications in a multi-cloud environment by understanding and applying these architectural and communication patterns.
@@
 Importance of Container Management and Multi-Cloud Strategies: Our findings report the significance of robust container management and multi-cloud strategies.
-Techniques like the Linux Container (LXC) project, Docker Container Images, and Lightweight Virtualization are frequently used, emphasizing the industry trend towards containerization due to its benefits in efficiency, scalability, and platform independence (Vaughn, 2020).
-Furthermore, the popularity of multi-cloud strategies like Hybrid/Multi-cloud Approach and Multi-cloud Load Balancing confirms the growing industry move towards using multiple cloud providers to enhance redundancy, flexibility, and avoid vendor lock-in (Gillam et al., 2013).
+Techniques like the Linux Container (LXC) project, Docker Container Images, and Lightweight Virtualization are frequently used, emphasizing the industry trend towards containerization due to its benefits in efficiency, scalability, and platform independence [153].
+Furthermore, the popularity of multi-cloud strategies like Hybrid/Multi-cloud Approach and Multi-cloud Load Balancing confirms the growing industry move towards using multiple cloud providers to enhance redundancy, flexibility, and avoid vendor lock-in [154].
+In particular, the reviewed studies illustrate how container technologies are adapted to meet challenges specific to multi-cloud deployments, such as orchestrating containers across provider-specific APIs, aligning container runtime environments across clouds, and ensuring policy compliance in cross-provider setups.
+These represent key peculiarities introduced by multi-cloud environments into the containerization landscape.
 Implications: Practitioners should focus on gaining expertise in container technologies and multi-cloud strategies to take advantage of their numerous benefits.
@@
 Security, Resilience, and Migration Patterns and Strategies: Our study also identifies patterns and strategies related to security, resilience, and efficient migration strategies in a multi-cloud context.
-Strategies such as File Encryption, Attribute-Based Encryption, and Secure Data sharing are often employed to ensure data security, an area of growing concern with the increase in cloud-based applications (Svantesson and Clarke, 2010).
-Moreover, resilience strategies such as Fault-tolerance with Redundant Engines and RAFT Consensus Algorithm are essential to ensure system reliability in distributed, multi-cloud environment (Ongaro and Ousterhout, 2014).
-Migration between different cloud environments, facilitated by strategies like Service-oriented Migration and Migration to a Virtualized Container, is another vital aspect, given the growth in cloud services and the need for interoperability (Jamshidi et al., 2013).
-Migrating applications and data between cloud environments is an important consideration in a multi-cloud scenario (Jamshidi et al., 2013).
-Strategies such as image-based migration and migration to virtualized containers, like Docker, are commonly used (Vaughn, 2020).
+Strategies such as File Encryption, Attribute-Based Encryption, and Secure Data sharing are often employed to ensure data security, an area of growing concern with the increase in cloud-based applications [155].
+Moreover, resilience strategies such as Fault-tolerance with Redundant Engines and RAFT Consensus Algorithm are essential to ensure system reliability in distributed, multi-cloud environment [156].
+Migration between different cloud environments, facilitated by strategies like Service-oriented Migration and Migration to a Virtualized Container, is another vital aspect, given the growth in cloud services and the need for interoperability [157].
+Migrating applications and data between cloud environments is an important consideration in a multi-cloud scenario [157].
+Strategies such as image-based migration and migration to virtualized containers, like Docker, are commonly used [153].
 These techniques facilitate portability and provide consistent environments across different cloud platforms.
 Additionally, they have significant implications for scaling and managing applications across different cloud service providers.
+Unlike traditional container security or resilience strategies in single-cloud contexts, the reviewed studies show that in multi-cloud scenarios, these approaches must handle cross-cloud identity management, consistency in fault detection and recovery across isolated infrastructures, and security rule federation across providers.
+Similarly, migration patterns uniquely address compatibility mismatches and compliance constraints introduced by the heterogeneity of cloud APIs and services.
 Implications: These findings suggest that security, resilience, and efficient migration should be key areas of focus for practitioners operating in multi-cloud environment.
 Researchers may also look into novel techniques for enhancing cloud security, improving system resilience, and easing the process of cloud migration.
-4.3. Quality Attributes and Tactics (RQ3)
+4.3 Quality Attributes and Tactics (RQ3)
 The exploration and implementation of quality attributes and associated tactics are critical to the successful deployment of containerized applications in multi-cloud environment, shaping performance, security, and compatibility characteristics.
@@
 Hashem et al.
-previously suggested similar efficiency gains from machine learning and container technology, validating the interpretive value of our results (Hashem et al., 2016).
+previously suggested similar efficiency gains from machine learning and container technology, validating the interpretive value of our results [158].
 Specifically, the application of machine learning aligns with a broader trend towards intelligent cloud resource management, a domain that Buyya et al.
-identified as significant in cloud computing research (Buyya et al., 2020).
+identified as significant in cloud computing research [159].
 Meanwhile, location-aware service brokering emerges as a novel and impactful strategy, stressing the role of geographical considerations in further optimizing containerized applications’ performance.
+Importantly, the reviewed studies emphasize that performance tactics must accommodate multi-cloud-specific challenges such as distributed latency, resource fragmentation across providers, and dynamic SLA enforcement, which are less relevant in single-cloud settings.
 Implications: For researchers, the results of our study offer opportunities to explore the integration of AI techniques with container-based cloud architectures further, especially the application of location-aware service brokering in multi-cloud environment.
@@
 It highlights the combination of encryption, network security protocols, machine learning, and user-specified security measures, while also encouraging deployment across different cloud providers.
-This finding complements Singh et al.’s argument for a comprehensive approach to cloud security, and aligns with the recent concept of distributed cloud security (Singh et al., 2016).
-Our results also report the importance of user-specified security measures, which are in agreement with the user-centric security model proposed by Shin and Dong-Hee (Shin, 2013).
+This finding complements Singh et al.’s argument for a comprehensive approach to cloud security, and aligns with the recent concept of distributed cloud security [160].
+Our results also report the importance of user-specified security measures, which are in agreement with the user-centric security model proposed by Shin and Dong-Hee [161].
+Notably, multi-cloud containerization introduces additional layers of complexity in managing security policies across heterogeneous platforms and enforcing access control in federated settings.
+Tactics such as decentralized IAM logging and policy-aware container placement directly respond to these unique multi-cloud constraints.
 Implications: From a research perspective, the interact between different security measures and their effect on overall system security in multi-cloud environment warrants further investigation.
@@
 Ensuring Compatibility Across Different Cloud Providers: The results of our study suggests strategies to ensure compatibility in multi-cloud environment through standardization of interfaces, componentization, and interoperability approaches.
-Emphasizing lightweight communication protocols and interoperability between clouds, the results related to compatibility mirrors Mell and Grance’s principles of service-oriented architecture, reinforcing the need for effective communication across different cloud platforms (Mell et al., 2011).
+Emphasizing lightweight communication protocols and interoperability between clouds, the results related to compatibility mirrors Mell and Grance’s principles of service-oriented architecture, reinforcing the need for effective communication across different cloud platforms [162].
 Petcu et al.
-have also previously stressed the importance of standardization and componentization for ensuring compatibility (Petcu et al., 2018).
+have also previously stressed the importance of standardization and componentization for ensuring compatibility [163].
+Our findings highlight that achieving compatibility in a containerized multi-cloud context often involves abstracting away provider-specific interfaces through standardized APIs and decoupling infrastructure logic via containerization.
+These are adaptations not typically required in homogeneous cloud environments.
 Implications: From a research standpoint, the dynamic between lightweight communication protocols, interface standardization, and overall system compatibility in multi-cloud deployments offers fertile ground for future work.
@@
 Future research could further explore the intersection of these quality attributes to develop more robust, efficient, and secure multi-cloud applications by using containerise technologies.
-4.4. Automation Challenges and Solution Framework (RQ4 - RQ5)
+4.4 Automation Challenges and Solution Framework (RQ4 - RQ5)
 The Automation Challenge-Solution Framework for Containerized Applications in multi-cloud environment (see Table LABEL:tab:Automation-solutions) suggests that multi-cloud management for containerized application is a complex task encompassing several dimensions.
@@
 This indicates that addressing one challenge may potentially impact the other areas, necessitating a holistic approach to tackle these issues.
+What makes these challenges unique to containerized multi-cloud setups is the dual complexity arising from container portability and cloud heterogeneity.
+Containers, while offering mobility and reproducibility, must now adapt to diverse runtime configurations, networking policies, and resource scheduling strategies across cloud providers.
 The multi-dimensionality of the challenges implies the need for an integrated approach that considers the interdependencies among these dimensions.
@@
 This suggests the growing maturity and applicability of these technologies in complex scenarios, a significant step towards achieving efficient multi-cloud management.
-Previous studies (e.g., (de Laat et al., 2020), (Chen et al., 2021), (Gill et al., 2022)) have recognized the potential of AI and ML in resolving complex cloud management issues, which is in line with the findings from the framework.
+Previous studies (e.g., [164], [165], [166]) have recognized the potential of AI and ML in resolving complex cloud management issues, which is in line with the findings from the framework.
 This further validates the significant role AI and ML have started to play in this domain.
+The reviewed studies specifically show how AI/ML techniques help in real-time placement of containers across heterogeneous clouds, optimizing latency and cost under dynamic workload and policy constraints—issues that are intensified by the distributed and federated nature of multi-cloud environments.
 The increasing adoption of AI/ML in multi-cloud environment signifies an essential evolution in the field.
@@
 Importance of Standardization and Interoperability: The framework underscores the crucial role of standardization and interoperability in multi-cloud environment, facilitating seamless integration and management across different cloud platforms.
-Recent literature validates the importance of standardization and interoperability in multi-cloud environment, emphasizing their role in enhancing the efficiency of cloud services (Alonso et al., 2023b).
+Recent literature validates the importance of standardization and interoperability in multi-cloud environment, emphasizing their role in enhancing the efficiency of cloud services [167].
 This agrees with our results, further confirming the critical need for these factors in the domain.
 The need for standardization and interoperability reiterates the necessity for a unified framework or protocol across different cloud environments.
+In containerized deployments, interoperability challenges become more pronounced due to variations in networking stacks, security policies, and orchestration tools between providers.
+Standardization at the container runtime level (e.g., Docker images, Kubernetes manifests) and orchestration abstraction (e.g., TOSCA, Helm charts) plays a vital role in enabling automation workflows that span diverse cloud ecosystems.
 It suggests that attention to these aspects can lead to more seamless, efficient, and secure management of multi-cloud services.
@@
 For practitioners, the findings suggest the need to adopt standards and focus on interoperability while selecting and implementing cloud services, ensuring more efficient management across different platforms.
-4.5. Security Challenges and Solutions Framework (RQ4 - RQ5)
+4.5 Security Challenges and Solutions Framework (RQ4 - RQ5)
 The framework proposed in our study provides a comprehensive and systematic approach to addressing the complex security issues associated with containerized applications in multi-cloud environment.
@@
 The selected studies suggest a variety of methods, including secure container orchestration, container isolation, and advanced deployment strategies that make use of deep reinforcement learning.
-This aligns with the existing literature (e.g., (Zhong et al., 2022), (Li et al., 2023) (Combe et al., 2016)) which highlights the unique security considerations brought forth by containerization.
+This aligns with the existing literature (e.g., [168], [169] [170]) which highlights the unique security considerations brought forth by containerization.
 Containers, being lighter than traditional virtual machines and offering process-level isolation, have revolutionized the way applications are packaged and run.
@@
 Implementing these will be key to leveraging the benefits of containerization in a secure manner.
+Implementing these will be key to using the benefits of containerization in a secure manner.
+In multi-cloud settings, these challenges are compounded by the need to secure container interactions across diverse platforms, enforce consistent policies under different providers, and protect workloads that dynamically move between clouds.
 Implications: For researchers, this highlights an area of potential study: the development of advanced container-specific security strategies, including those that use advanced techniques such as machine learning.
@@
 Implementing Multi-Layered Security and Compliance Measures: Implementing multi-layered security measures that encompass data protection, access control, and communication security, along with compliance with legal and regulatory requirements, is of utmost importance.
-Solutions range from data encryption, role-based access control, secure container orchestration, adherence to regulations like GDPR (Regulation, 2016).
-This is consistent with current cybersecurity frameworks, such as the NIST cybersecurity framework (of Standards and Technology, 2018), which emphasize a multi-layered approach to security.
-Furthermore, several studies point to the criticality of data security, user access control, and secure communication in cloud environments (Austin, 2018) (Singh et al., 2016).
+Solutions range from data encryption, role-based access control, secure container orchestration, adherence to regulations like GDPR [171].
+This is consistent with current cybersecurity frameworks, such as the NIST cybersecurity framework [172], which emphasize a multi-layered approach to security.
+Furthermore, several studies point to the criticality of data security, user access control, and secure communication in cloud environments [173] [160].
 In a multi-cloud, containerized environment, it is not enough to secure data; access and communication channels must also be secured, and all actions must comply with legal and regulatory requirements.
+In multi-cloud containerized setups, multi-layered security must also accommodate interoperability between CSPs, dynamic access management across domains, and distributed compliance enforcement for container workflows.
 Implications: For practitioners is to design and implement a comprehensive security strategy that encompasses multiple layers of protection.
@@
 Secure and standardized interfaces, such as OpenStack APIs, can be used to enhance security and performance across multiple cloud providers.
-This finding is aligned with existing literature, which emphasizes the role of standardization in achieving security and interoperability in multi-cloud environment (Buyya et al., 2010) (Petcu, 2011).
+This finding is aligned with existing literature, which emphasizes the role of standardization in achieving security and interoperability in multi-cloud environment [174] [175].
 By implementing standardized interfaces and practices, organizations can ensure a level of interoperability and security across cloud environments.
+For containerized applications, this also means ensuring consistent security postures across orchestrators, managing identity federation between providers, and enabling trust in federated environments that dynamically provision containers at runtime.
 Implications: For practitioners, this means choosing platforms and tools that support standardized interfaces and practices, or working to implement those standards within their own environments.
 Researchers, on the other hand, could focus on the development of new standards or the improvement of existing ones to better address security challenges in multi-cloud, containerized environments.
-4.6. Deployment Challenges and Solutions Framework (RQ4 - RQ5)
+4.6 Deployment Challenges and Solutions Framework (RQ4 - RQ5)
 Table LABEL:tab:monitoring-solutions offers a comprehensive overview of challenges and their corresponding solutions encountered during the deployment of containerized applications in multi-cloud environment.
@@
 As Raj et al.
-(Raj et al., 2018) highlighted, embracing automated deployment pipelines paired with standardized toolsets can significantly diminish complexity and mitigate human-induced errors within multi-cloud frameworks.
+[6] highlighted, embracing automated deployment pipelines paired with standardized toolsets can significantly diminish complexity and mitigate human-induced errors within multi-cloud frameworks.
 The challenges, captured under Performance Testing Deployment, Deployment Validation, and Cloud Management Platform Evaluation Deployment in the table, elucidate the nuanced complexities associated with deployments spanning multiple cloud platforms.
@@
 Multi-cloud deployments, inherently intricate, benefit profoundly from standardization, bringing forth predictability and coherence, while automation paves the way for diminishing human-driven inconsistencies and bolstering deployment speeds.
+Multi-cloud deployments, inherently intricate, benefit profoundly from standardization, bringing forth predictability and coherence, while automation paves the way for diminishing human-driven inconsistencies and bolstering deployment speeds.
+In containerized environments, these challenges are amplified due to the need to consistently deploy lightweight, encapsulated workloads across heterogeneous infrastructure with varying APIs and orchestration logic.
 Implications: For researchers, the quest for optimized frameworks championing standardization and automation within multi-cloud environment presents a promising research trajectory.
@@
 Fernandes et al.
-underscore that the labyrinth of diverse platforms and differing standards in multi-cloud configurations amplifies security complexities (Fernandes et al., 2014).
+underscore that the labyrinth of diverse platforms and differing standards in multi-cloud configurations amplifies security complexities [176].
 Challenges articulated as Security Deployment in Multi-container Environment and Malicious Service Management Deployment in Table LABEL:tab:Deployment-solutions accentuate the primacy of security considerations.
@@
 To attain comprehensive security across diverse platforms, strategies must encapsulate both depth of understanding and breadth of application.
+Containerization introduces new layers of isolation and deployment granularity, which—when combined with multi-cloud distribution—creates a unique surface of vulnerability that necessitates tailored controls at both the container and orchestration levels.
 Implications: For the academic community, pioneering avant-garde security solutions apt for multi-cloud environment offers a fertile domain of inquiry.
@@
 Integration, Monitoring, and Infrastructure Management: As highlighted by Ferrer et al.
-(Ferrer et al., 2016) the ability to operate seamlessly across disparate cloud providers and maintain flexible infrastructures is paramount in harnessing the full potential of multi-cloud deployments.
+[177] the ability to operate seamlessly across disparate cloud providers and maintain flexible infrastructures is paramount in harnessing the full potential of multi-cloud deployments.
 The presented challenges, including Network Latency Deployment and Vendor Lock-in Prevention Deployment, underscore the indispensability of agility and cross-platform compatibility.
 Proposed solutions such as VM Selection for Composite Applications and Interoperability Layer Above Cloud Infrastructure accentuate the importance of crafting adaptable infrastructure and formulating strategies to mitigate vendor entrenchment.
+Here, containerization plays a distinct role by enabling application components to be packaged once and executed across diverse cloud platforms, yet this same portability necessitates careful orchestration, consistent networking, and compatibility layers across varied cloud APIs and runtime environments.
 In the context of multi-cloud environment, flexibility and interoperability transcend mere advantageous attributes; they become pivotal determinants shaping adaptability, scalability, and the sustained success of deployment blackprints.
@@
 For industry professionals, anchoring strategies around flexibility and interoperability not only optimizes current deployments but also fortifies them against future technological evolutions and shifting business dynamics.
-4.7. Monitoring Challenges and Solutions (RQ4 - RQ5)
+4.7 Monitoring Challenges and Solutions (RQ4 - RQ5)
 Table LABEL:tab:monitoring-solutions presents a comprehensive view of the Monitoring Challenge-Solution Framework for Containerized Applications in multi-cloud environment.
@@
 Enhancing Performance Stability through Multi-Cloud Monitoring: In the context of multi-cloud environment, ensuring consistent and reliable performance monitoring for containerized applications can be challenging due to the dynamic nature of cloud resources.
-Existing studies (e.g., (Zhong et al., 2022) (Mungoli, 2023)) highlighted the impact of varying resource availability on the performance of containerized applications across multiple clouds.
+Existing studies (e.g., [168] [178]) highlighted the impact of varying resource availability on the performance of containerized applications across multiple clouds.
 They emphasized the need for performance benchmarking to track fluctuations and maintain a consistent user experience.
 The solutions presented in our framework, such as benchmarking performance variation and utilizing Kubernetes for monitoring, align with existing studies suggestion for benchmarking.
+Unlike traditional single-cloud deployments, containerized applications spanning multiple clouds must be monitored for platform-specific latency, orchestration timing mismatches, and fragmented telemetry data pipelines.
+Our proposed framework expands on this by incorporating Kubernetes for improved multi-cloud performance monitoring.
+Kubernetes, while originally designed for single-cluster orchestration, plays a central role in aggregating monitoring data across containerized services deployed in heterogeneous clouds, helping unify performance visibility.
 Our proposed framework expands on this by incorporating Kubernetes for improved multi-cloud performance monitoring.
@@
 This approach can effectively address performance variations and ensure a consistent user experience across diverse cloud environments.
-Implications: For practitioners, adopt benchmarking practices to track performance fluctuations, and leverage Kubernetes for unified multi-cloud performance monitoring.
+Implications: For practitioners, adopt benchmarking practices to track performance fluctuations, and use Kubernetes for unified multi-cloud performance monitoring.
 For researchers, explore further refinements in benchmarking methodologies and investigate the impact of dynamic multi-cloud environment on performance consistency.
@@
 discussed the complexities of managing network communications in multi-cloud environment.
-They emphasized the need for standardized approaches like service mesh to enhance network visibility and control (Sedghpour and Townend, 2022).
-Solutions in our proposed framework, such as using Docker Server for containerized services and Prometheus/Grafana for resource monitoring, align with Sedghpour et al.’s (Sedghpour and Townend, 2022) suggestion for standardized approaches.
+They emphasized the need for standardized approaches like service mesh to enhance network visibility and control [179].
+Solutions in our proposed framework, such as using Docker Server for containerized services and Prometheus/Grafana for resource monitoring, align with Sedghpour et al.’s [179] suggestion for standardized approaches.
+However, container-based deployments compound the difficulty by adding layers of ephemeral, distributed services that must be monitored across volatile container lifecycles.
+This increases the burden of synchronization and observability.
+Additionally, our framework extends these solutions to address broader resource management challenges.
+Containers, especially when orchestrated over multiple cloud platforms, generate high-churn environments that require lightweight, highly adaptive monitoring strategies to ensure timely insights and cost control.
 Additionally, our framework extends these solutions to address broader resource management challenges.
@@
 Strengthening Security, Compliance, and Trust: Ensuring data privacy, continuous security control, and compliance in multi-cloud containerized applications is essential but challenging due to the distributed nature of cloud resources.
-Existing studies (e.g., (Alonso et al., 2023b) (Raj et al., 2018)) discussed the challenges of ensuring security and compliance across multi-cloud environment.
+Existing studies (e.g., [167] [6]) discussed the challenges of ensuring security and compliance across multi-cloud environment.
 They highlighted the importance of continuous monitoring and enforcement mechanisms to address security and compliance gaps.
 Solutions presented in our proposed framework, such as using machine learning for performance supervision and employing security assurance platforms, are aligned with existing studies that emphasis on continuous monitoring and enforcement for security and compliance.
+Containerized deployments introduce distinct security risks, e.g., image poisoning, privilege escalation between co-resident containers, and cross-platform trust boundary misconfigurations that demand container-aware monitoring solutions.
 Implementing advanced techniques like machine learning for monitoring and adopting security assurance platforms can significantly enhance security, compliance, and trust management in multi-cloud containerized applications.
@@
 For researchers, explore the scalability and effectiveness of machine learning-driven monitoring techniques and investigate novel methods for enforcing security and compliance in multi-cloud setups.
-4.8. Tools and Frameworks (RQ6)
+4.8 Tools and Frameworks (RQ6)
 This study report 160 distinct tools and frameworks for container based applications in multi-cloud environment, all of which are systematically classified into 6 categories, further delineated into 46 subcategories.
@@
 Noteworthy cloud service providers highlighted in our study include Amazon EC2, Google Cloud App Engine, and Azure, further underscoring this finding.
-A pivotal insight gleaned from our research underscores the paramount importance of harmonizing cloud service models with specific application prerequisites, thereby facilitating efficient allocation of resources and optimization of costs.
+In multi-cloud environments, selecting services becomes more complex due to differing billing models, feature parity, and integration capabilities across providers—requiring tools that support cross-provider abstraction and orchestration.
 Implications: For practitioners, our outcomes underscore the critical necessity of meticulously evaluating and selecting the most fitting service model based on the unique demands of a given project.
@@
 These technologies receive direct validation through table references, serving as prime exemplars.
+However, our analysis also reveals that deploying Kubernetes in a multi-cloud environment demands adaptations, such as federation control planes, multi-zone DNS, and workload placement strategies tailored to cross-cloud latencies and SLAs.
 Our research yields a compelling revelation: Docker and Kubernetes hold pivotal roles in ensuring consistent application deployment and seamless scalability, effectively simplifying the complexities of multi-cloud environment.
 Implications: Practitioners stand to gain significantly by adopting Docker and Kubernetes, as this potent combination bolsters efficiency and cultivates heightened portability.
-Meanwhile, researchers have a promising avenue to explore deeper facets, including security, performance optimization, and the intricate domains of advanced use cases for Docker and Kubernetes within multi-cloud landscapes.
-This exploration promises valuable insights that can finely calibrate the potential of these cutting-edge technologies.
+Researchers can explore emerging gaps in areas such as secure cross-cloud scheduling, fault tolerance across zones, and policy-compliant orchestration.
 The Rise of DevOps: Embracing Automation Tools: Our exploration spotlights the important role of DevOps practices and tools such as Puppet, Chef, and Terraform within contemporary software development landscapes.
@@
 The table substantiates this notion through references to Puppet, Chef, and Terraform.
-Our study unearths a pivotal insight: DevOps practices wield the ability to augment the software development life cycle by automating tasks, ushering in a realm of efficient and dependable deployment.
+Within a multi-cloud context, these tools must be adapted to handle provider-specific APIs, authentication schemes, and compliance constraints, which introduces configuration drift and policy management complexity..
 Implications: Practitioners are advised to seamlessly integrate DevOps principles into their workflows by harnessing the prowess of these tools.
-Furthermore, a realm of exploration opens for researchers, inviting deeper analyses into the evolutionary trajectory of DevOps practices and their profound impact on elevating software quality and hastening development velocities.
-5. Threats to Validity
+Researchers may investigate how DevOps pipelines can be extended to support containerized workloads that span multiple clouds, ensuring traceability, consistency, and auditability across heterogeneous execution environments.
+5 Threats to Validity
 This systematic mapping study is susceptible to various threats that could influence its outcomes.
-To address these potential threats, we adhered to the established guidelines for conducting SMSs and SLRs as outlined in (Keele et al., 2007) and (Petersen et al., 2008b).
-We further categorized and analyzed the validity threats to our study based on the four distinct types of validity threats mentioned in (Zhou et al., 2016) and (Wohlin et al., 2012).
+To address these potential threats, we adhered to the established guidelines for conducting SMSs and SLRs as outlined in [17] and [180].
+We further categorized and analyzed the validity threats to our study based on the four distinct types of validity threats mentioned in [181] and [182].
 In the subsequent sections, we delve into the specific validity threats that pertain to the different phases of this SMS.
-5.1. Internal Validity
+5.1 Internal Validity
 Internal validity refers to the factors that could affect the analysis of the data extracted from the selected studies.
@@
 During this procedure, the lead two authors conducted the study screening based on the criteria explicitly detailed in Section 3.
-To assess the objectivity of the screening process and the level of agreement between the two authors, we applied Cohen’s Kappa (Cohen, 1968).
+To assess the objectivity of the screening process and the level of agreement between the two authors, we applied Cohen’s Kappa [22].
 In cases where the lead authors did not agree, the second and third authors independently reviewed the disputed studies to reach a consensus.
@@
 If any uncertainties arose regarding the extracted data, comprehensive discussions were convened among all authors.
-Following the recommendations in (Wohlin et al., 2012), a subset of the extracted data was cross-verified by the second and third authors.
+Following the recommendations in [182], a subset of the extracted data was cross-verified by the second and third authors.
   • Bias on Themes Classification: Misclassification of data and primary studies may result in biases stemming from subjective interpretation.
 To minimize this risk, we adhered to the thematic analysis guidelines established by Braun et al.
-(Braun and Clarke, 2006), and implemented a six-step process for thematic classification, as detailed in Section 2.4.
+[26], and implemented a six-step process for thematic classification, as detailed in Section 2.4.
   • Data Synthesis: We employed both qualitative and quantitative approaches to assess the gathered data.
 Potential biases in data synthesis could influence the interpretation of our findings.
-To address this concern, we used open coding and constant comparison techniques from Grounded Theory (Glaser and Strauss, 1967) to analyze the qualitative data extracted from the selected studies.
-5.2. External Validity
+To address this concern, we used open coding and constant comparison techniques from Grounded Theory [27] to analyze the qualitative data extracted from the selected studies.
+  • Transparency of Quality Assessment Process: While our replication package [11] provides a detailed breakdown of quality assessment scores for each study and each criterion, it does not include per-study justifications or annotations explaining how individual scores were assigned.
+This may present a limitation in terms of external reproducibility, as other researchers may not fully reconstruct the decision-making rationale behind each score.
+Although we followed a consistent and internally validated scoring logic (e.g., peer-reviewed venues scored higher, empirical data presence was required for full marks), we recognize that the absence of a fully documented scoring rubric or narrative rationale could reduce transparency.
+However, to partially mitigate this, we have explicitly described our scoring criteria and interpretation rules in Section 2.3.3.
+This clarification improves traceability by explaining the rationale used across all assessed studies.
+The limitation remains inherent to balancing the breadth and depth of quality assessment in a large-scale review and is acknowledged as a tradeoff in terms of replication of this study.
+5.2 External Validity
 Concerns about external validity relate to the applicability and generalizability of study findings.
@@
 To complement and enrich our findings, future work could include an industrial study (e.g., blogs, white papers) that would provide a broader view of the application of containerization in practice
-5.3. Construct Validity
+5.3 Construct Validity
 Construct validity concerns to the accuracy of the operational measures used for data collection in a study.
@@
 Additionally, we customized search string according to each database syntax.
-5.4. Conclusion Validity
+5.4 Conclusion Validity
 Threats to conclusion validity relate to factors, such as data inaccuracies, that can impede drawing accurate conclusions.
 To mitigate these threats, we adhered to best practices, including the search protocol, pilot search, and pilot data selection, as recommended by Kitchenham et al.
-(Keele et al., 2007) and Petersen et al.
-(Petersen et al., 2008b).
+[17] and Petersen et al.
+[180].
 Moreover, to further ensure the validity of our conclusions, the authors engaged in multiple brainstorming sessions to collaboratively interpret the results and finalize the conclusions.
-6. Related Work
+6 Related Work
 This section reviews the most relevant existing research in terms of secondary studies such as literature surveys, state-of-the-art analysis, systematic reviews, and mapping studies that consolidate published literature on containerization in multi-cloud environment.
 The review focuses on (i) existing challenges, proposed solutions, and emerging trends detailed in Section 6.1 and (ii) container-based deployment of multi-cloud systems in Section 6.2.
-6.1. Challenges, Solutions, and Performance of Multi-cloud Containers
-In recent years, several literature surveys have been published to analyze state-of-the-art on the applications (Pahl et al., 2017), tools and technologies (Casalicchio and Iannucci, 2020), potential and limitations (Bentaleb et al., 2022), as well as emerging trends (Watada et al., 2019) of container-based solutions for cloud computing systems, discussed below.
-6.1.1. Challenges, Applications, Tools, and Emerging Trends
-Containers in multi-cloud environment provide a standardized and portable way to package, deploy, and run applications, ensuring seamless deployment and management across diverse cloud platforms (Alonso et al., 2023c) (Saxena et al., 2021).
+6.1 Challenges, Solutions, and Performance of Multi-cloud Containers
+In recent years, several literature surveys have been published to analyze state-of-the-art on the applications [183], tools and technologies [184], potential and limitations [185], as well as emerging trends [186] of container-based solutions for cloud computing systems, discussed below.
+6.1.1 Challenges, Applications, Tools, and Emerging Trends
+Containers in multi-cloud environment provide a standardized and portable way to package, deploy, and run applications, ensuring seamless deployment and management across diverse cloud platforms [187] [188].
 One of the earliest SLR-based studies by Pahl et al.
-(Pahl et al., 2017) on containerized clouds reviewed a total of 46 studies to identify, taxonomically classify and systematically compare the existing research on containers and their application in the context of cloud-based systems.
+[183] on containerized clouds reviewed a total of 46 studies to identify, taxonomically classify and systematically compare the existing research on containers and their application in the context of cloud-based systems.
 The SLR classified and compared the selected research studies using a conceptual framework to highlight existing trends and needs for future research.
 SLR results indicate container orchestration, microservice delivery, continuous development and deployment as emerging trends of research on containerized clouds.
-The SLR in (Alonso et al., 2023c) reviewed a total of 88 studies (published from 2011 - 2021) on multi-cloud systems.
-Similar survey studies such as (Saxena et al., 2021) concepts, challenges, requirements and future directions for multi-cloud environment are discussed.
-A survey of existing approaches and solutions provided by different multi-cloud architectures is entailed along with analysis of the pros and cons of different architectures while comparing the same (Imran et al., 2020).
-6.1.2. Performance and Resource Utilization
+The SLR in [187] reviewed a total of 88 studies (published from 2011 - 2021) on multi-cloud systems.
+Similar survey studies such as [188] concepts, challenges, requirements and future directions for multi-cloud environment are discussed.
+A survey of existing approaches and solutions provided by different multi-cloud architectures is entailed along with analysis of the pros and cons of different architectures while comparing the same [189].
+6.1.2 Performance and Resource Utilization
 In a study by Bentaleb et al.
-(Bentaleb et al., 2022), the authors follow the taxonomical classification from (Pahl et al., 2017) to categorize container-based technologies for cloud systems.
+[185], the authors follow the taxonomical classification from [183] to categorize container-based technologies for cloud systems.
 The study argues for the need of performance metrics to objectively define and evaluate quality attributes such as resource virtualization, service elasticity, orchestration, and multi-tenancy in containerized cloud systems.
 The study also highlights the needs for future research in terms of best practices, i.e., patterns and tactics (enabling reuse) and tools (supporting automation) for container-based cloud deployment.
-In the context of performance, the studies (Casalicchio and Iannucci, 2020) (Watada et al., 2019) report literature review and experimental analysis of factors that influence performance of container-based cloud systems.
+In the context of performance, the studies [184] [186] report literature review and experimental analysis of factors that influence performance of container-based cloud systems.
 Specifically, Casalicchio et al.
-(Casalicchio and Iannucci, 2020) reviewed a total of 97 research studies investigating performance evaluation and run-time adaptation of container-based solution.
+[184] reviewed a total of 97 research studies investigating performance evaluation and run-time adaptation of container-based solution.
 The study highlighted several unsolved challenges such as I/O throughput optimization, performance prediction, and multi-layer monitoring performance bottlenecks.
 Moreover, Watada et al.
-(Watada et al., 2019) conducted an experimental study to compare the performance of VMs, containers and uni kernels in terms of and technological maturity using standard benchmarks and observed containers to optimise performance of container-based.
+[186] conducted an experimental study to compare the performance of VMs, containers and uni kernels in terms of and technological maturity using standard benchmarks and observed containers to optimise performance of container-based.
 The performance of containerized clouds is also determined by resource utilization, which includes but is not limited to CPU utilization, memory footprints, energy consumption, network bandwidth, and execution time.
 Kapil et al.
-(Netaji and Bhole, 2021) reviewed 64 research papers to gain insight into resource allocation, management, and scheduling.
+[190] reviewed 64 research papers to gain insight into resource allocation, management, and scheduling.
 Furthermore, limitations of existing resource allocation algorithms are discussed, indicating the need to investigate algorithms or techniques of performance optimization of containers for cloud systems.
 The study by Maenhaut et al.
-(Maenhaut et al., 2020) provided an overview of the current state of the art regarding resource management within the broad sense of cloud computing, complementary to existing surveys in the literature.
+[191] provided an overview of the current state of the art regarding resource management within the broad sense of cloud computing, complementary to existing surveys in the literature.
 The study investigated how research is adapting to the recent evolution within cloud solutions, including container technologies.
-6.2. Container-based Cloud Orchestration, Deployment, and Security
+6.2 Container-based Cloud Orchestration, Deployment, and Security
 This section reviews the most relevant research on the container-based deployment with a focus on orchestration and security issues of cloud-based systems, detailed below.
-6.2.1. Orchestration and Deployment
+6.2.1 Orchestration and Deployment
 Naweiluo et al.
-(Zhou et al., 2022) explored containerization in High Performance Computing (HPC) environments, contrasting it with cloud computing.
+[192] explored containerization in High Performance Computing (HPC) environments, contrasting it with cloud computing.
 The results of this study indicate that containers enhance application deployment efficiency, but face challenges in HPC due to high security levels and the need for extensive libraries and packages that affect cloud portability, often resulting in vendor lock-ins.
 The study also provides a survey and taxonomy on containerization and orchestration efforts in HPC, pointing out differences with cloud environments and identifying future research and engineering potentials.
-A systematic mapping study by Naylor et al.(Bachiega et al., 2018) explores virtualization in container-based cloud deployments.
+A systematic mapping study by Naylor et al.[193] explores virtualization in container-based cloud deployments.
 The mapping study, based on a comprehensive review of major databases, identifies a significant research gap in the performance evaluation of containers, underscoring the need for further investigation on cloud deployment.
 Carlos et al.
-(Awada, 2018) present an optimization strategy for deploying microservices in multi-cloud environment focusing on minimizing cloud service costs, network latency, and startup time for new microservices.
+[194] present an optimization strategy for deploying microservices in multi-cloud environment focusing on minimizing cloud service costs, network latency, and startup time for new microservices.
 Their approach uses a Non-dominated Sorting Genetic Algorithm II (NSGA-II) compared to a Greedy First-Fit algorithm, demonstrating a 300% improvement with NSGA-II.
@@
 Emiliano et al.
-(Paladi et al., 2018) survey container orchestration, proposing a reference architecture for autonomic orchestrators, and identifying research challenges in the field.
+[195] survey container orchestration, proposing a reference architecture for autonomic orchestrators, and identifying research challenges in the field.
 Their work emphasizes the importance of container technologies in cloud environments and the need for advanced orchestration solutions to manage complex multi-container applications effectively.
-Uchechukwu Awada (Guerrero et al., 2018b) reviews container orchestration tools and platforms, comparing the architectures, components, and capabilities of several Container Service Platforms (CSPs) and orchestration tools like Amazon ECS, Kubernetes, Docker Swarm, and Mesos.
+Uchechukwu Awada [196] reviews container orchestration tools and platforms, comparing the architectures, components, and capabilities of several Container Service Platforms (CSPs) and orchestration tools like Amazon ECS, Kubernetes, Docker Swarm, and Mesos.
 The study offers insights into the current state of container orchestration and suggests future research directions, serving as a guide for developers and organizations.
 Koustabh et al.
-(Dolui and Kiraly, 2018) explore multi-container deployment on IoT gateways to meet the stringent latency requirements of advanced IoT applications.
+[197] explore multi-container deployment on IoT gateways to meet the stringent latency requirements of advanced IoT applications.
 Through their study within the AGILE project, they highlight containerization’s role in overcoming the diversity and resource constraints of IoT gateways, showcasing containerization’s advantages for IoT gateway performance optimization.
@@
 Matteo et al.
-(Nardelli et al., 2018) introduce the Adaptive Container Deployment (ACD) model to optimize containerized application deployment in geo-distributed environments, focusing on IoT and fog computing resources.
+[198] introduce the Adaptive Container Deployment (ACD) model to optimize containerized application deployment in geo-distributed environments, focusing on IoT and fog computing resources.
 ACD, formulated as an Integer Linear Programming problem, aims to improve application performance by leveraging containers’ horizontal and vertical elasticity.
 The study evaluates ACD’s effectiveness against greedy heuristics, highlighting the need for advanced orchestration solutions to exploit emerging computing environments’ characteristics efficiently.
-6.2.2. Security of Containerized Clouds
+6.2.2 Security of Containerized Clouds
 Nicolae et al.
-(Casalicchio, 2019) analyze security challenges in cloud orchestration for multi-cloud deployments, proposing a security-enabled orchestration framework.
+[199] analyze security challenges in cloud orchestration for multi-cloud deployments, proposing a security-enabled orchestration framework.
 Their research identifies potential attack scenarios and security enforcement mechanisms, aiming to enhance security guarantees for cloud operators in a multi-cloud setting.
 Mohammad et al.
-(Bokhari et al., 2018) offer a detailed analysis of cloud computing, defining its essential characteristics, architecture, service models (SaaS, PaaS, IaaS), and deployment models.
+[200] offer a detailed analysis of cloud computing, defining its essential characteristics, architecture, service models (SaaS, PaaS, IaaS), and deployment models.
 They discuss the security requirements for public and private clouds across different service models, aiming to provide researchers with a comprehensive understanding of cloud computing’s potential and security challenges.
 Sari et al.
-(Sultan et al., 2019) survey the landscape of container security, addressing challenges and solutions across four generalized use cases within the host-container threat landscape.
+[201] survey the landscape of container security, addressing challenges and solutions across four generalized use cases within the host-container threat landscape.
 Their analysis covers both software-based solutions utilizing Linux kernel features and hardware-based solutions for enhanced security.
@@
 Hendrik et al.
-(Graupner et al., 2015b) tackle privacy, security, and trust issues in cloud computing through a multi-cloud architecture perspective.
+[202] tackle privacy, security, and trust issues in cloud computing through a multi-cloud architecture perspective.
 They propose a novel technique for enhancing security and unifying access control mechanisms across cloud providers, addressing the challenges of inadequate cross-provider APIs and non-unified access control.
 Rohan et al.
-(Gupta et al., 2016) propose a security-enhancing methodology for cloud data storage using container clustering and Docker instances for on-demand encryption.
+[203] propose a security-enhancing methodology for cloud data storage using container clustering and Docker instances for on-demand encryption.
 This approach aims to secure data while optimizing resource usage and cost, highlighting the benefits of container technology in improving cloud computing security and efficiency.
-6.3. Conclusive Summary
-Current survey-based studies and systematic reviews, discussed above (e.g., (Casalicchio and Iannucci, 2020) (Pahl et al., 2017)), aim to consolidate the latest findings in published research concerning container-based solutions for cloud computing systems.
+6.3 Conclusive Summary
+Current survey-based studies and systematic reviews, discussed above (e.g., [184] [183]), aim to consolidate the latest findings in published research concerning container-based solutions for cloud computing systems.
 These studies investigated various aspects of multi-cloud, including exploring applications, tools, and technologies, assessing potential and limitations, and identifying emerging trends.
 Table 16 presents an overview of the differences between this study and existing secondary studies.
-The symbols used in the table are as follows:  indicates “Yes”,  indicates “No”, and $\triangle$ indicates “Partially”.
-After thoroughly comparing our results with existing studies (e.g., (Pahl et al., 2017), (Casalicchio and Iannucci, 2020), (Netaji and Bhole, 2021)), it is clear that none of the results are fully identical with our findings.
-Our study distinguishes itself from prior research in several ways.
+The symbols used in the table are as follows: ✗ indicates “Yes”, ✓ indicates “No”, and $\triangle$ indicates “Partially”.
+This study distinguishes itself from prior secondary studies in multiple critical ways, thereby advancing the state of the art in the domain of containerization for multi-cloud environments.
+While existing surveys and mapping studies (e.g., [183], [184], [190]) provide valuable overviews of tools, technologies, performance, and resource management, they often focus on isolated aspects or lack integration across key quality and deployment dimensions.
+In contrast, our SMS presents a comprehensive, theme-based classification of 121 selected studies—making it the most extensive secondary analysis in this domain to date.
 First, our work systematically categorizes container utilization in multi-cloud environment into coherent categories and themes.
@@
 We identified and classified seventy-four patterns and strategies for containerization, which existing studies have not provided.
+This level of operational detail is absent in prior work.
 Furthermore, we explored the QAs and associated tactics for containerization in multi-cloud environment.
-This contribution offers practitioners actionable insights into tailoring their containerization strategies to meet desired quality standards.
-Lastly, our study introduces four distinct frameworks consisting of challenges and solutions related to security, automation, deployment, and monitoring within the context of containerization in multi-cloud environment.
-Table 16.
-Comparison of This SLR Results with Existing Secondary Studies.
-MS indicate “Mix Study”, LR indicate “Literature Review”  indicates “Yes”,  indicates “No,” and $\triangle$ indicates “Partially.”
-Contributions This Study (Pahl et al., 2017) (Casalicchio and Iannucci, 2020) (Netaji and Bhole, 2021) (Bentaleb et al., 2022) (Watada et al., 2019) (Saxena et al., 2021) (Imran et al., 2020) (Casalicchio, 2019)
+Unlike existing reviews that superficially mention quality attributes, we provide a dedicated framework linking QAs with implementation tactics, offering practical insights for engineering design decisions—an aspect previously unexplored.
+Our SMS also identifies and formalizes four distinct challenge-solution frameworks in the areas of security, automation, deployment, and monitoring.
+These frameworks synthesize fragmented research findings into cohesive, actionable knowledge structures, not previously consolidated in the literature.
+Lastly, our comparative analysis (Table 16) demonstrates that none of the previous studies collectively cover the breadth and depth addressed by our SMS, particularly in integrating performance, security, orchestration, and QA strategies across the entire lifecycle of container-based multi-cloud systems.
+By addressing these multiple dimensions holistically, our study not only fills significant gaps in the literature but also provides a practical roadmap for researchers and practitioners aiming to optimize container deployments in complex, heterogeneous cloud environments.
+Table 16: Comparison of this SLR results with existing secondary studies.
+MS indicates “Mix Study”, LR indicates “Literature Review”, ✓ indicates “Yes”, ✗ indicates “No”, and $\triangle$ indicates “Partially”
+Contributions This Study [183] [184] [190] [185] [186] [188] [189] [199]
 Number of selected studies 121 46 97 64 40 15 11
@@
 Tools and Frameworks (Section 3.9) $\triangle$ $\triangle$ $\triangle$
-7. Conclusions
+7 Conclusions
 This SMS presents the current state of research regarding containerization in multi-cloud environment, focusing on aims of selected studies, the roles of containers, containers implementation strategies, architectural patterns and strategies, quality attributes and corresponding tactics.
 Additionally, this SMS also explores deployment, monitoring, security challenges, accompanied by their respective solutions, tools, and frameworks used to implement containerized applications in multi-cloud environment.
-From an initial pool of extensive search results, 86 relevant studies were finally chosen to answer the research questions presented in Table 1.
-The key findings of this SMS are listed below.
+We investigated 121 relevant studies to answer the RQs in Table 1 with key findings of this SMS include:
   • The findings of this SMS reveal various insights regarding SMS demographics, publishers, publication types, authors’ affiliations, and research themes.
@@
   • This SMS presents a comprehensive security challenge-solution framework for containerized applications in multi-cloud environment.
-This framework offers valuable insights and actionable strategies for practitioners and researchers to enhance the security of containerized applications, address complex multi-cloud security concerns, and ensure compliance with various regulatory requirements, thereby enabling the development of more robust and secure containerized applications in multi-cloud environment.
+This framework offers valuable insights and actionable strategies for practitioners and researchers to enhance the security of containerized applications, address complex multi-cloud security concerns, and ensure compliance with various regulatory requirements, thereby enabling the development of robust and secure containerized applications for multi-clouds.
   • This SMS presents a catalogs for the automation challenge-solution catalogs for containerized applications in multi-cloud environment.
-The automation challenge-solution catalog offers a comprehensive guide for practitioners and researchers to address the complex challenges associated with automating various aspects of deploying, managing, scaling, testing, migrating, and ensuring interoperability of containerized applications in multi-cloud environment.
-By offering specific solutions for each challenge, this catalog aids in the development of more efficient, secure, and adaptable automation processes, empowering practitioners to streamline their operations and enabling researchers to advance the field of multi-cloud containerization.
+The automation challenge-solution catalog offers a comprehensive guide for practitioners and researchers to address the complex challenges associated with automating various aspects of deploying, managing, scaling, testing, migrating, and ensuring interoperability of containerized applications in multi-cloud environments.
   • The third catalog presented in this SMS is the deployment challenge-solution catalog for containerized applications in multi-cloud environment.
@@
 Data availability
-The dataset of this study has been provided online: https://zenodo.org/record/10732611.
+Link to our dataset is in the reference [11].
 Acknowledgments
-This research is funded by Business Finland through QLEAP (2022-24) project, and by the NSFC China under Grant No. 62172311.
-8. Appendices
-{xltabular}
-—c—p4.5cm—p4.5cm—c—c— List of selected studies for containerization in multi-cloud environments
-Study ID Title Venue Year Reference
+This research is funded by Business Finland through QLEAP (2022-24) project, the National Natural Science Foundation of China (NSFC) under Grant No.
+62172311, and the Major Science and Technology Project of Hubei Province under Grant No.
+2024BAA008.
+Declaration of AI Assistance
 […]
-Study ID Title Venue Year Reference
-[…]
-Continued on next page
-[…]
-[…]
-S01 A component-based adaptation approach for multi-cloud applications IEEE Conference on Computer Communications (INFOCOM) Workshops 2014 (Almeida et al., 2014)
-S02 A Cost-Efficient Multi-cloud Orchestrator for Benchmarking Containerized Web-Applications International Conference on Web Information Systems Engineering (WISE) 2019 (Jha et al., 2019)
-S03 A framework to support multi-cloud collaboration IEEE World Congress on Services (SERVICES) 2020 (Hua et al., 2020)
-S04 A Machine Learning Model for Detection of Docker-based APP Overbooking on Kubernetes International Conference on Communications (ICC) 2021 (Ramos et al., 2021)
-S05 A multi-cloud service mesh approach applied to Internet of Things Annual Conference of Industrial Electronics Society (IECON) 2022 (Gattobigio et al., 2022)
-S06 A Secure Container Placement Strategy Using Deep Reinforcement Learning in Cloud International Conference on Computer Supported Cooperative Work in Design (CSCWD) 2022 (Deng et al., 2022)
-S07 A Service Driver Based Application Execution and Development Method in Multi-cloud Context International Conference on Data Science and Computer Application (ICDSCA) 2021 (Zhang et al., 2021)
-S08 An adaptable framework to deploy complex applications onto multi-cloud platforms International Conference on Computing Communication Technologies - Research, Innovation, and Vision for Future (RIVF) 2015 (Pham et al., 2015)
-S09 An agile container-based approach to TaaS International Congress Future Telecommunications: Infrastructure and Sustainability (FITCE) 2016 (Verdugo et al., 2017)
-S10 An Efficient and Reliable Multi-Cloud Provider Monitoring Solution Conference on Global Communications (GLOBECOM) 2022 (Sabbioni et al., 2020)
-S11 An Efficient Multi-Keyword Search Scheme over Encrypted Data in Multi-Cloud Environment International Conference on Smart Cloud (SmartCloud) 2015 (He et al., 2022)
-S12 An Empirical Investigation of a Fault Tolerant Containerized Application Deployment International Conference on Informatics (ICI) 2015 (Bisht and Kaur, 2022)
-S13 Applying Computational Intelligence for enhancing the dependability of multi-cloud systems using Docker Swarm Symposium Series on Computational Intelligence (SSCI) 2017 (Naik, 2016a)
-S14 Automatically Design Heuristics for Multi-Objective Location-Aware Service Brokering in Multi-Cloud International Conference on Services Computing (SCC) 2020 (Chen et al., 2022)
-S15 Autonomic fine-grained migration and replication of component-based applications across multi-clouds National Foundation for Science and Technology Development Conference on Information and Computer Science (NICS) 2022 (Pham and Pham, 2015)
-S16 Band-Area Resource Management Platform and Accelerated Particle Swarm Optimization Alg. for Container Deployment IEEE Access 2016 (Ouyang et al., 2022)
-S17 Building a Multi-tenant Cloud Service from Legacy Code with Docker Containers International Conference on Cloud Engineering (IC2E) 2022 (Slominski et al., 2015)
-S18 Building cloud native application - Analysis for multi-component application deployment International Conference on Computer Communication and Informatics (ICCCI) 2021 (Pai and Kumar, 2021)
-S19 CLAMS: Cross-layer Multi-cloud Application Monitoring-as-a-Service Framework International Conference on Services Computing (SCC) 2014 (Alhamazani et al., 2014)
-S20 Cloud MF: Applying MDE to Tame the Complexity of Managing Multi-Cloud Applications International Conference on Utility and Cloud Computing (UCC) 2014 (Ferry et al., 2014)
-S21 Cluster as a Service: A Container Based Cluster Sharing Approach with Multi-user Support IEEE Symposium on Service-Oriented System Engineering (SOSE) 2016 (Cui et al., 2016)
-S22 Container Deployment Strategy for Edge Networking Workshop on Middleware for Edge Clouds and Cloudlets (MECC) 2019 (Wong et al., 2019)
-S23 Container-Based Performance Isolation for Multi-Tenant SaaS Applications in Micro-Service Architecture Journal of Physics: Conference Series 2020 (Wang et al., 2020)
-S24 Containerization and the PaaS Cloud IEEE Cloud Computing 2015 (Pahl, 2015)
-S25 ContainerLeaks: Emerging Security Threats of Information Leakages in Container Clouds International Conference on Dependable Systems and Networks (DSN) 2017 (Gao et al., 2017)
-S26 ContMEC: An Architecture of Multi-access Edge Computing for Offloading Container-Based Mobile Applications IEEE International Conference on Communications (ICC) 2022 (Watanabe et al., 2022)
-S27 DECIDE: An Extended DevOps Framework for Multi-Cloud Applications International Conference on Cloud and Big Data Computing (ICCBDC) 2022 (Alonso et al., 2019)
-S28 Delivering Elastic Containerized Cloud Applications to Enable DevOps International Workshop on Software Engineering for Adaptive and Self-Managing Systems (SEAMS) 2017 (Barna et al., 2017)
-S29 Deployment of Multi-Tier Application on Cloud and Continuous Monitoring using Kubernetes International Conference on Electrical, Electronics, Communication, Computer Technologies and Optimization Techniques (ICEECCOT) 2021 (Gupta et al., 2021)
-S30 Deployment-Time Multi-Cloud Application Security International Conference on Smart Computing (SMARTCOMP) 2017 (Sheridan et al., 2017)
-S31 DEVSML 3.0 Stack: Rapid deployment of DEVS Farm in distributed cloud environment using microservices and containers Symposium on Theory of Modeling and Simulation (TMS/DEVS) 2017 (Mittal and Risco-Martín, 2017)
-S32 Disaster Recovery and Business Continuity for Database Services in Multi-Cloud International Conference on Computer Applications & Information Security (ICCAIS) 2018 (Al-shammari and Alwan, 2018)
-S33 Dynamic On-Demand Fog Formation Offering On-the-Fly IoT Service Deployment IEEE Transactions on Network and Service Management 2020 (Sami and Mourad, 2020)
-S34 Dynamic security assurance in multi-cloud DevOps IEEE Conference on Communications and Network Security (CNS) 2017 (Rios et al., 2017)
-S35 Engineering and Experimentally Benchmarking a Container-based Edge Computing System International Conference on Communications (ICC) 2020 (Carpio et al., 2020)
-S36 FogDocker: Start Container Now, Fetch Image Later International Conference on Utility and Cloud Computing (UCC) 2019 (Lorenzo et al., 2019)
-S37 Importance of Application-Level Resource Management in Multi-Cloud Deployments International Conference on Cloud Engineering (IC2E) 2019 (Dimitrijevic et al., 2019)
-S38 Knowledge-Engineered Multi-Cloud Resource Brokering for Application Workflow Optimization IEEE Transactions on Network and Service Management 2022 (Pandey et al., 2022)
-S39 Label-affinity-Scheduler: Considering Business Requirements in Container Scheduling for Multi-Cloud and Multi-Tenant Brazilian Symposium on Computing System Engineering (SBESC) 2022 (Altran et al., 2022)
-S40 Leveraging kernel security mechanisms to improve container security: A survey International Conference on Availability, Reliability and Security (ARES) 2019 (Bélair et al., 2019)
-S41 Location-Aware and Budget-Constrained Service Deployment for Composite Applications in Multi-Cloud Environment IEEE Transactions on Parallel and Distributed Systems 2020 (Shi et al., 2020)
-S42 MicroCloud: A Container-Based Solution for Efficient Resource Management in the Cloud International Conference on Smart Cloud (SmartCloud) 2016 (Baresi et al., 2016)
-S43 Mobile Edge as Part of the Multi-Cloud Ecosystem: A Performance Study Euromicro International Conference on Parallel, Distributed and Network-Based Processing (PDP) 2019 (Dreibholz et al., 2019)
-S44 Modeling and Automated Deployment of Serverless Applications Using TOSCA Conference on Service-Oriented Computing and Applications (SOCA) 2018 (Wurster et al., 2018)
-S45 Monitoring Elastically Adaptive Multi-Cloud Services IEEE Transactions on Cloud Computing 2018 (Trihinas et al., 2015)
-S46 Monitoring-Aware Optimal Deployment for Applications Based on Microservices IEEE Transactions on Services Computing 2021 (Fadda et al., 2019)
-S47 MSSF: A Step towards User-Friendly Multi-cloud Data Dispersal International Conference on Cloud Computing (CLOUD) 2014 (Libardi et al., 2014)
-S48 Multi-agent QoS-aware autonomic resource provisioning framework for elastic BPM in containerized multi-cloud environment Journal of Ambient Intelligence and Humanized Computing 2022 (Saif et al., 2023)
-S49 Multicloud- auto scale with prediction and delta correction algorithm International Conference on Intelligent Computing, Instrumentation and Control Technologies (ICICICT) 2019 (Harwalkar et al., 2019b)
-50 Multi-cloud Distribution of Virtual Functions and Dynamic Service Deployment: Open ADN Perspective IEEE International Conference on Cloud Engineering (IC2E) 2015 (Bhamare et al., 2015)
-S51 Multi-cloud resource provisioning with Aneka: A unified and integrated utilization of Microsoft Azure and Amazon EC2International Conference on Computing and Network Communications (CoCoNet) 2015 (Buyya and Barreto, 2015)
-S52 MUSA Deployer: Deployment of Multi-cloud Applications IEEE International Workshop on Enabling Technologies: Infrastructure for Collaborative Enterprises (WETICE) 2017 (Casola et al., 2017)
-S53 NoMISHAP: A Novel Middleware Support for High Availability in Multicloud PaaS IEEE Cloud Computing 2017 (Acquaviva et al., 2017)
-S54 Object storage in the cloud and multi-cloud: State of the art and the research challengesIEEE International Conference on Automation, Quality and Testing, Robotics (AQTR) 2018 (Bucur et al., 2018)
-S55 On the Deployment of an Open-Source, 5G-Aware Evaluation Testbed International Conference on Mobile Cloud Computing, Services, and Engineering (MobileCloud) 2018 (Bolivar et al., 2018)
-S56 OpenStack Federation in Experimentation Multi-cloud Testbeds IEEE International Conference on Cloud Computing Technology and Science (CloudCom) 2013 (del Castillo et al., 2013)
-S57 Orchestrated Multi-Cloud Application Deployment in OpenStack with TOSCA International Conference on Smart Computing (SMARTCOMP) 2017 (Tricomi et al., 2017)
-S58 Orchestration Based Hybrid or Multi Clouds and Interoperability Standardization IEEE International Conference on Cloud Computing in Emerging Markets (CCEM) 2018 (Sitaram et al., 2018)
-S59 Performance Analysis of Zero-Trust multi-cloud IEEE International Conference on Cloud Computing (CLOUD) 2021 (Rodigari et al., 2021)
-S60 Policies Based Container Migration Using Cross-Cloud Management Platform International Conference on Information and Automation for Sustainability (ICIAfS) 2018 (Janarthanan et al., 2018)
-S61 Policy-based deployment in a hybrid and multi-cloud environment International Conference on Cloud Computing and Services Science (CLOSER) 2019 (Di Modica et al., 2019)
-S62 Predicting the End-to-End Tail Latency of Containerized Microservices in the Cloud International Conference on Cloud Engineering (IC2E) 2019 (Rahman and Lama, 2019)
-S63 Preventing vendor lock-ins via an interoperable multi-cloud deployment approach International Conference for Internet Technology and Secured Transactions (ICITST) 2017 (Pellegrini et al., 2017)
-S64 Private STaaS with OpenStack Cinder Volumes for Hybrid/Multi-cloud International Conference on Advances in Computing and Communication Engineering (ICACCE) 2019 (Harwalkar et al., 2019a)
-S65 Real-World Experience with a Multicloud Exchange IEEE Cloud Computing 2017 (Zeck and Bouroudjian, 2017)
-S66 Reliability Management for Blockchain-Based Decentralized Multi-Cloud International Symposium on Cluster Computing and the Grid (CCGRID) 2020 (Aral et al., 2020)
-S67 Reliability-Aware Cost-Efficient Scientific Workflows Scheduling Strategy on Multi-Cloud Systems IEEE Transactions on Cloud Computing 2022 (Tang, 2021)
-S68 Resource optimization of container orchestration: a case study in multi-cloud microservices-based applicationsThe Journal of Supercomputing 2018 (Guerrero et al., 2018a)
-S69 Scalable multi-cloud platform to support industry and scientific applications International Convention on Information and Communication Technology, Electronics and Microelectronics (MIPRO) 2018 (Kiss, 2018)
-S70 Search-based Methods for Multi-Cloud Configuration International Conference on Cloud Computing (CLOUD) 2022 (Lazuka et al., 2022)
-S71 Secure access control for multi-cloud resources IEEE Conference on ALocal Computer Networks Workshops (LCN Workshops) 2015 (Graupner et al., 2015a)
-S72 Secure and Dependable Multi-Cloud Network Virtualization International Workshop on Security and Dependability of Multi-Domain Infrastructures 2017 (Alaluna et al., 2017)
-S73 Secure and Parallel Expressive Search over Encrypted Data with Access Control in Multi-CloudIoT Cloudification of the Internet of Things (CIoT) 2018 (Guechi and Maamri, [n. d.])
-S74 Self-healing cloud services in private multi-clouds International Conference on High Performance Computing & Simulation (HPCS) 2018 (Mfula and Nurminen, 2018)
-S75 SpyStorage: A Highly Reliable Multi-Cloud Storage with Secure and Anonymous Data Sharing International Conference on Networking, Architecture, and Storage (NAS) 2017 (Shen et al., 2017)
-S76 Support Services for Applications Execution in Multi-clouds Environments International Conference on Autonomic Computing (ICAC) 2016 (Pop et al., 2016)
-S77 SynAPTIC: Secure and persistent connectivity for containers International Symposium on Cluster, Cloud and Grid Computing (CCGRID) 2017 (Ranjbar et al., 2017)
-S78 System Restore in a Multi-cloud Data Pipeline Platform International Conference on Dependable Systems and Networks – Industry Track 2019 (Wang et al., 2019)
-S79 Systematic and Recomputable Comparison of Multi-cloud Management Platforms International Conference on Cloud Computing Technology and Science (CloudCom) 2018 (Serhiienko and Spillner, 2018)
-S80 Towards a lightweight multi-cloud DSL for elastic and transferable cloud-native applications International Conference on Cloud Computing and Services Science (CLOSER) 2018 (Quint and Kratzke, 2018)
-S81 Towards Formal-Based Semantic Interoperability in Multi-Clouds: The FCLOUDS Framework IEEE International Conference on Cloud Computing (CLOUD) 2017 (Challita et al., 2017)
-S82 Towards Model-Driven Provisioning, Deployment, Monitoring, and Adaptation of Multi-cloud Systems IEEE International Conference on Cloud Computing (CLOUD) 2013 (Ferry et al., 2013)
-S83 Towards Trustworthy Multi-Cloud Services Communities: A Trust-Based Hedonic Coalitional Game IEEE Transactions on Services Computing 2018 (Wahab et al., 2016)
-S84 Multi-cloud provisioning of business processes Journal of Cloud Computing 2019 (Kritikos et al., 2019)
-S85 Evaluating the DevOps Reference Architecture for Multi-cloud IoT-Applications SN Computer Science 2021 (Bou Ghantous and Gill, 2021)
-S86 Hybrid IT and Multi Cloud an Emerging Trend and Improved Performance in Cloud Computing SN Computer Science 2020 (Gundu et al., 2020)